A Complete Bibliography of ACM Transactions on Design Automation of Electronic Systems

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

10 May 2024 Version 1.92

Title word cross-reference

1 [AGM01]. 2 [FWCL05, GH00, RL13, ZPLI23]. 2.5 [WCB15, WWCT18]. 3 [ADDM^{+}13, AJK^{+}21, CLT^{+}15, CBR^{+}22, CXR^{+}23, CWL^{+}22, DLC^{+}17, DLK24, DHZL23, JGM14, KK11, KKHK16, KLE18, LLKC13, LDD^{+}18, LDD^{+}19, LHZ^{+}06, LHC16, LW17, LS19, LS17, MAL23, OS03, OCK19, PSP24, PRKK21, PKC^{+}21, SKP21, SYX12, THM15, TMDF10, VILSL23, WYC10, WTW^{+}23, XGC^{+}20, YHH09, ZYS12, ZPLI23]. 4 [JCGP05]. ^{2} [BXG^{+}24, SJL23]. ^{3} [ZGB^{+}24]. dd [MLMM08]. DDX [SW04]. Fmax [PMB10]. g^{m} [LZ21]. GF(2^{m}) [RMPJ08]. H [CLT^{+}15]. I^{D} [LZ21]. k [CLH12, SSN22]. k/m [CHY05]. μ [DHZ^{+}11]. N [Pom16b, CLH12, Pom17a]. o(min(m, n)) [LM05]. t/t [CH13]. V_{t} [KOS09].

-Ary [CLH12]. -based [SW04]. -Cubes [CLH12]. -D [OS03, WYC10]. -Detection [Pom17a, Pom16b]. -Diagnosability [CH13]. -distinguishability [AGM01]. -domain [FWCL05]. -driven [MSD06]. -geometry [JCGP05]. -macrocell-based [CHY05]. -Matrix-Based [CLT^{+}15]. -VOR [SJL23].

/Nano [Kha23].

0.35V [ACF^{+}11]. 0.35V-Optimized

[ACF^{+}11].

1687 [IIEKS23].

2-stage [KSA^{+}10]. 2.0 [CLYP09, HWGY16, LLL^{+}18, ZZL^{+}23]. 2.5D [WTW^{+}23]. 2009 [GK09]. 252Kgates [CCC^{+}09a]. 252Kgates/4.9Kbytes [CCC^{+}09a].

36 [DHZ^{+}11]. 3D [LY24].

4.9Kbytes [CCC^{+}09a]. 40nm [ACF^{+}11]. 45-degree [CT13, TP08]. 45nm [BFL10].

71mW [CCC^{+}09a]. 7T [RM23a].

90nm [CFD^{+}16]. 9T [PS23].

A3MAP [JP12]. aberration [KPSW09]. absence [SPA^{+}03]. Abstraction [HZS^{+}19, LXGM23, CMNQ08, CLM^{+}10, HMB98]. abstraction/refinement [CLM^{+}10]. ABW [CIX15]. AC [MHA19]. Accelerated [CBR^{+}22, LD17, NHS23, RKKH24, SS24, XJF^{+}23, BHDS09, MLC08, RB19]. Accelerating [CXR^{+}23, CLX^{+}23, HW14, KZKAKP23, LS11, SKS12]. Acceleration [EJR22, GYZ^{+}22, LDP^{+}22, LY24, WFSS20, GPK^{+}09]. Accelerator [CBC22, FLG^{+}23, HLL^{+}24, HLW^{+}23, KP22, LCJ^{+}22, LYL^{+}19, LJJ^{+}22, LQL^{+}24, LPL^{+}21, OHA19, SKR^{+}22, SHBD21, SQL^{+}24, TWM^{+}23, WML^{+}24, WB^{+}24, AHL^{+}08]. Accelerator-rich [SHBD21]. Accelerators [CSO22, HJY23, LHC24, SYGC22, SV11, TL19, LSPC14, YLP^{+}13]. Access [BSP^{+}22, GSD^{+}18, HWDQ22, OKC08, PTPB22, RPR^{+}21, XYG^{+}16, Cha01, KLSP11, KCKG13]. Accesses [CLX^{+}23, KCKG16]. Accuracy [BH22, EAAK^{+}23, HSP^{+}22]. Accurate [DKZ^{+}15, LJ18, SV16, SKCM06, TWL16, TEK18, MFS09, RCD07, SGD10, XK97]. Accurately [CHA^{+}23]. Achieving [HSP^{+}22, KJT04, STL^{+}13]. ACM [GK09, BC08, CH10a, KLSZ09, QS11, SN10, CPX14]. acoustic [FIR^{+}97]. acquisition [NR03]. across [LBV^{+}06]. action [KC98]. Activation [WLM21]. Active [LKC^{+}18, VEO16]. Actively [PCT^{+}17]. Activity [GFJ16, KOO18, RG19, PR11, SXX^{+}06]. Actor [RGT^{+}14]. Actor-Oriented [RGT^{+}14]. Actuations [RB21]. acyclic [LKTD98]. AD [BXG^{+}24]. Adaptable [CRC15, KKK12, SHN12]. Adaptation [LYHL14, LCZ^{+}24, MDR15, RNA^{+}21, TZZH22]. Adapting [SSO16]. Adaptive [BM11, BYT22, CB17, CIX15, EW18b, JM14, KKHK16, LLKY13, LYSO19, LJJ^{+}22, LPY^{+}20, LIK22, SFM^{+}19, SJ23, SOS15, TZ17, WTR12, WQC^{+}16, ZLY^{+}15, CCYC14, CR12, CLQ12, DP04, FS13, HCK13, LMB^{+}12, LSL^{+}13, RL13, RAKK12, SCB01]. Adaptively [KLK^{+}17, DL11]. ADC [EO19]. ADCs [HWCL15, PKP^{+}03]. Add [LWZ^{+}19]. Adder [Brz24]. Adders [BH22, CXS^{+}23, EAAK^{+}23, KKK12]. Addition [BSP^{+}23]. Address [LP03, SR12]. addressing [SSP04]. Adjustable [LW21, KSA^{+}10, LLHT12]. Adjustment [MNMK^{+}21]. ADL [MSD06]. Admission [DZCD15]. ADMM [WTW^{+}23]. ADMM-based [WTW^{+}23]. Advanced [ATF^{+}23, MCY23, DDFR13]. Advances [CO18, JCPL23]. Advancing [LRHL24]. Adversarial [Ase23, BXG^{+}24, DZ24, FLG^{+}23, LYM^{+}20, LRHL24]. Aerial [HXB^{+}22]. Affine [WKL^{+}18, BC11]. after [XFJ^{+}16]. Against [ADB^{+}19, DZS^{+}18, LLQD23, LDX22, RNR^{+}21, AYS20, CYZL23, DFM15, GDTF17, HYK^{+}20, LQD22, ZLQ15]. AGENTS [dW97]. Agglomerative [LLLC13]. Agglomerative-based [LLLC13]. Aging [ADB^{+}19, DNT20, FYCT15, GC18, OT15, SJ23, TCW20, HTCP13]. Aging- [FYCT15]. Aging-Aware [OT15]. Aging-induced [TCW20]. Agnostic [BDBB19]. ahead [CSAHR07]. AI [APG24, CCY22]. Aided [HWF^{+}23, LJZ^{+}24]. AIMCU [ZXC^{+}23]. AIMCU-MESO [ZXC^{+}23]. Airgap [HS19]. algebra [GK07, GK09]. Algebraic [LAYZ23, ARLJH06]. Algorithm [DHVW18, GDPRG11, GYT12, HCRK11, HNS23, HLG^{+}15, JYHY21, KLSZ09, KLSZ11, MA16, MJB19, TZ17, YVC14, ZHC^{+}21, ZLG^{+}19, ZHJ^{+}23, BDB98, CD09, CT13, CSL^{+}07, CCW08, EK97, GBC07, JHL02, KT96, KL05, LM05, MBB01, MKBS05, MLMM08, MWG97, SCB01, SGJ96, VKKR02, XTW05, YMC^{+}13, YWW10, Zho08]. Algorithmic [AMO05, KRH18, LXWC20, RRHB21]. Algorithms [ACFM12, DK22, EWT23, GMN^{+}13, GdRJM21, SV16, SZB17, TCP97, WSY23, WCX^{+}24, Das04, Das09, EMO03, GMSSS02, JLF^{+}12, LKM04, LIA00, OWH08, PB14, PW99, TC98, YW09, YCHT00, ZSZ10, ZS02]. Aligned [LJJ^{+}22, SHL^{+}19, XYG^{+}16]. Allocating [KAKSP16, YHH09]. Allocation [ABC^{+}17, BK00, BM11, CET16, CARH18, KK14, KKLG15, LHC24, SCK18, ZYS12, AOC02, CLM^{+}10, CL99b, LCK^{+}09, SM00]. Alternative [KRL15, SYZ08]. among [DK08, LYSO19]. Amplifier [DMR23, RM23b]. Amps [AG22]. AMS [CVMP19, DDNAV04, MDM^{+}12, MPDG09, ZMS^{+}19]. Analog [ADB^{+}19, BBEM15, CFD^{+}16, CLC20, DZ18, GMS^{+}23, GPS^{+}24, HRC21, HSP^{+}22, LDP^{+}22, LYSO19, LS22, LLM^{+}23, LZ21, LHJ12, LCYN18, NL24, PTS^{+}20, SA24, SHD17, SCK^{+}23, STGR15, SOS15, TZ17, TZ20, WJYZ11, XAG^{+}20, ZGB^{+}24, ZSY18, BC05, DC07, DDNAV04, LON08, LFG^{+}09, LCKT12, LTPR^{+}13, ST99, SCJ01, WV02]. Analog-in-Memory [LDP^{+}22]. Analog/Mixed [GMS^{+}23, STGR15]. Analog/Mixed-Signal [STGR15, GMS^{+}23]. Analog/RF [BBEM15, PTS^{+}20, SA24]. Analyses [BFG17b, MCY23, YBM^{+}21]. Analysis [BS14b, Brz24, CZW^{+}03, CLT^{+}15, CB17, CXLL22, CH17, CYH19, CLMZ10, CYLC24, DKZ^{+}15, GD20, GLY^{+}12, HLZ^{+}22, HKL^{+}15, HHL14, HZJC23, JIR^{+}21, JM14, KM97, KOO18, Kha23, KC13, LJ18, LDLM20, LV14, MAS16, MHA19, NHS23, NSCM17, OM08, PLC24, PHKW12, Pie16, PEPP06, QBTM16, RRHB21, SMBT19, STWX12, SYH^{+}22, THT12, VTC20, WL12, XT16, ZFLS11, ZYW^{+}18, ZS16, ZKS^{+}16, ZMS^{+}19, ZBPF18, AC06, APB^{+}08, BWB14, BK10, CPR^{+}02, DCK10, Das04, DH06, FZKS11, GM08, GGBZ02, GDG^{+}08, IBMD07, JB98, JT98, KPR06, KVMH08, LWC07, LCHT02, LON08, LTPR^{+}13, MDG98, MFS09, MCMW08, NM13, QSK12, RMB10, ST99, VMP^{+}00, WYC10, YWGI09, ZHM07]. Analytic [AMM^{+}18, LFST21, JP12]. Analytical [HHL14, MA16, SV16, WTW^{+}23, XLL^{+}16, GG04, LON08]. Analyzing [CAP^{+}23, LH13]. Android [THC^{+}14]. Annealing [VLH04]. Annotating [BD05]. ANNs [ZMLH24]. Anomaly [LL19, VTC20]. ant [WGDK07]. anti [HTCP13]. anti-aging [HTCP13]. Any [JZG21]. Application [BH22, CGLH23, CYV^{+}14, DLK24, HKL^{+}15, HMMG^{+}20, HCZ^{+}16, HLL^{+}24, JBJ22, LPD^{+}17, LYHL14, LHF12, LF12, LIK22, MMM^{+}22, MDR15, RM23a, RCK^{+}15, STJG16, TCL14, VA17a, XLL^{+}16, XT16, YP10, ZYDP08, ZYPC17, CSC08, HLKN07, Hsi00, JCGP05, LM96, MMP00, MP07, SXZV13, WKR09, WSEA99, ZMTC13]. Application-aware [ZYDP08]. Application-Driven [YP10]. Application-level [HLL^{+}24]. application-oriented [Hsi00]. Application-Specific [HKL^{+}15, HMMG^{+}20, HCZ^{+}16, LPD^{+}17, LHF12, LF12, RCK^{+}15, TCL14, VA17a, CSC08, WKR09]. Applications [ACF^{+}11, BFV15, BLUS19, CLL^{+}22, EKEK22, ETAV18, EO19, HC17, HAB^{+}17, LFST21, LDLM20, MAS^{+}20, MS23, MLH^{+}17, NTSA18, PFHAH22, PMA24, RM23b, RS18, SBR^{+}17, SSK^{+}23, SVK17, SFM^{+}19, SLV^{+}22, SWT23, SESN15, WDZG16, WH20, ZLL^{+}16, CCC^{+}09a, DCK09, DCK10, DPNB02, DSH12, DVA02, HG07, KSS^{+}09, KCA04, KFH^{+}08, MHD^{+}04, NT05, PDN97, Ped96, SR12, VCLD03, VMP^{+}00, WLL^{+}11, WG11, ZHM07, ZAZ13]. Applying [CHBK15, WPR^{+}19]. Approach [CYZL23, CHK^{+}23, CJKS24, DY23, DZS^{+}18, DNT20, FG18, FMR23, GVJ15, HWL^{+}23a, HS19, JDLZ24, KRH18, LYL^{+}23, LHF12, LMA^{+}16, LTW^{+}16, MDR15, ORGD^{+}15, PGGD23, Pom18a, RRHB21, SHD17, SGGR14, SCK^{+}23, ZHJ^{+}23, ADS^{+}09, BD08, BMJ13, CBHK11, CHHL96, DDNAV04, DVA02, ETR07, GG04, GABP00, KSS^{+}09, KJKK03, LFG^{+}09, LCKT12, MSR09, MR96, NR01, SSP04, Vah02]. Approaches [HMMG^{+}20, KTKO13, LCOM07, Tes02, WAZ98]. approximability [BCC08]. Approximate [ADGSM22, EJR22, GT21, HWDQ22, JSS^{+}19, LKLC22, MED23, MHA19, NRDB19, OHA19, PMP17, WCX^{+}24, YBM^{+}21]. Approximating [GD20]. Approximation [BYT22, DHVW18, EKEK22, HWCL15, LNPL23, SYH^{+}22, HCS01, YWK^{+}03]. Arbiter [MMM^{+}22, NSCM17]. Arbitrary [WJG^{+}19]. Arbitration [AL19, IHM15]. Architecting [SABSA15]. Architectural [BRCS18, CXS^{+}23, KGS^{+}20, MA16, MLH^{+}17, APB^{+}08, CL99b, MSD06, VS12b]. Architecture [AJK^{+}21, BMdG17, CM20, CIB01, DK16, HSR^{+}24, HLG^{+}15, JP12, JYY^{+}22, LPLK22, LWZ^{+}19, LYL^{+}19, LJJ^{+}22, LYLW17, MD13, MSD06, MRL^{+}19, MS17, NGL^{+}21, PMT20, PCT^{+}17, SHBD21, SSL17, SJL23, TWM^{+}23, WKL^{+}18, WWCT18, WSY23, YKCG14, YMB15, YLP^{+}13, ZHL^{+}23, ZMLH24, CHY05, GM03, LCOM07, LTPT10, SCCH08, WTL^{+}13, XZC09, YBM^{+}21, ZYZ^{+}13, RJL^{+}09]. Architecture-aware [JP12]. Architecture-level [CIB01, LTPT10, WTL^{+}13]. Architectures [AMM^{+}18, CPS16, CBR^{+}22, CXR^{+}23, GADG19, GD22, HWX^{+}14, LM19, LLK^{+}14, RBWB20, VS12a, dONH23, ACT13, BD08, Cha01, CKAP07, CCL03, DP04, FS13, FRS97, GBK07, JBC^{+}10, JLF^{+}12, Kan06, KLSP11, LP03, LLKY13, LYCP13, OCRS07, PPDK09, QM12, WH05, ZM07, ZHTC09]. Area [EO19, HS18, HCW^{+}16, KKK12, KKLG15, SY07, SS14, TRM^{+}16, TCL14, Yan16, ZHJ^{+}23, DK08, GS00, HCS01, KL05, KNRK06, LC13, LCL08, MS00, SPMS02, SSP04, XPSE12, ZYZ^{+}13, ZHTC09]. area-array [LC13, LCL08]. Area-Aware [HCW^{+}16]. Area-Efficient [EO19, SS14]. Area-I [Yan16]. Area-I/O [Yan16]. Areas [WPR^{+}19]. Arithmetic [BSP^{+}23, PIK20, CCL03]. ARM [LLH^{+}17]. ARM-Based [LLH^{+}17]. ARM2 [HV98]. Array [CFD^{+}16, KCKG16, RBWB20, RB21, SPC^{+}15, AOC02, CZW00, LC13, LCL08, WV02, ZYZ^{+}13]. array-based [CZW00]. Array-Style [CFD^{+}16]. Arrays [HCW^{+}16, TRM^{+}16, WSY23, AC06, CH02, CD96, LMB^{+}12, PWY05, WAZ98]. Artificial [KAC^{+}23, WXH^{+}19]. Ary [CLH12]. ASIC [KLV15, THL^{+}13]. ASICs [PW99]. ASIPs [SM00]. ASP [YMB15]. ASP-Based [YMB15]. Aspect [HKJ^{+}23]. aspects [AMO05]. Assay [BTP^{+}20, LSCK20]. assembled [BC05]. assembly [AMR00]. assertion [BZ08, MPDG09, TBZ13]. assertion-based [TBZ13]. assertion-checker [BZ08]. Assertions [MDM^{+}12, WLM21]. Assessed [LLLL18]. Assessment [NPH^{+}20, RNR^{+}21]. Assignment [CK16, KLE18, LYCP17, LMS16, SV16, Yan16, Yan17, Yan20, BDB98, CCX06, CHH09, CPW04, CLYP09, KNDK96, Kuc03, LJV02, LCC11, LT11, VJBC07, WWG08, WLCJ09, XTW05, Yan11]. Assisted [CCMC20, GFJ16, HRC21, PTC^{+}15, SMBT19, SCK^{+}23, CSL^{+}07, MBB01]. Assistive [MVK^{+}18]. Assurance [XLY^{+}18]. Assured [JSS^{+}19]. Asymmetric [SBR^{+}17, RAKK12]. Asynchronous [PMS15, TB20, WWW^{+}12]. At-Speed [PTC^{+}15, TPC^{+}17, SXZV13]. ATM [RFYL98]. ATPG [HCC01, MT02, SGK08]. Attack [Ase23, BSP^{+}19, CYZL23, Che18, CYLC24, GLD^{+}22, JZG21, LTZ22, LLQD23, OK20, SQL^{+}24, YBM^{+}21, DDFR13]. Attacks [AYS20, CPK20, DZS^{+}18, DHB16, HYK^{+}20, JIR^{+}21, KD24, LSCK20, LYM^{+}20, LQD22, MLH^{+}17, PTPB22, RNR^{+}21, ZLQ15, LWK11]. Attempt [KR23]. Attention [YXG^{+}24]. Attestation [CRT19]. Attributed [PRCK08]. Augmented [VBP^{+}19]. Augmenting [TL19]. Authentication [HRK18, MPM^{+}17, YFT17]. Authorization [MPM^{+}17]. Auto [YCL^{+}23]. Auto-tuning [YCL^{+}23]. AutoDSE [SYGC22]. Autogenerated [APD^{+}11]. Automata [BZ08, PSD21, KT01]. Automata-based [BZ08]. Automated [BPTB17, IE12, KLV15, dONH23, GWR13]. Automatic [BFV15, CK96, CS22, CJLZ11, EWT23, GD20, GYZ^{+}22, MS08, SHD17, Shi20, SRTG19, WKR09, ADS^{+}09, KSS^{+}09, LFG^{+}09, TDE08, WWC04]. automating [HA05, RSR01]. Automation [ADB^{+}19, CH10a, CPX14, CO18, DZS^{+}18, DK22, FZL^{+}23, GHYR19, HHH^{+}21, JDD20, JCPL23, KLSZ09, KAC^{+}23, PSD21, SSK^{+}23, DTC^{+}09, LOC12]. Automotive [HK18, KPB19, LZZSV15, LMS16, MPM^{+}17, SRTG19, XLY^{+}18]. Autonomous [ML09, STL^{+}13]. Autotuning [MAL23]. Auxiliary [BDC08, CCQ98, Pie16]. Available [TEK18, dONH23]. AVB [DZK^{+}24]. Average [ZLW^{+}15]. Averaging [TWL16]. Avoid [WPR^{+}19]. Avoiding [AL19, HLG^{+}15, HGLC16, LLLL18, WSRH16, XPZ^{+}18, LYKW09]. award [GK09, QS11]. Aware [AKAKP18, BDBB19, BHY^{+}24, BLUS19, CMP10, CET16, CJKK19, DNT20, DZ18, FYCT15, GVJ15, HHK^{+}17, HC17, HXB^{+}22, HCW^{+}16, KPF16, KW16, KAC^{+}23, KPB19, LHW^{+}17, LLL^{+}18, LHK^{+}15, LZZSV15, LNG^{+}16, LMS16, MT15, OT15, PBZM19, RS18, RCK^{+}15, SBY^{+}20, SKP21, SCK^{+}23, SYX12, TBCH17, WSH^{+}18, WDD^{+}23, WLLH16, Yan20, YYG^{+}16, ZYPC17, ADP^{+}07, CHH09, CGV^{+}23, CLQ12, DHX^{+}23, DD02, ETR07, ENP20, FS13, GM08, GKM05, HJY23, JHL02, JDD20, JP12, JCS^{+}08, KPSW09, KJKK03, LC14, LKLC22, LWX^{+}23, LSZ^{+}21, LZ21, LG23, MAS^{+}20, MBD^{+}20, MJM11, MHQ07, MKW08, NWA^{+}24, OCK19, PSD21, PPDK09, PGGD23, RGM09, SSG12, SBC08, SRKS23, SMYH07, SKS12, SNL12, SWT23, TZ20, VGG19, WH05, WPHL08, WLL^{+}11, YB23, YXG^{+}24, YYLL09, ZYDP08, ZYP09, SGJN24]. awareness [RL13]. Ax [EJR22]. Ax-BxP [EJR22].

B* [WCC03]. B*-trees [WCC03]. back [CCK^{+}18, GABP00]. back-end [GABP00]. Backward [BS14b]. balanced [LLHT12]. Balancing [JIR^{+}21, MT15]. Band [WTR12]. Bandwidth [KLK^{+}17, BD08, GM03, LLKC13]. bank [CPW04, Kan06, SM00, Wu09]. banked [OK08]. Base [BSP^{+}19]. Based [APDC17, ALLE20, ANS^{+}20, ASAP17, AVG19, AKM^{+}22, AJK^{+}21, AAA15, Ase23, BHK17, BS14a, BD14, BHY^{+}24, CPS16, CCH^{+}15a, CAOM19, CLT^{+}15, CZZYW21, CXLL22, CYLC24, DLC^{+}17, ETAV18, EO19, GNGT21, GDTF17, GHYR19, HCL^{+}14, HWX^{+}14, HLG^{+}15, HC23, JHMGS18, JPHL16, JM14, KGS^{+}20, KC10, KLK^{+}17, KMO^{+}12, LZZ23, LLH^{+}17, LG18, LDLM20, LAYZ23, LZY^{+}23, LSZ^{+}24, LS11, LHK^{+}15, LLLL18, LH11, LPY^{+}20, LQD22, LGGJ14, LCC^{+}15, LKC^{+}18, LPL^{+}21, MNMK^{+}21, MCZ^{+}16, MA16, MS23, MCD12, NSP^{+}20, PIK20, PBH^{+}24, PSNC18, PG15, Pom17a, Pom18b, Pom20, PY20, QBTM16, QZZW24, RM23b, RS18, SV16, SMBT19, STGR15, TZ17, VEO16, WLZ^{+}19, WCB15, WQC^{+}16, WWCT18, WFSS20, WSY23, WC10, WL12, XS16, XCF18, YMB15, ZS16, ZHC^{+}18, AHAKP08, AM10, ADDM^{+}13, BLM00, BPRR98, BC11, BXG^{+}24, BBD00, BOC00, BH10, BZ08, CLM^{+}10, CNQ13, CGN96, CZW00, CFHM09]. based [CBR^{+}22, CH02, CBR^{+}05, CD96, CHY05, CFX09, CM13, CCL04, DP02, DCK09, DJP21, DDNAV04, DVA02, EMO03, EY12, FLG^{+}23, FS13, GK14, GG99, GPH^{+}09, GD20, GBC07, GDF09, GPK^{+}09, GH00, HWDQ22, HDZ^{+}20, HWF^{+}23, HZL^{+}22, HYK^{+}20, HZJC23, HCK13, HWCL13, HFMB20, HXZ^{+}23, HTC^{+}23, IIEKS23, IYF^{+}21, JZG21, JJH21, JLF^{+}12, KBN09, KZKAKP23, KK11, KLP^{+}24, KSD^{+}22, KNRK06, KSA^{+}10, LC13, LB00, LKM04, LWC07, LCC11, LWZ^{+}19, LJJ^{+}22, LJZ^{+}24, LHC24, LDK99, LZ21, LCHT02, LWG^{+}23, LOC12, LWK11, LLLC13, LXWC20, LYM^{+}20, LG23, MMM^{+}22, MP07, MS21, MLC08, NAK20, OM08, OHA19, OKC08, OK08, PSD21, PDN00, PRCK08, PMB10, PR09, Pom14b, RL13, RS98, SW04, SGK08, SWT23, SOC06, SC06, TFW24, TN99, TBZ13, VGG19, VILSL23, VKT02, WPR^{+}19, WH20, WTW^{+}23, WWC04, WC06, WPL23, WSEA99, XAG^{+}20, Yan00, Yan08]. based [YLY^{+}23, YYC09, ZHM07, ZZ24, ZGB^{+}24, ZHJ^{+}23, AA17, PBZM19, CCQ98, CH00, MW97, MHT14, MWG97, PBSV^{+}06]. Basic [AG22, VMP^{+}00]. Batch [LYL^{+}19]. Battery [MRL^{+}19, NSS^{+}16, Rak09, SKM^{+}16, CSAHR07, LCZ^{+}08]. battery-powered [CSAHR07]. Bayesian [BLR06, GPS^{+}24, PTS^{+}20, XJF^{+}23, ZGB^{+}24, ZGB^{+}23]. BDD [CCQ98, VKT02]. BDD-based [CCQ98, VKT02]. BDDs [BC16]. Beam [LZ17]. Behavior [CLMZ10, HXC^{+}18, RGT^{+}14, KRS06]. Behavior-Level [CLMZ10]. Behavioral [APD^{+}11, AA17, CLMZ10, KHP05, Sch17, TN99, WV02, WHRC12, Fuj05, HLKN07, KSS^{+}09, MRC06, VKKR02]. behaviors [BG01, KW02]. Benchmark [SA24, PSK08]. Benchmarking [JBC^{+}10]. Benders [ETAV18]. benefited [SLC^{+}22]. Best [GYZ^{+}22, GK09, QS11, SSCS10]. Best-Suited [GYZ^{+}22]. between [ATF^{+}23, CJKS24, Fuj05, YRH11]. Betweenness [SSN22]. Beyond [CPX14]. Biased [EKEK22, JCK^{+}18]. biasing [CFHM09]. BICS [RM09, RMB10]. BIFEST [LTH99]. Bifurcation [HHL14]. Binarized [BP23]. Binary [SV07, BCR^{+}08]. Binding [CET16, KK14, LHF12, ZLQ15, BD97, CLM^{+}10, CFX09, DS06, HLKN07, MKK13, MJM11, XK97]. Bio [BTP^{+}20]. Bio-chemical [BTP^{+}20]. Bio-IP [BTP^{+}20]. Biochemical [KGS^{+}20, RCK^{+}15]. Biochip [CPK20]. Biochips [CGLH23, GLD^{+}22, GHYR19, JYHY21, KGS^{+}20, KR23, LHC16, LSCK20, LCZ^{+}24, LKC^{+}18, MGR^{+}15, MWK21, PBWB21, PBF^{+}22, RCK^{+}15, RBWB20, RB21, SKS^{+}18, SOC06, SC06]. biomedical [APB^{+}08]. Bipartitioning [RTNL05, DPNB02]. bipolar [ZYZ^{+}13]. BIST [BBEM15, JNS^{+}17, LWC07, PKP^{+}03, PGB01, SSGS03]. Bit [HHK^{+}17, RM23a, LYCP13, NdLCR03, RMPJ08, RM09, RMB10, SBH^{+}06, VILSL23]. bit-width [LYCP13, SBH^{+}06]. Bits [SSO16]. Bitstream [HYK^{+}20, OK20]. black [LAS01]. BLAS [CCYC14]. BlOck [AG22, CM19, CCYC14, CCK^{+}18, DK16, ZLG^{+}19, KRS06, LPP00, MHD^{+}04, MS00, WCC03]. Block-level [CCYC14]. block-processing [LPP00]. Blockage [JD18]. Blockchain [CUA^{+}24, IK19, XRS^{+}19]. Blocked [EJR22]. Blocks [AFM14, JPM^{+}19, DK08, FLWW02, FLWC07, MHD^{+}04, MS00]. Blue [RGX^{+}24]. BNF [WWC04]. BNF-based [WWC04]. BoA [XJF^{+}23]. BoA-PTA [XJF^{+}23]. Board [MW97]. Board-level [MW97]. Boards [GDTF17, BPRR98, OW06]. body [CFHM09]. body-biasing [CFHM09]. bonding [WPL23]. BonnRoute [GMN^{+}13]. Boolean [PRCK08, BR12, BD97, BC11, CCQ98, GPK^{+}09, OK20, SGJ96]. BOOM [BSZ^{+}24]. BOOM-Explorer [BSZ^{+}24]. Boosting [CMNQ08, CSO22, XAG^{+}20, ZGB^{+}23]. borrowing [LCHT02]. Both [WH20]. bottleneck [NM13]. Bound [IIEKS23, JLJ15, HWF^{+}23, LC96, LTPR^{+}13, YWK^{+}03]. Boundary [Pom19a]. Boundary-Functional [Pom19a]. Bounded [CKKT98, LLLL18]. Bounded-skew [CKKT98]. bounds [TC98]. Boxes [HSR^{+}24, LAS01]. BoxRouter [CLYP09]. Brain [GNQ^{+}22, WSY23]. Brain-Inspired [WSY23]. Brain-network-inspired [GNQ^{+}22]. branch [CBHK11]. branch-and-cut [CBHK11]. Breaking [Che18, KSD^{+}22]. breakpoint [KRK98]. Breakpoints [KRK98]. bridges [LLQ^{+}03, EBR^{+}09]. bridging [LTH99, TCP97]. Broadside [Pom15a, Pom16a, Pom16c, Pom18b, Pom19a, Pom21a, Pom24b, Pom13, Pom14a, Pom14b]. BSP [SYHL14]. BTI [GC18]. BTI-Aging [GC18]. bubble [Yan00]. bubble-sorting-based [Yan00]. Budgeting [CXH^{+}16, STGR15, HLHT08, LCHT02]. Budgeting-Based [STGR15]. Buffer [LYLW17, MB04, SAL19, TCL14, WHRC12, CW01, FHHG12, JHL02, LLHT12, LT11, XTW05]. Buffered [OCK19, Yan16, CM08]. buffering [KRS06, KC13]. Buffers [CK16, SJN24]. Building [JDD20]. Buildings [ZHC^{+}18]. Built [EO19, IYF^{+}21, Pom13, Pom24a, SBB^{+}18, WCB15, LTH99].

Built-In [EO19, SBB^{+}18, WCB15, IYF^{+}21, Pom13, Pom24a, LTH99]. bump [DVA02]. bump-and-refit [DVA02]. Burst [CHBK15, CIX15]. Burst-Writes [CIX15]. Bus [GG99, Yan19, JWL^{+}03, LCOM07, LV02, OW06, SCJ01, YW09]. Bus-based [GG99]. Buses [Yan17, YGZ04]. Butterfly [LLQD23, ZYPC17]. BxP [EJR22]. Bypass [PMT20, YKCG14].

C [LWC18, RMPJ08]. C-Mine [LWC18]. C-testable [RMPJ08]. C2RTL [ZLL^{+}16]. Cache [AKM^{+}22, BFG^{+}19, CPS16, CAOM19, DJP21, GD22, GG04, HWX^{+}14, JZYZ15, JLK15, KLJ14, LYLW17, MACV14, Mit16, NTSA18, NAK20, SSS^{+}19, SABSA15, SMBT19, SJ23, SAL19, TYSF20, WDLD17, YPCF17, Giv06, JS13, LMW99, LSL^{+}13, PDN97, SLXZ12, TKVN07, TY97, VS12b, ZYDP08, NTSA18]. cache-coherence-enabled [LSL^{+}13]. Cacheline [PBL^{+}17]. Caches [CK19, CB17, SJN24, SYX12, CXK^{+}13, LSDV10, ZP08]. Caching [WQC^{+}16, HCK13]. CAD [BSP^{+}19, HAW20, KLSZ09, KLSZ11, LZR23, LYM^{+}20, NPH^{+}20, NSP^{+}20, PLH^{+}24, SB98, Vah02]. CAD-Base [BSP^{+}19]. CAD/EDA [LZR23]. calculation [RCD07]. Calibration [CCMC20, PMB10]. Call [Ano13, CH10a, Ped11, KLSZ09]. CALM [ZYPC17]. Cameras [YMB15]. Camouflaged [WCZ^{+}24]. Camouflaging [ISK21]. CAN [LMS16]. Cancellation [LTYW12, FIR^{+}97]. Cap [HC17, YLY^{+}23]. Capability [EW18b]. Capacitance [XLS15, YLY^{+}23]. capacitive [LXCH04]. Capacitor [HWCL15, HWCL13]. Capacitors [SCK18]. Capture [PTC^{+}15, XCW12, Xia24]. Carbon [WSH^{+}18]. Carbon-Nanotube [WSH^{+}18]. Care [DY23, TPC^{+}17]. cares [CBMM10, SGK08]. Carlo [FZL^{+}23, GLY^{+}12, ZFL22]. Carrying [IPWW17]. CASCA [DZS^{+}18]. Cascade [YYL^{+}15]. Case [APDC17, CH17, LLP^{+}16, LYM^{+}20, RCW22, RPR^{+}21, DCC^{+}23]. Cases [LWC18, KFH^{+}08]. Causal [CBC22]. Cause [PLC24]. caused [SHLL98]. Cayley [CCH15b]. CBDC [LLQD23]. CBDC-PUF [LLQD23]. CCM [TWL16]. CDTA [YFT17]. Cell [ACF^{+}11, CZZYW21, DBK^{+}18, JZYZ15, KRL15, RM23a, TRM^{+}16, WPR^{+}19, WC10, XNZ^{+}15, JCS^{+}08, KBN09, LCZ^{+}08, MRB^{+}11, MS00, RS03, SSCS10, dW97]. Cell-based [WPR^{+}19]. Cells [CYLC24, HWGY16, JCK^{+}18, MJB19, SKM^{+}16, GH00, TS96]. Cellular [PSD21, KT01]. CeMux [BH22]. Centralised [CK19]. Centrality [SSN22]. Centralized [ZHC^{+}23]. Centric [WGSH16, XLNB17, WB^{+}24, ZHOM08]. Centroid [WLLH16, HWCL13]. CGRA [KZKAKP23, WB^{+}24]. CGRCA [JDLZ24]. Chain [BSP^{+}19, CUA^{+}24, LHC16, LLQD23, Pom17b, RNR^{+}21, SLP^{+}19, XRS^{+}19, YFT17, YSF^{+}18, YFT18, YBS^{+}18, GKM05, RMKP03, TYH08, WPHL08]. chained [KC13]. Chains [Pom16b]. Challenges [BRCS18, MRL^{+}19, XLNB17, Ped11, RBA^{+}12]. Change [JSA18, LLP^{+}16]. changes [LG12]. Changing [MMM^{+}22]. Channel [BDBB19, CGLH23, DZS^{+}18, JM14, KD24, LSZ^{+}24, LQD22, PPP^{+}15, ZBPF18, CYZL23, FLWC07, HSA^{+}04, LLKY13, LM21, NPH^{+}20, Yan00, YCHT00]. Channels [BSP^{+}22, GNGT21, JLJ15, DSKB04]. Chaotic [CSC^{+}21]. Characteristics [CFD^{+}16, DHZL23, JLF^{+}12]. Characterization [KRL15, MMM^{+}22, SRC15, BW00, JCS^{+}08]. Charge [VA17b]. Chassis [APD^{+}11]. check [CL13, YCHT00]. checker [BZ08]. checkerboard [GC96]. Checking [AA17, KW16, ZZL^{+}23, AGM01, BK10, CNQ13, Fuj05, HMB98, KMS12, YWGI09]. Chemical [LTW^{+}16, BTP^{+}20]. Chief [Ano13, Hu20]. Chip [ADB^{+}19, ALL17, BHK17, BD14, BDBB19, CK19, CM20, FHL^{+}23, GADG19, GSD^{+}18, HAB^{+}17, HZS^{+}19, IHM15, JLJ15, JNS^{+}17, JZYZ15, JGM14, KBV^{+}15, LDD^{+}18, LDD^{+}19, LW17, PMT20, PGCB16, SCK18, SMBT19, STWX12, SGGR14, WLT08, XS16, XCF18, Yan16, YKCG14, ZHC^{+}21, ZYS12, ZYPC17, AYM05, APB^{+}08, ADS^{+}09, BMJ13, Cha01, CKAP07, CSC08, CXK^{+}13, CBR^{+}05, CCL04, DNT20, HDL^{+}12, JP12, KP13, KYN^{+}12, LCOM07, LLKY13, LLKC13, LH13, LC13, MD13, NR03, OM08, PLH^{+}24, PDN00, PPDK09, PTC05, Pom24a, TDE08, WM24, WDC^{+}22, WDLX21, Yan11, YLP^{+}13, ZSZ10, ZMTC13, ZM07, WLL^{+}11, AHL^{+}08]. Chip-Multiprocessors [HAB^{+}17]. chip-package [LC13]. Chip-to-Chip [GADG19]. Chipless [YBS^{+}18]. Chips [CCY22, HCZ^{+}16, LWX^{+}23, SOS15, GNQ^{+}22, HGBH09, VS12a]. Chisel [FMR23]. choice [SBGD13]. choose [DNA^{+}12]. CHSM [CUA^{+}24]. ciphers [LWK11]. circadian [GS13]. Circuit [ADB^{+}19, BBEM15, BZWZ17, BFL10, CM18, CM19, CZZYW21, FZL^{+}23, GBR07, GDTF17, GPS^{+}24, HS18, HRC21, HHX^{+}23, HS19, JK10, LYSO19, LH11, LQD22, RJBS09, SS24, SA24, SMYH07, Shi20, SCK^{+}23, TWL16, WSH^{+}18, WKC12, ZGB^{+}24, ZFL22, ADM^{+}13, AJM13, BDB98, CSC08, CBMM10, CSX^{+}05, DL11, GMSSS02, HRP00, LLQ^{+}03, OW06, RCD07, SPMS02, YH97, YMC^{+}13]. Circuit-Averaging [TWL16]. Circuit-simulated [SMYH07]. circuit-switched [CSC08]. Circuits [BJX15, GPS^{+}24, HDB22, HWL^{+}23a, HZL^{+}22, JZG21, KKS16, LD17, LSZ^{+}21, LS22, LLM^{+}23, LZ21, NL24, PB12, Pom16b, RGM15, SHD17, SCK^{+}23, WTR12, WCX^{+}24, XAG^{+}20, ZSY18, ZHJ^{+}23, BLM00, BLR06, BC05, BASB01, CSKR05, CLLK06, CACS05, Che96, CPR^{+}02, DC07, DD02, EMO03, HVF^{+}01, HH09, HWCL13, KJKK03, KOS09, KVMH08, LH09, LON08, LFG^{+}09, LTPR^{+}13, NS03, PL98, PSK08, PR98, PR09, RTNL05, SNH02, ST99, WV02, ZCG06, SSCS10]. Clamp [VEO16]. class [SB98]. Classification [GAT^{+}21, MS17, VNS19, RAKK12]. Classifiers [ALL17]. cleaning [JS13]. client [dW97]. client-server [dW97]. Clip [HWF^{+}23, GH00]. Clock [EK16, HN07, HYN15, KK14, KK11, KKS16, LLL^{+}18, LNG^{+}16, LT11, LS17, OCK19, TCW20, UE22, WCCC14, WKC12, WWW^{+}12, BDM^{+}99, BDB98, CGN96, CM08, CHH09, CKKT98, GHW^{+}12, GWR13, HTCP13, LLHT12, LLLC13, PL98, SSGS03, TDF^{+}09, wATkK02]. Clock-Aware [LLL^{+}18]. Clock-Gating [WKC12, BDM^{+}99]. Clock-Tree [KKS16]. Clock-Tree-Aware [LNG^{+}16]. clocked [BD00]. Clocking [BPTB17, MR05]. Cloning [JNCS19, Vah99]. Close [Pom18b]. Close-to-Functional [Pom18b]. Closed [CW01]. closure [LC14, YYC07]. Cloud [BD14]. Cluster [CM19, DD02, LJV02, SB98, KJR^{+}07, LWC07]. Cluster-aware [DD02]. Cluster-cover [SB98]. Clustered [CMP10, GBK07]. Clustering [HWF^{+}23, VILSL23, XLL^{+}16, CC06, HLCH07, MLMM08, SPMS02]. clusters [OWH08]. CMAPS [Hsi00]. CMOS [ACF^{+}11, ADB^{+}19, CFD^{+}16, GH00, LTH99, PHKW12, WSS^{+}18]. CMP [CXK^{+}13, WGSH16, ZHL^{+}23]. CmpCNN [ZHL^{+}23]. CMPs [CAOM19, SYX12]. CNN [LCJ^{+}22, LHC24, LYM^{+}20, MS23, TWM^{+}23, TZZH22, VFML23, YLY^{+}23, ZHL^{+}23]. CNN-Based [MS23, LYM^{+}20]. CNN-Cap [YLY^{+}23]. CNNFlow [NM23]. CNNs [PRKK21, WDD^{+}23]. Co [CVMP19, CBR^{+}22, Hua01, JSS^{+}19, LCG^{+}22, PLC24, PGGD23, SKM^{+}16, WWFT12, ZHC^{+}23].

Co-design [CBR^{+}22, ZHC^{+}23]. Co-optimization [LCG^{+}22]. Co-scheduling [PGGD23]. Co-Simulation [SKM^{+}16, WWFT12, CVMP19]. Co-synthesis [Hua01]. Co-Training [JSS^{+}19, PLC24]. coarse [KLSP11]. coarse-grained [KLSP11]. cocurrent [KI01]. Code [AMR00, AM98, CL99a, FHHR21, MLH^{+}17, TAP^{+}24, TY97, BH10, DHV^{+}00, KMS12, KNDK96, KH10, LP03, LB00, LKTD98, LDK99, OKC08, SR12, SBH^{+}06, SM00, VMP^{+}00, VLGG01]. Code-Injection [MLH^{+}17]. code-motion [DHV^{+}00]. codes [RM09, WHXZ13]. Codesign [BM11, CMM00, FIR^{+}97, GABP00, GGB97, HKL^{+}07, SCV06]. Coding [WZL^{+}21]. Coefficient [APDC17]. Coexistent [BDBB19]. Coffeee [RJL^{+}09]. Cognition [HXC^{+}18]. Coherence [GD22, HWX^{+}14, LSL^{+}13, ZYDP08]. coherency [VS12b]. Collection [GSD^{+}18, HCL^{+}14, ZLW^{+}15]. Collection-Induced [GSD^{+}18]. colony [WGDK07]. Coloring [ZLY^{+}15, CML98]. Combinational [CD96, HWL^{+}23a, LD17, EMO03, KT96, KOS09, PR98, RJBS09, TN99]. Combinatorial [AM05, VLH04]. Combining [ETAV18, LFST21, SPG^{+}08]. CoMETC [ANR13]. Commercial [LRHL24, MPDG09]. Commercial-quality [LRHL24]. Common [DHB16, LWC18, WLLH16, ZYZ^{+}13, HWCL13]. Common-Centroid [WLLH16]. common-centroid-based [HWCL13]. Common-source-line [ZYZ^{+}13]. Communication [CARH18, KPF16, SRTG19, YP10, ADS^{+}09, GBK07, GG99, LCOM07, MOZ06, PPDK09, PBSV^{+}06, ZM07]. Compact [LJ18, MAS16, PBH^{+}24, SYH^{+}22, WTR12, XCW12, HVF^{+}01, YHL07]. Compacting [PL03]. Compaction [Pom15a, Pom15b, Pom20, EMO03, MHD^{+}04, TBZ13, XLCL13]. Comparative [Brz24, Kha23, MLG12, PB14]. Comparing [VGG19]. Comparison [SA24]. Comparisons [PKC^{+}21]. compatible [SGK08, WWC04]. compensation [CFHM09]. Compilation [SFM^{+}19, SBH^{+}06, YHL07, KLSP11, MSR09, VLGG01]. Compile [KNRK06]. Compile-time [KNRK06]. compiled [PHM00]. Compiler [HTC^{+}23, LHS20, LPD^{+}17, LLHT03, SMBT19, SYHL14, WKL^{+}18, XPSE12, BD08, GGDN04, HG07, KRS06, SSG12]. Compiler-Assisted [SMBT19]. compiler-directed [HG07]. Compiler-in-the-loop [XPSE12]. Compilers [YLL06]. Compiling [Edw03]. Complementary [CYLC24, QSW^{+}15]. Complementation [Pom15a]. Complete [PDS12, AGM01]. complete- [AGM01]. completeness [LLYW10]. Complex [WTR12, TYH08]. Complex-Valued [WTR12]. Complexity [ASAP17, AL19, LTYW12, WYC10, BCC08, YCCG03]. Compliance [HC18, BGM04]. Component [HWL^{+}23b, LH14, PG15, RSR01]. Component-Based [PG15]. Component-Composition [LH14]. Composable [VGG19, WTL^{+}13, HGBH09]. Composition [LH14, AG22]. Compositions [NSCM17]. compound [FLWC07]. Comprehensive [DSHD23, GSFT16, JNS^{+}17, PTPB22, SA24, SSK^{+}23, YFT17, ZBPF18]. Compress [XCW12]. Compressed [PBL^{+}17]. Compression [BLNK14, EK16, Xia24, BH10, JCS^{+}08, LCT03, LDK99, NT05, OKC08]. Compressors [SMS22]. CoMPSoC [HGBH09]. Computation [BFG17a, CV17, CARH18, EJR22, FHL^{+}23, IIEKS23, KCKG16, KS23, MOZ06, Pom17a, BLM00, GMSSS02, HLCH07, HW00, Kag05, WYIG07, YH97]. Computational [BCC08].

Computations [CBR^{+}22, CXR^{+}23, ENP20, ARLJH06, LPP00, PGB01]. Compute [HJY23, LPL^{+}21, TCP97]. Compute-in-Memory [HJY23]. Compute-in-Memory-Based [LPL^{+}21]. Computer [MFHP12, CSL^{+}07, MBB01]. computer-assisted [CSL^{+}07, MBB01]. Computing [BMdG17, BXG^{+}24, CDB11, HHX^{+}23, HXZ^{+}23, JSS^{+}19, LY24, MHA19, NRDB19, SN10, WLH20, XGC^{+}20, YBM^{+}21, ZXC^{+}23, CLQ12, LC96, NR01]. Concept [AM10]. Concept-based [AM10]. Concolic [WJM24]. Concurrency [SSG12, Sen11]. Concurrency-aware [SSG12]. Concurrency-oriented [Sen11]. Concurrent [SOC06, WH20, Edw03, EY12, HCLC98, LC13, RBA^{+}12]. Conditional [CLH12, CCH15b, KW02]. Conditionally [CSC^{+}21]. conditions [HN07, YH97]. Confidence [JT98]. Configurable [EAAK^{+}23, LSPC14, LLQD23, BD08, LCD07, SPG^{+}08]. Configuration [WCZ^{+}24]. Configurations [HABS15, BHS11]. Conflict [GSD^{+}18]. Congestion [RGM15, SYL09, SAHF^{+}20, XXC^{+}24, YWK^{+}03, ZPLI23, LCJ^{+}10, RL13]. Congestion-Free [RGM15]. Connected [LJZ^{+}24]. connection [Yan11]. connections [YCCG03]. conquer [HPK99, SW12]. Conscious [LLP^{+}16]. Consecutive [Yan17]. Consideration [JD18, LYLW17, WPL23]. considered [HN07]. Considering [BHLG19, CCK^{+}18, GC18, JOH17, WCCC14, KPR06, LH13, LTPR^{+}13]. Consistency [CJKS24, YP10]. Consolidated [HC17]. Constant [CHC^{+}16, GYT12]. Constant-Cost [CHC^{+}16]. Constrained [LLM01, LLLL18, NRM^{+}24, PBF^{+}22, RKKH24, Yan18, BG01, GOC02, LSDV10, MMP00, NG06, NR01, OKC08, SCB01, WG11, WLH20, WLCJ09, XPX^{+}21, YWW10, ZHOM08]. Constraint [KKK12, MRMP08, RS18, VMP^{+}00, YRH11, Das09, PR96, TP08]. Constraint-Based [RS18]. Constraint-driven [MRMP08]. Constraints [CLC20, DBK^{+}18, Kuc03, MN17, Pom16a, Yan17, BD05, CSAHR07, Hua01, QS09, SSP04, wATkK02, VLH98, WWG08, ZAZ13, ZW98]. Constraints-driven [Kuc03]. Constructed [ZXC^{+}23]. Constructing [DSRV02, JZYZ15]. Construction [DLK24, EK16, HGLC16, LLLL18, CM08, LH09, LYKW09, Yan08, ZCG06]. Constructive [LYL^{+}23]. Consumption [Brz24, FG18, Kan06, TKVN07]. Contact [YLZ^{+}17]. Contact-Hole [YLZ^{+}17]. Containing [WWW^{+}12, LAS01]. Content [HHK^{+}17, RB19, MLC08]. Content-Aware [HHK^{+}17]. content-based [MLC08]. Contention [CHA^{+}23, DJP21, KLJ14, ZYPC17]. Contention-Aware [ZYPC17]. Context [RG19, BDC08, JHL02]. context-aware [JHL02]. context-triggered [BDC08]. Context-Varying [RG19]. Contiguous [KKLG15]. Control [AVG19, BDB12, BYT22, CGLH23, CS22, FHHR21, GDD21, JDD20, JK10, LDP^{+}22, LJJ^{+}22, MAS^{+}20, PIK20, PCT^{+}17, QSW^{+}15, SS24, VGG19, ADDM^{+}13, BMJ13, CXK^{+}13, CR12, FRS97, KSA^{+}10, MWG97, OM08, SHLL98, ZAJ^{+}12]. control-dominated [FRS97, MWG97]. Control-Flow [FHHR21]. Control-system [CGLH23]. Controlled [TRM^{+}16, DL11]. COntroller [KMR18, SSL17, GF06, HMLL11, LC14]. Controllers [LVS16, PDS12, BDM^{+}99, Fuj05, NCP01]. Controlling [KYL16]. controls [YHL07]. conversion [ZLL13]. Converter [FZL^{+}23, SGGR14, WDC^{+}22, ADS^{+}09]. Converters [SBB^{+}18, TWL16, WGT^{+}17, JR97]. Convolution [CLX^{+}23, HLW^{+}23].

Convolutional [CHK^{+}23, DCC^{+}23, MNMK^{+}21, NM23, NGL^{+}21, YLY^{+}23]. cooling [ANR13]. Cooperation [ATF^{+}23]. Cooperative [LHF12]. cooptimization [ZLL13]. Coordinated [ANR13, DJP21, GGDN04]. COPE [DJP21]. coprocessor [GDTG07]. coprocessors [SCV06]. Core [CAOM19, CYH19, ETAV18, KD24, LHLP16, SBY^{+}20, SESN15, WMT^{+}16, WDC^{+}22, WDLX21, CCL04, GD22, LBV^{+}06, LG23, RAKK12, SEN05, SZV^{+}12, XZC09]. core-based [CCL04]. core-external [XZC09]. Cores [RKKH24, SFM^{+}19, WGSH16, GG04, LV02, SSGS03, XZC09]. CoreSight [LLH^{+}17]. Corner [KQP^{+}19, MHD^{+}04, Meh98]. Corners [GPS^{+}24]. correct [ADS^{+}09]. Correcting [PGCB16]. Correction [DZ18, RM09, WHXZ13]. Correlated [SCL^{+}22, SXZV13]. Correlations [LYSO19]. cosimulation [FLPP09]. Cost [ABC^{+}17, CHC^{+}16, JPHL16, LSZ^{+}24, MHT14, MJB19, QS09, BPRR98, BWB14, Giv06, HCK13, JDLZ24, LG12]. Cost-Effective [JPHL16, MHT14]. cosynthesis [Hsi00, Wol96]. Counterfeit [YFT17]. Countermeasure [HYK^{+}20, OK20]. Countermeasures [CPK20, DZS^{+}18]. Counting [PB12]. coupled [LMB^{+}12]. Coupling [LDD^{+}19, KJKK03, LXCH04, SKCM06]. coupling-aware [KJKK03]. covariance [KPR06]. cover [SB98]. Coverage [AKAKP18, CYV^{+}14, CM13, IE12, Pom22, XAG^{+}20, DSH12, FZKS11, GF06, Sen11, SDP^{+}09, TCP97, WPHL08, WPR^{+}19]. Coverage-Directed [IE12, CM13]. Coverage-Driven [CYV^{+}14]. Covering [BZWZ17, Pom21a]. CoVerPlan [DSHD23]. Covert [GNGT21, KD24]. CPU [LG23, SEN05, ZBPF18]. CRA [LLH^{+}17]. Crash [WL12]. Creation [NRZ^{+}18]. criteria [CGN96]. Critical [AKAKP18, BSP^{+}22, FYCT15, GC18, IGN18, KMR18, LC14, STJG16, XGWL24, ETR07, HKB^{+}07]. Critical-path-aware [LC14, ETR07]. Criticality [BB17, CV17, CYH19, SZB17, ZABGZ17]. Cross [APG24, EKEK22, KD24, VBP^{+}19, WFT^{+}19, XNZ^{+}15]. Cross-Core [KD24]. Cross-layer [EKEK22]. Cross-level [VBP^{+}19]. Cross-Point [XNZ^{+}15, WFT^{+}19]. Cross-Stack [APG24]. Crossbar [BXG^{+}24, LHC24, XGC^{+}20, THL^{+}13]. Crossbar-based [BXG^{+}24]. crossbar-switch [THL^{+}13]. crossing [SW99]. Crosstalk [LWH06, LDX22, HR06, JPCJ06, LCC11, MCMW08, Mut09, ZW98]. crosstalk-driven [JPCJ06]. Crosstalk-Induced [LDX22]. CRP2.0 [ATF^{+}23]. Cryptographic [LQD22, DP04]. Cubes [CLH12, WC10]. Cubic [HWL^{+}23b]. cuboidal [WYC10]. Current [CH10b, MN17, PS23, WLLH16, HLCH07, HCN09]. Current-Ratio [WLLH16]. Custom [HRC21, KAKSP16, LW17, LSZ^{+}24, LHF12, LF12, TDF^{+}09, AMR00, HMVG13, TS96]. Customizable [LIK22, MPSJ07]. customization [CBMM10, MKK13, MSB^{+}09, YLP^{+}13]. Customized [PSP24]. Cut [SHL^{+}19, CBHK11]. Cutting [LVS16]. Cyber [CXLL22, SKM^{+}16]. Cyber-Physical [CXLL22, SKM^{+}16]. Cyberphysical [PGCB16]. Cycle [BHY^{+}24, LVS16, LS11, WZH^{+}23, Das04, Pom14a]. Cycle-Level [LS11]. cycled [JSG09]. Cycles [KAKSP16]. Cyclic [BR12, Che18].

D [CWL^{+}22, GH00, WCB15, ADDM^{+}13, AJK^{+}21, CLT^{+}15, CBR^{+}22, CXR^{+}23, DLC^{+}17, DLK24, DHZL23, JGM14, KK11, KKHK16, KLE18, LLKC13, LDD^{+}18, LDD^{+}19, LHZ^{+}06, LHC16, LW17, LS19, LS17, MAL23, OS03, OCK19, PSP24, PRKK21, PKC^{+}21, RL13, SA24, SKP21, SYX12, THM15, TMDF10, VILSL23, WYC10, WWCT18, WTW^{+}23, XGC^{+}20, YHH09, ZGB^{+}24, ZYS12, ZPLI23]. D-enabled [LDD^{+}19]. D-ICs [LS17]. D-NoC [ADDM^{+}13]. D-Stacked [SYX12]. DAG [SRKS23, WJG^{+}19]. DAGSizer [CHK^{+}23]. daisy [KC13]. daisy-chained [KC13]. DANCE [LCG^{+}22]. Dark [HAB^{+}17]. DARP [CRC15]. DARP-MP [CRC15]. Data [CPS16, CCMC20, DZCD15, FHHH22, JLK15, KW16, LWC18, LL19, NTSA18, NM23, OHA19, PCD^{+}01, Pom16c, PAV17, PA21, SPC^{+}15, SUC01, TYSF20, TZZH22, VTC20, WDD^{+}23, WB^{+}24, XCW12, XPZ^{+}18, BHW^{+}13, BK00, BWB14, BHS11, FWCL05, GFC^{+}09, GMN^{+}13, GDF09, IBMD07, JCS^{+}08, KMS12, KI01, KCA04, LS23, LSPC14, LCT03, Meh98, NR03, PDN97, PDN00, PGB01, RMKP03, SM00, VCLD03, YGZ04, LCG^{+}22]. Data-centric [WB^{+}24]. data-dominant [VCLD03]. Data-Driven [DZCD15, LL19]. Data-flow [FHHH22]. data-flow-driven [KMS12]. DAta-Network [LCG^{+}22]. Databases [HCL^{+}14]. Dataflow [ASAP17, BMdG17, BLUS19, BFG17b, BFG17a, CH17, HPB11, JJH21, JOH17, LFST21, SFM^{+}19, SS14, WML^{+}24, HKB^{+}07, MHF96, MB04]. Dataflows [LPLK22]. Datapath [JR97, PIK20, CL99b, GDTG07, MR05, XPSE12]. datapaths [Fuj05, GK07, GK09, NCP01]. DC [CFD^{+}16, SBB^{+}18, TWL16, WGT^{+}17, WDC^{+}22]. DC-DC [WGT^{+}17, WDC^{+}22]. DCM [TWL16]. DCW [WLZ^{+}19]. DDAM [WDD^{+}23]. Deadlines [ENP20, WJG^{+}19]. deadlock [LM05, TDE08]. deadspace [SY07]. Debug [EW18b, LHLP16, HW14]. Debugging [Ali12, BHK17, RPKC05]. Decade [XFJ^{+}16]. decap [LCL08]. Decision [CWL^{+}22, HZL^{+}22]. decode [TKVN07]. Decoder [CAP^{+}23, SJL23, CCC^{+}09a]. decoders [KHW06]. Decoding [CWL^{+}22].

Decomposition [ETAV18, GBR07, HWDQ22, HCW^{+}16, KHW06, LZ17, RFG20, YLZ^{+}17, ZLY^{+}15, ZGB^{+}24, CHHL96, CH00, EMO03, LM96, WSEA99]. Decomposition-based [ZGB^{+}24, EMO03]. Decompression [PBL^{+}17]. Decoupled [DMR23]. Decoupling [SCK18, XLS15]. deduction [DP02]. Deep [Ase23, CLL^{+}22, EJR22, HZJC23, HLX^{+}23, KZKAKP23, KLP^{+}24, LYL^{+}19, LJZ^{+}24, LCZ^{+}24, LPL^{+}21, NHS23, PSP24, RNA^{+}21, SKR^{+}22, SQL^{+}24, UPV23, ZHC^{+}23, ZBG^{+}23]. DeepFlow [APG24]. Defect [XAG^{+}20, ACT13, JT98]. defect-level [JT98]. Defective [PB12]. defects [XLCL13]. Defending [YFT18]. Defense [BXG^{+}24, GLD^{+}22, LDX22]. deficiency [ZCG06]. Defined [JHMGS18]. Definition [BC16, Pom15c, ZLG^{+}19, CCC^{+}09a, VCLD03]. Deflection [LLKC13]. Deformable [CLX^{+}23]. Degraded [SLC^{+}22]. degree [CT13, TP08]. Delay [CLC^{+}24, EAAK^{+}23, FYCT15, JLJ15, JK10, JOH17, LW21, LLQD23, MCD12, STJG16, XCW12, ZK15, BDB98, CFHM09, GS00, GMSSS02, HR06, KJKK03, LLHT12, MT02, MKW09, PT06, PMB10, PR98, PR96, RCD07, SC00, SSP04, TD03, WVYG99, XLCL13, XPSE12, YH97, YHL^{+}11]. Delay-Adjustable [LW21]. delay-area [XPSE12]. Delay-Fault [LW21]. delay-sensitivity-based [PMB10]. Delayed [SJ23]. Delivery [CAP^{+}23, XLS15, ZFLS11, ZLL13]. Demand [AAA15, PBF^{+}22, SKS^{+}18, WQC^{+}16]. Demand-Based [WQC^{+}16]. Demand-Driven [PBF^{+}22, SKS^{+}18]. demonstrable [JW08, LP07]. Dense [BYT22]. Density [RM23b, FLWC07, OWH08, ZYP09]. dependence [DH06]. Dependencies [BR12]. dependent [BLM00]. depth [CH00, LH09, ZCG06]. depth-optimal [CH00]. depth-size [LH09]. derive [GS00]. derived [CACS05, Zho08]. Describing [RHA08]. description [MSD06, PHM00, SSG12]. descriptions [Fuj05, MWG97]. Design [ADB^{+}19, ABC^{+}17, AFM14, BJX15, BSZ^{+}24, BH22, BS14a, BZWZ17, BS14c, BSP^{+}23, BHLG19, CK19, CD09, CH10a, CH10b, CPX14, CHC^{+}16, CYZL23, CSC^{+}21, CRC15, CGLH23, CO18, DZS^{+}18, DK22, DNT20, DHB16, EAP17, FZL^{+}23, FHL^{+}23, FLG^{+}23, FMR23, FCZ^{+}23, GACK22, GdRJM21, GCZ^{+}15, GHYR19, HCRK11, HXB^{+}22, HMMG^{+}20, HLG^{+}15, HHH^{+}21, HKJ^{+}23, ISK21, JCPL23, JDLZ24, JWL^{+}03, JLK15, KKLP15, KGS^{+}20, KO23, KP22, KLSZ09, KLSZ11, KLV15, KKS16, KAC^{+}23, KSD^{+}22, LLP^{+}16, LW17, LJJ^{+}22, LF12, LHK^{+}15, LZZSV15, LQD22, LLQD23, LPL^{+}21, MED23, MYSZ23, NWA^{+}24, OT15, OHA19, PSD21, PLH^{+}24, PMT20, PKC^{+}21, PDS12, PBH^{+}24, Pom14a, Pom16a, Pom18a, PS23, RFG20, RS18, SSK^{+}23, SMBT19, Sch17, SBY^{+}20, Shi20, SDP^{+}09, SGGR14, SHBD21, SYGC22, SHN12, SESN15, SYX12, STGR15, SCL^{+}22, TYSF20, TCL14, VGG19, VILSL23, VA17a, VEO16]. Design [WWCT18, WPR^{+}19, WS22, WDC^{+}22, WSS^{+}18, WPL23, XPX^{+}21, XLS15, XNZ^{+}15, YPCF17, YD16, ZLG^{+}19, ZGB^{+}23, ZYS12, ZZL^{+}23, ACT13, AHL^{+}08, APB^{+}08, AMM^{+}06, ADP^{+}07, BC05, BW00, BFP08, BASB01, CWW96, CIB01, CSL^{+}07, CBR^{+}22, DRG98, DTC^{+}09, EK97, FLWW02, FLWC07, FW00, FRS97, GPH^{+}09, GM03, GABP00, HV07, HA05, HJ08, HLCH07, JB98, JP08, KSS^{+}09, KG99, KCA04, LC13, LSL^{+}13, LFG^{+}09, LCL08, MOZ06, MBB01, MP07, MLG12, OCRS07, PB14, Ped96, Ped06, PBSV^{+}06, PW99, RFYL98, RS98, SW12, SGD10, SYL09, SSCS10, SUC01, SS11, SZV^{+}12, TW96, THL^{+}13, VAAH^{+}98, Voe01, WAZ98, WKR09, ZHM07, ZHC^{+}23]. Design-for-manufacturability [WPR^{+}19].

Design-for-Testability [Pom16a, Pom18a, Pom14a]. design-specific [ACT13]. Designed [KMO^{+}12, SPT^{+}17]. Designer [SS11]. Designing [BLNK14, DZS^{+}18, HBC^{+}08]. Designs [EK16, GD20, HLL^{+}24, LZY^{+}23, LTZ22, MACV14, PHKW12, WJM24, WWW^{+}12, WCX^{+}24, YVC14, Yan16, Yan17, ZK15, CH00, GM08, GOC02, HMB98, KI01, KK11, KHW06, LHW97, LCHT02, LLHT12, LAS01, LCKT12, MS00, MR96, RMKP03, Sen11, SSCS10, SNL12, WTL^{+}13, Yan11, ZMTC13]. Destination [RL13]. Destination-based [RL13]. Detailed [CJKS24, GdRJM21, HWL^{+}23a, MJB19, CBHK11, PWY05]. Detecting [DY23, DZ24]. Detection [CBO^{+}18, HDZ^{+}20, JYY^{+}22, KOO18, LXWC20, LYM^{+}20, LL19, LM21, PTPB22, Pom16b, Pom17a, VTC20, WH20, YFT17, ZHC^{+}18, CR12, DHZ^{+}11, FNP09, KI01, KRK98, KSA^{+}10, LM05, PR07, RM09, SCCH08, TDE08]. Determined [Pom18a]. Deterministic [EY12, KBV^{+}15, LB11, ZHC^{+}21, KT01]. Deterministic-Path [ZHC^{+}21]. detour [YW09]. Detours [Yan19]. developing [SMSB05]. Development [THT12]. developments [Lin97]. Device [BXG^{+}24, GHYR19, HXZ^{+}23, ZXC^{+}23, TZZH22]. Device-Based [GHYR19]. Devices [CLL^{+}22, GAT^{+}21, HSP^{+}22, KP22, Kha12, LPLK22, LKH19, LY24, PGGD23, PTPB22, SVK17, XPX^{+}21, YB23, JCS^{+}08, ZYZ^{+}13]. DFT [DDFR13, PTC^{+}15]. Diagnosability [CLH12, CCH15b, CH13, HWL^{+}23b, LH14]. Diagnosing [BDBB19]. Diagnosis [HFMB20, Pom17b, PA21, SBB^{+}18, WH19, WH20, XGWL24, CML98, KI01, TYH08]. Diagnostic [HVF^{+}01, HFMB20]. diagonal [DSKB04]. Diagram [HZL^{+}22]. Diagrams [CM19, KC98]. dictionaries [LCT03]. dictionary [HH09]. Diet [LS23]. difference [Das09]. differentiable [Con06].

Differential

[DMR23, HZJC23, JD18, LLP^{+}16, DDFR13]. differentiated [WHXZ13]. Digital [CM18, DZCD15, GLD^{+}22, JYHY21, LHC16, LCZ^{+}24, LKC^{+}18, MFHP12, MGR^{+}15, MWK21, PGCB16, PBF^{+}22, RB19, RCK^{+}15, RB21, SKS^{+}18, SOS15, VBP^{+}19, CPW04, RS03, SR12, SOC06]. Digitally [ZK15]. Dilution [GHYR19, KGS^{+}20]. Dimension [BC11, WPL23]. Dimension-reducible [BC11]. Dimensional [RGM15, SYH^{+}22, KQP^{+}19, Pom24b, WXH^{+}19, YYC07, YYC09]. Directed [CHK^{+}23, IE12, QM12, WLM21, CM13, HLCH07, HG07, LKTD98, MD08]. Direction [Yan18]. Direction-Constrained [Yan18]. Directives [SCL^{+}22]. discharging [HLCH07]. Discovering [NGL^{+}21]. Discrete [CHK^{+}23, HLG^{+}15, LGGJ14, MLG12, SV16]. Disjunctive [WYIG07]. disk [CD09, SLXZ12]. Dispatching [WHRC12]. Displacement [BFG^{+}19]. Dissipative [ZMS^{+}19]. Distance [HRK18, LKLC22, LDLM20, NAK20]. Distance-aware [LKLC22]. Distance-based [NAK20]. distinguishability [AGM01]. Distributed [APG24, CGLH23, EAP17, HXC^{+}18, JJH21, MVK^{+}18, SCK18, SRKS23, WLZ^{+}19, YMB15, CFX09, LC14, PEPP06, Wol96, dW97]. Distribution [JCK^{+}18, SSO16, WDD^{+}23, KSA^{+}10, SW99]. Distribution-Aware [WDD^{+}23]. Distributions [KYL16, STJG16]. Disturb [LHS^{+}21]. Disturbance [SBB^{+}18]. Disturbance-Free [SBB^{+}18]. Diversity [DZ24]. Divide [SW12, HPK99]. divide-and-conquer [HPK99]. Divided [TMDF10]. divider [EKEK22]. Division [PY20, LWG^{+}23]. DME [wATkK02]. DNN [CSO22, GYZ^{+}22, HWDQ22]. DNUCA [DK16]. Domain [ZGB^{+}24, FWCL05, IAI^{+}09, JBC^{+}10, LTPR^{+}13, SCV06]. domain-specific [SCV06]. Domains [WWW^{+}12, LBV^{+}06]. dominant [VCLD03]. dominated [FRS97, KI01, LDLM20, MWG97]. domino [KJKK03, ZS02, CLLK06, NTSA18]. Don’t [DY23, TPC^{+}17, CBMM10, SGK08]. don’t-cares [CBMM10, SGK08]. Dot [RBWB20, RB21]. Double [HWDQ22, HNS23, PMA24, SHL^{+}19, XYG^{+}16]. Double-row [HNS23]. Double-Shift [HWDQ22]. DPRTM [ADDM^{+}13]. DRAGON [HLW^{+}23]. DRAM [BLNK14, CJKK19, LYLW17, LMA^{+}16, PKJK20, PSP24, SSS^{+}19, SAL19, ZZCY17]. DRAM/PCM [BLNK14, LYLW17]. DRAMs [LS19]. DRC [ZZL^{+}23]. DRC-SG [ZZL^{+}23]. DRDU [IBMD07]. DReAM [LMA^{+}16]. Drive [CCS15, VA17b]. Driven [AMM^{+}18, CYV^{+}14, DKT^{+}16, DZCD15, EAP17, GDD21, HWGY16, HWCL15, LVS16, LHJ12, LNG^{+}16, PBF^{+}22, SKS^{+}18, Yan16, YP10, ZFLS11, ZSY18, CSAHR07, CZW00, CXS^{+}23, DRG98, EK97, GK14, HC23, HW00, JPCJ06, KMS12, Kuc03, KSA^{+}10, LLM^{+}23, LOC12, LL19, MPSJ07, MD08, MRMP08, NM23, PBH^{+}24, WY06, WLC02, XK97, Yan08, ZSZ10, MSD06]. drives [CCYC14]. Driving [dONH23]. Droplet [LKC^{+}18, RBWB20]. DSA [YLZ^{+}17]. DSP [AFM14, CL99a, LP03, SXX^{+}06, SESN15]. DSPs [AM98]. Dual [BLNK14, BPTB17, HS18, KKS16, CT13, HLHT08, MLMM08, SM00, WGDK07, WYC10]. Dual-Edge [BPTB17]. Dual-Edge-Triggered [HS18]. Dual-Mode [KKS16]. Dual-Phase [BLNK14]. dual-scanline [CT13]. dual-Vdd [HLHT08]. duplication [CC06, WY06]. During [TPC^{+}17, EW18b, HR06, MRC06, PTC^{+}15, RGM09, XPSE12, YWK^{+}03, YWW10, ZMTC13]. Duty [BHY^{+}24, JSG09]. Duty-Cycle [BHY^{+}24]. duty-cycled [JSG09]. DVFS [CXK^{+}13, SQL^{+}24]. DVFS-induced [SQL^{+}24]. Dynamic [ADDM^{+}13, BMJ13, BLUS19, BHS11, Brz24, CLX^{+}23, HKL^{+}15, HRP00, HLX^{+}23, HLW^{+}23, IAI^{+}09, LDP^{+}22, LHW^{+}17, LCZ^{+}24, LV14, MNMK^{+}21, MDR15, NDA^{+}23, ORGD^{+}15, PSP24, PBL^{+}17, RNA^{+}21, SKP21, SV11, WMT^{+}16, WML^{+}24, WGSH16, WZL^{+}21, XPX^{+}21, ZGB^{+}24, AHAKP08, ADM^{+}13, AMM^{+}06, BLR06, CMNQ08, GK14, GPH^{+}09, KJT04, KSA^{+}10, LTPT10, LLHT12, MR05, VJBC07, KMR18]. Dynamical [CS22]. Dynamically [CRC15, DHX^{+}23, DHW^{+}23, JPHL16, Pom18a, RNR^{+}21, ARLJH06, WLC02, YYLL09]. dynamics [WHXZ13]. DYNASCORE [KMR18].

E-Beam [LZ17]. E/E [dONH23]. E2HRL [SKR^{+}22]. Early [KO23, LTZ22, PBL^{+}17, SZB17, MKBS05, SYL09]. Early-Release [SZB17]. Easy [VS12a]. EBL [YYG^{+}16]. ECC [KRH18]. ECDSA [DHB16]. ECG [APB^{+}08]. echo [FIR^{+}97]. ECO [DVA02, LG12, LNPL23]. ECO-GNN [LNPL23]. ECR [LTYW12]. EDA [JHMGS18, LZR23]. EDF [GDG^{+}08, SZB17, WDZG16]. Edge [BPTB17, HS18, KP22, LQL^{+}24, LY24, MS23, PGGD23, WB^{+}24, YXG^{+}24, RS98]. Edge-aware [YXG^{+}24]. edge-based [RS98]. Editor [Ano13, Hu20, MYSZ23]. Editor-in-Chief [Ano13, Hu20]. Editorial [CH10b, CPX14, Dut05, Dut06, Dut07, Dut08c, Dut08a, Dut08b, Hu20, Irw00, MD13, Ped08, TK18, SJ02, Mar00]. EF [TZZH22]. EF-Train [TZZH22]. Effect [LHW^{+}17, NSS^{+}16, WCCC14, WSH^{+}18, WSRH16, LTH99]. Effective [DS06, JPHL16, LCJ^{+}10, LTW^{+}16, LCL08, NAK20, PCT^{+}17, XLY^{+}18, YVC14, YLZ^{+}17, YLY^{+}23, LPP00, LSPC14, MHT14, SBC08, WSV^{+}14, XLCL13]. effectiveness [WAZ98]. Effects [BDB98, BFL10, GC18, JIR^{+}21, VFML23, MRB^{+}11, RJBS09]. Efficiency [HSP^{+}22, KKLG15, LWC18, RB19, TCL14, WH19, WCX^{+}24, KJT04, ZAZ13]. Efficient [AKAKP18, BS14a, BHDS09, BW00, CK19, CCY22, CAOM19, CBC22, CYV^{+}14, CSO22, CLC^{+}24, DMR10, EO19, FHL^{+}23, GADG19, GT21, GFJ16, HLZ^{+}22, HSR^{+}24, HMB98, HAB^{+}17, HKB^{+}07, HCS01, HMMG^{+}20, HG07, HWX^{+}14, JSS^{+}19, JYY^{+}22, JLK15, KBN09, KC10, KW02, LHLP16, LJ18, LDD^{+}18, LCJ^{+}22, LHZ^{+}06, LWZ^{+}19, LAYZ23, LZ21, LF12, LHCT05, LM96, LB11, MWS^{+}20, MNMK^{+}21, MWK21, NTSA18, PMP17, QZZW24, RM09, RGM15, SV16, SMS22, SMBT19, SPC^{+}15, SPMS02, SS14, SYGC22, SCK^{+}23, SJL23, SRC15, TLCF16, TYSF20, TZZH22, VNS19, WKL^{+}18, WS22, WJY^{+}07, WWFT12, YPCF17, YCHT00, YP10, ZYW^{+}18, ZLG^{+}19, ZZL^{+}23, ARLJH06, BP23, CD09, Das09, EKEK22, FNP09, GM03, GBC07, IBMD07, JS13, JP08, KL05, LCD07, LH13, MR96, MR05, MP07, MWG97, SGD10, SLXZ12, SKR^{+}22, SHN12, SZV^{+}12, VILSL23, VKKR02, Wu09, ZSZ10]. efficient [ZYZ^{+}13, ZMLH24, Zho08, LCG^{+}22]. Efficiently [RCG^{+}08, TY19, ADM^{+}13]. Eh [DKT^{+}16, DBK^{+}18]. Elastic [LYL^{+}19, SZB17]. Electric [AKM^{+}22, VA17b]. Electrical [BHLG19, WM24]. Electrode [RBWB20]. Electromagnetic [JIR^{+}21, WFSS20]. Electromigration [DNT20, HZJC23]. Electron [HCW^{+}16]. Electronic [CH10a, HHH^{+}21, KLSZ09, Kha23, KAC^{+}23, SSK^{+}23, HV07]. Electronics [BSP^{+}19, CPX14, XRS^{+}19, CH10a]. Electrostatic [LDD^{+}19]. Electrostatics [LCC^{+}15]. Electrostatics-Based [LCC^{+}15]. Element [CLT^{+}15, ZK15]. elements [HMVG13]. eliminate [Mut09]. Eliminating [SHLL98]. Elimination [LHF12]. Elite [ZKS^{+}16]. Embedded [BMdG17, BD14, BS14c, BM11, BYT22, CHA^{+}23, DFM15, EAP17, GAT^{+}21, HCL^{+}14, IK19, IGN18, JJH21, KC10, LS23, LL15, LHLP16, LHK^{+}15, LL19, NSH^{+}16, OHA19, PG15, RFG20, SPT^{+}17, SL18, SJN24, SLV^{+}22, VBP^{+}19, WHRC12, XPZ^{+}18, XPX^{+}21, YP10, AM10, BPRR98, BH10, CSAHR07, CMM00, CSL^{+}07, CM13, DCK07, DCK09, DRG98, GDTG07, GPH^{+}09, GG04, GABP00, HKL^{+}07, HV07, HCK13, IAI^{+}09, JS13, KNDK96, LJV02, LCZ^{+}08, LSDV10, LB00, LMW99, LDK99, MBB01, MDG98, ML09, NG06, NR03, PDN97, PDN00, PCD^{+}01, PHM00, PEPP06, QS09, RSR01, SR12, SUC01, TKVN07, WAZ98, Wol96, XZC09, ZYDP08, ZP08]. Embedding [CM18, ZGB^{+}23]. Embeddings [CM19]. Emerging [BRCS18, SN10, YPCF17, BC08]. Employing [GS13, ZK15]. emulated [THC^{+}14]. Emulation [ALLE20, LTZ22, ADP^{+}07, HMVG13, KRK98, MW97]. En/Decoder [SJL23]. Enable [CLL^{+}22, TZZH22]. Enabled [CXR^{+}23, XRS^{+}19, YSF^{+}18, LDD^{+}19, LSL^{+}13, SLC^{+}22, YFT18]. Enabling [BSP^{+}22, IK19, JS13, SYGC22, ZHOM08]. Encoder [CAP^{+}23, QSW^{+}15, SLV^{+}22]. Encoder-Decoder [CAP^{+}23]. Encoding [CGV^{+}23, MDR15, OT15, PMP17, YMB15, ZLG^{+}19, KJT04, LCD07, LWC07, NT05, RTNL05, YGZ04]. Encryption [Che18]. End [ENP20, SJL23, GABP00]. End-to-End [SJL23, ENP20]. Ended [RM23b]. Endurance [CHC^{+}16, CCK^{+}18, HHK^{+}17]. Energy [BP23, BFL10, CCY22, CBC22, CSO22, DMR10, EKEK22, ENP20, GADG19, GT21, GFJ16, HXB^{+}22, HXC^{+}18, HSP^{+}22, JDD20, JSS^{+}19, JPHL16, KC10, LDD^{+}18, LWX^{+}23, LF12, LWC18, LMA^{+}16, MNMK^{+}21, MBD^{+}20, MR05, NTSA18, NRM^{+}24, PMP17, RB19, SMS22, SPC^{+}15, SKR^{+}22, TLCF16, TYSF20, TBCH17, VILSL23, WH05, WKL^{+}18, WCX^{+}24, XPZ^{+}18, XPX^{+}21, YB23, YPCF17, YP10, ZMLH24, ZHTC09, ZMS^{+}19, ANR13, CSAHR07, CLQ12, GBC07, HG07, HW00, JS13, JCS^{+}08, KSK^{+}05, KRS06, Kan06, KC13, KJR^{+}07, LSL^{+}13, LC07, MED23, MRC06, OK08, SLXZ12, SHN12, WLL^{+}11, Wu09, ZAZ13]. Energy- [LWX^{+}23, YP10]. Energy-Aware [HXB^{+}22, TBCH17, ENP20, JDD20, MBD^{+}20, WH05, JCS^{+}08]. Energy-Constrained [NRM^{+}24, XPX^{+}21]. Energy-Efficient [CCY22, CBC22, DMR10, GT21, GFJ16, JSS^{+}19, KC10, LDD^{+}18, LF12, MNMK^{+}21, NTSA18, PMP17, SPC^{+}15, TLCF16, TYSF20, WKL^{+}18, YPCF17, BP23, EKEK22, MR05, SKR^{+}22, VILSL23, ZMLH24, SLXZ12, SHN12, Wu09]. energy/thermal/cooling [ANR13]. Enforcing [EWT23]. Engine [LLL^{+}18, TMDF10, CNQ13, DP02, DP04]. Engineering [AYS20, CM18, EAP17, GDTF17, WSS^{+}18]. Engines [HKL^{+}15, VFML23]. Enhance [DLC^{+}17, GS13]. Enhanced [CYH19, CGV^{+}23, DZK^{+}24, LKH19, Pom15a, PS23, TWL16, FWCL05]. Enhancement [CYLC24, HWL^{+}23a, HWCL13, LCKT12]. Enhancements [Che18, PKC^{+}21, ZAZ13]. Enhancing [CCK^{+}18, GPS^{+}24, NRDB19, PPP^{+}15, SJN24]. Enlarged [ZS16]. Ensemble [WB16, WH19, WLH20]. Ensure [SLC^{+}22]. Enterprise [DKZ^{+}15]. entries [LCT03]. enumerative [STJG16]. Environment [RHN00, HKL^{+}07, Hsi01, SCV06]. Environmental [GPS^{+}24]. Environmentally [YBS^{+}18]. EPGAs [YTHC97]. EPHA [ZMLH24]. EPIC [AMR00]. ePlace [LCC^{+}15]. Equal [Pom21b]. Equation [Shi20, WTW^{+}23]. Equations [HZJC23]. Equipment [GCL^{+}16]. Equivalence [AA17, Fuj05, AGM01, HMB98, HCC01, KMS12]. Equivalent [Pom21b, MCMW08]. Era [HAB^{+}17]. ERfair [NSH^{+}16]. Error [CS22, DHZL23, HWL^{+}23a, LTYW12, LD17, LWC18, LW21, PB12, PHKW12, PGCB16, SMS22, TLCF16, WH20, KI01, KSA^{+}10, RM09, SCCH08, VAAH^{+}98, WHXZ13]. Error-Correcting [PGCB16]. Errors [DFM15, RJBS09]. Escape [JD18, Yan17, Yan18]. ESD [PLH^{+}24]. ESL [KSS^{+}09]. ESPSim [LAYZ23]. Establishing [GSFT16]. establishment [AJM13]. Estimate [LMA^{+}16]. Estimates [CM19, GS00]. Estimating [Meh98]. Estimation [APDC17, APS18, BZWZ17, Kha23, LD17, LZY^{+}23, NSP^{+}20, PB12, SNH02, SSN22, TC98, WXH^{+}19, ZLG^{+}19, ZPLI23, CIB01, DTC^{+}09, FLPP09, HKV^{+}07, JT98, KCA04, KNRK06, LMW99, MHF96, ZSZ10]. estimators [XK97]. Ethernet [MAS^{+}20]. evaluating [JBC^{+}10]. Evaluation [BBEM15, EBR^{+}09, GD20, GQW19, HBPW14, IYF^{+}21, LFST21, LTZ22, QBTM16, RGX^{+}24, CHY05, JLF^{+}12, LCOM07, PB14, SGJ96, WSV^{+}14]. Event [KRL15, MCD12, PMA24, RCD07, YH97, ZKS^{+}16, CBR^{+}05, HW00]. event-based [CBR^{+}05]. event-driven [HW00]. Evolution [PSK08]. Evolutionary [EWT23, JYHY21, WSY23]. Evolvable [SS24]. EWD [MPSJ07]. Exact [EAAK^{+}23, EKS^{+}14, Sch17, FLWC07, FNMS01, NR01]. Examples [DZ24]. Excitation [SOS15]. exclusive [DK08]. Execution [APDC17, GDD21, HLZ^{+}22, LSCK20, NRDB19, VGG19]. EXFI [BPRR98]. exhaustive [CMB07]. Expansion [MS17]. Experience [YXG^{+}24]. experiment [FIR^{+}97]. Experimental [Das04, AYM05]. Experiments [LHK^{+}15, BCC08, CIB01]. Experts [TEK18]. Explaining [YYL^{+}15]. explicit [EK97]. exploitation [GFC^{+}09]. Exploiting [GSD^{+}18, JLK15, OT15, WKC12, WHXZ13, DSRV02, FW00, Kan06]. Exploration [BSZ^{+}24, FLG^{+}23, FMR23, FCZ^{+}23, GACK22, HMMG^{+}20, LLLL18, MA16, RFG20, RS18, Sch17, WS22, APB^{+}08, CSL^{+}07, EK97, JP08, KSS^{+}09, LCOM07, MBB01, MSD06, PB14, PPDK09, RJL^{+}09, SW12, SUC01, VCLD03, XPSE12]. Explorer [BSZ^{+}24]. Exploring [CK19, QZZW24, TLCF16, WGDK07, YPCF17]. Exponential [APS18]. Express [JSA18]. expressions [SGJ96]. Extended [WWFT12, CK96, YTHC97]. Extensibility [SGC^{+}14]. Extensible [KAKSP16, MP07]. Extension [LF12, YCL^{+}23]. extensions [WKR09]. extensive [CBMM10]. External [KG09, CBMM10, XZC09]. Extra [CVMP19, KAKSP16]. Extra-Functional [CVMP19]. Extracting [Pom24b]. Extraction [BHBS22, HDZ^{+}20, YLY^{+}23, ZZL^{+}23]. Extreme [HKJ^{+}23, Pom15b].

fabric [MSB^{+}09]. fabrication [WLT08]. factorization [BOC00]. Factory [DZCD15]. FACTS [VMP^{+}00]. Fail [PAV17, PA21, BWB14]. Failure [XNZ^{+}15]. Failures [YYL^{+}15]. False [AKAKP18, AL19, GGBZ02, SHLL98]. False-noise [GGBZ02]. family [BD05]. fan [LH09]. fan-out [LH09]. Fast [ATF^{+}23, CPW04, DK16, DNT20, GdRJM21, GLY^{+}12, HNS23, HGLC16, IHM15, JZYZ15, KKLG15, LZY^{+}23, LH11, SMBT19, SGD10, STWX12, Tes02, TZ17, ZHJ^{+}23, CCW08, GMN^{+}13, GBC07, JHL02, KT96, LC14, LCKT12, NR01, SBGD13, SGJ96, YTHC97, HHX^{+}23, LCC^{+}15, OS03, QSK12]. FastCFI [FHHR21]. Faster [SSN22]. fastest [Das04]. Fault [CYH19, CGV^{+}23, EKS^{+}14, GT21, GVJ15, HDB22, HWL^{+}23b, IYF^{+}21, JIR^{+}21, JPM^{+}19, LW17, LW21, LXWC20, LTZ22, NGL^{+}21, Pom22, RRHB21, SQL^{+}24, XCF18, Xia24, XGWL24, YYL^{+}15, BPRR98, BH03, CEB06, DNA^{+}12, HH09, JLF^{+}12, LTH99, LLQ^{+}03, SC06, TCP97, TD03]. Fault-Aware [GVJ15]. Fault-based [IYF^{+}21]. Fault-Induced [RRHB21]. Fault-Tolerant [CYH19, GT21, LW17, XCF18, NGL^{+}21, SC06]. FaultDroid [RRHB21]. Faults [BDBB19, HDB22, MCD12, Pom17b, Pom19b, Pom20, Pom21b, ZHC^{+}21, HVF^{+}01, LTH99, LIA00, MT02, PT06, PR98, PR09, TYH08, XZC09]. Faulty [JCK^{+}18, JPM^{+}19]. FBGA [WPL23]. Feature [HDZ^{+}20, VTC20]. Features [LL19]. featuring [EK97]. Federated [ZHC^{+}23]. Feed [Ase23, LHS20]. Feed-Forward [Ase23, LHS20]. feedback [LWK11]. FeFET [LSZ^{+}24]. FeFET-Based [LSZ^{+}24]. FET [AKM^{+}22]. fetches [KTKO13]. FFT [HDZ^{+}20, TMDF10]. FFT-based [HDZ^{+}20]. FH [HGLC16]. FH-OAOS [HGLC16]. Fidelity [WFSS20, SCL^{+}22]. Field [JDLZ24, WSH^{+}18, CH02, CD96, PWY05, WV02]. field-programmable [CH02, PWY05]. FIFO [BK00, ZLL^{+}16]. File [TLCF16, CFX09, GF10, ZYP09]. Files [WKL^{+}18]. Fill [LTW^{+}16, LIA00]. Filling [TPC^{+}17]. Filter [BH22, EO19, MED23, PCT^{+}17, FS13, TKVN07]. filtering [CL13, ZYDP08]. Filters [RB19]. finding [KL05]. Fine [BYT22, LG18, LPY^{+}20, RCW22]. Fine-Grain [LG18]. Fine-Grained [BYT22, RCW22, LPY^{+}20]. FinFET [PS23, WLLH16]. Finite [CLT^{+}15, SRC15, CK96, CHHL96, GK07, GK09]. Finite-Element-Based [CLT^{+}15]. Finite-Point [SRC15]. Firmware [KC10, RGT^{+}14]. first [MR96]. first-time-right [MR96]. Fixed [ALL17, WDZG16, YCL^{+}23, ZHJ^{+}23, AM98, CPW04, LCT03, MHQ07]. fixed-length [LCT03]. Fixed-Point [ALL17, YCL^{+}23, AM98, CPW04]. Fixed-Priority [WDZG16, MHQ07]. Fixing [LSZ^{+}21]. Flash [CCK^{+}18, CWL^{+}22, DHZL23, HCL^{+}14, KC10, MWS^{+}20, PPP^{+}15, WQC^{+}16, WL12, WZL^{+}21, ZLW^{+}15, HCK13, JCS^{+}08, Wu09]. Flash-Based [HCL^{+}14, KC10]. flash-memory [Wu09]. Flattened [ZYPC17]. flexibility [JDLZ24]. Flexible [BHK17, FMR23, IGN18, LKC^{+}18, RS18, CL99b, MS00]. FlexRay [SGC^{+}14]. Flip [HS18, HKJ^{+}23, Kha23, KMO^{+}12, LW21, VILSL23, XCW12, Yan16, KOS09, KSA^{+}10, LLLC13, Yan11, ZMTC13, WB^{+}24]. Flip-Chip [Yan16, Yan11, ZMTC13]. Flip-Flop [Kha23, KMO^{+}12, LW21, XCW12, HKJ^{+}23, VILSL23, LLLC13]. Flip-Flops [HS18, KOS09, KSA^{+}10]. Floating [BS14a, BSP^{+}23, SKCM06, WG11]. Floating-point [BSP^{+}23, WG11]. Floorplan [KQP^{+}19, YVC14, YCCG03, HCS01, LCL08, MRMP08, SY07]. Floorplan-Guided [YVC14]. Floorplanning [DHX^{+}23, DHW^{+}23, HCRK11, HCZ^{+}16, KLE18, LJL^{+}23, YXG^{+}24, HMLL11, LHZ^{+}06, LCC11, LLM01, SYZ08, WLCJ09, YYC07, YYC09]. floorplanning-based [LCC11]. floorplans [DSK01, MSKBD07, MS00, WYC10]. Flop [Kha23, KMO^{+}12, LW21, XCW12, HKJ^{+}23, LLLC13, VILSL23]. Flops [HS18, KOS09, KSA^{+}10]. Flow [FHHR21, HMO^{+}14, IGN18, KGS^{+}20, KW16, LJJ^{+}22, MJB19, NPH^{+}20, NM23, PKC^{+}21, PDS12, QSW^{+}15, RJ14, XPX^{+}21, ZGB^{+}23, BFP08, DTC^{+}09, FHHH22, GDF09, KMS12, LC13, OM08, WC06]. Flow-Based [KGS^{+}20]. Flows [DZK^{+}24, JLJ15, VGG19]. Fluid [GHYR19, KR23]. Fluids [KGS^{+}20, RCK^{+}15]. Flux [LSZ^{+}21]. Fly [VFML23]. FOLD [Pom15b]. Folded [AFM14, HS18]. Folding [Pom15b, BHS11, TS96]. footprint [AMM^{+}06]. Forced [RSR01]. Forecasting [LG23]. form [CW01, PR09, Shi20]. Formal [Ali12, BGM04, EW18a, HLL^{+}24, KMS12, KG99, SSS^{+}19, SGGR14, VS12a, ADS^{+}09, CMM00, MR96, RFYL98, SMSB05, VS12b, Zho08]. Formally [KRH18]. formats [AMR00]. Forming [PR07]. FORTIS [GSFT16]. Forward [Ase23, GSFT16, GS00, LHS20]. Foundation [TB20]. Four [HGLC16]. Four-Step [HGLC16]. Fourier [LCC^{+}15]. FPGA [AMM^{+}18, ACT13, ALLE20, BS14c, BHS11, CWW96, CZW^{+}03, CH00, CLC^{+}24, DP02, EW18b, FW00, FHHR21, GPK^{+}09, GVJ15, HSR^{+}24, HABS15, HYK^{+}20, HLHT08, HW14, JLF^{+}12, KT96, KL05, KFH^{+}08, LKM04, LLL^{+}18, LM19, LWG^{+}23, LZA^{+}21, LDX22, MMM^{+}22, MW97, MA16, MP07, MS21, OK20, PSD21, PL98, PMT20, PBH^{+}24, PSNC18, PFHAH22, PY20, SLV^{+}22, SYGC22, SAHF^{+}20, TZZH22, TW96, XXC^{+}24, ZLQ15, ZHTC09]. FPGA-based [MW97, ALLE20, PSNC18, DP02, GPK^{+}09]. FPGA/FPIC [CZW^{+}03]. FPGAPRO [LDX22]. FPGAs [CZW00, CEB06, CHY05, DVA02, GNGT21, GDG^{+}08, KNRK06, LZY^{+}23, LQL^{+}24, LB11, MCZ^{+}16, MLMM08, SPMS02, Tes02, VKT02, WG11, WS22, WTW^{+}23, WLC02, WSEA99, YGH^{+}10, YYLL09]. FPIC [CZW^{+}03]. Framework [APG24, CSC^{+}21, DK16, DSHD23, FMR23, GACK22, GDTF17, HWDQ22, HLZ^{+}22, HRC21, HZJC23, HTC^{+}23, JJH21, JSS^{+}19, JPHL16, KPB19, LL15, LZY^{+}23, LHC24, LIK22, LTZ22, LLQD23, LDX22, MBD^{+}20, NPH^{+}20, NDA^{+}23, QZZW24, RG19, RB21, SKM^{+}16, THT12, WLZ^{+}19, WWFT12, XPX^{+}21, YP10, ZLL^{+}16, ZF23, ZFL22, ADP^{+}07, HR06, HV07, KKJ^{+}08, KH10, MPSJ07, MP07, RPKC05, SB98, SBH^{+}06, SS11, ZM07]. Free [RGM15, SBB^{+}18, ZBG^{+}23, BLR06]. frequencies [PL03]. Frequency [GC18, JPHL16, WTR12, WGSH16, GM08, JDT^{+}08, LTPR^{+}13, ML09]. frequency- [LTPR^{+}13]. Frequent [YGZ04]. FSM [AGM01, CGV^{+}23]. FSMs [CK16, EWT23]. FTT [NGL^{+}21]. FTT-NAS [NGL^{+}21]. FUBOCO [AG22]. fuel [LCZ^{+}08]. fuel-cell-battery [LCZ^{+}08]. Full [STWX12, HDL^{+}12]. Full-Chip [STWX12]. fully [FW00]. Function [BHY^{+}24, CSC^{+}21, LYL^{+}23, LLQD23]. Functional [CVMP19, DCK07, FMR23, FRS97, PR98, Pom15b, Pom15c, Pom16a, Pom16c, Pom18a, Pom18b, Pom19a, Pom21a, Pom24b, VLH98, WSEA99, XLY^{+}18, CMB07, CK96, EWT23, LOC12, MT02, Pom13, Pom14b, Vah99, AG22]. Functionality [BFV15, HLCH07]. functionality-directed [HLCH07]. Functions [Ase23, BC11, CCQ98, TW96]. Fundamental [SBY^{+}20, XLNB17, Voe01]. FUNI [LIA00]. Future [HAB^{+}17, Kha23, KBV^{+}15, ZZCY17]. FuzzRoute [RGM15].

Galois [JDLZ24]. GALS [SS11]. GALS-Designer [SS11]. game [HR06, RJL^{+}09]. game-theoretic [HR06]. GAN [LRHL24]. GAN-Place [LRHL24]. GANDSE [FLG^{+}23]. Garbage [GSD^{+}18, HCL^{+}14, ZLW^{+}15]. Gate [CM19, CDB11, Che96, CHK^{+}23, HDB22, HMO^{+}14, KKS16, LGGJ14, SV16, SRC15, VTC20, CCW08, CH02, CD96, CH00, HH09, LG12, LLYW10, PWY05, RGM09, SC00, WY06]. Gate-Level [CDB11, HMO^{+}14, VTC20, CM19, Che96]. gated [CM08]. Gates [WSS^{+}18, KOS09]. Gateway [HXC^{+}18, JSG09]. Gating [CMP10, CLMZ10, KKHK16, WKC12, XLS15, BDM^{+}99, ETR07, HTCP13, KBN09, SSCS10, YHL07]. Gaussian [ZYW^{+}18]. GBDD [YTHC97]. GEMM [CSO22, WML^{+}24]. General [CH02, HWF^{+}23, LSZ^{+}24, wATkK02]. Generalized [GMS^{+}23, Pom15c, DS06].

Generated [CCH15b]. Generating [MFS09, MN17, PKJK20, KT01]. Generation [BKW15, BFV15, CYV^{+}14, GMS^{+}23, IE12, Kha23, LCY12, LV14, LCYN18, MFHP12, MCD12, NPH^{+}20, PCT^{+}17, Pom17a, Pom17b, Pom18b, Pom24a, SHD17, Shi20, STJG16, SOS15, TAP^{+}24, VFML23, WLM21, WWW^{+}12, YLZ^{+}17, YD16, ZZL^{+}23, AM98, CK96, Che96, CL99a, CCW08, GF06, HRP00, KKMB02, KJR^{+}07, KNDK96, KH10, LTH99, LP03, LKTD98, MMP00, MSD06, MD08, PFHAH22, PR98, PR07, Pom13, QM12, SR12, SNL12, SM00, TBZ13, VMP^{+}00, dW97]. Generative [FLG^{+}23, LRHL24]. generator [BCR^{+}08, WWC04]. Generic [SA24, FLWW02, FLWC07]. Genetic [MA16]. Genetic-Algorithm-Based [MA16]. Geometric [CM18, HWF^{+}23, WJYZ11]. geometry [JCGP05]. Global [AOC02, BM11, CJKS24, DHNR23, GD22, RGM15, WSH^{+}18, ZPLI23, CLYP09, DHV^{+}00, SPA^{+}03, ZHTC09]. Global/Local [BM11]. Globally [PMS15]. Gmax [BP23]. GMDF [FIR^{+}97]. Gmin [BP23]. Gmin-Gmax [BP23]. GNN [LNPL23, VILSL23]. GNN-based [VILSL23]. Good [GdRJM21, GMN^{+}13, YWK^{+}03]. GP [APS18]. GPGPU [SBR^{+}17]. GPGPUs [HIW15, TLCF16]. GPlace3.0 [AMM^{+}18]. GPU [CDB11, CBR^{+}22, HCRK11, LLK^{+}14, LH11, NHS23, SSN22, TYSF20]. GPU-Based [LH11]. GPUs [BYT22, SABSA15, SQL^{+}24, TY19, WKL^{+}18, ZWD11]. Gradient [SV16, GBC07]. gradient-based [GBC07]. grading [PT06]. Grain [LG18]. Grained [BYT22, RCW22, KLSP11, LPY^{+}20]. Grammar [JHMGS18]. Granularity [RBWB20]. Graph [CHK^{+}23, CH17, CBR^{+}22, CXR^{+}23, FCZ^{+}23, HRC21, HLW^{+}23, JHMGS18, JOH17, LB00, LJL^{+}23, LNPL23, OKJH22, SSK^{+}23, SS14, YXG^{+}24, WYC10, WC06]. Graph-based [LB00]. Graph-Grammar-Based [JHMGS18]. graphene [YMC^{+}13]. graphical [BLR06]. GraphPlanner [LJL^{+}23]. Graphs [ASAP17, BFG17b, CM18, CCH15b, CHK^{+}23, ENP20, HPB11, LH14, CH13, DSK01, HKB^{+}07, LKTD98, MHF96]. Gravity [OS03]. Grid [DNT20, HXC^{+}18, LAYZ23, MN17, SCK18, ZS16, MFS09]. gridless [LCC11]. Grids [BS14b]. GRIP [JHMGS18]. Ground [HC23, LHJ12, YHH09]. Grouping [XCW12, KSA^{+}10]. Guarantee [MN17]. Guaranteed [PMS15]. Guest [CH10b, Mar00, SJ02, MYSZ23]. Guidance [ZKS^{+}16]. Guided [YVC14, RNR^{+}21]. Guidelines [WPR^{+}19]. Guiding [EW18a].

H3D [LY24]. H3D-Transformer [LY24]. Hamming [HRK18]. Handling [DH06, GdRJM21]. Hard [CHBK15, CWL^{+}22, NRM^{+}24, WDZG16, PW99, QS09]. hard/soft [QS09]. Hardened [BS14c]. hardness [WYC10]. Hardware [ANS^{+}20, BS14a, BSP^{+}23, BM11, CMM00, CBR^{+}22, DY23, DZS^{+}18, GFJ16, GQW19, HJY23, HLL^{+}24, IPWW17, KTKO13, KP22, LG18, LHF12, LF12, LPL^{+}21, MED23, MRL^{+}20, MFHP12, MCY23, MRL^{+}19, NWA^{+}24, PTPB22, RGX^{+}24, RB19, SS24, SGJN24, SKR^{+}22, TY19, VTC20, WSY23, XFJ^{+}16, YSF^{+}18, YCL^{+}20, YBM^{+}21, YGH^{+}10, ZHC^{+}23, ZLG^{+}19, AMO05, BHDS09, BGM04, FNP09, GGB97, GPK^{+}09, HKL^{+}07, HBC^{+}08, JW08, KSK^{+}05, KG99, LP07, LVL03, MSB^{+}09, MLC08, ML09, RHA08, SSG12]. Hardware-accelerated [RB19, MLC08]. Hardware-Assisted [GFJ16]. Hardware-aware [HJY23]. Hardware-Based [BS14a]. Hardware-Efficient [ZLG^{+}19].

Hardware-Enabled [YSF^{+}18]. Hardware-Software [BM11, GGB97, HKL^{+}07, LVL03]. Hardware/Software [LHF12, CMM00, KTKO13, YGH^{+}10, AMO05, ML09]. Harmonic [Kha23]. Harnessing [RBWB20]. Hartley [HHX^{+}23]. Harvest [YB23]. Harvesting [SAL19, XPZ^{+}18]. hash [YTHC97]. Hashing [CJKK19, JCK^{+}18]. hazards [HA05]. HBM [PRKK21]. HBM-like [PRKK21]. healing [SS24]. Heap [JPM^{+}19]. Heaps [KLK^{+}17]. heartbeat [DHZ^{+}11]. heartbeat-detection [DHZ^{+}11]. Height [CZZYW21]. HeM [AJK^{+}21]. Heterogeneous [AJK^{+}21, CUA^{+}24, DHW^{+}23, ETAV18, GADG19, LY24, MBD^{+}20, RKKH24, RS18, SPT^{+}17, SVK17, SRKS23, SSL17, SAL19, SWT23, TBCH17, WTW^{+}23, XPX^{+}21, BWB14, CL99a, HV07, KJR^{+}07, LLKY13, PTC05, QS09, SCB01, SKS12]. Heterogeneously [ZP08]. Heuristic [AKAKP18, HGLC16, CLM^{+}10, LCKT12, OCRS07, SBGD13]. heuristics [TN99]. HEVC [SLV^{+}22]. Hidden [HYK^{+}20]. Hierarchical [CV17, HWL^{+}23b, JDD20, LMB^{+}12, LJ18, MSKBD07, OKJH22, SKR^{+}22, TZ17, WMT^{+}16, WLH20, XT16, BG01, HKV^{+}07, VKKR02, ZM07]. Hierarchy [CM19, FW00]. High [AKAKP18, Ali12, CYZL23, CSC^{+}21, CET16, CS22, CK16, DKT^{+}16, DBK^{+}18, DLC^{+}17, EKEK22, FCZ^{+}23, GHW^{+}12, HIW15, HSP^{+}22, ISK21, JD00, JDLZ24, Kha23, LLL^{+}18, LYKW09, LQD22, LQL^{+}24, MACV14, NSP^{+}20, PSD21, PRKK21, PTC05, PFHAH22, RCW22, RJ14, RM23a, RM23b, Sch17, SYH^{+}22, SS14, SLV^{+}22, VAAH^{+}98, WMT^{+}16, WS22, ZYW^{+}18, ZZ24, ZLG^{+}19, ACT13, AYM05, BHW^{+}13, BD00, CCC^{+}09a, GDTG07, GF06, GGDN04, GWR13, HJ08, JP08, KW02, KJT04, LJV02, LC14, Lin97, LFG^{+}09, MKBS05, MJM11, MLMM08, NS03, OW06, OWH08, PB14, RFYL98, SW12, SLXZ12, TC98, VKKR02, XK97, YWW10]. high-density [OWH08]. High-Dimensional [SYH^{+}22]. High-flexibility [JDLZ24]. High-Level [CET16, CS22, FCZ^{+}23, ISK21, RCW22, RJ14, Sch17, SS14, SLV^{+}22, JD00, NSP^{+}20, PTC05, PFHAH22, VAAH^{+}98, WS22, AYM05, BD00, GGDN04, HJ08, JP08, KW02, LC14, Lin97, MKBS05, MJM11, MLMM08, PB14, RFYL98, SW12, TC98, VKKR02, XK97, YWW10]. High-order [CYZL23]. High-Performance [DKT^{+}16, DLC^{+}17, LLL^{+}18, LQL^{+}24, WMT^{+}16, CYZL23, GHW^{+}12, LYKW09, GDTG07, GWR13, LJV02, LFG^{+}09, NS03, SLXZ12]. high-quality [BHW^{+}13]. High-Security [LQD22]. High-speed [PSD21, OW06]. High-Throughput [HIW15, EKEK22, PRKK21]. Higher [BS14a, LYSO19, XPSE12]. Highly [dONH23]. Hindsight [YXG^{+}24]. History [JM14]. History-Based [JM14]. Hits [SAL19]. HLS [SCL^{+}22]. Hmap [YTHC97]. HMP [SPT^{+}17]. Hold [LSZ^{+}21, KSA^{+}10]. hold-driven [KSA^{+}10]. holding [Pom14a]. Hole [YLZ^{+}17]. Holes [Pom21a]. Holistic [RGT^{+}14]. Hop [AL19]. HoPE [PBL^{+}17]. Hot [PBL^{+}17]. Hot-Cacheline [PBL^{+}17]. Hotspot [HDZ^{+}20, JYY^{+}22, LYM^{+}20]. HPC [LZA^{+}21]. Huffman [BH10, NT05, WZL^{+}21]. Huffman-based [BH10]. huge [HCK13]. huge-scale [HCK13]. Human [BHBS22]. Human-Readable [BHBS22]. HVAC [JDD20]. HW [ADP^{+}07, FLPP09, WWFT12]. HW-SW [ADP^{+}07]. HW/SW [FLPP09, WWFT12]. Hybrid [BLNK14, GD22, GCL^{+}16, HRC21, KKK12, LFST21, LZ17, LZ21, LYLW17, LV14, LGGJ14, MACV14, NAK20, PA21, SLXZ12, WSS^{+}18, ZMLH24, CLYP09, KT01, KKMB02, LCZ^{+}08]. Hypercube [TMDF10].

I/O [LC13, SLC^{+}22, Wu09, Yan16]. IC [ABC^{+}17, AYS20, BHLG19, DLK24, EK97, IK19, KK11, KKHK16, LCJ^{+}10, LTZ22, Ped96, WCB15, WXH^{+}19, WSS^{+}18, XGC^{+}20, ZLL13]. IC/MCM [EK97]. ICOS [HCLC98]. ICP [XGWL24]. ICP-RL [XGWL24]. ICs [CM18, CM19, CLT^{+}15, GSFT16, LHJ12, LS17, PKC^{+}21, THM15, VILSL23, WWCT18, YHH09]. IDDQ [TCP97]. Identification [LYL^{+}23, VTC20, DNA^{+}12, JDT^{+}08]. identify [LIA00]. Identifying [XGWL24]. Idle [LC07]. Idleness [GSD^{+}18]. IDs [SOS15]. IEEE [IIEKS23]. II [JW08, SA24]. IIoT [PTPB22]. ILA [HZS^{+}19]. illegal [LIA00]. ILP [GBK07, MRC06, MWG97, OCRS07, OK08, SR12, WPL23]. ILP-based [MWG97, OK08, WPL23]. Image [GAT^{+}21, RB19, WYIG07]. Imbalanced [HDZ^{+}20]. IMC [CYLC24]. Imitation [RKKH24]. Impact [GBK07, LDD^{+}19, MDR15, RB19, TY19, TWM^{+}23, XNZ^{+}15, KTKO13]. Impacts [LHS^{+}21]. implement [ADM^{+}13]. Implementation [ANS^{+}20, ALL17, BP23, HCRK11, JM14, KKLP15, LS22, LXGM23, MMM^{+}22, MAS16, ORGD^{+}15, SLV^{+}22, ZABGZ17, CD09, JWL^{+}03, KYN^{+}12]. Implementing [HKL^{+}15, KBA08]. Implication [LPLK22, WH20, WC06]. Implication-based [WH20]. implications [BLM00, DNA^{+}12, GGBZ02, ZLL13]. Implicit [PT06]. Imprecise [ENP20, PKP^{+}03]. Improve [KKLG15, Pom19b, WHXZ13]. Improved [DMR23, HWGY16, KKLP15, LWC18, Giv06, LV02, PDN97, Vah99, KO23]. Improvement [JGM14, KMO^{+}12, THM15, DD02]. Improvements [KAKSP16, VLH98]. Improving [ALLE20, CL13, CHC^{+}16, CJKS24, CWL^{+}22, KRS06, KYL16, RAKK12, TWM^{+}23, WDLD17, WSH^{+}18, WH19].

In-Cache [BFG^{+}19]. In-Memory [ZXC^{+}23, HHX^{+}23]. In-network [CXK^{+}13]. In-Order [ZBPF18]. in-place [KCKG13, YWW10]. In-Scratchpad [DFM15]. In-Situ [SL18]. inclusive [TZ20]. Incomplete [Pom19b]. Inconsistency [XPZ^{+}18]. Increase [KMR18]. Increasing [HW14, Pom22]. Incremental [BS14b, DNT20, EO19, HKV^{+}07, LYCP17, LNG^{+}16, SGGR14, WJM24, DVA02, LG12, LLM01, SMSB05]. Independent [Pom16b, VEO16]. Index [BC16, HCL^{+}14, HCK13]. index-based [HCK13]. Index-Resilient [BC16]. indexed [AC06]. indexing [Giv06]. indices [LCT03]. indirectly [AC06]. Indoor [MVK^{+}18]. Induced [CIX15, GSD^{+}18, LS19, LDX22, RRHB21, DHZL23, TCW20, SQL^{+}24]. Inductive [IPWW17, HMLL11, LXCH04]. Inductor [WDC^{+}22]. Industry [MCY23]. Inference [CBC22, HTC^{+}23, KZKAKP23, KLP^{+}24, LCG^{+}22, LPL^{+}21, MNMK^{+}21]. Inferencing [PGGD23]. Information [HMO^{+}14, NPH^{+}20, RRHB21, ZZL^{+}23, ZBPF18]. Informative [TEK18]. Initializability [CPR^{+}02]. Initialization [WL12]. Injection [CGV^{+}23, JIR^{+}21, LTZ22, MLH^{+}17, SQL^{+}24, BPRR98]. Input [JK10, LV14, PIK20, Pom16a, Pom16c, Pom21b, SRC15, BD05, BH03, CCW08, KM97]. Inputs [Pom18a]. Insertion [GMS^{+}23, HS19, LTW^{+}16, PSD21, SHL^{+}19, WZH^{+}23, CW01, JHL02, LXCH04, LLHT12, LCL08]. insertion/sizing [CW01]. Inspired [WSY23, GNQ^{+}22]. Instinctive [MVK^{+}18]. Instruction [HKL^{+}15, HZS^{+}19, KKMB02, LPD^{+}17, LCD07, LHF12, LF12, LXGM23, OT15, SEN05, TYSF20, AMR00, Hua01, KSK^{+}05, KTKO13, KHW06, LP03, LLHT03, LYCP13, LMW99, WH05]. Instruction-Level [HZS^{+}19, LXGM23, TYSF20, SEN05]. Instruction-Set [HKL^{+}15, LP03].

Instructions [KAKSP16]. Instrumentation [FHHR21]. Instrumenting [MPDG09]. Integer [ETAV18, TFW24, TZ17, GH00]. integer-programming-based [GH00]. Integrate [LLH^{+}17]. Integrated [HMLL11, HWX^{+}14, HS19, JNCS19, KK14, KO23, KLE18, LLM^{+}23, LZ21, NCP01, PLC24, RGM15, SHD17, BWB14, LFG^{+}09, XTW05]. Integrating [BMdG17]. Integration [APD^{+}11, AJK^{+}21, BPTB17, BRCS18, CUA^{+}24, IGN18, JHMGS18, TMDF10, YD16, DL11, LHZ^{+}06, SSP04]. Integrity [CUA^{+}24, DCC^{+}23, FHHR21, FHHH22, XRS^{+}19, ZF23, XZC09, YHH09]. intellectual [KHP05]. Intelligence [KAC^{+}23, MVK^{+}18]. Intelligent [KP22, LJZ^{+}24, HCLC98]. intensive [KCA04]. intent [SDP^{+}09]. Inter [DJP21]. Inter-tile [DJP21]. interacting [NCP01]. interactive [SCV06]. intercluster [GBK07]. Interconnect [DHNR23, HCZ^{+}16, LKLC22, MSB^{+}09, WTR12, XS16, YLY^{+}23, HR06, HLHT08, JPCJ06, SY07]. Interconnection [GADG19, CFX09]. Interconnections [GNQ^{+}22, KM97]. Interconnects [WM24, CML98, CH96, XZC09]. Interface [HLL^{+}24, LZZ23, LHLP16]. Interfaces [PMP17]. Interference [CIX15]. Interleaving [SPC^{+}15]. intermediate [LTH99]. Internal [BDB12, Yan19]. Internet [DP04, TK18]. interpolation [CMNQ08, YHL^{+}11]. Interposer [WCB15, WWCT18]. Interposer-Based [WCB15, WWCT18]. Interrupt [JP08]. Interrupts [Ali12]. Interval [PIK20, ST99]. Intra [SLV^{+}22]. intrasignal [KCKG13]. Intrinsic [HRK18, SCJ01]. Introducing [PGB01]. Introduction [ADGSM22, BC08, BJX15, CCY22, CO18, CLQ12, Har05, HAW20, HJ08, JCPL23, JW08, LP07, LZR23, MYSZ23, NWA^{+}24, Ped06, PFHAH22, RW03, RBA^{+}12].

Introspection [KI01]. Intrusive [LL15, SL18]. Invariant [Pom18b, PL03]. Invariants [IPWW17]. Inversion [LHW^{+}17]. Inversion-Aware [LHW^{+}17]. inverted [DH06]. Inverter [VEO16]. Investigating [RB19]. Investigation [XLNB17]. IO [Yan11]. IoT [BSP^{+}22, CCMC20, CARH18, MMM^{+}22, PTPB22, XLNB17, YB23, YFT17, YFT18]. IoT/IIoT [PTPB22]. IP [BTP^{+}20, BFV15, ISK21, JHMGS18, RGX^{+}24, SLP^{+}19, SSGS03]. IP-Integration [JHMGS18]. IPs [GSFT16, LLH^{+}17, LG18, Sch17, VBP^{+}19]. Irregular [CLX^{+}23, KCKG16, KCKG13]. ISAs [SBH^{+}06]. Ising [MS21]. Ising-FPGA [MS21]. Island [GMS^{+}23, LCY12, GM08]. Islands [JPHL16]. Isolation [CCS15]. Issue [ADGSM22, BJX15, HAW20, LZR23, NWA^{+}24, TK18, BC08, LP07, Ped06, Ped11]. Iteration [CZZYW21]. Iterative [KLV15, Yan20, DD02]. iTimerM [LJ18].

JAMS [KPB19]. JAMS-SG [KPB19]. Java [BHDS09, PSL^{+}98]. JETC [BC08]. JETC/TODAES [BC08]. Jitter [KPB19, ZZ24]. Jitter-Aware [KPB19]. joint [BC08]. Jointly [CCK^{+}18, GYT12, XXC^{+}24, ZLW^{+}15]. Journal [SN10]. JPEG2000 [GFC^{+}09].

kEP [BCC08]. kEP-SOPs [BCC08]. kernel [EKEK22, WKR09]. Kernels [MLH^{+}17]. Key [ISK21, JZG21, ZZL^{+}23]. Key-based [JZG21]. Key-Obfuscated [ISK21]. knapsack [SBGD13]. Knowledge [EO19, ZHC^{+}23]. Knowledge- [EO19].

L [LM96, Meh98, OKJH22]. L-shaped [Meh98]. L-shapes [LM96]. L0 [KJR^{+}07]. L2 [SYX12, TYSF20]. Lab [PGCB16]. Lab-on-Chip [PGCB16]. Lagrangian [LGGJ14, PY20, ZBG^{+}23]. Language [TAP^{+}24, MSD06, MLC08, PHM00, RHN00]. languages [BGM04, Edw03, SSG12]. Large [CK19, CSX^{+}05, DNT20, GNQ^{+}22, JZYZ15, LYL^{+}19, NDA^{+}23, TAP^{+}24, WTW^{+}23, YVC14, ZHC^{+}21, ZGB^{+}24, AM10, DD02, HH09, MRB^{+}11, SCB01]. Large-Scale [LYL^{+}19, YVC14, CSX^{+}05, GNQ^{+}22, WTW^{+}23, ZGB^{+}24]. Last [KLJ14, SABSA15, SAL19, CXK^{+}13]. Last-Level [KLJ14, SABSA15, SAL19]. Latch [JNCS19, Kha23, LCHT02]. latch-based [LCHT02]. late [LG12]. Latencies [Sch17]. Latency [LWX^{+}23, QBTM16, YKCG14, ZYPC17, PMT20, WHXZ13]. Latency-aware [LWX^{+}23]. Latency-Minimal [ZYPC17]. Lattices [GSS14, HMO^{+}14]. Launch [Pom21b, PTC^{+}15, WWW^{+}12, XCW12, Xia24, WPHL08]. launch-off-shift [WPHL08]. Launch-on-Capture [XCW12, Xia24]. Launch-On-Shift [PTC^{+}15, Pom21b, WWW^{+}12]. Launch-to-Capture [PTC^{+}15]. Layer [DHZL23, LYCP17, MWS^{+}20, WL12, Yan17, Yan20, CLYP09, DDNAV04, EKEK22, OW06, Yan00, Yan19]. Layer-induced [DHZL23]. Layout [CFD^{+}16, DZ18, HWF^{+}23, JYY^{+}22, LZ17, LCYN18, RCK^{+}15, SPC^{+}15, TZ20, WPHL08, WPR^{+}19, XK97, YLZ^{+}17, ZLY^{+}15, GS00, GH00, KG09, WJYZ11]. Layout-Aware [RCK^{+}15, WPHL08]. Layout-driven [XK97]. Layouts [GMS^{+}23, GFC^{+}09, LM96]. Lazy [ZLW^{+}15, ZLW^{+}15]. Lazy-RTGC [ZLW^{+}15]. LBNoC [PMT20]. LDE [TZ20]. LDE-aware [TZ20]. LDOs [SCK18]. LDPC [CWL^{+}22, DHZL23]. leaf [dW97]. Leak [PCT^{+}17]. Leakage [CFHM09, DHB16, HYN15, JK10, LDX22, PIK20, PS23, RRHB21, STWX12, SYHL14, SKP21, XT16, YYLL09, ZBPF18, CS07, CCW08, KOS09, MLG12, YLL06]. Leakage-Aware [SKP21, YYLL09]. Learn [RG19]. Learned [XFJ^{+}16]. Learning [ALLE20, CLL^{+}22, CAOM19, CCMC20, CJKS24, DNT20, EW18a, GT21, HDZ^{+}20, HAW20, HZJC23, HMMG^{+}20, HXC^{+}18, HFMB20, HHH^{+}21, HC23, IE12, JBJ22, KP22, KLP^{+}24, LG18, LYHL14, LZY^{+}23, LJZ^{+}24, LCZ^{+}24, LZR23, LG23, LPL^{+}21, LRHL24, MBD^{+}20, MYSZ23, NL24, NDA^{+}23, NSP^{+}20, PJL14, QZZW24, RKKH24, RNA^{+}21, RPR^{+}21, SKR^{+}22, SCK^{+}23, SWT23, SQL^{+}24, SAHF^{+}20, TEK18, WH19, WLH20, WS22, WDLX21, XAG^{+}20, XXC^{+}24, XGWL24, ZHL^{+}23, ZHC^{+}23, ZKS^{+}16, ZHC^{+}18, ZPLI23, CXS^{+}23, STL^{+}13]. Learning-Based [LG18, HFMB20, LJZ^{+}24, LG23, SWT23, XAG^{+}20]. Learning-to-Search [NDA^{+}23]. Least [JLJ15]. Legalization [CZZYW21, HNS23]. Legalizer [DBK^{+}18, DBK^{+}18]. length [CCC09b, Con06, LCT03]. Lens [KPSW09]. Lessons [XFJ^{+}16]. LET [WLZ^{+}19]. LET-Based [WLZ^{+}19]. Level [CDB11, CET16, CS22, CLMZ10, DKZ^{+}15, FCZ^{+}23, HKL^{+}15, HMO^{+}14, HZS^{+}19, ISK21, KLJ14, LL15, LG18, LS11, LXGM23, MNMK^{+}21, PDS12, Pie16, RCW22, RJ14, SABSA15, Sch17, SS14, SLV^{+}22, SAL19, TYSF20, VTC20, WDLD17, WCZ^{+}24, WJM24, AYM05, BdM00, BD00, CM19, CCYC14, CIB01, CXK^{+}13, Che96, GM08, GG99, GS00, GGDN04, HJ08, HLL^{+}24, JD00, JR97, JP08, JT98, KI01, KRK98, KW02, LC14, LLQ^{+}03, LTPT10, Lin97, MW97, MOZ06, MKBS05, MT02, MJM11, MLMM08, NSP^{+}20, OCRS07, PB14, PPDK09, PTC05, Ped06, PFHAH22, PBSV^{+}06, RFYL98, RFG20, SW12, Sen11, SEN05, TC98, TJ99, Vah99, VAAH^{+}98, VKKR02, VS12b, VBP^{+}19, WTL^{+}13, WS22, XK97, YWW10, ZHM07, ZLL13]. Leveling [CCH^{+}15a, CHC^{+}16, Kha12, CD09]. levelized [KPR06]. Levels [BFL10]. Leveraging [CS22, DSHD23, SQL^{+}24]. LFSR [KJT04, Pom17a, Pom18b].

LFSR-Based [Pom17a, Pom18b]. Libraries [ACF^{+}11]. Library [KRH18, KKS16, MCZ^{+}16, BD97, DDNAV04, JD00]. Library-Based [MCZ^{+}16, DDNAV04]. lifecycle [HDL^{+}12]. Lifetime [AAA15, DLC^{+}17, SJN24, WDLD17, MHT14]. Lightning [SQL^{+}24]. Lightweight [MPM^{+}17, NSCM17, MMM^{+}22]. like [PRKK21]. limitations [Voe01]. limited [LLKC13]. line [SNH02, ZYZ^{+}13]. Linear [ACFM12, CGV^{+}23, ETAV18, MFHP12, TZ17, DSRV02, KC98, LWK11, ST99]. Linking [HRC21]. Links [KQP^{+}19]. list [HCS01, MHD^{+}04]. list-approximation [HCS01]. lists [HVF^{+}01]. Lithographic [LYM^{+}20]. Lithography [HDZ^{+}20, LZ17, ZLY^{+}15]. liveness [MS08]. LLC [PBZM19, SJ23]. LLCs [PBL^{+}17]. LLR [CWL^{+}22]. Load [CLC20, LLHT12, Pom19a, Pom14b]. Load-balanced [LLHT12]. Local [BM11, KC13]. Locality [LDLM20, MT15, TYSF20, ZFLS11, GFC^{+}09, Kan06]. Locality-Aware [MT15]. Locality-Driven [ZFLS11]. Localization [HDB22, YYL^{+}15]. localized [CMNQ08]. Locally [PMS15, KC13]. Locked [IYF^{+}21, JZG21]. Locking [BTP^{+}20, Mit16]. Lockout [ISK21]. Logic [ALLE20, AYS20, BFL10, CBMM10, Che18, CZW19, CXS^{+}23, ETAV18, EKS^{+}14, HS18, HIW15, JZG21, KKH^{+}02, KMO^{+}12, LWZ^{+}19, LSZ^{+}21, LWC18, PA21, QZZW24, SLP^{+}19, WB16, WCZ^{+}24, WKC12, ZHJ^{+}23, ZWD11, ARLJH06, BLM00, BDM^{+}99, BOC00, CSKR05, CD96, GGBZ02, KJKK03, KMC97, KVMH08, LWH06, MW97, RJBS09, TW96, TN99, TJ99, VKT02, WVYG99, ZS02, PRCK08]. Logic-Based [ETAV18]. Logical [SJ23]. logics [BD05]. long [SSP04]. long-path [SSP04]. Longevity [KBV^{+}15]. Look [KSD^{+}22]. Look-up-table-based [KSD^{+}22]. Lookahead [PMT20]. lookup [CH02, WSEA99]. Loop [AA17, EO19, GDD21, LDLM20, SXX^{+}06, HKV^{+}07, PCC09, XPSE12]. Loop-dominated [LDLM20]. Loops [IYF^{+}21, BG01, CL99a, KNDK96, SHLL98]. Lose [KBV^{+}15]. Loss [WSRH16, KC13]. Losses [ZMS^{+}19]. Low [ACF^{+}11, AYS20, ALL17, BPTB17, CH10b, CM08, CHHL96, CLMZ10, DMR23, GBR07, GAT^{+}21, HWDQ22, HLKN07, HTCP13, JDLZ24, KP22, Kha23, KLP^{+}24, LTYW12, LS23, LSL^{+}13, LSZ^{+}24, LQD22, LS17, MED23, MKK13, MACV14, PMT20, PMB10, Pom14b, RFB10, RM23a, SMS22, SYH^{+}22, SCK^{+}23, SESN15, TWL16, TMDF10, WGT^{+}17, WPR^{+}19, YKCG14, ZK15, BD00, BPRR98, CH10a, CCX06, DS06, GOC02, HLCH07, HCK13, JWL^{+}03, KBN09, KKH^{+}02, KJR^{+}07, KHW06, KYN^{+}12, LLHT03, LYCP13, LHW97, ML09, RTNL05, SUC01, TJ99, YGZ04, ZYDP08, ZP08]. Low-Complexity [LTYW12]. Low-Cost [LSZ^{+}24, JDLZ24, BPRR98, HCK13]. Low-coverage [WPR^{+}19]. Low-data [LS23]. Low-energy [LSL^{+}13, MED23]. Low-Latency [YKCG14, PMT20]. Low-Overhead [KLP^{+}24, LQD22, PMB10]. Low-Power [ALL17, BPTB17, CH10b, CLMZ10, GBR07, GAT^{+}21, HWDQ22, LS17, TWL16, TMDF10, WGT^{+}17, ZK15, CM08, HTCP13, KP22, MKK13, Pom14b, RFB10, SMS22, BD00, CH10a, DS06, GOC02, HLCH07, JWL^{+}03, KBN09, KKH^{+}02, KHW06, KYN^{+}12, LYCP13, ML09, RTNL05, SUC01, ZYDP08, ZP08]. Low-Rank [SYH^{+}22]. Low-Voltage [DMR23, SCK^{+}23]. Lower [HWF^{+}23, LC96, TC98]. Lower-bound [HWF^{+}23, LC96]. Lowering [JLK15]. LSTM [CBC22]. LUT [CD96, CH00, CYLC24, KNRK06, LKM04, VKT02]. LUT-based [CH00, KNRK06, LKM04, VKT02]. LVS [LBV^{+}06].

MAC [BS14a]. Machine

[ALLE20, BHBS22, CAOM19, CCMC20, CXS^{+}23, CJKS24, DNT20, EW18a, HAW20, HMMG^{+}20, HXC^{+}18, HHH^{+}21, HC23, IE12, KP22, LYHL14, LZY^{+}23, LZR23, MYSZ23, NSP^{+}20, RPR^{+}21, SCK^{+}23, SAHF^{+}20, XAG^{+}20, ZHC^{+}18, ZPLI23, CK96, KMC97, MMP00, PHM00, MSR09]. Machine-Learning [ZHC^{+}18]. Machine-learning-driven [CXS^{+}23]. Machines [DMR10, BDC08, CHHL96, MS08, BHDS09]. Macro [LJ18]. macrocell [CHY05]. Macromodel [SHD17]. MAESTRO [RGT^{+}14]. Magnetic [WDC^{+}22]. Magneto [AKM^{+}22]. Magneto-Electric [AKM^{+}22]. Main [AAA15, BLNK14, NAK20, PBZM19]. Makespan [SRKS23]. Making [TCW20, XLNB17]. Managed [KLK^{+}17]. Management [ABC^{+}17, BM11, CHBK15, DLC^{+}17, DMR10, GCL^{+}16, HC17, HXC^{+}18, JPM^{+}19, KKLG15, LHW^{+}17, LZA^{+}21, MBD^{+}20, MDR15, NDA^{+}23, PJL14, PSP24, PBZM19, SKP21, SAHF^{+}20, VA17b, WMT^{+}16, WXH^{+}19, YB23, AHAKP08, ADDM^{+}13, AMM^{+}06, ANR13, BHDS09, BMJ13, CLQ12, DS05, FHHG12, GK14, HCK13, IBMD07, LMB^{+}12, STL^{+}13]. Managing [TY19, BD08]. Manhattan [DSKB04]. Manhattan-diagonal [DSKB04]. manipulation [CCQ98, Zho08]. manufacturability [WPR^{+}19]. Manufacturing [MCY23, YCL^{+}20]. Many [CAOM19, GD22, SA24, SESN15, WMT^{+}16, WDLX21, ZHC^{+}21]. Many-Core [CAOM19, SESN15, WMT^{+}16, GD22, WDLX21]. Manycore [AJK^{+}21, KLK^{+}17, NDA^{+}23]. Manycore-Based [KLK^{+}17]. mapper [YTHC97]. Mapping [CPS16, CGLH23, ETAV18, GT21, GYZ^{+}22, HABS15, HAB^{+}17, HJY23, JBJ22, LFST21, SWT23, VNS19, WDD^{+}23, XGC^{+}20, ZYPC17, CSL^{+}07, CH02, CH00, CHY05, JP12, JD00, KL05, LKM04, MBB01, PL98, SKS12, WY06, WSEA99, ZS02]. Marching [CCH^{+}15a]. Marching-Based [CCH^{+}15a]. Markov [CB17]. Masking [CYZL23]. Massively [ZWD11]. Matched [LCYN18]. Matching [CLC20, HWF^{+}23, MS17, THM15, WLLH16, ZLG^{+}19, BD97]. Matching-based [HWF^{+}23]. Mathematical [LHC24]. MATLAB [LPD^{+}17]. matrices [KVMH08]. Matrix [CLT^{+}15, CZZYW21, LCJ^{+}22, LXWC20, LKC^{+}18]. Matrix-Based [LKC^{+}18]. Maximization [LM21]. Maximize [CS22]. Maximizing [BH22, HHK^{+}17]. MaxSense [LM21]. Maze [LLLL18, JCGP05]. MCC [YYG^{+}16]. MCEmu [THT12]. MCM [EK97]. MCMM [EK16]. McPAT [LLK^{+}14]. MCUs [MRB^{+}11]. MDE [ORGD^{+}15]. mean [Das04]. Measurement [APDC17, CRT19, JB98, XAG^{+}20, LG12]. Measurement-Based [APDC17]. Measurements [LFST21, LYSO19]. Measuring [CHA^{+}23, WAZ98]. MEC [LJZ^{+}24]. Mechanical [BHLG19, LTW^{+}16]. Mechanism [QSW^{+}15, SVK17, WQC^{+}16, ZLW^{+}15, ZK15, Wu09]. Mechanisms [CBO^{+}18, PTPB22, GBK07]. MEDA [KR23, LSCK20, PBWB21]. Media [SLV^{+}22]. Medium [MED23]. MEDUSA [ZPLI23]. MeF [AKM^{+}22]. MeF-RAM [AKM^{+}22]. memetic [LFG^{+}09]. Memories [AAA15, DFM15, DHZL23, JSA18, LS23, LSZ^{+}24, SKP21, JD00, MRB^{+}11, NR03, OK08, RMB10, SPG^{+}08]. Memory [AKM^{+}22, BLNK14, BD14, CPS16, CCK^{+}18, CIX15, CLX^{+}23, DFM15, DHX^{+}23, HJY23, JCK^{+}18, JPM^{+}19, KLSP11, KKLG15, LHS20, LDP^{+}22, LZZ23, LLP^{+}16, LCJ^{+}22, LWZ^{+}19, LPL^{+}21, MWS^{+}20, MS23, NAK20, NM23, PDN97, PPP^{+}15, PRKK21, PBZM19, RPR^{+}21, SHBD21, SSL17, TLCF16, TRM^{+}16, TMDF10, VFML23, WQC^{+}16, WDZG16, WFT^{+}19, WDD^{+}23, WGSH16, WZL^{+}21, XNZ^{+}15, ZXC^{+}23, ZLW^{+}15, ZZCY17, AMM^{+}06, BD08, BHDS09, BGN^{+}07, CPW04, CJLZ11, HHX^{+}23, HKV^{+}07, IBMD07, JCS^{+}08, Kan06, KG09, LSPC14, MB04, NdLCR03, OKC08, PDN00, PCD^{+}01, SUC01, SM00, WH05, Wu09, ZYZ^{+}13, ZP08]. Memory-aware [DHX^{+}23]. Memory-Based [BD14, CPS16, LWZ^{+}19]. memory-constrained [OKC08]. Memory-driven [NM23]. Memory-Throughput [MS23]. Memristive [BXG^{+}24, KZKAKP23, WSY23, XGC^{+}20]. Memristive-based [KZKAKP23]. Memristor [LS22]. MEMS [BHLG19, Kha12]. MEMS-IC [BHLG19]. Merging [ASAP17, CZW19, TCL14, LLLC13, MB04]. Mesh [CHA^{+}23, JM14, KK14, GHW^{+}12, RL13]. MESO [ZXC^{+}23]. Message [Hu20, KPB19, DSH12, EY12]. message-passing-based [EY12]. metamodeling [MPSJ07]. Metastability [PBH^{+}24]. Method [AKAKP18, BZWZ17, CZZYW21, JSS^{+}19, KO23, LCC^{+}15, MNMK^{+}21, RGM15, SYH^{+}22, SRC15, STGR15, WTR12, WMT^{+}16, WZL^{+}21, YLZ^{+}17, ZYW^{+}18, ZPLI23, CGN96, CL99b, HW00, Kag05, LH13, LDK99]. Methodologies [PLH^{+}24, BW00, CEB06, MD13, SSCS10]. Methodology [BFV15, DK22, EKEK22, EAP17, GMS^{+}23, HXB^{+}22, KKLP15, KJR^{+}07, KMO^{+}12, LW17, LSZ^{+}21, LZ21, LZZSV15, LLLL18, NSP^{+}20, SWT23, VA17a, VEO16, VBP^{+}19, WCZ^{+}24, XPX^{+}21, AMM^{+}06, DRG98, FLPP09, HDL^{+}12, HCLC98, Hsi00, KYN^{+}12, NR03, PW99, SEN05, SMSB05, SZV^{+}12]. Methods [CLL^{+}22, EW18a, GDF09, KRL15, ZHC^{+}18, FZKS11, SW04, ZAJ^{+}12]. Metric [YRH11]. Metrics [LIK22]. Micro [Kha23, RBWB20, YBM^{+}21]. Micro- [Kha23]. Micro-/Nano [Kha23]. Micro-architecture [YBM^{+}21]. Micro-Electrode-Dot-Array [RBWB20]. Microarchitectural [GOC02, LS11, HMLL11]. Microarchitecture [BSZ^{+}24, ZBPF18, CFX09]. microcontrollers [CD09]. MicroElectrode [RB21]. MicroFix [YHL^{+}11]. Microfluidic [CPK20, CGLH23, GLD^{+}22, GHYR19, JYHY21, KGS^{+}20, LHC16, LCZ^{+}24, LKC^{+}18, MGR^{+}15, MWK21, PGCB16, PBF^{+}22, RCK^{+}15, RB21, SKS^{+}18]. microfluidics [SOC06, SC06]. microfluidics-based [SOC06, SC06]. Microgrid [VA17a]. Microprocessor [OT15, BPRR98, HV98, LBV^{+}06, WAZ98, WWC04]. microprocessor-based [BPRR98]. Microprocessors [Ali12, WMT^{+}16, LTPT10, MKW09, VAAH^{+}98, WTL^{+}13]. Migration [DK16, Kha12, TZ20]. Migration-Resistant [Kha12]. million [HH09]. million-gate [HH09]. Min [HS18, SSP04]. Min-Area [HS18, SSP04]. min-delay [SSP04]. Mine [LWC18]. Minimal [MCD12, ZYPC17, KL05]. minimal-area [KL05]. Minimally [EKEK22, RNA^{+}21]. Minimization [HYN15, KR23, PIK20, WB16, AMR00, CSAHR07, CGN96, CCC09b, HPK99, HCS01, HCN09, KC13, LXCH04, LKM04, LDK99, LWH06, LC07, MRC06, OK08, Ped96, PR96, QS09, SXX^{+}06, TJ99, ZYP09]. Minimizing [GSD^{+}18, KOS09, PKJK20, SRKS23, TPC^{+}17, WDZG16, WC10, KT96]. Minimum [BFL10, HYN15, JLK15, KJKK03, FNMS01, MS00, ZCG06]. minimum-area [MS00]. Minimum-Energy [BFL10]. Mining [LJZ^{+}24, LWC18]. Mismatched [WPL23]. miss [TY97]. Missing [HDB22]. Mission [BSP^{+}22]. Mistakes [DHB16]. Mitigate [JIR^{+}21, MDR15, RJBS09]. Mitigating [KS23, LHS^{+}21, MRB^{+}11, VFML23]. Mitigation [BFL10, HWL^{+}23a, KD24, KRL15, MRL^{+}20, HMLL11]. Mixed [BB17, CZZYW21, CYH19, HRC21, IGN18, KZKAKP23, KMR18, SZB17, TFW24, YVC14, ZABGZ17, ZSY18, AM05, KOS09, MS00, YWGI09]. mixed- [KOS09]. Mixed-Cell-Height [CZZYW21]. Mixed-Critical [IGN18, KMR18]. Mixed-Criticality [BB17, CYH19, SZB17, ZABGZ17]. Mixed-Signal [HRC21, STGR15, ZSY18, GMS^{+}23, KZKAKP23]. Mixed-Size [YVC14, AM05]. Mixing [KR23]. Mixture [RCK^{+}15, SKS^{+}18]. ML [LYM^{+}20]. ML-based [LYM^{+}20]. MLC [JSA18, KYL16, MWS^{+}20, PPP^{+}15, SJN24]. MM* [LH14]. MNFTL [MWS^{+}20]. Mobile [CLL^{+}22, GYZ^{+}22, JZYZ15, LPLK22, LKH19, YPCF17, ISE08, JBC^{+}10]. MoC [MPSJ07]. Mode [EAAK^{+}23, EK16, JOH17, KKS16, KS23, SLC^{+}22, UE22, LC07]. Mode-benefited [SLC^{+}22]. Mode-Reconfigurable [UE22]. Mode-switch [KS23]. Model [AVG19, Ase23, CLH12, CCH15b, CB17, EAP17, GFJ16, GGB97, JJH21, KW16, KLP^{+}24, LH14, LJ18, LCG^{+}22, LOC12, LY24, MS21, SZB17, TFW24, TAP^{+}24, XLNB17, YWGI09, YMB15, BLR06, BK10, BH03, CNQ13, CH13, CK96, LLQ^{+}03, MP07, MCMW08, PWY05, RS98]. Model-based [JJH21, MP07]. Model-Centric [XLNB17]. Model-Driven [EAP17, LOC12]. modeled [ARLJH06]. Modeling [BKW15, BLUS19, CVMP19, GS00, GCZ^{+}15, LG18, LLK^{+}14, LLQD23, PSL^{+}98, QBTM16, RGT^{+}14, RPR^{+}21, SSS^{+}19, TWL16, WTR12, WGT^{+}17, ZHL^{+}23, BBD00, JP08, LMW99, LON08, LVL03, MPSJ07, PTC05, RHN00, RFYL98, Rak09, SKCM06, VAAH^{+}98, VLGG01, WTL^{+}13, WJY^{+}07, ZM07]. Models [APD^{+}11, APS18, BBEM15, BFG17a, HHL14, LS23, LFST21, LS22, MA16, RG19, WLM21, YLY^{+}23, YBM^{+}21, ZABGZ17, GMSSS02, LTPT10, MRC06, SGD10, SMSB05]. Modern [DKT^{+}16, NTSA18]. Modification [JK10, PAV17]. Modified [DMR23]. Modular [GAT^{+}21, ZMS^{+}19]. Module [HRC21, LCYN18, SC06, WCZ^{+}24, CCX06, SCJ01, TW96]. Module-Level [WCZ^{+}24]. Module-Linking [HRC21]. Modules [DHW^{+}23, CWW96, CZW^{+}03, KT96, OWH08]. Modulo [PG15]. Modulus [CZZYW21]. Modulus-Based [CZZYW21]. MOEA [SA24]. MOEA/D [SA24]. Monitoring [FYCT15, LL15, LHLP16, LLH^{+}17, SL18, APB^{+}08, CXK^{+}13, CBR^{+}05, KP13, WJY^{+}07]. Monitors [VBP^{+}19]. Monolithic [AJK^{+}21, DLK24, LDD^{+}18, LDD^{+}19, PKC^{+}21]. Monotone [DPNB02]. Monster [FHHH22]. Monte [FZL^{+}23, GLY^{+}12, ZFL22]. morphing [RAKK12]. MOS [ZK15]. MOSFET [BFL10]. motes [RFB10]. Motion [FG18, ZLG^{+}19, DHV^{+}00, KMS12]. Movable [TFW24]. Movement [HWGY16]. MP [CRC15]. MPSoC [BGN^{+}07, GK14, KKJ^{+}08, KH10, SGD10]. MPSoCs [ADP^{+}07, DJP21, EWT23, LFST21, MRL^{+}20, MHT14, RGT^{+}14, SKS12, SSL17, SWT23, YP10]. MRAM [JZYZ15, SMBT19]. MSG [WY06]. MTCMOS [HLCH07]. Muller [ZHJ^{+}23]. Multi [BS14c, CYH19, EKEK22, ETAV18, EWT23, GACK22, HC17, JOH17, KGS^{+}20, KLE18, KR23, LFST21, LWG^{+}23, LG23, PBWB21, PBF^{+}22, PY20, RKKH24, SA24, SFM^{+}19, SBY^{+}20, SCL^{+}22, VILSL23, WFSS20, WZH^{+}23, WDLX21, XXC^{+}24, ZLY^{+}15, ZGB^{+}23, ZHJ^{+}23, ZPLI23, dONH23, CNQ13, HGBH09, HMB98, KOS09, MPSJ07, PB14, Pom14a, RAKK12, SZV^{+}12, Wu09]. multi- [KOS09]. multi-bank [Wu09]. Multi-bit [VILSL23]. Multi-chip [WDLX21]. Multi-Core [CYH19, ETAV18, SBY^{+}20, LG23, RAKK12, SZV^{+}12]. Multi-Cores [RKKH24, SFM^{+}19]. Multi-Cycle [WZH^{+}23, Pom14a]. multi-engine [CNQ13]. Multi-Fidelity [WFSS20, SCL^{+}22]. Multi-FPGA [BS14c, LWG^{+}23, PY20]. Multi-kernel [EKEK22]. multi-MoC [MPSJ07]. Multi-Mode [JOH17]. Multi-Objective [GACK22, KLE18, SFM^{+}19, dONH23, EWT23, LFST21, SCL^{+}22, ZGB^{+}23, PB14]. multi-phase [HMB98]. multi-processor [HGBH09]. Multi-Resolution [ZPLI23]. Multi-Start [ZLY^{+}15]. Multi-strategy [ZHJ^{+}23]. Multi-Target [KGS^{+}20, PBWB21, PBF^{+}22, KR23]. Multi-Task [XXC^{+}24]. Multi-threaded [HC17]. Multi/Many [SA24]. multibank [WH05]. Multicast [WWCT18, XS16, XCF18]. multichip [OWH08]. Multicore [BM11, CRC15, DFM15, HWX^{+}14, JPHL16, KLSZ11, LS11, LHK^{+}15, LMA^{+}16, QBTM16, SPT^{+}17, SAL19, THT12, WDZG16, XPX^{+}21, BHW^{+}13, CNQ13, DSH12, HDL^{+}12, KP13, LTPT10, Ped11, QM12, SNL12, WTL^{+}13]. Multicycle [Pom15a, Pom20, Pom13]. multidimensional [SBGD13]. multidomain [AM10, BMJ13]. multifunctional [AM10]. Multigrid [LAYZ23]. Multiharmonic [WGT^{+}17]. Multilayer [DLK24, KKHK16, LLLL18]. Multilevel [HBPW14, JZYZ15, PJL14, ZF23, JCS^{+}08, SGK08]. multilevel-cell [JCS^{+}08]. multimedia [HKL^{+}07, ZHM07, ZHOM08]. multimetric [HR06, RGM09]. Multimode [SSGS03]. multiplane [AJM13]. Multiple [BM11, GYT12, GPS^{+}24, KRL15, OKJH22, Pom16b, SRC15, WC06, YLZ^{+}17, CH96, GM08, JR97, KFH^{+}08, LBV^{+}06, LLHT12, MRB^{+}11, MR05, NdLCR03, PT06, PMB10, RMKP03, RM09, SBGD13, WLT08, WLCJ09, WSEA99]. multiple-bit [RM09]. multiple-choice [SBGD13]. multiple-output [WSEA99]. multiple-project [WLT08]. Multiple-Supply [BM11]. Multiple-Transient [KRL15]. Multiplexed [LHC16, LM19]. Multiplexer [Pom18a]. Multiplexing [LWG^{+}23, PY20]. Multiplication [GYT12, JDLZ24]. Multiplier [EKEK22, SMS22, WCX^{+}24]. Multiplier-divider [EKEK22]. Multiplierless [ACFM12, AFM14]. Multipliers [CXS^{+}23, RMPJ08]. multiprocessing [ZM07]. Multiprocessor [CHBK15, CH17, JOH17, KFH^{+}08, NSH^{+}16, APB^{+}08, DCK07, DCK09, DCK10, HCLC98, Kan06, MOZ06, WLL^{+}11, WG11, ZAJ^{+}12]. Multiprocessors [HAB^{+}17, JGM14, KBV^{+}15, PJL14, IAI^{+}09, PTC05, ZYDP08]. Multirate [ZABGZ17]. Multistage [Shi20, LON08]. multistandard [CCC^{+}09a]. Multitarget [SKS^{+}18]. multitasking [NG06, PW99]. Multiterminal [UPV23, JCGP05, MW97]. Multithread [SYHL14]. Multithreaded [HPB11]. Multiversion [HCL^{+}14]. multivoltage [CCX06]. Multiway [FW00]. mutually [DK08]. Mux [BH22]. MVP [LCJ^{+}22].

n [RG19, PR07]. N-detection [PR07]. NAND [CWL^{+}22, MWS^{+}20, PPP^{+}15, WQC^{+}16, WZL^{+}21, ZLW^{+}15]. Nano [Kha23]. Nanometer [BFL10, BPTB17, STWX12]. Nanophotonic [LKLC22]. nanoribbon [YMC^{+}13]. Nanotube [WSH^{+}18]. NAS [NGL^{+}21]. Native [LS22]. Navigation [MVK^{+}18]. NBTI [BDB12, CMP10]. NBTI-Aware [CMP10]. Near [KCKG13, LCJ^{+}22, PRKK21, SHN12]. Near-Memory [PRKK21]. Near-optimal [KCKG13]. near/sub [SHN12]. near/sub-threshold [SHN12]. Nearest [PSD21]. Negative [LHS^{+}21]. Negatives [AL19]. Negligible [EAAK^{+}23]. Neighborhood [PSD21].

Neighborhood-aware [PSD21]. Nested [AA17, CL99a]. Nesterov [LCC^{+}15]. Net [Yan19, LXCH04, MW97]. nets [JCGP05]. Network [Ase23, CM20, CHK^{+}23, CARH18, DJP21, DNT20, DCC^{+}23, EJR22, FLG^{+}23, HZL^{+}22, HCZ^{+}16, HXC^{+}18, HC23, KZKAKP23, KLK^{+}17, LDD^{+}18, LDD^{+}19, LW17, LJJ^{+}22, LJL^{+}23, MT15, NHS23, PMT20, WXH^{+}19, WDLX21, XS16, XCF18, YKCG14, YLY^{+}23, YXG^{+}24, ZHC^{+}21, ZYS12, ZBG^{+}23, CSC08, CL13, CM08, CXK^{+}13, CCL04, GNQ^{+}22, HW14, KMC97, LCOM07, LLKY13, LLKC13, OCRS07, RFB10, LCG^{+}22]. Network-Based [Ase23, FLG^{+}23, YLY^{+}23]. Network-on-Chip [CM20, LDD^{+}18, LW17, PMT20, XS16, XCF18, YKCG14, ZHC^{+}21, ZYS12, CSC08, LCOM07, LLKY13, LLKC13]. Network-on-Chips [HCZ^{+}16, GNQ^{+}22]. Networked [KC10]. Networking [DZK^{+}24]. Networks [BKW15, BP23, BDBB19, CZW19, CAP^{+}23, CLX^{+}23, FCZ^{+}23, GAT^{+}21, GPS^{+}24, HWL^{+}23b, HLX^{+}23, IIEKS23, IHM15, JLJ15, KPB19, LHS20, LDP^{+}22, LYL^{+}19, LNPL23, LRHL24, MAS^{+}20, MNMK^{+}21, MPM^{+}17, NM23, PSP24, SSK^{+}23, SRTG19, UPV23, UE22, XLS15, YMB15, ZFLS11, ZYPC17, ZMP16, BLR06, CXK^{+}13, CBR^{+}05, GWR13, HMVG13, JP12, JSG09, MD13, MDM07, OM08, RL13, TDE08, VS12a]. Networks-on-Chip [BDBB19, IHM15, JLJ15, CXK^{+}13, JP12, OM08]. Networks-on-Chips [VS12a]. Neumann [KT01]. NeuPow [NSP^{+}20]. Neural [Ase23, BP23, DCC^{+}23, EJR22, FLG^{+}23, FCZ^{+}23, GAT^{+}21, GPS^{+}24, HXZ^{+}23, HLX^{+}23, HTC^{+}23, JYY^{+}22, KZKAKP23, LHS20, LDP^{+}22, LPLK22, LYL^{+}19, LJJ^{+}22, LJL^{+}23, LNPL23, MNMK^{+}21, NHS23, NM23, NGL^{+}21, PSP24, SSK^{+}23, UPV23, WXH^{+}19, WDLX21, YLY^{+}23, ZBG^{+}23]. NeuroCool [PSP24]. Neuromorphic [BXG^{+}24, GT21, LS22, XGC^{+}20]. Neuron [ZK15]. Neuron-MOS [ZK15]. Next [PFHAH22, YD16]. Next-generation [PFHAH22]. NMOS [RM23b]. NoC [ADDM^{+}13, CAOM19, CBR^{+}22, CXR^{+}23, DJP21, HWX^{+}14, JBJ22, MHT14, QBTM16, SGJN24, TCL14, SPT^{+}17]. NoC-based [MHT14, CAOM19, HWX^{+}14, QBTM16, CBR^{+}22, DJP21]. NoC-Enabled [CXR^{+}23]. Noc-HMP [SPT^{+}17]. NoCs [AJM13, AL19, CHA^{+}23, DLC^{+}17, HMMG^{+}20, JM14, KPF16, MT15]. Node [BDB12, CZW19, PDS12, PMA24, DHZ^{+}11, JSG09, ZHOM08]. node-centric [ZHOM08]. Nodes [ATF^{+}23, BPTB17, LZA^{+}21, NSS^{+}16]. noise [GGBZ02, HR06, HMLL11]. nominations [Ano13]. Non [AKM^{+}22, EWT23, GLY^{+}12, HSP^{+}22, HKJ^{+}23, LL15, SL18, STJG16, WDLD17, ZYW^{+}18, KCKG13]. Non-enumerative [STJG16]. Non-functional [EWT23]. Non-Gaussian [ZYW^{+}18]. Non-Intrusive [LL15, SL18]. Non-Monte-Carlo [GLY^{+}12]. non-overlapping [KCKG13]. Non-uniform [HKJ^{+}23]. Non-Volatile [AKM^{+}22, HSP^{+}22, WDLD17]. noncomplementary [RS03]. Nonfunctional [HBPW14, RGT^{+}14]. Nonideal [TWL16, WFT^{+}19]. noniterative [MCMW08]. nonlinear [CCC09b, Con06]. nonManhattan [Yan00]. nonpreemptive [GDG^{+}08]. nonslicing [LCC11]. Nonspecified [WC10]. nonstationary [AHAKP08]. nonuniform [VCLD03]. nonvolatile [SLXZ12, ZYZ^{+}13]. note [CSL^{+}07]. Notions [SGC^{+}14]. Novel [GD22, KKHK16, LWZ^{+}19, LJJ^{+}22, LLQD23, MS17, VNS19, DDFR13, SCCH08, Ped06]. NP [DK22]. NP-Separate [DK22]. NPU [LPLK22, RKKH24]. NPU-Accelerated [RKKH24]. NSGA [SA24]. NSGA-II [SA24]. number [HPK99]. NVM [BRCS18, SJ23, SJN24].

NVMe [HC18].

O [LC13, SLC^{+}22, Wu09, Yan16]. OAOS [HGLC16]. OBDD [FWCL05]. Obfuscated [ISK21, LMS16, RNR^{+}21]. Obfuscation [AYS20, GDTF17, HYK^{+}20, KSD^{+}22, OK20, RGX^{+}24, SLP^{+}19]. Obfuscation-Based [GDTF17, HYK^{+}20]. Object [SJL23, Wol96, HCLC98, Hsi01]. Object-oriented [Wol96, HCLC98, Hsi01]. Objective [GACK22, KLE18, SA24, SFM^{+}19, dONH23, EWT23, LFST21, PB14, SCL^{+}22, ZGB^{+}23]. Observability [CLMZ10, CM13]. observability-based [CM13]. Observation [LL15, HW14, Pom13]. Observing [DBK^{+}18]. Obstacle [HLG^{+}15, HGLC16, LLLL18, WSRH16, Yan20, LYKW09, SMYH07]. Obstacle-Avoiding [HLG^{+}15, HGLC16, LLLL18, WSRH16, LYKW09]. Obstacle-Aware [Yan20, SMYH07]. obtain [MS00]. Obviating [PBWB21]. Occupancy [ZHC^{+}18]. Octilinear [HGLC16, Yan08]. Off [FG18, BHY^{+}24, KSD^{+}22, MS23, PDN00, RJL^{+}09, WPHL08]. off-chip [PDN00]. Office [GCL^{+}16]. Offline [MGR^{+}15]. Offlining [JPM^{+}19]. Offloading [LJZ^{+}24]. offs [FHHG12, PCC09, WVYG99, WGDK07, XPSE12]. OLED [LKH19]. On-Chip [ALL17, JNS^{+}17, JZYZ15, SCK18, SMBT19, ZYPC17, DNT20, LCOM07, PLH^{+}24, PDN00, Pom24a, WM24, WDC^{+}22, ZSZ10, ADS^{+}09, CCL04, KP13, LH13, NR03, PPDK09, YLP^{+}13, ZM07]. On-Demand [AAA15]. On-device [TZZH22]. On-the-Fly [VFML23]. Once [CHBK15]. One [MWK21, XFJ^{+}16]. One-pass [MWK21]. Ones [PB12]. Online [BYT22, HLW^{+}23, MBD^{+}20, TZZH22, ZAJ^{+}12, ADDM^{+}13, CSAHR07, RAKK12]. Only [CHBK15]. onto [OKJH22, SWT23]. Op [AG22]. Op-Amps [AG22]. Opamp [Shi20]. OPC [TZ20]. OPC-inclusive [TZ20]. Open [LRHL24, BCR^{+}08, BD05]. open-source [BCR^{+}08]. OpenCL [TL19]. Operating [EAAK^{+}23, TWL16, TL19, PMB10]. Operation [BPTB17, CLMZ10, GDTF17, MACV14, KJR^{+}07]. Operations [BC16, LWZ^{+}19, LXWC20, ARLJH06, BG01, HPK99]. operators [BD05]. OPportunistic [SGJN24]. opportunities [VCLD03]. Opposite [HCN09]. Opposite-phase [HCN09]. Optical [DZ18, WM24]. Optimal [ABC^{+}17, BKW15, BASB01, Cha01, CCX06, CARH18, CH96, FG18, GSS14, HNS23, HWCL13, IIEKS23, KLP^{+}24, KNDK96, LCHT02, OWH08, PL98, SCK18, TS96, TPC^{+}17, ZW98, BW00, BMJ13, CACS05, CGN96, CH00, DSK01, GH00, KCKG13, LH09, MKW08]. Optimization [ACFM12, BZWZ17, BHLG19, CZW19, CYH19, CWL^{+}22, CK16, DHVW18, DHNR23, DZCD15, GLY^{+}12, GK07, GPS^{+}24, HRC21, HWF^{+}23, HLG^{+}15, HC23, HS19, HKJ^{+}23, JBJ22, JPHL16, JNCS19, KKK12, KKS16, LFST21, LHC16, LZZSV15, LWG^{+}23, LH11, LYCP17, NL24, NM23, PTS^{+}20, PPP^{+}15, PY20, RKKH24, SA24, SFM^{+}19, SYHL14, SHBD21, SRTG19, SHL^{+}19, SCK^{+}23, SCL^{+}22, TRM^{+}16, VILSL23, WHRC12, WFSS20, WTW^{+}23, WDC^{+}22, WKC12, WSRH16, WDLX21, XJF^{+}23, ZGB^{+}24, ZGB^{+}23, ZHJ^{+}23, dONH23, BLM00, BDM^{+}99, BdM00, BCC08, BDB98, BFP08, BOC00, BGN^{+}07, CLLK06, CSC08, CCC09b, CFX09, CJLZ11, Con06, DP02, GG04, GBC07, GDF09, GHW^{+}12, HR06, HPK99, HG07, JPCJ06, KJKK03, KLSP11, KCKG13, KSA^{+}10, LLHT03, LCG^{+}22, LCHT02, LC07, LLLC13, MKBS05, MHT14, MKW09, MLG12, OM08, PCD^{+}01, PEPP06, RGM09, RJBS09, SB98, SPA^{+}03, THL^{+}13, VKKR02, VLH04]. optimization [WGDK07, WLL^{+}11, XZC09, GK09]. optimizations [GGDN04, KRS06, SSG12, SC00, ZHTC09]. Optimized [ACF^{+}11, BC05, HCRK11, MJB19, VA17b, ZABGZ17, ZYS12, KCA04, SY07]. Optimizer [LDLM20]. Optimizing [GYT12, KSK^{+}05, LPP00, LPLK22, LHC24, LAS01, RBWB20, SYZ08, ZLW^{+}15]. optimum [Das04]. OR-based [ZHJ^{+}23]. Oracle [RNR^{+}21]. Oracle-guided [RNR^{+}21]. Orchestrated [SAL19]. Orchestration [EW18a]. Order [DZCD15, KQP^{+}19, LYSO19, SXZV13, ZBPF18, CYZL23]. Ordered [JD18]. Ordering [AJM13, GKM05, LXCH04, MKW08]. organization [PDN97]. Oriented [CLC20, RGT^{+}14, HCLC98, Hsi00, Hsi01, LHZ^{+}06, Sen11, Wol96]. Orthogonal [GLY^{+}12]. Oscillator [CLC^{+}24]. outbreak [FNP09]. Outcome [HFMB20]. Output [JM14, LJJ^{+}22, WSEA99]. Outputs [LHS20]. Overhead [AYS20, EAAK^{+}23, FHHH22, KLP^{+}24, LQD22, PKJK20, WLL^{+}11, MHQ07, PMB10]. Overhead-aware [WLL^{+}11]. Overlapping [KCKG16, YYG^{+}16, KCKG13]. Overlay [EW18b, LM19]. Overscaling [CS22]. Overview [SLP^{+}19].

P3 [HK18]. Pack [ZHJ^{+}23]. Package [WPL23, BC05, LC13, LCJ^{+}10]. packaging [VLH98]. Packed [YCL^{+}23]. Packet [MS17, VNS19, CL13]. packings [SYZ08]. Packs [SKM^{+}16]. pad [IBMD07]. padding [SSP04]. Page [AAA15]. Pair [JD18]. Pairing [AAA15]. Pairwise [ZLY^{+}15]. paper [GK09, QS11]. papers [CH10a, KLSZ09, Ped11]. paradigm [DS05, TYH08]. paradigms [Ped06, PBSV^{+}06]. Parallel [Brz24, DL11, EBR^{+}09, EAP17, FZL^{+}23, GDPRG11, JJH21, KLSZ11, KLK^{+}17, KMC97, LAYZ23, LB11, Sch17, ZFLS11, ZS16, ZGB^{+}24, ZMLH24, ZWD11, CBHK11, CT13, Hsi00, Hsi01, KKJ^{+}08, KH10, LM05, LH09, RMPJ08, TW96, ZCG06, KLSZ09]. parallel-programming [KKJ^{+}08]. Parallelism [HC18, DSRV02]. Parallelization [LH11, ZLL^{+}16]. parallelizing [GGDN04]. Parameter [HRC21, MAL23, ZKS^{+}16, ZGB^{+}23, ST99]. Parameterised [HABS15]. parameterizable [BHS11]. Parameterized [LTPT10, CT13, TP08]. Parameters [BBEM15, BHLG19, KPR06]. Parametric [BFG17a, LON08, LCKT12]. Parasitic [LZ21, WLLH16]. Parasitic-Aware [WLLH16, LZ21]. Pareto [NL24]. parity [RMB10]. PARR [XYG^{+}16]. parser [MLC08]. Part [HLZ^{+}22]. ParTBC [SSN22]. Partial [HZJC23, KQP^{+}19, MCZ^{+}16, ETR07, GDG^{+}08, KBN09, KJT04]. Partially [DHX^{+}23, DHW^{+}23, Pom16c, Pom18b, SSC17, LSDV10, YYLL09]. Particle [HLG^{+}15, FS13]. Partition [WDLD17, ZLL^{+}16, CFHM09, WY06]. partition-based [CFHM09]. Partition-Level [WDLD17]. Partitioned [WDZG16, FWCL05]. Partitioning [CPS16, CXLL22, DHX^{+}23, DHW^{+}23, KD24, KLP^{+}24, LSDV10, SS14, SRTG19, TBCH17, TP08, Vah02, AM10, AMO05, CT13, CJLZ11, DCK07, DD02, FW00, GF10, LLKY13, LVL03, MSKBD07, ML09, PDN00, VLH98, Vah99, WH05, YGH^{+}10]. Partitions [ZS16]. pass [BWB14, MWK21]. pass-fail [BWB14]. passing [DSH12, EY12]. Passive [DHB16, EO19]. Past [WS22]. Path [AKAKP18, CV17, FYCT15, KPF16, LVS16, LLLL18, MMM^{+}22, MCD12, PSD21, STJG16, TD03, ZHC^{+}21, ZBG^{+}23, ETR07, LC14, PT06, PMB10, SHLL98, SSP04, XLCL13, Yan08]. Path-Assessed [LLLL18]. Path-Aware [AKAKP18]. Path-Driven [LVS16]. Pathfinding [APG24, UPV23].

Paths [GC18, XGWL24, BK00, PGB01]. PATRON [CGV^{+}23]. Pattern [BKW15, CCK^{+}18, HWF^{+}23, IIEKS23, NPH^{+}20, BH03, FNMS01, OKC08]. pattern-based [OKC08]. Patterning [LZ17, SHL^{+}19, XYG^{+}16, YLZ^{+}17, ZLY^{+}15]. Patterns [LM21, Pom18b, ZMTC13]. Pay [CHBK15]. Pay-Burst-Only-Once [CHBK15]. PBO [ZGB^{+}24]. PCB [Yan17]. PCM [AAA15, BLNK14, CCH^{+}15a, CHC^{+}16, HHK^{+}17, LYLW17, PBZM19]. PCM-Based [PBZM19, AAA15, CCH^{+}15a]. PeaCE [HKL^{+}07]. Peak [JGM14, PTC^{+}15, TPC^{+}17, HCN09]. PeaPaw [TBCH17]. Pegged [IK19]. Penalty [JK10]. per-Task [LMA^{+}16]. per-word [RMB10]. Performance [Ali12, BG01, BDBB19, CCS15, CWL^{+}22, DKT^{+}16, DBK^{+}18, DLC^{+}17, DHZL23, DKZ^{+}15, FG18, GK14, GDD21, HWCL15, Kha23, KYL16, LHS20, LDD^{+}18, LMW99, LLL^{+}18, LS19, LLM^{+}23, LTPR^{+}13, LQL^{+}24, NRZ^{+}18, QBTM16, SJN24, SYX12, TWM^{+}23, TBCH17, TRM^{+}16, TK18, THT12, THC^{+}14, WY06, WMT^{+}16, WLC02, WLCJ09, Yan16, YP10, ZLW^{+}15, CL13, CYZL23, DP02, EK97, FLPP09, GDTG07, Giv06, GOC02, GHW^{+}12, GWR13, HDL^{+}12, LC96, LJV02, LYKW09, LFG^{+}09, LV02, NS03, PDN97, RAKK12, SLXZ12, VLH98, WWG08, ZHM07]. Performance-Aware [BDBB19]. Performance-constrained [BG01, WLCJ09, GOC02]. Performance-Driven [GDD21, HWCL15, Yan16, GK14, LLM^{+}23, WY06, WLC02, EK97]. Performance-Efficient [YP10]. performance/power [ZHM07]. Performance/Thermal [SYX12]. Performance/Thermal-Aware [SYX12]. Period [HYN15, BDB98, CGN96, PL98]. Periodic [CHBK15, Pom16c, SBY^{+}20]. Perpendicular [RPR^{+}21].

Personalization [TZZH22]. Perspective [KAC^{+}23, RJ14, SS14, MOZ06, ZHOM08]. Perspectives [YBM^{+}21]. Perturbation [LYM^{+}20]. Pharmaceutical [YSF^{+}18]. Phase [BLNK14, IYF^{+}21, JSA18, KSA^{+}10, LLP^{+}16, LQD22, LG23, CR12, HMB98, HCN09, Kag05, RAKK12]. Phase-adjustable [KSA^{+}10]. Phase-aware [LG23]. Phase-Change [LLP^{+}16]. Phenomena [ADB^{+}19]. Physical [Ase23, BHY^{+}24, CXLL22, CO18, HLHT08, JCPL23, LLQD23, MYSZ23, PKC^{+}21, PTPB22, SKM^{+}16, YD16, GWR13, HMVG13, MLG12, SYL09]. Physically [CSC^{+}21]. Piecewise [HBPW14]. PIM [CXR^{+}23, KLP^{+}24]. PIM-based [KLP^{+}24]. Pin [XYG^{+}16, Yan20, OWH08, XTW05]. Pin-Access [XYG^{+}16]. Pins [TFW24]. Pipeline [CRC15, RPKC05]. Pipelined [CHBK15, LF12, MED23, MRL^{+}20, Hua01, MS08, MD08, NS03, RTNL05, YGH^{+}10]. pipelines [HA05]. Pipelining [AA17, KLV15, BG01, BASB01, CACS05, CL99a, HV98]. Pixel [DZ24]. place [KCKG13, YWW10, LRHL24]. Placement [ATF^{+}23, DK16, HWGY16, HWL^{+}23a, HWCL15, HKJ^{+}23, JYHY21, JNCS19, KRL15, LLL^{+}18, LNG^{+}16, LCC^{+}15, LB11, MCZ^{+}16, MJB19, MAL23, PSD21, SAHF^{+}20, TFW24, TRM^{+}16, WSH^{+}18, WTW^{+}23, WSRH16, WLLH16, WDLX21, YVC14, ZSY18, AM05, ACT13, CBHK11, CACS05, CC06, CSX^{+}05, EK97, KPSW09, LCK^{+}09, OS03, RS03, SC06, Tes02, TY97, VLH04, WLC02, WCC03, WLT08, YWK^{+}03]. placements [HWCL13]. Placer [AMM^{+}18, DKT^{+}16, DKT^{+}16]. Placers [LRHL24, MAL23]. Plaintext [HYK^{+}20]. planar [DPNB02]. Planning [DSHD23, XYG^{+}16, YYG^{+}16, LC13, LHZ^{+}06, MKBS05, SBC08, XTW05]. PLAs [LWH06]. Plasticine [EKEK22]. Platform [APD^{+}11, IGN18, KLP^{+}24, LY24, VGG19, FNP09, JCS^{+}08, RFB10, ZHM07, PBSV^{+}06]. Platform-aware [VGG19]. platform-based [ZHM07, PBSV^{+}06]. Platforms [BS14c, ETAV18, LS11, LMS16, MBD^{+}20, RS18, TBCH17, VGG19, WDZG16, YPCF17, BMJ13, CNQ13, JW08, LP07, MPDG09]. Playing [RJL^{+}09]. PMC [CLH12, CCH15b, CH13]. PMU [APD^{+}11]. Point [ALL17, BS14a, BFL10, SRC15, WZH^{+}23, XNZ^{+}15, AM98, BSP^{+}23, CPW04, DPNB02, LCOM07, WG11, WFT^{+}19, YCL^{+}23, Yan08]. point-to-point [LCOM07]. Pointer [RCW22]. points [PMB10, Pom13, TD03]. Poisson [QSK12, WTW^{+}23]. Polar [JNS^{+}17]. Polarity [ZHJ^{+}23, CHH09, LT11]. Policies [DZCD15, Kha12]. policy [CXK^{+}13]. Polishing [LTW^{+}16]. Polling [LZZ23]. Polling-Based [LZZ23]. Pollution [DJP21]. polygon [LLM01]. polygons [CT13, LM96, TP08]. Polymerase [LHC16]. polymorphic [LLYW10]. polynomial [GK07, GK09]. Polynomials [GLY^{+}12]. port [CL13, SBC08]. port-scalable [SBC08]. portable [LCZ^{+}08, Rak09]. Portion [GD20]. POSE [Hsi01]. Positioning [HK18]. Post [GDD21, PTS^{+}20, VILSL23]. Post-clustering [VILSL23]. Post-Processing [GDD21]. Post-silicon [PTS^{+}20]. Postlayout [CLLK06]. Postplacement [CMB07, LCY12, WWG08, XLL^{+}16]. PostRouting [KO23]. Postscheduling [FHHG12]. postsilicon [MKK13]. Power [ACF^{+}11, ALL17, BLM00, BS14b, BM11, BPTB17, Brz24, CMP10, CH10b, CHBK15, CXH^{+}16, CAP^{+}23, CGV^{+}23, CLMZ10, DLC^{+}17, DNT20, DCC^{+}23, FG18, FZL^{+}23, GBR07, GCL^{+}16, GAT^{+}21, HWDQ22, HPK99, HYN15, HC23, JIR^{+}21, JLK15, Kha23, KKHK16, LG18, LKM04, LYHL14, LAYZ23, LSZ^{+}24, LLK^{+}14, LHJ12, LHK^{+}15, LKH19, LS17, LNPL23, MAS16, MKW09, MN17, NPH^{+}20, NDA^{+}23, NSP^{+}20, PJL14, Ped96, PTC^{+}15, RM23a, SCK18, SC00, SBC08, SYHL14, SSCS10, SESN15, TWL16, TRM^{+}16, TMDF10, TCL14, VNS19, WVYG99, WGT^{+}17, WZH^{+}23, WC10, WSRH16, XXC^{+}24, XLS15, ZFLS11, ZK15, ZS16, ZMTC13, ZF23, AHAKP08, BDM^{+}99, BdM00, BD00, BMJ13, BBD00, CS07, CH10a, CM08, CIB01, CCX06, CCW08, CHHL96, CCC09b, CJLZ11, CLQ12, DS06, DTC^{+}09, ETR07, GOC02, GDF09, GF10, GS13, HR06, HLCH07, HLHT08, HTCP13, JWL^{+}03, KBN09, KP22]. power [KKH^{+}02, KOS09, KC13, KHW06, KYN^{+}12, LMB^{+}12, LLHT03, LYCP13, LHW^{+}17, LBV^{+}06, LHW97, MKK13, MRC06, MKW08, MLG12, MFS09, ML09, NT05, PPDK09, Pom14b, PWY05, PR96, RFB10, RTNL05, SMS22, STL^{+}13, SUC01, SPMS02, SNL12, SZV^{+}12, TKVN07, TJ99, THC^{+}14, WJY^{+}07, YHL^{+}11, YGZ04, YLL06, YHL07, YHH09, ZHM07, ZLL13, ZYDP08, ZP08, ZYP09]. Power-Aware [LHK^{+}15, CGV^{+}23, SBC08, SNL12]. Power-delay [MKW09, SC00, WVYG99]. power-density [ZYP09]. Power-Efficient [JLK15, SZV^{+}12]. Power-Gating [KKHK16, YHL07]. Power-On [WZH^{+}23]. power-optimal [MKW08]. Power-safe [ZMTC13]. power-transmission [KC13]. Power/Ground [HC23, LHJ12]. Power/Thermal [ZF23]. Powered [XPZ^{+}18, CSAHR07]. Powerful [LTYW12, MB04]. PowerPC [WAZ98]. PPA [LS23, MAL23]. Practical [CPK20, Pie16, UPV23, VJBC07]. Practice [MDM^{+}12, RGX^{+}24, SSCS10]. PRAM [KYL16]. PREASC [GD20]. precedence [ZAZ13]. Precise [Ali12, RCW22, ZZ24]. Precision [EJR22, HLX^{+}23, YCL^{+}23]. Precision-reconfigurable [EJR22]. predefined [PSK08]. Predict [KO23].

Predictability [NSCM17]. Predictable [VGG19, WLZ^{+}19, HGBH09]. Predicting [LHS20, XXC^{+}24]. Prediction [CS07, CBC22, DNT20, DCC^{+}23, DKZ^{+}15, FG18, HWX^{+}14, JGM14, LPY^{+}20, LNPL23, PBL^{+}17, SAHF^{+}20, YB23, CR12, OM08, SYL09]. prediction-based [OM08]. Predictive [AVG19, HW00, TKVN07]. Preemptive [IHM15, SSC17, GDG^{+}08]. Preface [YD16]. Preferred [Pom18a]. Prefetching [DJP21, LV02, PSP24]. Prefix [Brz24, LH09, ZCG06]. Preparation [PGCB16, PBWB21, PBF^{+}22, RCK^{+}15, SKS^{+}18]. prescribed [DSRV02]. Presence [EKS^{+}14, MCMW08]. Preserving [HK18, HTC^{+}23]. Prevent [WSS^{+}18]. Preventing [YCL^{+}20]. Previewer [HFMB20]. Primary [Pom16a, Pom21b]. Primitive [MMM^{+}22]. Principle [CHBK15]. Principles [SBY^{+}20, Ped96]. Print [DZCD15]. Printed [GDTF17, OW06]. Priority [IHM15, KPF16, LMS16, WDZG16, MHQ07]. Priority-Aware [KPF16]. Priority-Preemptive [IHM15]. Privacy [HTC^{+}23, HK18]. Privacy-preserving [HTC^{+}23]. Proactive [KBV^{+}15]. Probabilistic [APS18, CKAP07, CB17, GQW19, KW16, KVMH08, BLR06, FZKS11]. Probe [Kha12, BC05]. Probe-Wear [Kha12]. Problem [Ase23, DPNB02, DS06, FNMS01, LVL03, NR01, PDN00, SW99, YWW10]. Problem-tailored [Ase23]. problems [SB98, WGDK07]. Procedure [Vah99]. Process [AKAKP18, BHY^{+}24, BHLG19, GC18, LWZ^{+}19, RJ14, TWM^{+}23, VEO16, CS07, GM08, KTKO13, KPR06, LG12, LH13, LTPR^{+}13]. Process-in-memory [LWZ^{+}19]. processes [JB98]. Processing [BM11, GFJ16, GDD21, HXB^{+}22, LCJ^{+}22, LYL^{+}19, LS22, MFHP12, PRKK21, WDD^{+}23, HMVG13, JSG09, LPP00, NM13, TYH08, ZHOM08].

Processing-In-Memory [WDD^{+}23]. Processing-Near-Memory [LCJ^{+}22]. Processor [HKL^{+}15, ISE08, LHLP16, LYHL14, LF12, NSH^{+}16, NRZ^{+}18, OHA19, SPT^{+}17, VLGG01, DHZ^{+}11, GG04, Giv06, HGBH09, KBA08, LMB^{+}12, OCRS07, PDN97, PDN00, RFB10, SGD10, WKR09]. processor-based [PDN00]. Processors [CRC15, JZYZ15, KAKSP16, KLK^{+}17, KLJ14, LPD^{+}17, LHF12, OKJH22, TY19, BH10, CL99a, CPW04, Edw03, Hua01, KJR^{+}07, LJV02, LCD07, LB00, MD08, PHM00, RAKK12, SR12, TKVN07, LSV06]. product [DK08]. Production [PBWB21, PKP^{+}03]. profile [ZSZ10]. Profiling [KLP^{+}24, SMBT19, THC^{+}14]. Profiling-Based [SMBT19]. Program [HKL^{+}15, BGN^{+}07, RAKK12, WWC04]. Programmable [CLC^{+}24, GHYR19, HHX^{+}23, KP22, WCZ^{+}24, WSS^{+}18, ZK15, CH02, CD96, LSPC14, MSD06, PTC05, PWY05, WV02]. Programmers [SYGC22]. Programming [CGV^{+}23, ETAV18, KLSZ11, TFW24, TZ17, WLZ^{+}19, ADDM^{+}13, GH00, KLSZ09, KKJ^{+}08, TP08, WJYZ11]. programming-based [ADDM^{+}13]. Programs [PMS15, SYHL14, EY12, Vah02, YWGI09]. Progressive [KC10]. project [WLT08]. projective [DL11]. Prolonging [AAA15]. Proof [CCMC20, IPWW17]. Proof-Carrying [IPWW17]. Propagation [AL19, MCD12, KPR06, RCD07, YH97]. Properties [CVMP19, HBPW14, RGT^{+}14, WFT^{+}19, BDC08, BH03, BFP08, BZ08]. property [KHP05]. Protect [MLH^{+}17]. protected [LSDV10, RMB10]. Protecting [DFM15, GSFT16, YBS^{+}18]. Protection [GDTF17, LSZ^{+}24, PLH^{+}24, SLP^{+}19, KHP05]. ProtFe [LSZ^{+}24]. Protocol [LXGM23, ADS^{+}09, BGM04, DP04]. prototype [APB^{+}08]. Prototyping [ARLJH06, ORGD^{+}15, JDT^{+}08]. Provably [ADS^{+}09, Das09, YWK^{+}03]. Provide [KKLG15]. Providing [HC18]. Proximity [DZ18]. Pruning [GYZ^{+}22, MNMK^{+}21, ZBG^{+}23, DHV^{+}00]. Pseudo [KD24, MAL23, PKC^{+}21]. Pseudo- [MAL23, PKC^{+}21]. Pseudo-Partitioning [KD24]. PSL [BZ08]. PSS [DSHD23]. PTA [XJF^{+}23]. PTM [LLH^{+}17]. PUF [CCMC20, CLC^{+}24, IK19, LLQD23, MMM^{+}22, NSCM17]. PUFs [HRK18]. Pulse [HWL^{+}23a, LQD22]. Push [KMO^{+}12]. PV [DZ18]. PV-Aware [DZ18]. PVT [PPDK09]. PWM [TWL16, WGT^{+}17].

QoS [LYLW17, RKKH24].

QoS-Constrained [RKKH24]. quad [LBV^{+}06]. quad-core [LBV^{+}06]. Quadratic [AL19]. Quadruple [JIR^{+}21]. QuadSeal [JIR^{+}21]. Quality [BZWZ17, JSS^{+}19, LKH19, LPY^{+}20, LIK22, Pom19b, BHW^{+}13, LRHL24, XPSE12]. Quality-Assured [JSS^{+}19]. Quality-Enhanced [LKH19]. QuanDA [NHS23]. Quantifying [SGC^{+}14, YRH11]. Quantitative [NHS23, LCOM07]. Quantization [GYT12, HWDQ22, HJY23, HLX^{+}23, LDP^{+}22, ZZ24]. Quantization/Mapping [HJY23]. Quantum [HZL^{+}22, LSZ^{+}21, ZFL22]. Quenching [HWL^{+}23a]. Queuing [SSL17].

Race [BK10, HN07]. Radio

[JDT^{+}08, JSG09]. Radix [BS14a]. RAID [SLC^{+}22]. RAID-enabled [SLC^{+}22]. Rail [LQD22, VEO16]. RAM [AKM^{+}22, LSL^{+}13, SABSA15]. ramp [KM97]. Random [BZWZ17, BS14b, RPR^{+}21, ZGB^{+}23, JT98, KPR06, SXZV13, SNL12]. Range [LDP^{+}22, MS17, CL13, LSPC14]. Rank [SYH^{+}22]. Rapid [EW18b, ORGD^{+}15]. Rare [ZKS^{+}16]. Rare-Event [ZKS^{+}16]. RASCv2 [BSP^{+}22]. Rate [CJKK19, HDZ^{+}20, LD17, MDG98, PB12, PHKW12, TY97]. rates [ACT13]. Ratio [HKJ^{+}23, WLLH16, Das04]. RC [KM97, VEO16]. RDL [Yan11]. Reachable [XLNB17]. React [ADB^{+}19]. Reaction [LHC16]. Reactive [WLZ^{+}19, ZABGZ17, PSL^{+}98]. Read [DHZL23, JSA18, LHS^{+}21, PPP^{+}15, WHXZ13]. Readable [BHBS22]. Real [CHBK15, CBC22, CH17, DZK^{+}24, FG18, FHHR21, GYZ^{+}22, HXC^{+}18, KPF16, LSCK20, LQL^{+}24, NSH^{+}16, NRM^{+}24, PKJK20, PSNC18, SSC17, SBY^{+}20, SLV^{+}22, SWT23, WLZ^{+}19, WDZG16, WJG^{+}19, YRH11, ZLW^{+}15, APB^{+}08, DRG98, HMVG13, MHQ07, PEPP06, PW99, WLL^{+}11, ZAZ13]. Real-Time [CHBK15, CBC22, CH17, FG18, GYZ^{+}22, HXC^{+}18, KPF16, LSCK20, LQL^{+}24, NSH^{+}16, NRM^{+}24, PSNC18, SSC17, SBY^{+}20, SLV^{+}22, WLZ^{+}19, WDZG16, WJG^{+}19, YRH11, ZLW^{+}15, DZK^{+}24, FHHR21, SWT23, APB^{+}08, DRG98, HMVG13, MHQ07, PEPP06, PW99, WLL^{+}11, ZAZ13]. realistic [MFS09]. Reality [XLNB17]. Realization [ACFM12, CHHL96]. reallocation [ZYP09]. realtime [HG07]. Reassignment [Yan20, Yan08]. ReChannel [RHA08]. Recognition [GFJ16, RG19, SJL23]. recompilation [GF10]. Reconfigurable [ADB^{+}19, AVG19, BKW15, CPS16, CM20, DHX^{+}23, DHW^{+}23, EK16, JDLZ24, JPHL16, LPL^{+}21, MS21, MLC08, MRL^{+}19, ORGD^{+}15, RM23a, SSC17, SVK17, UE22, ZLQ15, ZMS^{+}19, ARLJH06, EJR22, GDG^{+}08, HBC^{+}08, HW14, JBC^{+}10, KKMB02, KLSP11, LCK^{+}09, RHA08, WKR09, WLC02, YLP^{+}13, YGH^{+}10, YYLL09]. Reconfiguration [CAOM19, MCZ^{+}16]. reconfigurations [RCG^{+}08]. reconnections [WC06]. reconstruction [Yan08]. Recover [BFV15]. Recovering [JCK^{+}18]. Recovery [NSS^{+}16, WL12, ZAZ13]. Rectangle [Yan18]. rectangular [DSK01, Meh98]. Rectilinear [GC96, LLLL18, WCC03, LYKW09, MHD^{+}04, MS00, OWH08]. Recurrent [HLW^{+}23]. recursive [LC96]. Recycling [TCW20]. Red [RGX^{+}24]. Reduce [CIX15, JK10, Pom16c]. Reduced [PAV17, Pom24a, AMM^{+}06, SBH^{+}06]. reducible [BC11]. Reducing [ASAP17, BFG^{+}19, BWB14, CJKK19, DJP21, HH09, Kan06, KLJ14, LYCP13, PR11, SYHL14, KTKO13, MB04, PGB01, TKVN07]. Reduction [ABC^{+}17, BDB12, FLWW02, PTC^{+}15, PS23, Shi20, WB16, WDLD17, WH19, WLH20, CFHM09, CCW08, DK08, ETR07, GF10, HLHT08, KYN^{+}12, LCC11, LLHT12, LCJ^{+}10, NT05, RMKP03, SY07, SBH^{+}06, SPMS02, TY97, WVYG99, YHL^{+}11, YWK^{+}03, YLL06]. Redundancy [CJKK19, JLK15, CMNQ08]. Redundant [KMO^{+}12, SHL^{+}19, PGB01]. Reed [ZHJ^{+}23]. reference [AOC02, SM00]. Refinement [TFW24, CLM^{+}10, GGB97, MS08, MOZ06]. refit [DVA02]. Refresh [CJKK19, LSL^{+}13]. Region [BZWZ17, ZGB^{+}23]. Regions [JCK^{+}18]. Register [GF10, HWCL15, LHF12, LQD22, MHF96, TLCF16, WKL^{+}18, WJM24, XLL^{+}16, CACS05, CFX09, HCN09, KI01, KNDK96, LWK11, VKKR02, ZYP09]. register-file [CFX09]. Register-Transfer [WJM24]. Registers [PBH^{+}24, CL99a]. Regression [BBD00, GD20]. Regression-based [BBD00]. Regular [XYG^{+}16, CH13]. regulation [ZLL13]. Reinforced [MAL23]. Reinforcement [JBJ22, LJZ^{+}24, LCZ^{+}24, NL24, PJL14, QZZW24, SKR^{+}22, WDLX21, XGWL24, STL^{+}13]. Related [dONH23]. Relaxation [LGGJ14, PY20, ZBG^{+}23]. Relaxation-Based [PY20]. Release [SZB17, YP10]. Reliability [APS18, BHY^{+}24, CSC^{+}21, CET16, CCK^{+}18, CXLL22, GPS^{+}24, KMO^{+}12, LHJ12, NWA^{+}24, PPP^{+}15, RMB10, TK18, WXH^{+}19, XLY^{+}18, GS13, JS13, KVMH08, LH13, ZAZ13]. Reliability-Aware [BHY^{+}24, CET16]. Reliability-Driven [LHJ12]. Reliable [BJX15, GC18, JPCJ06, MACV14, WZL^{+}21, XCF18, XNZ^{+}15]. Relocation [HWF^{+}23, LLLC13]. Remote [BSP^{+}22, CRT19, KOO18, KC10]. Removal [MGR^{+}15, CMNQ08]. reorder [WPHL08]. Reordering [WC10, GFC^{+}09, Hua01, PR96]. Reorganizing [JCK^{+}18]. Repair [CJKK19, KMO^{+}12, PSNC18, MRMP08, NR03]. Repairable [KMO^{+}12]. repeating [LWC07]. Replacement [CZW19, JCK^{+}18, CCW08]. Replay [YXG^{+}24, ZLQ15, EY12]. Replication [DFM15]. Representation [HZL^{+}22, CCQ98, YYC09]. Representations [KQP^{+}19, YCCG03]. Representative [FYCT15, PKJK20]. Reprogramming [ANS^{+}20]. Request [AL19, Wu09]. Requests [CIX15, AHAKP08]. Requirement [XLY^{+}18, KCA04]. Requirements [EWT23, Pie16, SL18, Meh98, MB04]. ReRAM [BP23, HXZ^{+}23, LJJ^{+}22, LHC24]. ReRAM-based [HXZ^{+}23, LJJ^{+}22, LHC24]. ReSC [YFT18]. rescheduling [GK14]. Rescuing [HXZ^{+}23]. Research [BRCS18, MRL^{+}19, XFJ^{+}16]. reseeding [KJT04]. Reservation [HC18]. Reserved [KKLG15]. reset [SPA^{+}03]. Reshaping [TZZH22]. Residential [VA17a]. Residue [MGR^{+}15]. Resilience [GD20, LWC18]. Resilient [BJX15, BC16, CRC15, KKLP15, SMS22]. Resistance [CYLC24, KYL16]. Resistant [Kha12]. Resistive [CYLC24, EBR^{+}09, LWZ^{+}19, TLCF16, WFT^{+}19, XNZ^{+}15, LLQ^{+}03, SKCM06]. Resolution [LQL^{+}24, ZPLI23]. resolving [Das09]. Resource [CET16, CS22, DK08, FS13, HC17, KK14, LZY^{+}23, LF12, MBD^{+}20, PBF^{+}22, TCL14, WG11, WLH20, WGSH16, BDB98, CFX09, HLKN07, Kuc03, LSDV10, MKK13, MJM11, NR01, WGDK07, YWW10, ZHOM08, KMR18]. Resource-aware [FS13]. Resource-Constrained [PBF^{+}22, WG11, WLH20, LSDV10, NR01, ZHOM08]. Resources [DHW^{+}23, JNS^{+}17, PGB01]. Response [CH17, KS23, PMS15, SSO16, DC07, SCJ01]. Responses [XCW12]. Responsiveness [SLC^{+}22]. Restore [ZZCY17]. Restricted [HSR^{+}24, KD24]. results [AYM05]. Resynthesis [WPR^{+}19]. Retargetable [PHM00, AMR00, KKJ^{+}08, VLGG01]. Retargeting [DZ18, IIEKS23, WJYZ11]. Retention [CJKK19]. reticle [WLT08]. Retiming [BOC00, HMB98, HLHT08, SSP04, Zho08]. Retiming-based [BOC00]. Retracing [LLLL18]. Retrain [ZBG^{+}23]. Retrain-Free [ZBG^{+}23]. Reuse [AC06, BFP08, CSO22, LDLM20, NAK20, OHA19, IBMD07, LSPC14, RSR01, VCLD03]. Reuse-based [OHA19]. Reusing [CCL04]. Revealing [CM19]. Reverse [AYS20, CM18, GDTF17, WSS^{+}18]. Reversible [HDB22, PS23, MDM07]. Review [IE12]. revisited [RS98, SDP^{+}09]. Revisiting [GWR13, ZSY18]. Revitalized [PCT^{+}17]. Rewarding [TEK18]. Rewiring [LTYW12, CMB07]. rewriting [ARLJH06]. rewriting-logic [ARLJH06]. RF [BBEM15, HCZ^{+}16, LYSO19, LZ21, PTS^{+}20, SA24]. RF-Interconnect [HCZ^{+}16]. RF/Analog [LYSO19]. RFID [DTC^{+}09, YFT18, YBS^{+}18]. RFID-Enabled [YFT18]. RGMU [JDLZ24]. rhythms [GS13]. rich [SHBD21]. right [MR96]. Ring [CLC^{+}24, GK07, GK09]. Ripple [HWGY16]. rISAs [SBH^{+}06]. RISC [BSZ^{+}24, HV98, YCL^{+}23, ZBPF18]. RISC-V [BSZ^{+}24, YCL^{+}23]. risk [DS05].

Risks [MCY23]. river [ZW98]. RL [NT05, XGWL24]. RL-Huffman [NT05]. RLC [MN17]. RO [PBH^{+}24]. RO-driven [PBH^{+}24]. Robust [ATF^{+}23, BJX15, BP23, CZZYW21, DZ18, GCZ^{+}15, MCD12, PBWB21, STGR15, TLCF16, ZK15, ZHC^{+}23, CLYP09, ST99]. Robustness [BHLG19]. Role [CK19]. Root [PLC24]. Root-Cause [PLC24]. rotary [TDF^{+}09]. Routability [AMM^{+}18, HWGY16, HC23, HKJ^{+}23, SAHF^{+}20, THL^{+}13, ZSY18, CLYP09, HSA^{+}04, SYZ08, WSV^{+}14, YCHT00]. Routability-Driven [AMM^{+}18, HWGY16, ZSY18, HC23]. Routable [LCYN18]. Route [CJKS24]. Router [PMT20, TCL14, XS16, CLYP09, JCGP05, MLC08, TDF^{+}09, wATkK02]. Routers [JM14]. Routing [ATF^{+}23, CLC20, DLK24, GdRJM21, GKM05, JD18, LHJ12, LLLL18, LWG^{+}23, LKC^{+}18, MAS^{+}20, MCZ^{+}16, RGM15, RBWB20, SGJN24, TZ17, TZ20, WLLH16, WPL23, XYG^{+}16, Yan18, Yan19, Yan20, ZHC^{+}21, ZPLI23, CZW00, CKKT98, DSKB04, DVA02, GMN^{+}13, LLKC13, LCC11, LCJ^{+}10, MW97, OW06, OWH08, RL13, SMYH07, Yan00, YW09, Yan11, YMC^{+}13, YCHT00, ZW98, ZHTC09]. Routing-aware [GKM05]. Routing-Based [LLLL18, LWG^{+}23]. Row [SAL19, HNS23, LC13]. row-based [LC13]. Row-Buffer [SAL19]. RRAM [LXWC20]. RRAM-based [LXWC20]. RSMT [TFW24]. RSMTs [DLK24]. RSPP [KD24]. RTGC [ZLW^{+}15]. RTL [BK00, BBD00, BFP08, BFV15, Fuj05, GS00, ISK21, LZY^{+}23, LV14, PGB01, PSK08, PIK20, WLM21, XK97]. Rule [GdRJM21, KMO^{+}12, MS17, VNS19, ZZL^{+}23, RS98]. Run [DP02, KS23, HMLL11]. Run-time [DP02, KS23, HMLL11]. Runtime [BHW^{+}13, LL15, LPL^{+}21, NRZ^{+}18, VTC20, WXH^{+}19, ADDM^{+}13, GFC^{+}09, GDG^{+}08, HW14, RCG^{+}08, SKS12, WJY^{+}07, YGH^{+}10]. runtime-reconfigurable [GDG^{+}08].

Saber [CYZL23]. safe [ZMTC13]. Safety [MN17, XLY^{+}18, dONH23, MS08]. Safety-Related [dONH23]. Salsa20 [MAS16]. Sample [PGCB16, PBWB21, PBF^{+}22, ZKS^{+}16]. Sampling [WTR12, ZYW^{+}18]. SAT [CLM^{+}10, Che18, CYV^{+}14, DP02, IIEKS23, RCD07, SGK08]. SAT-based [CLM^{+}10, IIEKS23, SGK08]. Satisfiability [BR12, GMSSS02, OK20, PG15, GPK^{+}09, HSA^{+}04]. satisfying [QS09]. saturation [CCL03]. Saving [RM23a, HW00]. Savings [LKH19]. Scalable [AA17, KLK^{+}17, LAYZ23, PJL14, SS24, SESN15, SKM^{+}16, ZF23, HG07, KCKG13, SBC08, SBGD13, WSV^{+}14]. Scalable-Throughput [SESN15]. Scale [DNT20, HC17, LYL^{+}19, YVC14, ZHC^{+}21, CSX^{+}05, GNQ^{+}22, HCK13, WTW^{+}23, ZGB^{+}24, KBA08]. Scaled [PHKW12]. Scaling [GC18, HC17, HHL14, LV14, WGSH16, IAI^{+}09, KSA^{+}10, ML09]. Scaling-Aware [HC17]. Scan [BKW15, KMO^{+}12, LWC07, LWK11, PSD21, Pom16b, Pom16c, Pom17b, RNR^{+}21, WC10, WWW^{+}12, XCW12, DDFR13, GKM05, KBN09, NT05, PR09, PR11, RMKP03, SSGS03, TYH08, WPHL08]. Scan-based [LWK11, KBN09, PR09]. Scan-BIST [LWC07]. Scan-Cell [WC10]. Scan-In [Pom16c]. Scan-Shift [WC10]. scanline [CT13]. Scenario [BLUS19, DCK09, EK16, HLZ^{+}22, KW16, SWT23, GPH^{+}09]. Scenario-Aware [BLUS19, KW16, SWT23]. Scenario-based [DCK09]. Scenarios [NRZ^{+}18, SPG^{+}08]. Schedulability [GDG^{+}08]. Schedule [SGC^{+}14]. Scheduler [NSH^{+}16, SRKS23, JP08]. Schedules [GDD21, DSRV02, LC96]. Scheduling [ABC^{+}17, BB17, BDBB19, CACS05, CIX15, DZK^{+}24, DHX^{+}23, DHW^{+}23, ENP20, JOH17, KPB19, LHW97, MAS^{+}20, NRM^{+}24, OKJH22, PMS15, SSC17, SLC^{+}22, SAL19, SZB17, WCB15, WDZG16, WWCT18, WJG^{+}19, XPX^{+}21, CLM^{+}10, CJLZ11, DS05, DHV^{+}00, GBC07, HN07, JR97, KW02, Kuc03, LLHT03, MKBS05, MJM11, MHQ07, MR05, MWG97, NR01, PGGD23, RCG^{+}08, SXX^{+}06, TC98, WH05, WGDK07, YWW10, YGH^{+}10, YYLL09]. schematic [KG09]. Scheme [BM11, CWL^{+}22, HDB22, JDD20, KKLG15, KLK^{+}17, LTYW12, LJZ^{+}24, WHRC12, WH20, XS16, HCK13, KSA^{+}10, XLCL13]. Schemes [GYZ^{+}22, MGR^{+}15, CSC08, KCKG13]. Scoping [dONH23]. Score [XLL^{+}16]. scratch [IBMD07]. scratch-pad [IBMD07]. Scratchpad [CPS16, DFM15, BD14]. Script [ZZL^{+}23, NPH^{+}20]. Scrubbing [SVK17]. SDF [OKJH22]. SDF/L [OKJH22]. Search [FZL^{+}23, JYY^{+}22, LPLK22, NDA^{+}23, Pom24b, RFG20, VCLD03, ZFL22, CMB07, DVA02, YWW10]. search-based [DVA02]. Search-space [RFG20]. Searching [DK16, SYZ08]. Secret [LDX22]. Section [BMdG17, CCY22, CO18, JCPL23, KLSZ11, PFHAH22, YD16, CH10a, CLQ12, HJ08, JW08, KLSZ09, MD13, RBA^{+}12]. Secure [BHK17, LSZ^{+}24, LSCK20, YCL^{+}20, HBC^{+}08, ISE08, HRK18]. SecureTVM [HTC^{+}23]. Security [CM20, CPK20, CYLC24, GQW19, GLD^{+}22, HMO^{+}14, KAC^{+}23, KSD^{+}22, LHLP16, LZZSV15, LQD22, LMS16, MMM^{+}22, MAS^{+}20, MCY23, MPM^{+}17, NSCM17, NWA^{+}24, RNR^{+}21, RGX^{+}24, SLP^{+}19, TK18, WM24, WLM21, YSF^{+}18, YBM^{+}21, DP04, IAI^{+}09]. Security-Aware [KAC^{+}23, LZZSV15, LMS16, MAS^{+}20, NWA^{+}24]. SEDONUT [PMA24]. Seeds [Pom17a, Pom24a].

Segment [WL12]. Segment-Based [WL12]. Segmentation [LCG^{+}22]. Segmented [HSA^{+}04, JWL^{+}03, YCHT00]. Select [Pom18a]. Selection [AKAKP18, CXS^{+}23, CV17, FYCT15, GC18, JM14, KPF16, STJG16, ZKS^{+}16, CGN96, CCC09b, LB00, PMB10, VLGG01, XLCL13]. Selective [HTC^{+}23, HKJ^{+}23, Mut09, NRDB19, LCT03, WY06]. selectively [BD00]. selectively-clocked [BD00]. Self [CRT19, EO19, IYF^{+}21, LW21, PIK20, SS24, SBB^{+}18, SJ23, SHL^{+}19, WCB15, WZH^{+}23, XYG^{+}16, SEN05, SZV^{+}12]. Self-Aligned [SHL^{+}19, XYG^{+}16]. Self-healing [SS24]. Self-Measurement [CRT19]. Self-Similarity [PIK20]. Self-Test [EO19, SBB^{+}18, WCB15, WZH^{+}23, IYF^{+}21]. Self-Testable [LW21]. self-testing [SEN05]. self-tuning [SZV^{+}12]. Semantic [Pie16]. Semantics [KC98]. Semi [PLC24]. Semi-Supervised [PLC24]. Sense [ADB^{+}19, DMR23, RM23b]. Sensing [DMR23, LSCK20, LTH99, WJY^{+}07]. Sensitive [CHA^{+}23, DZK^{+}24, YBS^{+}18]. Sensitivity [LM21, LON08, PMB10, ST99]. Sensor [CCMC20, NSS^{+}16, PDS12, ZHC^{+}18, DHZ^{+}11, JSG09, LCK^{+}09, RFB10, ZSZ10]. sensor-driven [ZSZ10]. Sensors [FG18, RG19, YHL^{+}11]. Separate [DK22]. Separation [EK16]. sequence [GF06, LC07, MMP00]. Sequences [PKJK20, Pom15b, Pom15c, Pom17b, Pom18a, Pom24b, KT01, LWC07, PL03, PR11]. Sequential [LVS16, LD17, LWG^{+}23, SPA^{+}03, WKC12, BLR06, BOC00, Che96, CPR^{+}02, Edw03, HVF^{+}01, HRP00, HCC01, JB98, KT96, KOS09, MMP00, PL98, SNH02, Vah02, YWGI09]. sequentially [LIA00]. SER [LD17]. Serial [PMP17]. Serialized [KH10]. Series [TW96]. Series-parallel [TW96]. server [dW97]. servers [ANR13]. Service [DKZ^{+}15, AHAKP08, CBR^{+}05]. Service-Level [DKZ^{+}15]. Set [HKL^{+}15, LPD^{+}17, LHF12, LF12, MCD12, OT15, Pom19b, Pom22, DPNB02, Hua01, LP03, LCD07, LLYW10]. Sets [Pom16b, YRH11, PR07, TCP97]. Settings [ZHC^{+}23]. setup [KO23]. SEU [JLF^{+}12]. SG [KPB19, ZZL^{+}23]. SHAIP [HRK18]. Shannon [GBR07]. shaped [Meh98]. shapes [LM96]. Shaping [KLK^{+}17]. Share [RG19]. Share-n-Learn [RG19]. Shared [KLJ14, SHBD21, ZAZ13]. Sharing [CS22, LF12, RG19, TCL14, WGSH16, BDB98, DK08, SHLL98]. Sherlock [GACK22]. shield [LXCH04]. shielding [Mut09]. Shift [HWDQ22, PBH^{+}24, Pom21b, PTC^{+}15, WC10, WWW^{+}12, LWK11, WPHL08]. shifter [Kag05]. Shifts [LS19]. short [SSP04]. short-path [SSP04]. Shuffling [HHK^{+}17, KJR^{+}07]. shutdown [HW00]. SID [LHK^{+}15]. SID-Based [LHK^{+}15]. Side [BSP^{+}22, CYZL23, DZS^{+}18, LSZ^{+}24, LQD22, LM21, NPH^{+}20, ZBPF18]. Side-Channel [DZS^{+}18, LSZ^{+}24, LQD22, ZBPF18, CYZL23, LM21, NPH^{+}20]. Side-Channels [BSP^{+}22]. sided [Yan19]. Sigma [ZYW^{+}18]. Signal [HRC21, LS22, MFHP12, STGR15, WGT^{+}17, ZSY18, CPW04, GMS^{+}23, KZKAKP23, LLLC13, SR12, TYH08, XZC09]. signal-integrity [XZC09]. Signals [Yan16, MKW08]. Significance [LJJ^{+}22, MHA19]. Signoff [LNPL23]. Silicon [ANS^{+}20, HAB^{+}17, PTS^{+}20]. SIMD [EKEK22, YCL^{+}23]. Similarity [PIK20, TYSF20, YRH11]. Simplifying [HA05]. Simulated [ZYS12, SMYH07]. simulating [RHA08]. Simulation [BLUS19, CDB11, EKS^{+}14, EO19, GDPRG11, HBPW14, HIW15, HPB11, IHM15, LS22, MDM^{+}12, PLH^{+}24, PRCK08, ST99, SKM^{+}16, WFSS20, WWFT12, XJF^{+}23, ZWD11, CVMP19, DCK10, DL11, HVF^{+}01, HKB^{+}07, KMC97, LOC12, PTC05, PHM00, RSR01, WTL^{+}13].

Simulation-Based

[EO19, PRCK08, LOC12]. Simulations [LS11]. Simulator [LAYZ23, LHK^{+}15, FWCL05, EBR^{+}09]. simulators [RPKC05]. Simultaneous [CC06, CYV^{+}14, CFX09, JK10, LXCH04, SM00, CCX06, CCW08, CW01, MRC06, YHH09]. simultaneously [HLCH07, SSP04]. Single [BD14, HCW^{+}16, KRL15, LSZ^{+}21, LQD22, PMA24, RM23b, SKS^{+}18, SSL17, VEO16, Yan19, Yan20, PTC05, VJBC07, YW09]. Single- [SKS^{+}18]. Single-Chip [BD14, PTC05]. single-detour [YW09]. Single-Electron [HCW^{+}16]. Single-Event [KRL15]. Single-Inverter-Based [VEO16]. Single-Layer [Yan20, Yan19]. Single-Rail [LQD22]. Single-Tier [SSL17]. Situ [HSP^{+}22, SL18]. Size [KCKG16, YVC14, ZLG^{+}19, AMR00, AM05, FNMS01, HH09, HKV^{+}07, LDK99, LH09, SBH^{+}06]. Sizing [CHK^{+}23, DZ18, KKS16, LLM^{+}23, LZ21, LGGJ14, SV16, SCK^{+}23, ZLL^{+}16, ZGB^{+}24, CW01, HR06, LG12, MLG12, RGM09, SC00]. Skew [CHH09, TCW20, CKKT98, HN07, HTCP13, LLHT12, LT11, wATkK02]. Skew-aware [CHH09]. Skewed [Pom19a, CSKR05, Pom14b]. Skewed-Load [Pom19a, Pom14b]. Slack [ASAP17, NRZ^{+}18, CGN96, KSA^{+}10]. Slack-Based [ASAP17, KSA^{+}10]. Slacks [PSNC18]. SLAM [BYT22]. Sleeping [TEK18]. Slew [WCCC14]. Slicible [DSK01]. SLO [HC18]. slow [NS03]. slow-speed [NS03]. Small [WGT^{+}17, XLCL13]. small-delay [XLCL13]. Small-Signal [WGT^{+}17]. Smart [AL19, FHL^{+}23, HXC^{+}18, HK18, JDD20, SKM^{+}16, YMB15, ZHC^{+}18, JS13, AL19]. Smart-Gateway [HXC^{+}18]. Smart-Grid [HXC^{+}18]. Smart-Hop [AL19]. SmartCap [LYHL14]. SmartDR [GdRJM21]. Smarter [HFMB20]. Smartphone [LYHL14].

Smartphones [LYLW17]. SMs [SBR^{+}17]. SMT [AA17]. SMT-Based [AA17]. SNNs [ZMLH24]. Snoop [PCT^{+}17, ZYDP08]. Snooping [GD22]. SoC [HZS^{+}19, GM03, GDF09, XZC09, BHW^{+}13, DCK10, Kan06, LLH^{+}17, LCL08, LXGM23, MOZ06, SBC08, TCL14, WLCJ09]. SOC-based [GDF09]. SoCDAL [AHL^{+}08]. SOCs [MSD06, BM11, JHMGS18, JPHL16, ZM07]. Soft [CWL^{+}22, DFM15, EKEK22, HWL^{+}23a, LD17, LW21, PHKW12, SWT23, TLCF16, QS09, RJBS09, ANS^{+}20]. Soft-Error [HWL^{+}23a, LW21, TLCF16]. Soft-Error-Rate [LD17]. Soft-HaT [ANS^{+}20]. Software [ANS^{+}20, BM11, CBR^{+}22, HLL^{+}24, JHMGS18, JJH21, KMR18, LLP^{+}16, LHF12, SYGC22, THT12, YYL^{+}15, ZHC^{+}23, AMO05, BASB01, CMM00, CACS05, CM13, FHHG12, GGB97, HKL^{+}07, JW08, KSK^{+}05, KTKO13, LMW99, LP07, LVL03, MSD06, ML09, NG06, SS11, WYIG07, WJY^{+}07, YWGI09, YGH^{+}10]. Software-Based [ANS^{+}20]. Software-Defined [JHMGS18]. Software/Hardware [CBR^{+}22, HLL^{+}24]. Solid [CCS15, CD09, CCYC14]. Solid-State [CCS15, CCYC14]. solid-state-disk [CD09]. Solution [GSFT16, JNS^{+}17, YFT17, YFT18, FNMS01, SR12]. Solutions [WFT^{+}19, CW01, NR01]. Solver [MS21, XJF^{+}23]. solvers [DP02, QSK12]. Solving [CYV^{+}14, HZJC23, WGDK07]. Some [KAKSP16]. SOPs [BCC08]. Sorting [ZMP16, Yan00]. Source [LRHL24, YKCG14, BCR^{+}08, KRK98, ZYZ^{+}13]. source-level [KRK98]. Source-Synchronous [YKCG14]. Sources [DHB16, CH96]. Space [AKAKP18, BSZ^{+}24, FLG^{+}23, FMR23, FCZ^{+}23, GACK22, GCZ^{+}15, HMMG^{+}20, PGGD23, Pom24b, RGX^{+}24, RS18, Sch17, SHBD21, WS22, APB^{+}08, ARLJH06, BW00, EK97, JP08, KSS^{+}09, RFG20, SW12, VCLD03]. Space-aware [PGGD23]. space-efficient [ARLJH06]. spaces [BC11]. spacing [MKW09]. spare [ACT13]. SparGD [WML^{+}24]. Sparing [NRM^{+}24]. Sparse [WML^{+}24]. Spatial [GFC^{+}09, RB19, Das09]. Spatio [SSC17]. Spatio-Temporal [SSC17]. Special [ADGSM22, BJX15, BMdG17, CCY22, CO18, HAW20, JCPL23, KLSZ11, LZR23, NWA^{+}24, PFHAH22, TK18, YD16, BC08, CH10a, CLQ12, HJ08, JW08, KLSZ09, LP07, MD13, Ped06, RBA^{+}12]. specialization [ADM^{+}13]. specialized [BC08]. Specific [HKL^{+}15, HMMG^{+}20, HCZ^{+}16, LPD^{+}17, LHF12, LF12, RCK^{+}15, TCL14, VA17a, ACT13, CSC08, SCV06, WKR09]. Specification [HZS^{+}19, HV98, MD08, VS12a, BD00, BGM04, HV07]. Specification-driven [MD08]. Specifications [DSHD23, LXGM23, Pie16, CMM00, DDNAV04, MB04, VKKR02]. Spectral [KOO18, ZF23, TN99]. spectral-based [TN99]. Speculative [NRDB19]. Speed [CK16, DMR23, Kha23, PTC^{+}15, RM23a, TPC^{+}17, NS03, OW06, PSD21, SXZV13]. Speeding [CLM^{+}10]. Speeding-up [CLM^{+}10]. Speedup [Che18, KAKSP16]. Speedups [GDTG07]. SPICE [LS22, XJF^{+}23]. Spill [LHF12]. Spin [RPR^{+}21]. Spin-Transfer-Torque [RPR^{+}21]. Spintronics [MS21]. Spintronics-based [MS21]. Split [SJ23, YCL^{+}20]. Splitting [BHY^{+}24, CZZYW21]. SPMCloud [BD14]. Spread [MJB19]. SQLite [LLP^{+}16]. SRAM [CCC^{+}09a, DMR23, HHL14, JLF^{+}12, NdLCR03, PMA24, PS23, RM23a, RM23b, ZYW^{+}18]. SRAM-based [JLF^{+}12]. SRAM/71mW [CCC^{+}09a]. SRAMs [RM09]. SSA [MHA19]. SSA-AC [MHA19]. SSAGA [SBR^{+}17]. SSD [WHXZ13]. SSDs [GSD^{+}18, HC18, LHS^{+}21, SLC^{+}22]. SSER [PHKW12]. Stability [HHL14]. Stack [APG24, WDZG16]. Stacked [SYX12, THM15, LHZ^{+}06]. Stacking [HKJ^{+}23]. Stage [LZ17, Shi20, KSA^{+}10]. Stage-form [Shi20]. Stages [KO23, SYL09]. staircases [MSKBD07]. Stairway [MHD^{+}04]. Standard [ACF^{+}11, DBK^{+}18, KRL15, TRM^{+}16, PR09, SSCS10, TS96]. Standard-Cell [DBK^{+}18, SSCS10]. standard-scan [PR09]. Standby [NRM^{+}24]. Standby-Sparing [NRM^{+}24]. Start [ZLY^{+}15]. State [AVG19, BHBS22, CCS15, CK16, Pom15a, RGX^{+}24, BDC08, CD09, CCYC14, CK96, CHHL96, HRP00, Pom14a, SNH02]. State-Based [AVG19]. States [Pom16c, LIA00]. Static [BDB12, ETAV18, KD24, LV14, MHA19, Pom15b, XPX^{+}21, ZFLS11, DH06, EMO03]. Statically [KKLG15]. Statistical [BBEM15, CV17, JGM14, KPR06, LM21, PHKW12, RPR^{+}21, SV16, STWX12, XT16, ZKS^{+}16]. statistics [SNH02, SXZV13]. steering [HKV^{+}07]. Steiner [CKKT98, GC96, HGLC16, LLLL18, LYKW09, SMYH07, Yan08]. Steiner-point [Yan08]. Stencil [YYG^{+}16]. Step [HGLC16, Vah02]. stimuli [MFS09]. Stimulus [CYV^{+}14, LV14, BLR06, PKP^{+}03]. stimulus-free [BLR06]. stitching [Meh98]. Stochastic [BH22, GLY^{+}12, MMP00, GBC07, NM13]. Stopper [PCT^{+}17]. Storage [BD14, CCH^{+}15a, CGLH23, HWDQ22, Kha12, KCA04, Pom24a, WQC^{+}16, ZLW^{+}15, ZMS^{+}19, BD08, Meh98, Wu09]. storages [HCK13]. STR [ZZ24]. STR-based [ZZ24]. Straightforward [LH09]. Strategies [HJY23, JM14, XLS15]. Strategy [KKHK16, ADDM^{+}13, ZHJ^{+}23]. stream [LWK11, NM13]. Streaming [LWX^{+}23, RS18, TY19, ZLL^{+}16, ZMP16, FHHG12, KSS^{+}09, WLL^{+}11]. Streamlining [LWX^{+}23]. Stress [HZJC23, LS19, WXH^{+}19]. Stress-based [HZJC23]. Stress-Induced [LS19]. striping [CCYC14]. Strong [AYS20]. Structural [CML98, CH00, AYM05, CL99a, HA05, VLH98]. Structure [AG22, KKHK16, FWCL05]. Structured [HLX^{+}23, THL^{+}13]. Structures [TB20, BK00, DDFR13, GMN^{+}13, Hua01, Meh98]. STT [JZYZ15, LSL^{+}13, SABSA15, SMBT19, WSS^{+}18]. STT-MRAM [SMBT19]. STT-RAM [SABSA15]. Stuck [TPC^{+}17, HVF^{+}01, PR09]. Stuck-At [TPC^{+}17, HVF^{+}01, PR09]. Study [LLP^{+}16, LYM^{+}20, MAL23, LC13, MLG12]. Style [CFD^{+}16]. Styles [LCYN18]. Sub [BFL10, PS23]. Sub-45nm [BFL10]. Sub-threshold [PS23, SHN12]. Subgraph [LNPL23, YYC07]. subnetworks [TDF^{+}09]. Substrate [WPL23, Yan20, LCJ^{+}10, SKCM06]. substrates [SKCM06]. subsystems [JSG09]. Subthreshold [BFL10]. Subtraction [BSP^{+}23]. Successive [HWCL15]. Successive-Approximation-Register [HWCL15]. Suited [GYZ^{+}22]. sum [DK08]. sum-of-product [DK08]. Super [LQL^{+}24]. Super-Resolution [LQL^{+}24]. SUPERB [EBR^{+}09]. Superposing [ZZ24]. Supervised [PLC24, RNA^{+}21]. Supply [BSP^{+}19, BM11, CUA^{+}24, JLK15, SLP^{+}19, WCCC14, XRS^{+}19, YFT17, YSF^{+}18, YFT18, YBS^{+}18, JR97, LLHT12, WLCJ09]. Support [MCZ^{+}16, WKL^{+}18, ZP08]. Supporting [LYL^{+}19, ZLL^{+}16]. Supports [MLH^{+}17]. Suppressed [BC16]. Surrogate [WFSS20, ZBG^{+}23]. Surrogate-Based [WFSS20]. Survey [BFG17a, BRCS18, GLD^{+}22, HHH^{+}21, KAC^{+}23, LM19, Mit16, MRL^{+}19, PTPB22, RJ14, SSK^{+}23, WM24, WCX^{+}24, BD97, CEB06, KG99, KP13, SW04]. survivability [ACT13]. suspect [DNA^{+}12]. Suspension [NSH^{+}16]. Sustainable [CXH^{+}16]. SW [ADP^{+}07, BFV15, FLPP09, WWFT12]. Swarm [HLG^{+}15]. Switch [CYLC24, HSR^{+}24, MMM^{+}22, CWW96, CZW^{+}03, FLWW02, FLWC07, KS23, RFYL98, THL^{+}13, ZHTC09]. switchboxes [DSKB04]. switched [CSC08, HWCL13]. switched-capacitor [HWCL13]. Switching [AVG19, BP23, GSS14, RM23b, SRC15, BLR06, HCN09, PR11, SXX^{+}06]. switching-activity [SXX^{+}06]. SwitchX [BP23]. Symbolic [BDM^{+}99, BFG17b, DY23, MCD12, SHD17, BLM00, FWCL05, KVMH08, YWGI09]. Symbolic-Event-Propagation-Based [MCD12]. symmetric [IAI^{+}09]. Symmetrical [OCK19, CZW00]. symmetries [CMB07]. Synaptic [HSP^{+}22]. Synchronizing [MDM^{+}12]. Synchronous [CH17, HPB11, PMS15, TB20, WWW^{+}12, YKCG14, ZABGZ17, BDM^{+}99, BASB01, CACS05, CPR^{+}02, HKB^{+}07, MB04]. SynergyFlow [LYL^{+}19]. Synthesis [AG22, AA17, BR12, BD00, BSP^{+}23, CSKR05, CET16, CXS^{+}23, CS22, CLMZ10, CCL03, EO19, EWT23, FCZ^{+}23, GBR07, HS18, HRC21, HMVG13, HCZ^{+}16, ISK21, JJH21, KK14, KKK12, KKS16, LS17, MWK21, NG06, OCK19, PDS12, PG15, PFHAH22, QZZW24, QSW^{+}15, RCW22, RJ14, Sch17, SGC^{+}14, SS14, SGGR14, SLV^{+}22, SV11, SCCH08, UE22, WCCC14, WS22, YMB15, ADS^{+}09, BDM^{+}99, BZ08, CLLK06, CMM00, CBMM10, CL99b, CD96, DDNAV04, FHHG12, GG99, GOC02, GH00, GGDN04, GWR13, HLKN07, HCLC98, Hsi01, HLHT08, Hua01, JLF^{+}12, KSS^{+}09, KKH^{+}02, KK11, KW02, KHP05, KFH^{+}08, LCD07, LC14, Lin97, LLHT12, LWH06, MMP00, MDM07, MKBS05, MJM11, MRC06, PBSV^{+}06, RFYL98, RS03, SW12, SCB01, SV07, TN99, TC98, VLH98, VKT02, VKKR02, WV02, WG11, WKR09, XK97, XPSE12, YWW10]. Synthesis-time [BSP^{+}23]. Synthesized [RB21, SBR^{+}17]. Synthesizing [GSS14, GNQ^{+}22]. synthetic [PSK08]. System [BdM00, CH17, DMR10, GM08, GPH^{+}09, HKL^{+}15, HZS^{+}19, LL15, LG18, NAK20, NRZ^{+}18, PDS12, PPDK09, Pie16, PBSV^{+}06, RFG20, SL18, SGGR14, TK18, WL12, YYG^{+}16, ZHM07, APB^{+}08, BPRR98, BMJ13, Cha01, CKAP07, CSC08, CGLH23, DC07, GG99, GABP00, HGBH09, HMVG13, HW00, LTH99, LCC11, MOZ06, MPSJ07, OCRS07, Ped06, SPG^{+}08, Sen11, Vah99, ZLL13, dW97, AHL^{+}08, LVL03, WLL^{+}11]. System-Level [HKL^{+}15, LL15, LG18, PDS12, Pie16, BdM00, GM08, PPDK09, RFG20, ZHM07, MOZ06, OCRS07, Ped06, Sen11, Vah99, ZLL13]. system-on-a-chip [Cha01, CKAP07]. System-on-Chip [HZS^{+}19, SGGR14, APB^{+}08, BMJ13, CSC08, WLL^{+}11, AHL^{+}08]. System-scenario-based [GPH^{+}09]. Systematic [AMM^{+}06, SLP^{+}19, KPR06, RPKC05]. SystemC [BK10, CVMP19, GD20, HV07, WWFT12, ZMS^{+}19, RHA08]. SystemC-AMS [CVMP19, ZMS^{+}19]. SystemC-based [GD20]. SystemCoDesigner [KSS^{+}09]. Systemization [ZHC^{+}23]. SystemJ [MSR09, SPT^{+}17]. Systems [ALLE20, ADGSM22, APG24, BHK17, BLNK14, BJX15, BSP^{+}22, BB17, BXG^{+}24, BS14c, CLL^{+}22, CHA^{+}23, CH10a, CCH^{+}15a, CHBK15, CXLL22, CYH19, DFM15, DHX^{+}23, DHW^{+}23, EAP17, GT21, HXZ^{+}23, HK18, IGN18, JJH21, KLSZ09, Kha23, KC10, KMR18, LL15, LWX^{+}23, LHK^{+}15, LZZSV15, LWG^{+}23, LMA^{+}16, LL19, LZA^{+}21, MRL^{+}19, NSH^{+}16, NDA^{+}23, NRM^{+}24, ORGD^{+}15, PLC24, PPP^{+}15, PSNC18, PG15, PBZM19, PY20, QBTM16, RFG20, RG19, RNA^{+}21, SSC17, SPT^{+}17, SRKS23, SBY^{+}20, STWX12, SS14, SHBD21, SAL19, TB20, THT12, TL19, UPV23, WLZ^{+}19, WHRC12, WQC^{+}16, WDD^{+}23, WDLX21, XPZ^{+}18, XGC^{+}20, YBM^{+}21, YRH11, ZLW^{+}15, ZMS^{+}19, ADM^{+}13, AM10, ADDM^{+}13, ARLJH06, BD00, BWB14, CSAHR07, CMM00, CSL^{+}07, Con06, CLQ12, CCL04, DCK07, DRG98, DDNAV04, DTC^{+}09, GDTG07, GPH^{+}09, GDF09, HKL^{+}07, HV07, HDL^{+}12, HCLC98, Hsi00, HBC^{+}08, JS13]. systems [JWL^{+}03, JW08, KKMB02, KC13, KP13, KFH^{+}08, LCZ^{+}08, LCK^{+}09, LSDV10, LDK99, LP07, MBB01, MDG98, MHQ07, ML09, OKC08, PDN00, PCD^{+}01, PSL^{+}98, Ped11, PEPP06, QS09, Rak09, RSR01, SCB01, SLXZ12, SUC01, SHN12, SS11, SZV^{+}12, THC^{+}14, Wol96, Wu09, ZAJ^{+}12, ZP08, SN10, CPX14]. Systems-on-Chip [BHK17, HDL^{+}12, KP13]. Systems-on-Chips [LWX^{+}23]. SystemVerilog [CYV^{+}14].

T [YYC09]. T-trees [YYC09]. TAAL [JZG21]. table [KSD^{+}22, WSEA99]. table-based [WSEA99]. tables [CH02, YTHC97]. Tag [YBS^{+}18]. tagged [ZP08]. tailored [Ase23]. Tailoring [CSC08]. Taming [FHHH22]. Tampering [HYK^{+}20, JZG21]. Tandem [MSR09]. tap [GMS^{+}23]. Tapered [BSP^{+}23, KKHK16]. Target [KGS^{+}20, KYL16, PBWB21, PBF^{+}22, Pom20, FS13, KR23]. Targeted [SNL12]. Targeting [LPD^{+}17, LZY^{+}23, PTPB22, JBC^{+}10, MLMM08]. Task [DHW^{+}23, ENP20, LJZ^{+}24, LMA^{+}16, SZB17, XXC^{+}24, DCK07, GK14, GBC07, YYLL09]. Tasks [CH17, SSC17, WJG^{+}19]. taxonomy [KP13]. TCAM [VNS19]. TCONMAP [HABS15]. tdf [ZMTC13]. TDM [VGG19]. TDM-based [VGG19]. Team [RGX^{+}24]. Technique [CV17, JK10, JPM^{+}19, LGGJ14, SBB^{+}18, DHV^{+}00, HLCH07, IBMD07, KI01, LC96, MB04, Mut09, RSR01]. Techniques [GD20, GdRJM21, MDM07, Mit16, PTC^{+}15, SJ23, TWL16, WSV^{+}14, YD16, AM05, BD97, BdM00, BH10, BASB01, CLM^{+}10, CSAHR07, CACS05, CFHM09, DS06, DD02, HPK99, HCS01, HCC01, KSK^{+}05, KMS12, KHP05, LSDV10, LB00, LHW97, LHCT05, LVL03, OCRS07, OK08, PCD^{+}01, RJBS09, TY97, TBZ13, TYH08, VMP^{+}00, XK97, ZHOM08]. Technologies [PFHAH22, SN10, BC08]. Technology [ATF^{+}23, BFL10, CHY05, DKT^{+}16, DBK^{+}18, GLD^{+}22, HABS15, JZYZ15, PS23, SABSA15, YD16, ZS02, BLM00, CH02, CH00, KL05, LKM04, PL98, WY06, WSEA99, ZLL13]. technology-dependent [BLM00]. Technology-Driven [DKT^{+}16]. TEI [LHW^{+}17]. TEI-power [LHW^{+}17]. Temperature [BHY^{+}24, JGM14, LHW^{+}17, SRKS23, ZYP09, ADP^{+}07, CLQ12, DH06, WJY^{+}07]. Temperature-aware [SRKS23, ZYP09, ADP^{+}07, CLQ12]. template [HGBH09]. Temporal [Pie16, SSC17, YYC07, BD05, Das09, YYC09]. Temporally [PRCK08]. Tensor [HZL^{+}22, SYH^{+}22]. terminals [ISE08]. Terrestrial [PMA24]. Test [AYM05, BDBB19, EMO03, EO19, FHL^{+}23, GF06, IE12, LCT03, LYSO19, LM21, MCD12, NSCM17, PKJK20, Pom15a, Pom15b, Pom15c, Pom16b, Pom16c, Pom17a, PAV17, Pom18a, Pom19b, Pom20, Pom21a, Pom22, Pom24a, Pom24b, RJ14, SBB^{+}18, TBZ13, WCB15, WWCT18, WH19, WH20, WZH^{+}23, WLM21, WC10, WWW^{+}12, XCW12, XLCL13, Xia24, BC05, BWB14, Cha01, Che96, CCL04, ETR07, FNMS01, GM03, HLKN07, HRP00, HJ08, IYF^{+}21, KT01, LTH99, MD08, NCP01, NT05, PR98, PR07, PR11, QM12, RMKP03, SW04, SBC08, SEN05, SNL12, TCP97, TD03, WPHL08, WWC04, XZC09, ZMTC13, SSGS03]. Test-Architecture [WWCT18, XZC09]. Testability [LW21, NWA^{+}24, Pom16a, Pom18a, FRS97, PSK08, Pom14a, SCJ01].

Testable [GBR07, LW21, RMPJ08]. testbenches [BFP08]. testers [NS03, SBC08]. Testing [LPY^{+}20, NS03, PTC^{+}15, TPC^{+}17, WWCT18, WJM24, WWW^{+}12, XCW12, XS16, XCF18, Xia24, JT98, KBN09, LHCT05, PKP^{+}03, SEN05, SXZV13, SCJ01, SOC06, TD03, XZC09]. Tests [Pom15a, Pom16a, Pom16c, Pom18b, Pom19a, Pom19b, Pom20, Pom21a, Pom21b, Pom24b, DNA^{+}12, PR09, Pom13, Pom14a, Pom14b]. text [LDK99]. text-compression-based [LDK99]. Theft [BTP^{+}20]. Their [MLH^{+}17, PTPB22, DSK01]. theoretic [HR06]. Theoretical [TB20, SB98]. Theories [PG15, YW09]. Theory [CXLL22, KR23, MDM^{+}12, SSK^{+}23, JWL^{+}03]. Thermal [CK19, CLT^{+}15, CXH^{+}16, CVMP19, CAP^{+}23, CR12, DCK10, JGM14, LCK^{+}09, LHW^{+}17, LDD^{+}18, LZA^{+}21, MDR15, OCK19, PSP24, RKKH24, SBY^{+}20, SKP21, WMT^{+}16, ZHC^{+}18, ZF23, ADDM^{+}13, ANR13, GK14, LH13, LHZ^{+}06, LTPT10, QSK12, WTL^{+}13, WJY^{+}07, YHH09, ZAJ^{+}12, ZSZ10]. Thermal-Aware [SBY^{+}20, SYX12, OCK19]. thermal-oriented [LHZ^{+}06]. Thermal-Sensor-Based [ZHC^{+}18]. Thermally [RGM15]. thermodynamic [VLH04]. Things [TK18]. Thread [CNQ13, SV11, KBA08]. Thread-based [CNQ13]. threaded [HC17]. Threat [MCY23, YBM^{+}21]. Three [KQP^{+}19, LQD22, RGM15, WXH^{+}19, Yan00, Vah02, YYC07, YYC09]. Three-Dimensional [RGM15, KQP^{+}19, WXH^{+}19, YYC07, YYC09]. Three-layer [Yan00]. Three-Phase [LQD22]. three-step [Vah02]. Threshold [CZW19, DHVW18, LYL^{+}23, SV16, PS23, SHN12]. Throughput [HCRK11, HIW15, KLJ14, MS23, SESN15, ZZ24, CJLZ11, EKEK22, GM08, PRKK21, SKS12, SHN12]. throughput-aware [SKS12].

Throughput-Optimized [HCRK11].

Thwart [BTP^{+}20, LSCK20]. Tier [SSL17]. TIGFET [LQD22]. TIGFET-Based [LQD22]. tightly [LMB^{+}12]. tightly-coupled [LMB^{+}12]. Tightness [APS18]. tile [DJP21]. Tiled [DK16]. Tiled-DNUCA [DK16]. Time [APDC17, BB17, CHA^{+}23, CHBK15, CBC22, CH17, CJKK19, FG18, GYZ^{+}22, HXC^{+}18, IGN18, KPF16, KPB19, LM19, LSZ^{+}21, LSCK20, LWG^{+}23, LQL^{+}24, NSH^{+}16, NRM^{+}24, PSNC18, PGGD23, PY20, SSC17, SBY^{+}20, SLV^{+}22, WLZ^{+}19, WDZG16, WJG^{+}19, YRH11, ZLW^{+}15, ZZCY17, APB^{+}08, ARLJH06, BSP^{+}23, CSAHR07, DP02, DRG98, DZK^{+}24, FHHR21, HMLL11, HLKN07, HMVG13, KS23, KNRK06, LCHT02, LTPR^{+}13, MR96, MHQ07, NG06, PEPP06, PW99, SCB01, SWT23, WGDK07, WLL^{+}11, ZAZ13]. Time- [PGGD23, ARLJH06]. time-constrained [NG06, SCB01]. time-constraints [CSAHR07]. Time-Division [PY20, LWG^{+}23]. time-domain [LTPR^{+}13]. Time-Multiplexed [LM19]. Time-Sensitive [CHA^{+}23, DZK^{+}24]. Time-Triggered [BB17, IGN18, KPB19]. time/resource [WGDK07]. Times [PMS15]. Timing [CZW00, CB17, CJKS24, HIW15, HS19, JNCS19, KKK12, LVS16, LJ18, LWC18, LYCP17, LNG^{+}16, LL19, MJM11, MKW08, TB20, VBP^{+}19, WSH^{+}18, WKC12, WL12, Yan08, YRH11, DCK09, DRG98, DH06, KPSW09, KPR06, KC98, LC14, LCHT02, MCMW08, QS09, SXX^{+}06, SCCH08, YHL^{+}11]. Timing-aware [MKW08]. Timing-Driven [LNG^{+}16, CZW00, Yan08, DRG98]. timing-error [SCCH08]. Timing-Yield [WSH^{+}18]. TinyOS [RFB10]. TLB [KSK^{+}05]. TLC [CWL^{+}22, WZL^{+}21]. TLM [BFP08, ZMS^{+}19]. TLM-to-RTL [BFP08]. TMDS [SRKS23]. TODAES [CH10a, KLSZ09, BC08, GK09, QS11, TK18]. Toffoli [MDM07]. Toggles [TPC^{+}17]. Tolerability [LW21]. Tolerance [GVJ15, JPM^{+}19, BXG^{+}24]. Tolerant [CYH19, GT21, LW17, PMA24, XCF18, CEB06, NdLCR03, NGL^{+}21, SC06]. tolerate [SPG^{+}08]. Tolerating [ZHC^{+}21]. Tool [BBEM15, JHMGS18, TDE08, VLH98]. Toolchain [GVJ15]. toolkit [MSD06]. tools [BdM00, GS00, MD13, MT02]. Top [SSN22]. Top- [SSN22]. Topological [SHD17]. Topologies [Kha23]. Topology [BDBB19, HCZ^{+}16, UE22, TDF^{+}09]. Topology-Agnostic [BDBB19]. Torque [RPR^{+}21]. Trace [BHK17, SJN24, BHW^{+}13]. Trace-Based [BHK17]. Traceability [IK19, YFT17]. track [LCC11]. Tracking [HMO^{+}14, NPH^{+}20, FS13]. Trade [BHY^{+}24, KSD^{+}22, MS23, PCC09, FHHG12, RJL^{+}09, WVYG99, WGDK07, XPSE12]. Trade-off [BHY^{+}24, KSD^{+}22, MS23, RJL^{+}09]. Trade-offs [PCC09, FHHG12, WVYG99, WGDK07, XPSE12]. Tradeoff [RS18]. Tradeoff-Aware [RS18]. Tradeoffs [LDD^{+}18]. Trading [FG18]. Traffic [QBTM16]. Train [TZZH22]. Training [ALL17, HSP^{+}22, JSS^{+}19, LS23, LCG^{+}22, PLC24, TZZH22]. Transactions [CH10a, CPX14, KLSZ09]. Transceivers [JNS^{+}17]. Transfer [LRHL24, RNA^{+}21, RPR^{+}21, WJM24, ZHL^{+}23, KI01, KVMH08]. Transform [HHX^{+}23, LCC^{+}15]. Transformation [SPC^{+}15, ZFL22, BGN^{+}07, KKH^{+}02, Vah99, VJBC07]. transformational [Voe01]. transformations [HKV^{+}07, LLM01, PCC09, WVYG99]. Transformer [LY24]. Transforms [ACFM12, MFHP12]. Transient [KRL15, SQL^{+}24, DC07, MRC06]. Transistor [CFD^{+}16, HCW^{+}16, PR96, RS03, WSH^{+}18].

Transition [JOH17, MHQ07, Xia24, LHCT05, PL03, PR09, WPHL08]. Transition-overhead-aware [MHQ07]. Transitions [DY23, Mut09]. transitive [YYC07]. Translation [MWS^{+}20, WL12]. transmission [KC13]. Transmissions [CBO^{+}18]. TransNet [RNA^{+}21]. Transparency [WHRC12]. Transparent [Pom17b, SV11, PR11]. Transparent-Scan [Pom17b, PR11]. Transposition [CCH15b]. traversal [HRP00]. Tree [FZL^{+}23, HGLC16, KK11, KKS16, LLLL18, LNG^{+}16, LS17, OCK19, PSD21, WCCC14, ZFL22, CHH09, LLHT12, LYKW09, LLLC13, TDF^{+}09, wATkK02, Yan08, YYC09]. Tree-based [PSD21, YYC09]. Trees [CCH15b, EK16, GC96, WCC03, YYC09]. Trends [CH10b, HHL14]. Triggered [BB17, DY23, HS18, IGN18, KPB19, BDC08]. Triggering [EW18b, HW14]. Triple [LZ17, ZLY^{+}15]. Tristate [CK16]. TRNG [PBH^{+}24, ZZ24]. Trojan [ANS^{+}20, LM21, MRL^{+}20, YCL^{+}20]. Trojans [DY23, SGJN24, VTC20, XFJ^{+}16]. TROP [SGJN24]. True [MAL23]. True- [MAL23]. Truncated [Pom22]. Trust [GSFT16, ZGB^{+}23, SGJN24]. TRust-aware [SGJN24]. Trust-region [ZGB^{+}23]. Trustworthy [CCMC20]. TSN [MAS^{+}20]. TSocket [CXH^{+}16]. TSV [KK11, KKHK16, WDC^{+}22]. TSV-based [KK11]. TSV-Inductor [WDC^{+}22]. Tunable [OK20, CFHM09]. tuned [RFB10]. Tuning [PTS^{+}20, ZGB^{+}23, LT11, SZV^{+}12, YCL^{+}23]. Turbine [WSRH16]. Turn [HSR^{+}24]. Turn-Restricted [HSR^{+}24]. Tutorial [Edw03]. TVM [HTC^{+}23, YCL^{+}23]. TVM-based [HTC^{+}23]. twisted [YW09]. Two [HLZ^{+}22, LZ17, OW06, Pom24b, TJ99, Yan19, CSC08, DDNAV04, LHZ^{+}06]. Two-dimensional [Pom24b]. Two-layer [OW06, DDNAV04]. Two-level [TJ99]. Two-Part [HLZ^{+}22]. Two-sided [Yan19]. two-stacked-die [LHZ^{+}06]. Two-Stage [LZ17].

UAV [LJZ^{+}24]. UAV-aided [LJZ^{+}24]. UCR [YBS^{+}18]. Ultra [ACF^{+}11, CK16, GBC07, Kha23, MACV14, SESN15, ZLG^{+}19]. Ultra-fast [GBC07]. Ultra-High [Kha23]. Ultra-High-Definition [ZLG^{+}19]. Ultra-High-Speed [CK16]. Ultra-Low [ACF^{+}11, MACV14, SESN15]. UltraScale [AMM^{+}18]. Unauthorized [CBO^{+}18, GDTF17, KOO18]. Unbounded [VS12a]. Uncertain [CXLL22, KW16]. uncertainties [CS07]. Uncertainty [CXLL22, GC18, STGR15, YB23]. Uncertainty-aware [YB23]. Unclonable [Ase23, BHY^{+}24, CSC^{+}21, LLQD23, YBS^{+}18]. Uncore [WGSH16]. Understanding [HHL14]. Undetectable [Pom19b]. Unicast [XS16, XCF18]. Unicast-Based [XS16, XCF18]. unified [Kag05]. Uniform [HZS^{+}19, KCKG16, HKJ^{+}23]. Unique [SOS15]. UNISIM [LS11]. UNISIM-Based [LS11]. Unison [SGJ96]. Unit [BM11, HWCL15, JDLZ24, ZXC^{+}23, HWCL13]. Unit-Capacitor [HWCL15]. Units [CLC^{+}24, LCJ^{+}22]. Universal [CWW96, CJKK19, JCK^{+}18, FLWW02, FLWC07]. universality [RHN00]. Unknown [SSO16]. Unknowns [EKS^{+}14]. Unmanned [HXB^{+}22]. Unnecessary [Pom15c]. unpredictabilities [DS05]. unpredictability [SPG^{+}08]. unscheduled [MHF96]. Unstructured [VTC20]. Untangling [Yan19, YW09]. untestable [LIA00]. UPaK [WKR09]. Update [KC10]. Upper [IIEKS23, JLJ15]. Upper-Bound [IIEKS23]. Upset [PMA24, NdLCR03, RM09]. upsets [MRB^{+}11]. Use [KBV^{+}15, KFH^{+}08, MS00]. use-cases [KFH^{+}08]. Useful [TCW20]. Using [APDC17, APD^{+}11, ASAP17, AVG19, AGM01, BBEM15, BDB12, BS14b, BM11, BLUS19, CM19, CAOM19, CYV^{+}14, CJKK19, CLC^{+}24, DCC^{+}23, DNA^{+}12, EW18a, EW18b, EWT23, EK16, FZL^{+}23, FWCL05, FHHR21, FYCT15, GFJ16, GBR07, GNGT21, GD20, GHYR19, HS18, HWF^{+}23, HWL^{+}23a, HLL^{+}24, JBJ22, JNS^{+}17, JSS^{+}19, KQP^{+}19, LHS20, LLH^{+}17, LFST21, LYHL14, LYSO19, LSCK20, LCZ^{+}24, LLK^{+}14, LCC^{+}15, LNPL23, LXGM23, LRHL24, LM21, MA16, NL24, NPH^{+}20, NRM^{+}24, PJL14, PMT20, PG15, PR09, Pom15a, SMS22, SS24, SKS^{+}18, SJN24, TB20, TYSF20, THM15, TMDF10, TCL14, WKL^{+}18, WXH^{+}19, WSS^{+}18, XXC^{+}24, XGWL24, YHL^{+}11, ZHC^{+}18, ZYS12, ZMS^{+}19, BLR06, BWB14, BK10, BGN^{+}07, BASB01, CACS05, CBMM10, CFHM09, CK96, GGBZ02, GK07, GK09, HVF^{+}01, HMB98, HPK99, HCC01, HW14, KSK^{+}05, KRS06, KPR06, KMS12, KMC97, LCT03, LSL^{+}13, LON08, MHD^{+}04, MSR09, MS08]. using [MR05, MP07, MLC08, MVK^{+}18, NRZ^{+}18, PRCK08, PKP^{+}03, PMB10, PHM00, RJL^{+}09, RCD07, SGK08, SABSA15, SFM^{+}19, STL^{+}13, SYH^{+}22, SBH^{+}06, SCJ01, TLCF16, TWL16, TN99, TD03, TYH08, Vah02, WVYG99, WJYZ11, WCC03, XLCL13, XK97, YTHC97, YYC07, ZHOM08, ZHC^{+}23]. UST [wATkK02]. UST/DME [wATkK02]. Utilisation [NAK20]. utility [BCR^{+}08]. Utilization [HKJ^{+}23, KKLG15, KMR18, MT15, GM03, SBC08, SY07]. Utilizing [BLNK14, CK16, DZ24, EBR^{+}09, LQD22, LLQD23]. UTPlaceF [LLL^{+}18].

V [BSZ^{+}24, MLMM08, YCL^{+}23].

Validation [HLL^{+}24, RB21, VS12a, CM13, DRG98, FLPP09, HJ08, MD08, QM12, RPKC05, WAZ98]. Value [DZ24, YGZ04]. Valued [WTR12]. Values [Pom18a]. Variability [CFD^{+}16, JIR^{+}21, NRZ^{+}18, TY19, LON08].

Variable [PSNC18, ZLG^{+}19, LHW97, WH05]. Variables [Pie16, CCQ98, Pom14a, SXZV13]. Variation [APDC17, AKAKP18, BXG^{+}24, FYCT15, GPS^{+}24, HXZ^{+}23, LSZ^{+}21, RGM09, SCK^{+}23, WCCC14, WDLD17, WSH^{+}18, GM08, KTKO13, MJM11, PPDK09]. Variation-Aware [FYCT15, SCK^{+}23, WSH^{+}18, LSZ^{+}21, RGM09, MJM11, PPDK09]. Variation-tolerance [BXG^{+}24]. Variations [BHY^{+}24, GC18, TWM^{+}23, XAG^{+}20, ZZCY17, KPR06, LH13, LTPR^{+}13, ST99]. various [WAZ98]. Varying [RG19, SSO16]. VBR [JLJ15]. Vdd [HLHT08]. Vector [BSP^{+}19, JK10, LCJ^{+}22, PIK20, CCW08, EMO03, KBA08]. vector-thread [KBA08]. Vectorized [BSP^{+}23]. Vectorizing [LPD^{+}17]. Vectorless [ZF23]. Vectors [Pom15c, Pom21b, CK96]. Vehicle [VA17b]. Vehicles [HXB^{+}22, LJZ^{+}24, dONH23]. Verification [Ali12, BKW15, DSHD23, DSH12, EW18a, HZS^{+}19, KYN^{+}12, LXGM23, PKJK20, Ped11, SSS^{+}19, VBP^{+}19, ZF23, BHW^{+}13, BDC08, BGM04, DCK07, DCK09, DCK10, DC07, GF06, HA05, HDL^{+}12, HV98, KMS12, KG99, KC98, LBV^{+}06, LOC12, MS08, MPDG09, PRCK08, RFYL98, RBA^{+}12, Sen11, VAAH^{+}98, VS12b, WYIG07, WWC04]. Verify [KRH18]. Verifying [APD^{+}11, HCC01]. VeriGen [TAP^{+}24]. Verilog [TAP^{+}24]. versatile [TYH08]. Vertical [AJK^{+}21, LLKC13]. Vertices [SSN22]. Very [ZHC^{+}21]. Very-Large-Scale [ZHC^{+}21]. VFI [DLC^{+}17]. VFI-Based [DLC^{+}17]. vGreen [DMR10]. VHDL [DDNAV04, GDPRG11, MR96, MWG97]. VHDL-AMS [DDNAV04]. Via [SHL^{+}19, WPL23, BZWZ17, CRT19, CSO22, CCC09b, FHL^{+}23, GPS^{+}24, HHL14, HSA^{+}04, IPWW17, IK19, JYHY21, JYY^{+}22, KOO18, KRL15, KLK^{+}17, LHZ^{+}06, PB12, PTS^{+}20, RAKK12, SAL19, VAAH^{+}98, WB16, WHXZ13, Yan20, YWGI09, ZZL^{+}23]. vias [YHH09]. Victim [NAK20, SSS^{+}19]. Video [MDR15, SJL23, ZLG^{+}19, CCC^{+}09a, ZHOM08]. Videos [LWX^{+}23]. viewpoint [LKTD98]. Violations [KO23, Das09]. Virtual [BHDS09, DMR10, JLJ15, MSR09, SSL17, Fuj05, KMC97, LLKY13, ZP08]. virtualization [ISE08]. Virtuoso [LWX^{+}23]. visibility [HW14]. visual [FS13]. VLAN [SRTG19]. VLIW [AMR00, GBK07, KJR^{+}07, LJV02, LLHT03, LYCP13, SXX^{+}06]. VLSI [CHK^{+}23, DPNB02, DK22, DD02, GMN^{+}13, GOC02, HLG^{+}15, JT98, LM96, MSKBD07, MYSZ23, MKW09, OS03, RS03, STWX12, SB98, SSCS10, UPV23, ZGB^{+}23]. VLSI-CAD [SB98]. VNCS [BXG^{+}24]. Volatile [AKM^{+}22, HSP^{+}22, WDLD17, LSL^{+}13]. Voltage [BHY^{+}24, CS22, DHVW18, DMR23, DS05, GNGT21, JPHL16, JLK15, KLE18, LCY12, MACV14, RM23a, SV16, SCK^{+}23, WCCC14, WGSH16, ZLL13, GM08, GBC07, KSA^{+}10, LHW97, LLHT12, MHQ07, ML09, Rak09, SHN12, WWG08, WLCJ09]. Voltage-Based [GNGT21]. Voltage-Frequency [JPHL16, GM08]. voltage/frequency [ML09]. voltages [JR97, MR05]. Volume [Pom16c, RMKP03]. Volumes [PAV17]. VOR [SJL23]. vs [KG09, MAL23, PDN00, SA24]. VSSD [CCS15]. Vulnerabilities [GQW19, MAS16, PTPB22]. Vulnerability [NPH^{+}20].

W [DHZ^{+}11]. Wafer

[THM15, BC05, WLT08, ZMTC13]. wafer-probe [BC05]. Wafer-to-Wafer [THM15]. Wake [WSRH16]. Walks [BS14b]. Wall [VFML23]. Warp [LSV06]. Warping [SV11]. Washing [MGR^{+}15]. Waste [KR23, PBWB21]. watt [RAKK12]. waveform [MCMW08]. Wavelet [AHAKP08, GFC^{+}09]. Wavelet-based [AHAKP08]. WaveSync [YKCG14]. WCET [APS18]. WCPNet [XXC^{+}24]. WCRT [CYH19]. Weakly [NRM^{+}24]. Wear [CCH^{+}15a, CHC^{+}16, Kha12, CD09]. Wear-Leveling [CCH^{+}15a, CD09]. Wearable [FG18, RG19, RNA^{+}21]. Wearables [GFJ16]. WEB [MS08]. Weights [HWDQ22, VFML23]. Well [GMS^{+}23]. Well-tap [GMS^{+}23]. while [QS09]. Wide [WTR12]. width [LYCP13, SBH^{+}06]. Wind [WSRH16]. Wire [CZW19, CLC20, LLM^{+}23, WPL23, Yan20, CW01, HR06, MKW09, WC06]. Wire-bonding [WPL23]. Wirelength [XXC^{+}24, LLLC13, SYZ08]. Wireless [CBO^{+}18, GADG19, NSS^{+}16, PDS12, WM24, DHZ^{+}11, JSG09, RFB10]. wiresizing [CH96]. with14nm [PS23]. within [SCK18, SAHF^{+}20]. Without [MS17, FHHR21, KKLG15, PTPB22, PR07]. Wolf [ZHJ^{+}23]. Word [CCC09b, Con06, WDLD17, RMB10]. Word- [WDLD17]. Word-length [CCC09b, Con06]. work [KYN^{+}12]. Workload [CSAHR07, GC18, LG23, PKJK20, TBCH17, CR12, WHXZ13]. Workload-ahead-driven [CSAHR07]. workstations [KMC97]. world [RBA^{+}12]. worm [FNP09]. wormhole [TDE08]. Worst [APDC17, CH17, DCC^{+}23, ZLW^{+}15]. Worst-Case [APDC17, CH17, DCC^{+}23]. wrapper [LV02]. Write [CCK^{+}18, CIX15, KYL16, LLP^{+}16, WDLD17]. Write-back [CCK^{+}18]. Write-Conscious [LLP^{+}16]. Write-Induced [CIX15]. Writeback [PBZM19]. Writeback-Aware [PBZM19]. Writebacks [BFG^{+}19]. Writes [CIX15].

X [HLG^{+}15]. X-Architecture [HLG^{+}15].

Xbars [BP23]. XFM [SMSB05]. XNOR [ZHJ^{+}23]. XNOR/OR [ZHJ^{+}23]. XNOR/OR-based [ZHJ^{+}23].

Yield [GLY^{+}12, GPS^{+}24, JGM14, KAKSP16, KMO^{+}12, SV16, SYH^{+}22, THM15, WSH^{+}18, ZYW^{+}18, HWCL13, KPSW09, LCKT12, MHT14].

Zero [BC16, Giv06, JK10, HTCP13, ZCG06]. zero-deficiency [ZCG06]. Zero-Suppressed [BC16]. Zoom [EO19]. Zoom-ADC [EO19].

References

Azarbad:2017:SSB

[AA17] Mohammad Reza Azarbad and Bijan Alizadeh. Scalable SMT-Based equivalence checking of nested loop pipelining in behavioral synthesis. ACM Transactions on Design Automation of Electronic Systems, 22(2):22:1–22:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Asadinia:2015:PLP

[AAA15] Marjan Asadinia, Mohammad Arjomand, and Hamid Sarbazi Azad. Prolonging lifetime of PCM-based main memories through on-demand page pairing. ACM Transactions on Design Automation of Electronic Systems, 20(2):23:1– 23:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Agrawal:2017:OSA

[ABC^{+}17] Prabhav Agrawal, Mike Broxterman, Biswadeep Chatterjee, Patrick Cuevas, Kathy H. Hayashi, Andrew B. Kahng, Pranay K. Myana, and Siddhartha Nath. Optimal scheduling and allocation for IC design management and cost reduction. ACM Transactions on Design Automation of Electronic Systems, 22(4): 60:1–60:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Absar:2006:RAI

[AC06] Javed Absar and Francky Catthoor. Reuse analysis of indirectly indexed arrays. ACM Transactions on Design Automation of Electronic Systems, 11(2):282–305, April 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abouzeid:2011:COS

[ACF^{+}11] Fady Abouzeid, Sylvain Clerc, Fabian Firmin, Marc Renaudin, Tiempo Sas, and Gilles Sicard. 40nm CMOS 0.35Voptimized standard cell libraries for ultra-low power applications. ACM Transactions on Design Automation of Electronic Systems, 16(3): 35:1–35:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Aksoy:2012:OAM

[ACFM12] Levent Aksoy, Eduardo Costa, Paulo Flores, and Jose Monteiro. Optimization algorithms for the multiplierless realization of linear transforms. ACM Transactions on Design Automation of Electronic Systems, 17(1):3:1–3:??, January 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Agarwal:2013:SDS

[ACT13] Amit Agarwal, Jason Cong, and Brian Tagiku. The survivability of design-specific spare placement in FPGA architectures with high defect rates. ACM Transactions on Design Automation of Electronic Systems, 18(2):33:1–33:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Afacan:2019:CRC

[ADB^{+}19] Engín Afacan, Günhan Dündar, Faík Baskaya, Alí Emre Pusane, and Mustafa Berke Yelten. On chip reconfigurable CMOS analog circuit design and automation against aging phenomena: Sense and react. ACM Transactions on Design Automation of Electronic Systems, 24(4):44:1– 44:22, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3325069.

Al-Dujaily:2013:DPB

[ADDM^{+}13] Ra’ed Al-Dujaily, Nizar Dahir, Terrence Mak, Fei Xia, and Alex Yakovlev. Dynamic programming-based runtime thermal management (DPRTM): an online thermal control strategy for 3D-noc systems. ACM Transactions on Design Automation of Electronic Systems, 19(1):2:1–2:??, December 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Alaghi:2022:ISI

[ADGSM22] Armin Alaghi, Eva Darulova, Andreas Gerstlauer, and Phillip Stanley-Marbell. Introduction to the special issue on approximate systems. ACM Transactions on Design Automation of Electronic Systems, 27(2):10:1– 10:2, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3488726.

Abouelella:2013:HEI

[ADM^{+}13] Fatma Abouelella, Tom Davidson, Wim Meeus, Karel Bruneel, and Dirk Stroobandt. How to efficiently implement dynamic circuit specialization systems. ACM Transactions on Design Automation of Electronic Systems, 18(3):35:1– 35:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Atienza:2007:HSE

[ADP^{+}07] David Atienza, Pablo G. Del Valle, Giacomo Paci, Francesco Poletti, Luca Benini, Giovanni De Micheli, Jose M. Mendias, and Roman Hermida. HW-SW emulation framework for temperature-aware design in MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 12(3): 26:1–26:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Avnit:2009:PCC

[ADS^{+}09] K. Avnit, V. D’silva, A. Sowmya, S. Ramesh, and S. Parameswaran. Provably correct on-chip communication: a formal approach to automatic protocol converter synthesis. ACM Transactions on Design Automation of Electronic Systems, 14(2): 19:1–19:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Aksoy:2014:MDF

[AFM14] Levent Aksoy, Paulo Flores, and Jose Monteiro. Multiplierless design of folded DSP blocks. ACM Transactions on Design Automation of Electronic Systems, 20(1):14:1– 14:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abel:2022:FSS

[AG22] Inga Abel and Helmut Graeb.

FUBOCO: Structure synthesis of basic op-amps by FUnctional BlOck COmposition. ACM Transactions on Design Automation of Electronic Systems, 27(6):63:1–63:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522738.

Ashar:2001:UCD

[AGM01] Pranav Ashar, Aarti Gupta, and Sharad Malik. Using complete-1-distinguishability for FSM equivalence checking. ACM Transactions on Design Automation of Electronic Systems, 6(4):569–590, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abbasian:2008:WBD

[AHAKP08] A. Abbasian, S. Hatami, A. Afzali-Kusha, and M. Pedram. Wavelet-based dynamic power management for nonstationary service requests. ACM Transactions on Design Automation of Electronic Systems, 13(1):13:1–13:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ahn:2008:SSC

[AHL^{+}08] Yongjin Ahn, Keesung Han, Ganghee Lee, Hyunjik Song, Junhee Yoo, Kiyoung Choi, and Xingguang Feng. SoC-DAL: System-on-chip design AcceLerator. ACM Transactions on Design Automation of Electronic Systems, 13(1):17:1– 17:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Arka:2021:HHM

[AJK^{+}21] Aqeeb Iqbal Arka, Biresh Kumar Joardar, Ryan Gary Kim, Dae Hyun Kim, Janardhan Rao Doppa, and Partha Pratim Pande. HeM3D: Heterogeneous manycore architecture based on monolithic 3D vertical integration. ACM Transactions on Design Automation of Electronic Systems, 26(2):16:1– 16:21, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3424239.

Abousamra:2013:OCE

[AJM13] Ahmed Abousamra, Alex K. Jones, and Rami Melhem. Ordering circuit establishment in multiplane NoCs. ACM Transactions on Design Automation of Electronic Systems, 18(4): 49:1–49:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abolmaali:2018:EFP

[AKAKP18] Sheis Abolmaali, Mehdi Kamal, Ali Afzali-Kusha, and Massoud Pedram. An efficient false path-aware heuristic critical path selection method with high coverage of the process variation space. ACM Transactions on Design Automation of Electronic Systems, 23(3): 32:1–32:??, April 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Angizi:2022:MRN

[AKM^{+}22] Shaahin Angizi, Navid Khoshavi, Andrew Marshall, Peter Dowben, and Deliang Fan. MeF-RAM: a new non-volatile cache memory based on magnetoelectric FET. ACM Transactions on Design Automation of Electronic Systems, 27(2):18:1– 18:18, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3484222.

Asgarieh:2019:SHA

[AL19] Yashar Asgarieh and Bill Lin. Smart-hop arbitration request propagation: Avoiding quadratic arbitration complexity and false negatives in SMART NoCs. ACM Transactions on Design Automation of Electronic Systems, 24(6):64:1– 64:25, November 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3356235.

Alizadeh:2012:FVD

[Ali12] Bijan Alizadeh. Formal verification and debugging of precise interrupts on high performance microprocessors. ACM Transactions on Design Automation of Electronic Systems, 17(4): 37:1–37:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Albalawi:2017:TFP

[ALL17] Hassan Albalawi, Yuanning Li, and Xin Li. Training fixedpoint classifiers for on-chip lowpower implementation. ACM Transactions on Design Automation of Electronic Systems, 22(4):69:1–69:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Agnesina:2020:IFB

[ALLE20] Anthony Agnesina, Sung Kyu Lim, Etienne Lepercq, and Jose Escobedo Del Cid. Improving FPGA-based logic emulation systems through machine learning. ACM Transactions on Design Automation of Electronic Systems, 25(5):46:1– 46:20, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3399595.

Araujo:1998:CGF

[AM98] Guido Araujo and Sharad Malik. Code generation for fixed-point DSPs. ACM Transactions on Design Automation of Electronic Systems, 3(2):136–161, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p136-araujo/p136-
araujo.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-2/p136-araujo/

.

Adya:2005:CTM

[AM05] S. N. Adya and I. L. Markov.

Combinatorial techniques for mixed-size placement. ACM Transactions on Design Automation of Electronic Systems, 10(1):58–90, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ahmed:2010:CBP

[AM10] Waseem Ahmed and Douglas Myers. Concept-based partitioning for large multidomain multifunctional embedded systems. ACM Transactions on Design Automation of Electronic Systems, 15(3): 22:1–22:??, May 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Atienza:2006:SDM

[AMM^{+}06] David Atienza, Jose M. Mendias, Stylianos Mamagkakis, Dimitrios Soudris, and Francky Catthoor. Systematic dynamic memory management design methodology for reduced memory footprint. ACM Transactions on Design Automation of Electronic Systems, 11(2): 465–489, April 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abuowaimer:2018:GRD

[AMM^{+}18] Ziad Abuowaimer, Dani Maarouf, Timothy Martin, Jeremy Foxcroft, Gary Gréwal, Shawki Areibi, and Anthony Vannelli. GPlace3.0: Routability-driven analytic placer for UltraScale FPGA architectures. ACM Transactions on Design Automation of Electronic Systems, 23(5):66:1–66:??, October 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Arato:2005:AAH

[AMO05] Péter Arató, Zoltán Ádám Mann, and András Orbán. Algorithmic aspects of hardware/ software partitioning. ACM Transactions on Design Automation of Electronic Systems, 10(1):136–156, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Aditya:2000:CSM

[AMR00] Shail Aditya, Scott A. Mahlke, and B. Ramakrishna Rau. Code size minimization and retargetable assembly for custom EPIC and VLIW instruction formats. ACM Transactions on Design Automation of Electronic Systems, 5(4):752–773, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-4/p752-aditya/p752-
aditya.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-4/p752-aditya/

.

Anonymous:2013:CNE

[Ano13] Anonymous. Call for nominations for Editor-in-Chief. ACM Transactions on Design Automation of Electronic Systems, 18(4):44:1–44:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ayoub:2013:CCM

[ANR13] Raid Ayoub, Rajib Nath, and Tajana Simunic Rosing. CoMETC: Coordinated management of energy/thermal/ cooling in servers. ACM Transactions on Design Automation of Electronic Systems, 19(1): 1:1–1:??, December 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Alam:2020:SSB

[ANS^{+}20] Md Mahbub Alam, Adib Nahiyan, Mehdi Sadi, Domenic Forte, and Mark Tehranipoor. Soft-HaT: Softwarebased silicon reprogramming for hardware Trojan implementation. ACM Transactions on Design Automation of Electronic Systems, 25(4): 35:1–35:22, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3396521.

Araujo:2002:GAR

[AOC02] Guido Araujo, Guilherme Ottoni, and Marcelo Cintra. Global array reference allocation. ACM Transactions on Design Automation of Electronic Systems, 7(2):336–357, April 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

AlKhatib:2008:MSC

[APB^{+}08] Iyad Al Khatib, Francesco Poletti, Davide Bertozzi, Luca Benini, Mohamed Bechara, Hasan Khalifeh, Axel Jantsch, and Rustam Nabiev. A multiprocessor system-on-chip for real-time biomedical monitoring and analysis: ECG prototype architectural design space exploration. ACM Transactions on Design Automation of Electronic Systems, 13(2): 31:1–31:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ain:2011:CPV

[APD^{+}11] Antara Ain, Debjit Pal, Pallab Dasgupta, Siddhartha Mukhopadhyay, Rajdeep Mukhopadhyay, and John Gough. Chassis: a platform for verifying PMU integration using autogenerated behavioral models. ACM Transactions on Design Automation of Electronic Systems, 16(3):33:1–33:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Abella:2017:MBW

[APDC17] Jaume Abella, Maria Padilla, Joan Del Castillo, and Francisco J. Cazorla. Measurementbased worst-case execution time estimation using the coefficient of variation. ACM Transactions on Design Automation of Electronic Systems, 22(4):72:1–72:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ardalani:2024:DCS

[APG24] Newsha Ardalani, Saptadeep Pal, and Puneet Gupta. Deep-Flow: a cross-stack pathfinding framework for distributed AI systems. ACM Transactions on Design Automation of Electronic Systems, 29(2):30:1– 30:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3635867.

Arcaro:2018:RTG

[APS18] Luís Fernando Arcaro, Karila Palma Silva, and Rômulo Silva De Oliveira. On the reliability and tightness of GP and exponential models for probabilistic WCET estimation. ACM Transactions on Design Automation of Electronic Systems, 23(3):39:1–39:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ayala-Rincon:2006:PTS

[ARLJH06] M. Ayala-Rincón, C. H. Llanos, R. P. Jacobi, and R. W. Hartenstein. Prototyping time- and space-efficient computations of algebraic operations over dynamically reconfigurable systems modeled by rewriting-logic. ACM Transactions on Design Automation of Electronic Systems, 11(2): 251–281, April 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ali:2017:RCD

[ASAP17] Hazem Ismail Ali, Sander Stuijk, Benny Akesson, and Luís Miguel Pinho. Reducing the complexity of dataflow graphs using slack-based merging. ACM Transactions on Design Automation of Electronic Systems, 22(2):24:1– 24:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Aseeri:2023:PTA

[Ase23] Ahmad O. Aseeri. A problemtailored adversarial deep neural network-based attack model for feed-forward physical unclonable functions. ACM Transactions on Design Automation of Electronic Systems, 28(4): 59:1–59:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3557742.

Aghaeekiasaraee:2023:CFR

[ATF^{+}23] Erfan Aghaeekiasaraee, Aysa Fakheri Tabrizi, Tiago Augusto Fontana, Renan Netto, Sheiny Fabre Almeida, Upma Gandhi, José Luís Güntzel, David Westwick, and Laleh Behjat. CRP2.0: a fast and robust cooperation between routing and placement in advanced technology nodes. ACM Transactions on Design Automation of Electronic Systems, 28(5): 79:1–79:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590962.

Amir:2019:SPC

[AVG19] Maral Amir, Frank Vahid, and Tony Givargis. Switching predictive control using reconfigurable state-based model. ACM Transactions on Design Automation of Electronic Systems, 24(1):2:1–2:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Al-Yamani:2005:TCE

[AYM05] Ahmad A. Al-Yamani and Edward J. McCluskey. Test chip experimental results on highlevel structural test. ACM Transactions on Design Automation of Electronic Systems, 10(4):690–701, October 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Alasad:2020:SLO

[AYS20] Qutaiba Alasad, Jiann-Shuin Yuan, and Pramod Subramanyan. Strong logic obfuscation with low overhead against IC reverse engineering attacks. ACM Transactions on Design Automation of Electronic Systems, 25(4): 34:1–34:31, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3398012.

Boyer:2001:ODS

[BASB01] François R. Boyer, El Mostapha Aboulhamid, Yvon Savaria, and Michel Boyer. Optimal design of synchronous circuits using software pipelining techniques. ACM Transactions on Design Automation of Electronic Systems, 6(4): 516–532, October 2001. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Behera:2017:TTS

[BB17] Lalatendu Behera and Purandar Bhaduri. Time-triggered scheduling of mixed-criticality systems. ACM Transactions on Design Automation of Electronic Systems, 22(4):74:1– 74:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bogliolo:2000:RBR

[BBD00] Alessandro Bogliolo, Luca Benini, and Giovanni De Micheli. Regression-based RTL power modeling. ACM Transactions on Design Automation of Electronic Systems, 5(3):337–372, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p337-bogliolo/p337-
bogliolo.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p337-bogliolo/

.

Beznia:2015:TAR

[BBEM15] Kamel Beznia, Ahcene Bounceur, Reinhardt Euler, and Salvador Mir. A tool for analog/RF BIST evaluation using statistical models of circuit parameters. ACM Transactions on Design Automation of Electronic Systems, 20(2): 31:1–31:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bhattacharya:2005:OWP

[BC05] Soumendu Bhattacharya and Abhijit Chatterjee. Optimized wafer-probe and assembled package test design for analog circuits. ACM Transactions on Design Automation of Electronic Systems, 10(2): 303–329, April 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bahar:2008:IJA

[BC08] R. Iris Bahar and Krishnendu Chakrabarty. Introduction to joint ACM JETC/TODAES special issue on new, emerging, and specialized technologies. ACM Transactions on Design Automation of Electronic Systems, 13(2):36:1–36:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bernasconi:2011:DRB

[BC11] Anna Bernasconi and Valentina Ciriani. Dimension-reducible Boolean functions based on affine spaces. ACM Transactions on Design Automation of Electronic Systems, 16(2): 13:1–13:??, March 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bernasconi:2016:IRZ

[BC16] Anna Bernasconi and Valentina Ciriani. Index-resilient zerosuppressed BDDs: Definition and operations. ACM Transactions on Design Automation of Electronic Systems, 21(4):72:1– 72:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bernasconi:2008:OKS

[BCC08] Anna Bernasconi, Valentina Ciriani, and Roberto Cordone. The optimization of kEP-SOPs: Computational complexity, approximability and experiments. ACM Transactions on Design Automation of Electronic Systems, 13(2): 35:1–35:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Baldassin:2008:OSB

[BCR^{+}08] Alexandro Baldassin, Paulo Centoducatte, Sandro Rigo, Daniel Casarotto, Luiz C. V. Santos, Max Schultz, and Olinto Furtado. An opensource binary utility generator. ACM Transactions on Design Automation of Electronic Systems, 13(2):27:1–27:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Benini:1997:SBM

[BD97] Luca Benini and Giovanni De Micheli. A survey of Boolean matching techniques for library binding. ACM Transactions on Design Automation of Electronic Systems, 2(3):193–226, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1997-2-3/p193-benini/p193-
                                        benini.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1997-2-3/p193-benini/

.

Benini:2000:SLPb

[BD00] L. Benini and G. De Micheli. Synthesis of low-power selectivelyclocked systems from highlevel specification. ACM Transactions on Design Automation of Electronic Systems, 5(3):311–321, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p311-benini/p311-
                                        benini.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/2000-5-3/p311-benini/

.

Banerjee:2005:OFT

[BD05] Ansuman Banerjee and Pallab Dasgupta. The open family of temporal logics: Annotating temporal operators with input constraints. ACM Transactions on Design Automation of Electronic Systems, 10(3):492–522, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Baradaran:2008:CAM

[BD08] Nastaran Baradaran and Pedro C. Diniz. A compiler approach to managing storage and memory bandwidth in configurable architectures. ACM Transactions on Design Automation of Electronic Systems, 13(4):61:1–61:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bathen:2014:STS

[BD14] Luis Angel D. Bathen and Nikil D. Dutt. SPMCloud: Towards the single-chip embedded ScratchPad memory-based storage cloud. ACM Transactions on Design Automation of Electronic Systems, 19(3): 22:1–22:??, June 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bhattacharya:1998:ERS

[BDB98] Subhrajit Bhattacharya, Sujit Dey, and Franc Breglez. Effects of resource sharing on circuit delay: an assignment algorithm for clock period optimization. ACM Transactions on Design Automation of Electronic Systems, 3(2):285–307, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p285-bhattacharya/
p285-bhattacharya.pdf;
http://www.acm.org/pubs/
citations/journals/todaes/
1998-3-2/p285-bhattacharya/

.

Bild:2012:SNR

[BDB12] David R. Bild, Robert P. Dick, and Gregory E. Bok. Static NBTI reduction using internal node control. ACM Transactions on Design Automation of Electronic Systems, 17(4): 45:1–45:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bhowmik:2019:PAT

[BDBB19] Biswajit Bhowmik, Jatindra Kumar Deka, Santosh Biswas, and Bhargab B. Bhattacharya. Performance-aware test scheduling for diagnosing coexistent channel faults in topology-agnostic networkson-chip. ACM Transactions on Design Automation of Electronic Systems, 24(2):17:1– 17:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3291532.

Banerjee:2008:ASM

[BDC08] Ansuman Banerjee, Pallab Dasgupta, and P. P. Chakrabarti. Auxiliary state machines + context-triggered properties in verification. ACM Transactions on Design Automation of Electronic Systems, 13(4):62:1– 62:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Benini:1999:SSC

[BDM^{+}99] L. Benini, G. De Micheli, E. Macii, M. Poncino, and R. Scarsi. Symbolic synthesis of clock-gating logic for power optimization of synchronous controllers. ACM Transactions on Design Automation of Electronic Systems, 4(4):351–375, October 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-4/p351-benini/p351-
benini.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1999-4-4/p351-benini/

.

Benini:2000:SLPa

[BdM00] Luca Benini and Giovanni de Micheli. System-level power optimization: techniques and tools. ACM Transactions on Design Automation of Electronic Systems, 5(2):115–192, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-2/p115-benini/p115-
benini.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-2/p115-benini/

.

Bouakaz:2017:SPD

[BFG17a] Adnan Bouakaz, Pascal Fradet, and Alain Girault. A survey of parametric dataflow models of computation. ACM Transactions on Design Automation of Electronic Systems, 22(2): 38:1–38:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bouakaz:2017:SAD

[BFG17b] Adnan Bouakaz, Pascal Fradet, and Alain Girault. Symbolic analyses of dataflow graphs. ACM Transactions on Design Automation of Electronic Systems, 22(2):39:1–39:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bakhshalipour:2019:RWT

[BFG^{+}19] Mohammad Bakhshalipour, Aydin Faraji, Seyed Armin Vakil Ghahani, Farid Samandi, Pejman Lotfi-Kamran, and Hamid Sarbazi-Azad. Reducing writebacks through in-cache displacement. ACM Transactions on Design Automation of Electronic Systems, 24(2):16:1– 16:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3289187.

Bol:2010:NME

[BFL10] David Bol, Denis Flandre, and Jean-Didier Legat. Nanometer MOSFET effects on the minimum-energy point of sub-45nm subthreshold logic— mitigation at technology and circuit levels. ACM Transactions on Design Automation of Electronic Systems, 16(1):2:1– 2:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bombieri:2008:ROT

[BFP08] Nicola Bombieri, Franco Fummi, and Graziano Pravadelli. Reuse and optimization of testbenches and properties in a TLM-to-RTL design flow. ACM Transactions on Design Automation of Electronic Systems, 13(3):47:1–47:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bombieri:2015:MRR

[BFV15] Nicola Bombieri, Franco Fummi, and Sara Vinco. A methodology to recover RTL IP functionality for automatic generation of SW applications. ACM Transactions on Design Automation of Electronic Systems, 20(3):36:1–36:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bakshi:2001:PCH

[BG01] Smita Bakshi and Daniel D.

Gajski. Performance-constrained hierarchical pipelining for behaviors, loops, and operations. ACM Transactions on Design Automation of Electronic Systems, 6(1):1– 25, April 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/citations/journals/
todaes/2001-6-1/p1-bakshi/

.

Bunker:2004:FHS

[BGM04] Annette Bunker, Ganesh Gopalakrishnan, and Sally A. Mckee. Formal hardware specification languages for protocol compliance verification. ACM Transactions on Design Automation of Electronic Systems, 9(1):1– 32, January 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bouchebaba:2007:MMO

[BGN^{+}07] Youcef Bouchebaba, Bruno Girodias, Gabriela Nicolescu, El Mostapha Aboulhamid, Bruno Lavigueur, and Pierre Paulin. MPSoC memory optimization using program transformation. ACM Transactions on Design Automation of Electronic Systems, 12(4):43:1– 43:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Blanton:2003:PIP

[BH03] R. D. (Shawn) Blanton and John P. Hayes. On the properties of the input pattern fault model. ACM Transactions on Design Automation of Electronic Systems, 8(1):108– 124, January 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bonny:2010:HBC

[BH10] Talal Bonny and Jörg Henkel.

Huffman-based code compression techniques for embedded processors. ACM Transactions on Design Automation of Electronic Systems, 15(4):31:1– 31:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Baker:2022:CMA

[BH22] Timothy J. Baker and John P.

Hayes. CeMux: Maximizing the accuracy of stochastic mux adders and an application to filter design. ACM Transactions on Design Automation of Electronic Systems, 27(3): 27:1–27:26, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3491213.

Brunner:2022:THR

[BHBS22] Michaela Brunner, Alexander Hepp, Johanna Baehr, and Georg Sigl. Toward a human-readable state machine extraction. ACM Transactions on Design Automation of Electronic Systems, 27(6): 58:1–58:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3513086.

Bertels:2009:EMM

[BHDS09] Peter Bertels, Wim Heirman, Erik D’Hollander, and Dirk Stroobandt. Efficient memory management for hardware accelerated Java Virtual Machines. ACM Transactions on Design Automation of Electronic Systems, 14(4):48:1– 48:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Backer:2017:SFT

[BHK17] Jerry Backer, David Hely, and Ramesh Karri. Secure and flexible trace-based debugging of systems-on-chip. ACM Transactions on Design Automation of Electronic Systems, 22(2): 31:1–31:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Burcea:2019:MIR

[BHLG19] Florin Burcea, Andreas Herrmann, Bing Li, and Helmut Graeb. MEMS-IC robustness optimization considering electrical and mechanical design and process parameters. ACM Transactions on Design Automation of Electronic Systems, 24(4):43:1– 43:24, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3325068.

Bruneel:2011:DDF

[BHS11] Karel Bruneel, Wim Heirman, and Dirk Stroobandt. Dynamic data folding with parameterizable FPGA configurations. ACM Transactions on Design Automation of Electronic Systems, 16(4):43:1–43:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Backasch:2013:RVM

[BHW^{+}13] Rico Backasch, Christian Hochberger, Alexander Weiss, Martin Leucker, and Richard Lasslop. Runtime verification for multicore SoC with highquality trace data. ACM Transactions on Design Automation of Electronic Systems, 18(2): 18:1–18:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bian:2024:RAS

[BHY^{+}24] Jingchang Bian, Zhengfeng Huang, Peng Ye, Zhao Yang, and Huaguo Liang. A reliability-aware splitting dutycycle physical unclonable function based on trade-off process, voltage, and temperature variations. ACM Transactions on Design Automation of Electronic Systems, 29(1):5:1– 5:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3594667.

Bahar:2015:ISI

[BJX15] R. Iris Bahar, Alex K. Jones, and Yuan Xie. Introduction to the special issue on reliable, resilient, and robust design of circuits and systems. ACM Transactions on Design Automation of Electronic Systems, 20(4): 59:1–59:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Balakrishnan:2000:AFS

[BK00] M. Balakrishnan and Heman Khanna. Allocation of FIFO structures in RTL data paths. ACM Transactions on Design Automation of Electronic Systems, 5(3):294–310, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p294-balakrishnan/
                                        p294-balakrishnan.pdf;
                                        http://www.acm.org/pubs/
                                        citations/journals/todaes/
                                        2000-5-3/p294-balakrishnan/

.

Blanc:2010:RAS

[BK10] Nicolas Blanc and Daniel Kroening. Race analysis for SystemC using model checking. ACM Transactions on Design Automation of Electronic Systems, 15(3):21:1–21:??, May 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Baranowski:2015:RSN

[BKW15] Rafal Baranowski, Michael A. Kochte, and Hans-Joachim Wunderlich. Reconfigurable scan networks: Modeling, verification, and optimal pattern generation. ACM Transactions on Design Automation of Electronic Systems, 20(2):30:1– 30:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bahar:2000:POT

[BLM00] R. Iris Bahar, Ernest T.

Lampe, and Enrico Macii. Power optimization of technologydependent circuits based on symbolic computation of logic implications. ACM Transactions on Design Automation of Electronic Systems, 5(3):267– 293, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
2000-5-3/p267-bahar/p267-
bahar.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p267-bahar/

.

Baek:2014:DHD

[BLNK14] Seungcheol Baek, Hyung Gyu Lee, Chrysostomos Nicopoulos, and Jongman Kim. Designing hybrid DRAM/PCM main memory systems utilizing dual-phase compression. ACM Transactions on Design Automation of Electronic Systems, 20(1):11:1–11:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bhanja:2006:SFG

[BLR06] Sanjukta Bhanja, Karthikeyan Lingasubramanian, and N. Ranganathan. A stimulus-free graphical probabilistic switching model for sequential circuits using dynamic Bayesian networks. ACM Transactions on Design Automation of Electronic Systems, 11(3):773–796, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bonna:2019:MSD

[BLUS19] Ricardo Bonna, Denis S. Loubach, George Ungureanu, and Ingo Sander. Modeling and simulation of dynamic applications using scenario-aware dataflow. ACM Transactions on Design Automation of Electronic Systems, 24(5):58:1– 58:29, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3342997.

Bondade:2011:HSC

[BM11] Rajdeep Bondade and Dongsheng Ma. Hardware-software codesign of an embedded multiple-supply power management unit for multicore SoCs using an adaptive global/local power allocation and processing scheme. ACM Transactions on Design Automation of Electronic Systems, 16(3): 31:1–31:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Basten:2017:SSI

[BMdG17] Twan Basten, Orlando Moreira, and Robert de Groote. Special section: Integrating dataflow, embedded computing and architecture. ACM Transactions on Design Automation of Electronic Systems, 22(2): 35:1–35:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bogdan:2013:DPM

[BMJ13] Paul Bogdan, Radu Marculescu, and Siddharth Jain. Dynamic power management for multidomain system-on-chip platforms: an optimal control approach. ACM Transactions on Design Automation of Electronic Systems, 18(4): 46:1–46:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bommu:2000:RBF

[BOC00] Surendra Bommu, Niall O’Neill, and Maciej Ciesielski. Retimingbased factorization for sequential logic optimization. ACM Transactions on Design Automation of Electronic Systems, 5(3):373–398, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/

2000-5-3/p373-bommu/p373-
bommu.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p373-bommu/

.

Bhattacharjee:2023:SGG

[BP23] Abhiroop Bhattacharjee and Priyadarshini Panda. SwitchX: Gmin-gmax switching for energy-efficient and robust implementation of binarized neural networks on ReRAM xbars. ACM Transactions on Design Automation of Electronic Systems, 28(4):60:1–60:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3576195.

Benso:1998:ELC

[BPRR98] A. Benso, P. Prinetto, M. Rebaudengo, and M. Sonza Reorda. EXFI: a low-cost fault injection system for embedded microprocessor-based boards. ACM Transactions on Design Automation of Electronic Systems, 3(4):626–634, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p626-benso/p626-
benso.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-4/p626-benso/

.

Bonetti:2017:AID

[BPTB17] Andrea Bonetti, Nicholas Preyss, Adam Teman, and Andreas Burg. Automated integration of dual-edge clocking for low-power operation in nanometer nodes. ACM Transactions on Design Automation of Electronic Systems, 22(4): 62:1–62:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Backes:2012:SCD

[BR12] John D. Backes and Marc D.

Riedel. The synthesis of cyclic dependencies with Boolean satisfiability. ACM Transactions on Design Automation of Electronic Systems, 17(4):44:1– 44:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Boukhobza:2018:ENS

[BRCS18] Jalil Boukhobza, Stéphane Rubini, Renhai Chen, and Zili Shao. Emerging NVM: a survey on architectural integration and research challenges. ACM Transactions on Design Automation of Electronic Systems, 23(2):14:1–14:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Brzozowski:2024:CAD

[Brz24] Ireneusz Brzozowski. Comparative analysis of dynamic power consumption of parallel prefix adder. ACM Transactions on Design Automation of Electronic Systems, 29(3): 49:1–49:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3651984.

BasiriM:2014:EHB

[BS14a] Mohamed Asan Basiri M. and Noor Mahammad Sk. An efficient hardware-based higher radix floating point MAC design. ACM Transactions on Design Automation of Electronic Systems, 20(1):15:1– 15:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Boghrati:2014:IAP

[BS14b] Baktash Boghrati and Sachin S. Sapatnekar. Incremental analysis of power grids using backward random walks. ACM Transactions on Design Automation of Electronic Systems, 19(3):31:1–31:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bolchini:2014:DHE

[BS14c] Cristiana Bolchini and Chiara Sandionigi. Design of hardened embedded systems on multi-FPGA platforms. ACM Transactions on Design Automation of Electronic Systems, 20 (1):16:1–16:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Basu:2019:CBA

[BSP^{+}19] Kanad Basu, Samah Mohamed Saeed, Christian Pilato, Mohammed Ashraf, Mohammed Thari Nabeel, Krishnendu Chakrabarty, and Ramesh Karri. CAD-Base: an attack vector into the electronics supply chain. ACM Transactions on Design Automation of Electronic Systems, 24(4): 38:1–38:30, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3315574.

Bai:2022:RER

[BSP^{+}22] Yunkai Bai, Andrew Stern, Jungmin Park, Mark Tehranipoor, and Domenic Forte. RASCv2: Enabling remote access to side-channels for mission critical and IoT systems. ACM Transactions on Design Automation of Electronic Systems, 27(6):65:1–65:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3524123.

Bommana:2023:DST

[BSP^{+}23] Ashish Reddy Bommana, Susheel Ujwal Siddamshetty, Dhilleswararao Pudi, Arvind Thumatti K. R., Srinivas Boppu, M. Sabarimalai Manikandan, and Linga Reddy Cenkeramaddi. Design of synthesistime vectorized arithmetic hardware for tapered floatingpoint addition and subtraction. ACM Transactions on Design Automation of Electronic Systems, 28(3):32:1–32:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3567423.

Bai:2024:BER

[BSZ^{+}24] Chen Bai, Qi Sun, Jianwang Zhai, Yuzhe Ma, Bei Yu, and Martin D. F. Wong. BOOM-Explorer: RISC-V BOOM microarchitecture design space exploration. ACM Transactions on Design Automation of Electronic Systems, 29(1):20:1– 20:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3630013.

Bhattacharjee:2020:BCA

[BTP^{+}20] Sukanta Bhattacharjee, Jack Tang, Sudip Poddar, Mohamed Ibrahim, Ramesh Karri, and Krishnendu Chakrabarty. Bio-chemical assay locking to thwart bio-IP theft. ACM Transactions on Design Automation of Electronic Systems, 25(1):5:1–5:20, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3365579.

Blythe:2000:EOD

[BW00] Stephen A. Blythe and Robert A. Walker. Efficient optimal design space characterization methodologies. ACM Transactions on Design Automation of Electronic Systems, 5(3):322–336, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p322-blythe/p322-
blythe.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p322-blythe/

.

Biswas:2014:RTC

[BWB14] Sounil Biswas, Hongfei Wang, and R. D. (Shawn) Blanton. Reducing test cost of integrated, heterogeneous systems using pass-fail test data analysis. ACM Transactions on Design Automation of Electronic Systems, 19(2):20:1– 20:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bi:2024:AVA

[BXG^{+}24] Yongtian Bi, Qi Xu, Hao Geng, Song Chen, and Yi Kang. AD^{2}VNCS: Adversarial defense and device variationtolerance in memristive crossbarbased neuromorphic computing systems. ACM Transactions on Design Automation of Electronic Systems, 29(1):8:1– 8:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3600231.

Bu:2022:TFG

[BYT22] Tiancong Bu, Kaige Yan, and Jingweijia Tan. Towards finegrained online adaptive approximation control for dense SLAM on embedded GPUs. ACM Transactions on Design Automation of Electronic Systems, 27(2):11:1–11:19, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3486612.

Boule:2008:ABA

[BZ08] Marc Boulé and Zeljko Zilic. Automata-based assertion-checker synthesis of PSL properties. ACM Transactions on Design Automation of Electronic Systems, 13(1):4:1–4:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Bi:2017:OQE

[BZWZ17] Zhaori Bi, Dian Zhou, Sheng-Guo Wang, and Xuan Zeng. Optimization and quality estimation of circuit design via random region covering method. ACM Transactions on Design Automation of Electronic Systems, 23(1):1:1–1:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chabini:2005:SOR

[CACS05] Noureddine Chabini, El Mostapha Aboulhamid, Ismaïl Chabini, and Yvon Savaria. Scheduling and optimal register placement for synchronous circuits derived using software pipelining techniques. ACM Transactions on Design Automation of Electronic Systems, 10(2):187–204, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Charles:2019:ECR

[CAOM19] Subodha Charles, Alif Ahmed, Umit Y. Ogras, and Prabhat Mishra. Efficient cache reconfiguration using machine learning in NoC-based manycore CMPs. ACM Transactions on Design Automation of Electronic Systems, 24(6):60:1– 60:23, November 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3350422.

Chhabria:2023:EDN

[CAP^{+}23] Vidya A. Chhabria, Vipul Ahuja, Ashwath Prabhu, Nikhil Patil, Palkesh Jain, and Sachin S. Sapatnekar. Encoderdecoder networks for analyzing thermal and power delivery networks. ACM Transactions on Design Automation of Electronic Systems, 28(1):3:1– 3:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3526115.

Chopra:2018:OAC

[CARH18] Abhimanyu Chopra, Hakan Aydin, Setareh Rafatirad, and Houman Homayoun. Optimal allocation of computation and communication in an IoT network. ACM Transactions on Design Automation of Electronic Systems, 23(6):78:1– 78:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2017:AMM

[CB17] Chao Chen and Giovanni Beltrame. An adaptive Markov model for the timing analysis of probabilistic caches. ACM Transactions on Design Automation of Electronic Systems, 23(1):12:1–12:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2022:EEL

[CBC22] Zhe Chen, Hugh T. Blair, and Jason Cong. Energyefficient LSTM inference accelerator for real-time causal prediction. ACM Transactions on Design Automation of Electronic Systems, 27(5): 44:1–44:19, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3495006.

Cauley:2011:PBC

[CBHK11] Stephen Cauley, Venkataramanan Balakrishnan, Y. Charlie Hu, and Cheng-Kok Koh.

A parallel branch-and-cut approach for detailed placement. ACM Transactions on Design Automation of Electronic Systems, 16(2):18:1–18:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2010:LSC

[CBMM10] Kai-Hui Chang, Valeria Bertacco, Igor L. Markov, and Alan Mishchenko. Logic synthesis and circuit customization using extensive external don’t-cares. ACM Transactions on Design Automation of Electronic Systems, 15(3):26:1–26:??, May 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2018:DMU

[CBO^{+}18] Doohwang Chang, Ganapati Bhat, Umit Ogras, Bertan Bakkaloglu, and Sule Ozev. Detection mechanisms for unauthorized wireless transmissions. ACM Transactions on Design Automation of Electronic Systems, 23(6):70:1–70:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ciordas:2005:EBM

[CBR^{+}05] Calin Ciordas, Twan Basten, Andrei Rădulescu, Kees Goossens, and Jef Van Meerbergen. An event-based monitoring service for networks on chip. ACM Transactions on Design Automation of Electronic Systems, 10(4): 702–723, October 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Choudhury:2022:SHC

[CBR^{+}22] Dwaipayan Choudhury, Reet Barik, Aravind Sukumaran Rajam, Ananth Kalyanaraman, and Partha Pratim Pande. Software/hardware co-design of 3D NoC-based GPU architectures for accelerated graph computations. ACM Transactions on Design Automation of Electronic Systems, 27 (6):61:1–61:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3514354.

Chen:2006:SPC

[CC06] Gang Chen and Jason Cong. Simultaneous placement with clustering and duplication. ACM Transactions on Design Automation of Electronic Systems, 11(3):740–772, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chien:2009:SMV

[CCC^{+}09a] Chih-Da Chien, Cheng-An Chien, Jui-Chin Chu, Jiun-In Guo, and Ching-Hwa Cheng. A 252Kgates/4.9Kbytes SRAM/ 71mW multistandard video decoder for high definition video applications. ACM Transactions on Design Automation of Electronic Systems, 14(1): 17:1–17:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Clarke:2009:WLS

[CCC09b] Jonathan A. Clarke, George A.

Constantinides, and Peter Y. K. Cheung. Word-length selection for power minimization via nonlinear optimization. ACM Transactions on Design Automation of Electronic Systems, 14(3):39:1–39:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2015:MBW

[CCH^{+}15a] Hung-Sheng Chang, Yuan-Hao Chang, Pi-Cheng Hsiu, Tei-Wei Kuo, and Hsiang-Pang Li. Marching-based wearleveling for PCM-based storage systems. ACM Transactions on Design Automation of Electronic Systems, 20(2):25:1– 25:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2015:CDC

[CCH15b] Naiwen Chang, Eddie Cheng, and Sunyuan Hsieh. Conditional diagnosability of Cayley graphs generated by transposition trees under the PMC model. ACM Transactions on Design Automation of Electronic Systems, 20(2):20:1– 20:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2018:EFM

[CCK^{+}18] Tseng-Yi Chen, Yuan-Hao Chang, Yuan-Hung Kuan, Ming-Chang Yang, Yu-Ming Chang, and Pi-Cheng Hsiu. Enhancing flash memory reliability by jointly considering write-back pattern and block endurance. ACM Transactions on Design Automation of Electronic Systems, 23(5): 64:1–64:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Constantinides:2003:SSA

[CCL03] G. A. Constantinides, P. Y. K. Cheung, and W. Luk. Synthesis of saturation arithmetic architectures. ACM Transactions on Design Automation of Electronic Systems, 8(3):334–354, July 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cota:2004:RCN

[CCL04] Érika Cota, Luigi Carro, and Marcelo Lubaszewski. Reusing an on-chip network for the test of core-based systems. ACM Transactions on Design Automation of Electronic Systems, 9(4):471–499, October 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chatterjee:2020:MLA

[CCMC20] Urbi Chatterjee, Soumi Chatterjee, Debdeep Mukhopadhyay, and Rajat Subhra Chakraborty. Machine learning assisted PUF calibration for trustworthy proof of sensor data in IoT. ACM Transactions on Design Automation of Electronic Systems, 25(4): 32:1–32:21, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3393628.

Cabodi:1998:AVB

[CCQ98] Gianpiero Cabodi, Paolo Camurati, and Stefano Quer. Auxiliary variables for BDDbased representation and manipulation of Boolean functions. ACM Transactions on Design Automation of Electronic Systems, 3(3):309–340, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-3/p309-cabodi/p309-
cabodi.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-3/p309-cabodi/

.

Chang:2015:VPI

[CCS15] Da-Wei Chang, Hsin-Hung Chen, and Wei-Jian Su. VSSD: Performance isolation in a solid-state drive. ACM Transactions on Design Automation of Electronic Systems, 20(4): 51:1–51:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cheng:2008:FSI

[CCW08] Lei Cheng, Deming Chen, and Martin D. F. Wong. A fast simultaneous input vector generation and gate replacement algorithm for leakage power reduction. ACM Transactions on Design Automation of Electronic Systems, 13(2): 34:1–34:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2006:OSM

[CCX06] Deming Chen, Jason Cong, and Junjuan Xu. Optimal simultaneous module and multivoltage assignment for low power. ACM Transactions on Design Automation of Electronic Systems, 11(2):362–386, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chandra:2022:ISS

[CCY22] Vikas Chandra, Yiran Chen, and Sungjoo Yoo. Introduction to the special section on energy-efficient AI chips. ACM Transactions on Design Automation of Electronic Systems, 27(5):41:1–41:2, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3538502.

Chang:2014:BBL

[CCYC14] Da-Wei Chang, Hsin-Hung Chen, Dau-Jieu Yang, and Hsung-Pin Chang. BLAS: Block-level adaptive striping for solid-state drives. ACM Transactions on Design Automation of Electronic Systems, 19(2):21:1–21:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cong:1996:CLS

[CD96] Jason Cong and Yuzheng Ding. Combinational logic synthesis for LUT based field programmable gate arrays. ACM Transactions on Design Automation of Electronic Systems, 1(2):145–204, April 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-2/p145-cong/p145-
cong.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1996-1-2/p145-cong/

.

Chang:2009:DIE

[CD09] Li-Pin Chang and Chun-Da Du. Design and implementation of an efficient wear-leveling algorithm for solid-state-disk microcontrollers. ACM Transactions on Design Automation of Electronic Systems, 15(1):

6:1–6:??, December 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chatterjee:2011:GLS

[CDB11] Debapriya Chatterjee, Andrew Deorio, and Valeria Bertacco. Gate-level simulation with GPU computing. ACM Transactions on Design Automation of Electronic Systems, 16(3): 30:1–30:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cheatham:2006:SFT

[CEB06] Jason A. Cheatham, John M. Emmert, and Stan Baumgart. A survey of fault tolerant methodologies for FP-GAs. ACM Transactions on Design Automation of Electronic Systems, 11(2):501–533, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2016:RAR

[CET16] Liang Chen, Mojtaba Ebrahimi, and Mehdi B. Tahoori. Reliabilityaware resource allocation and binding in high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 21(2):30:1–30:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2016:DCV

[CFD^{+}16] Gong Chen, Toru Fujimura, Qing Dong, Shigetoshi Nakatake, and Bo Yang. DC characteristics and variability on 90nm CMOS transistor array-style analog layout. ACM Transactions on Design Automation of Electronic Systems, 21(3): 45:1–45:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2009:LRD

[CFHM09] Po-Yuan Chen, Chiao-Chen Fang, Tingting Hwang, and Hsi-Pin Ma. Leakage reduction, delay compensation using partition-based tunable body-biasing techniques. ACM Transactions on Design Automation of Electronic Systems, 14(4):53:1–53:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cong:2009:SRB

[CFX09] Jason Cong, Yiping Fan, and Junjuan Xu. Simultaneous resource binding and interconnection optimization based on a distributed register-file microarchitecture. ACM Transactions on Design Automation of Electronic Systems, 14(3): 35:1–35:??, May 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2023:AMC

[CGLH23] Zhisheng Chen, Wenzhong Guo, Genggeng Liu, and Xing Huang. Application mapping and control-system design for microfluidic biochips with distributed channel storage. ACM Transactions on Design Automation of Electronic Systems, 28(2):29:1–29:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3564288.

Chang:1996:OCP

[CGN96] En-Shou Chang, Daniel D. Gajski, and Sanjiv Narayan. An optimal clock period selection method based on slack minimization criteria. ACM Transactions on Design Automation of Electronic Systems, 1(3):352–370, July 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-3/p352-chang/p352-
chang.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-3/p352-chang/

.

Choudhury:2023:EPF

[CGV^{+}23] Muhtadi Choudhury, Minyan Gao, Avinash Varna, Elad Peer, and Domenic Forte. Enhanced PATRON: Fault injection and power-aware FSM encoding through linear programming. ACM Transactions on Design Automation of Electronic Systems, 28(6): 93:1–93:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611669.

Cong:1996:OWI

[CH96] Jason Cong and Lei He. Optimal wiresizing for interconnects with multiple sources. ACM Transactions on Design Automation of Electronic Systems, 1(4):478–511, October 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-4/p478-cong/p478-
cong.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1996-1-4/p478-cong/

.

Cong:2000:SGD

[CH00] Jason Cong and Yean-Yow Hwang. Structural gate decomposition for depth-optimal technology mapping in LUTbased FPGA designs. ACM Transactions on Design Automation of Electronic Systems, 5(2):193–225, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-2/p193-cong/p193-
cong.pdf;   http://www.acm.

                                        org/pubs/citations/journals/
                                        todaes/2000-5-2/p193-cong/

.

Chowdhary:2002:GTM

[CH02] Amit Chowdhary and John P. Hayes. General technology mapping for field-programmable gate arrays based on lookup tables. ACM Transactions on Design Automation of Electronic Systems, 7(1):1–32, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2010:CPA

[CH10a] Naehyuck Chang and Jörg Henkel. Call for papers: ACM Transactions on Design Automation of Electronic Systems (TODAES) special section on low-power electronics and design. ACM Transactions on Design Automation of Electronic Systems, 15(2):20:1– 20:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2010:GEC

[CH10b] Naehyuck Chang and Jörg Henkel. Guest editorial: Current trends in low-power design. ACM Transactions on Design Automation of Electronic Systems, 16(1):1:1–1:??, November 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2013:DRG

[CH13] Chun-An Chen and Sun-Yuan Hsieh. t/t-diagnosability of regular graphs under the PMC model. ACM Transactions on Design Automation of Electronic Systems, 18(2):20:1– 20:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Choi:2017:WCR

[CH17] Junchul Choi and Soonhoi Ha.

Worst-case response time analysis of a synchronous dataflow graph in a multiprocessor system with real-time tasks. ACM Transactions on Design Automation of Electronic Systems, 22(2):36:1–36:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chakrabarty:2001:OTA

[Cha01] Krishnendu Chakrabarty. Optimal test access architectures for system-on-a-chip. ACM Transactions on Design Automation of Electronic Systems, 6(1):26–49, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-1/p26-chakrabarty/
p26-chakrabarty.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/2001-6-1/
p26-chakrabarty/.

Cardona:2023:AMC

[CHA^{+}23] Jordi Cardona, Carles Hernández, Jaume Abella, Enrico Mezzetti, and Francisco J. Cazorla. Accurately measuring contention in mesh NoCs in time-sensitive embedded systems. ACM Transactions on Design Automation of Electronic Systems, 28(3):43:1–43:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582006.

Chen:2015:APB

[CHBK15] Gang Chen, Kai Huang, Christian Buckl, and Alois Knoll. Applying pay-burst-only-once principle for periodic power management in hard real-time pipelined multiprocessor systems. ACM Transactions on Design Automation of Electronic Systems, 20(2):26:1– 26:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2016:IPE

[CHC^{+}16] Yu-Ming Chang, Pi-Cheng Hsiu, Yuan-Hao Chang, Chi-Hao Chen, Tei-Wei Kuo, and Cheng-Yuan Michael Wang. Improving PCM endurance with a constant-cost wear leveling design. ACM Transactions on Design Automation of Electronic Systems, 22(1): 9:1–9:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cheng:1996:GLT

[Che96] Kwang-Ting Cheng. Gatelevel test generation for sequential circuits. ACM Transactions on Design Automation of Electronic Systems, 1(4):405–442, October 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-4/p405-cheng/p405-
cheng.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-4/p405-cheng/

.

Chen:2018:ESA

[Che18] Yung-Chih Chen. Enhancements to SAT attack: Speedup and breaking cyclic logic encryption. ACM Transactions on Design Automation of Electronic Systems, 23(4):52:1– 52:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2009:SAP

[CHH09] Po-Yuan Chen, Kuan-Hsien Ho, and Tingting Hwang. Skew-aware polarity assignment in clock tree. ACM Transactions on Design Automation of Electronic Systems, 14(2): 31:1–31:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chow:1996:LPR

[CHHL96] Sue-Hong Chow, Yi-Cheng Ho, TingTing Hwang, and C. L. Liu. Low power realization of finite state machines — a decomposition approach. ACM Transactions on Design Automation of Electronic Systems, 1(3):315–340, July 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-3/p315-chow/p315-
chow.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1996-1-3/p315-chow/

.

Cheng:2023:DDG

[CHK^{+}23] Chung-Kuan Cheng, Chester Holtz, Andrew B. Kahng, Bill Lin, and Uday Mallappa. DAG-Sizer: a directed graph convolutional network approach to discrete gate sizing of VLSI graphs. ACM Transactions on Design Automation of Electronic Systems, 28(4):52:1– 52:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3577019.

Cong:2005:TMA

[CHY05] Jason Cong, Hui Huang, and Xin Yuan. Technology mapping and architecture evaluation for k/m-macrocell-based FPGAs. ACM Transactions on Design Automation of Electronic Systems, 10(1):3–23, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2001:ALP

[CIB01] Rita Yu Chen, Mary Jane Irwin, and Raminder S. Bajwa. Architecture-level power estimation and design experiments. ACM Transactions on Design Automation of Electronic Systems, 6(1):50–66, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-1/p50-chen/p50-chen.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/2001-6-1/p50-chen/.

Cheng:2015:ABW

[CIX15] Hsiang-Yun Cheng, Mary Jane Irwin, and Yuan Xie. Adaptive burst-writes (ABW): Memory requests scheduling to reduce write-induced interference. ACM Transactions on Design Automation of Electronic Systems, 21(1):7:1–7:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Choi:2019:RDR

[CJKK19] Kyu Hyun Choi, Jaeyung Jun, Minseong Kim, and Seon Wook Kim. Reducing DRAM refresh rate using retention time aware universal hashing redundancy repair. ACM Transactions on Design Automation of Electronic Systems, 24(5):53:1– 53:31, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3339851.

Chhabria:2024:MLA

[CJKS24] Vidya A. Chhabria, Wenjing Jiang, Andrew B. Kahng, and Sachin S. Sapatnekar. A machine learning approach to improving timing consistency between global route and detailed route. ACM Transactions on Design Automation of Electronic Systems, 29(1):18:1– 18:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3626959.

Cong:2011:AMP

[CJLZ11] Jason Cong, Wei Jiang, Bin Liu, and Yi Zou. Automatic memory partitioning and scheduling for throughput and power optimization. ACM Transactions on Design Automation of Electronic Systems, 16(2):15:1–15:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cheng:1996:AGF

[CK96] Kwang-Ting Cheng and A. S. Krishnakumar. Automatic generation of functional vectors using the extended finite state machine model.

ACM Transactions on Design Automation of Electronic Systems, 1(1):57–79, January 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-1/p57-cheng/p57-cheng.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1996-1-1/p57-cheng/

.

Czerwinski:2016:SAO

[CK16] Robert Czerwinski and Dariusz Kania. State assignment and optimization of ultrahigh-speed FSMs utilizing tristate buffers. ACM Transactions on Design Automation of Electronic Systems, 22(1):3:1– 3:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chakraborty:2019:ERL

[CK19] Shounak Chakraborty and Hemangee K. Kapoor. Exploring the role of large centralised caches in thermal efficient chip design. ACM Transactions on Design Automation of Electronic Systems, 24(5):52:1– 52:28, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3339850.

Chakrapani:2007:PSC

[CKAP07] Lakshmi N. Chakrapani, Pinar Korkmaz, Bilge E. S. Akgul, and Krishna V. Palem. Probabilistic system-on-a-chip architectures. ACM Transactions on Design Automation of Electronic Systems, 12(3): 29:1–29:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cong:1998:BSC

[CKKT98] Jason Cong, Andrew B. Kahng, Cheng-Kok Koh, and C.- W. Albert Tsao. Boundedskew clock and Steiner routing. ACM Transactions on Design Automation of Electronic Systems, 3(3):341–388, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-3/p341-cong/p341-
cong.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1998-3-3/p341-cong/

.

Cheng:1999:CGN

[CL99a] Wei-Kai Cheng and Youn-Long Lin. Code generation of nested loops for DSP processors with heterogeneous registers and structural pipelining. ACM Transactions on Design Automation of Electronic Systems, 4(3):231–256, July 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-3/p231-cheng/p231-

cheng.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1999-4-3/p231-cheng/

.

Choi:1999:FDA

[CL99b] Kyumyung Choi and Steven P.

Levitan. A flexible datapath allocation method for architectural synthesis. ACM Transactions on Design Automation of Electronic Systems, 4(4):376–404, October 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-4/p376-choi/p376-
choi.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1999-4-4/p376-choi/

.

Chang:2013:IPP

[CL13] Yen-Jen Chang and Hsiang-Yu Lu. Improving the performance of port range check for network packet filtering. ACM Transactions on Design Automation of Electronic Systems, 19(1): 3:1–3:??, December 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chi:2020:WLO

[CLC20] Hao Yu Chi, Chien Nan Jimmy Liu, and Hung Ming Chen. Wire load oriented analog routing with matching constraints. ACM Transactions on Design Automation of Electronic Systems, 25(6):55:1–55:26, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3403932.

Cui:2024:ERO

[CLC^{+}24] Yijun Cui, Jiang Li, Yunpeng Chen, Chenghua Wang, Chongyan Gu, Máire O’neill, and Weiqiang Liu. An efficient ring oscillator PUF using programmable delay units on FPGA. ACM Transactions on Design Automation of Electronic Systems, 29(1):2:1– 2:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3593807.

Chang:2012:CDA

[CLH12] Nai-Wen Chang, Tzu-Yin Lin, and Sun-Yuan Hsieh. Conditional diagnosability of k-ary ncubes under the PMC model. ACM Transactions on Design Automation of Electronic Systems, 17(4):46:1–46:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cai:2022:EDL

[CLL^{+}22] Han Cai, Ji Lin, Yujun Lin, Zhijian Liu, Haotian Tang, Hanrui Wang, Ligeng Zhu, and Song Han. Enable deep learning on mobile devices: Methods, systems, and applications. ACM Transactions on Design Automation of Electronic Systems, 27(3):20:1–20:50, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486618.

Cao:2006:POS

[CLLK06] Aiqun Cao, Ruibing Lu, Chen Li, and Cheng-Kok Koh. Postlayout optimization for synthesis of Domino circuits. ACM Transactions on Design Automation of Electronic Systems, 11(4):797–821, October 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cabodi:2010:SHA

[CLM^{+}10] Gianpiero Cabodi, Luciano Lavagno, Marco Murciano, Alex Kondratyev, and Yosinori Watanabe. Speeding-up heuristic allocation, scheduling and binding with SATbased abstraction/refinement techniques. ACM Transactions on Design Automation of Electronic Systems, 15(2):12:1– 12:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cong:2010:BLO

[CLMZ10] Jason Cong, Bin Liu, Rupak Majumdar, and Zhiru Zhang. Behavior-level observability analysis for operation gating in low-power behavioral synthesis. ACM Transactions on Design Automation of Electronic Systems, 16(1):

4:1–4:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Coskun:2012:ISS

[CLQ12] Ayse Kivilcim Coskun, Yung-Hsiang Lu, and Qinru Qiu. Introduction to the special section on adaptive power management for energy and temperature-aware computing systems. ACM Transactions on Design Automation of Electronic Systems, 18(1):1:1–1:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2015:MBF

[CLT^{+}15] Hai-Bao Chen, Ying-Chi Li, Sheldon X.-D. Tan, Xin Huang, Hai Wang, and Ngai Wong. H-matrix-based finite-element-based thermal analysis for 3D ICs. ACM Transactions on Design Automation of Electronic Systems, 20(4):47:1– 47:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chu:2023:ADC

[CLX^{+}23] Cheng Chu, Cheng Liu, Dawen Xu, Ying Wang, Tao Luo, Huawei Li, and Xiaowei Li. Accelerating deformable convolution networks with dynamic and irregular memory accesses. ACM Transactions on Design Automation of Electronic Systems, 28(4):67:1–67:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597431.

Cho:2009:BHR

[CLYP09] Minsik Cho, Katrina Lu, Kun Yuan, and David Z. Pan. BoxRouter 2.0: a hybrid and robust global router with layer assignment for routability. ACM Transactions on Design Automation of Electronic Systems, 14(2):32:1– 32:??, March 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chao:2008:LPG

[CM08] Wei-Chung Chao and Wai-Kei Mak. Low-power gated and buffered clock network construction. ACM Transactions on Design Automation of Electronic Systems, 13(1):20:1– 20:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Costa:2013:CDO

[CM13] José C. Costa and José C.

Monteiro. Coverage-directed observability-based validation for embedded software. ACM Transactions on Design Automation of Electronic Systems, 18(2):19:1–19:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cakir:2018:RED

[CM18] Burcin Cakir and Sharad Malik. Reverse engineering digital ICs through geometric embedding of circuit graphs. ACM Transactions on Design Automation of Electronic Systems, 23(4):50:1–50:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cakir:2019:RCH

[CM19] Burcin Cakir and Sharad Malik. Revealing cluster hierarchy in gate-level ICs using block diagrams and cluster estimates of circuit embeddings. ACM Transactions on Design Automation of Electronic Systems, 24(5):50:1– 50:19, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3329081.

Charles:2020:RNC

[CM20] Subodha Charles and Prabhat Mishra. Reconfigurable network-on-chip security architecture. ACM Transactions on Design Automation of Electronic Systems, 25(6):53:1– 53:25, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3406661.

Chang:2007:PRE

[CMB07] Kai-Hui Chang, Igor L. Markov, and Valeria Bertacco. Postplacement rewiring by exhaustive search for functional symmetries. ACM Transactions on Design Automation of Electronic Systems, 12(3):32:1– 32:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:1998:SDI

[CML98] X. T. Chen, F. J. Meyer, and F. Lombardi. Structural diagnosis of interconnects by coloring. ACM Transactions on Design Automation of Electronic Systems, 3(2):249–271, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p249-chen/p249-
chen.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1998-3-2/p249-chen/

.

Carchiolo:2000:HSS

[CMM00] Vincenza Carchiolo, Michele Malgeri, and Giuseppe Mangioni. Hardware/software synthesis of formal specifications in codesign of embedded systems. ACM Transactions on Design Automation of Electronic Systems, 5(3):399–432, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p399-carchiolo/
p399-carchiolo.pdf;   http:
//www.acm.org/pubs/citations/

                                        journals/todaes/2000-5-3/
                                        p399-carchiolo/.

Cabodi:2008:BID

[CMNQ08] Gianpiero Cabodi, Marco Murciano, Sergio Nocco, and Stefano Quer. Boosting interpolation with dynamic localized abstraction and redundancy removal. ACM Transactions on Design Automation of Electronic Systems, 13(1): 3:1–3:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Calimera:2010:NAC

[CMP10] Andrea Calimera, Enrico Macii, and Massimo Poncino. NBTI-aware clustered power gating. ACM Transactions on Design Automation of Electronic Systems, 16(1):3:1–3:??, November 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cabodi:2013:TBM

[CNQ13] Gianpiero Cabodi, Sergio Nocco, and Stefano Quer. Thread-based multi-engine model checking for multicore platforms. ACM Transactions on Design Automation of Electronic Systems, 18(3):36:1– 36:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chu:2018:ISS

[CO18] Chris Chu and Mustafa Ozdal. Introduction to the special section on advances in physical design automation. ACM Transactions on Design Automation of Electronic Systems, 23(4): 41:1–41:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Constantinides:2006:WLO

[Con06] George A. Constantinides.

Word-length optimization for differentiable nonlinear systems. ACM Transactions on Design Automation of Electronic Systems, 11(1):26– 43, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2020:SMB

[CPK20] Huili Chen, Seetal Potluri, and Farinaz Koushanfar. Security of microfluidic biochip: Practical attacks and countermeasures. ACM Transactions on Design Automation of Electronic Systems, 25(3): 27:1–27:29, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3382127.

Corno:2002:IAS

[CPR^{+}02] F. Corno, P. Prinetto, M. Rebaudengo, M. Sonza Reorda, and G. Squillero. Initializability analysis of synchronous sequential circuits. ACM Transactions on Design Automation of Electronic Systems, 7(2): 249–264, April 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chakraborty:2016:PDM

[CPS16] Prasenjit Chakraborty, Preeti Ranjan Panda, and Sandeep Sen. Partitioning and data mapping in reconfigurable cache and scratchpad memory-based architectures. ACM Transactions on Design Automation of Electronic Systems, 22(1):12:1– 12:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cho:2004:FMB

[CPW04] Jeonghun Cho, Yunheung Paek, and David Whalley. Fast memory bank assignment for fixed-point digital signal processors. ACM Transactions on Design Automation of Electronic Systems, 9(1):52– 74, January 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2014:EBT

[CPX14] Naehyuck Chang, David Z. Pan, and Yuan Xie. Editorial: ACM Transactions on Design Automation of Electronics Systems and beyond. ACM Transactions on Design Automation of Electronic Systems, 20(1): 1:1–1:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cochran:2012:TPA

[CR12] Ryan Cochran and Sherief Reda. Thermal prediction and adaptive control through workload phase detection. ACM Transactions on Design Automation of Electronic Systems, 18(1):7:1–7:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2015:DMD

[CRC15] Hu Chen, Sanghamitra Roy, and Koushik Chakraborty. DARP-MP: Dynamically adaptable resilient pipeline design in multicore processors. ACM Transactions on Design Automation of Electronic Systems, 21(1):3:1–3:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Carpent:2019:RAS

[CRT19] Xavier Carpent, Norrathep Rattanavipanon, and Gene Tsudik. Remote attestation via self-measurement. ACM Transactions on Design Automation of Electronic Systems, 24(1): 11:1–11:??, January 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2007:PLP

[CS07] Hongliang Chang and Sachin S.

Sapatnekar. Prediction of leakage power under process uncertainties. ACM Transactions on Design Automation of Electronic Systems, 12(2): 12:1–12:??, April 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chowdhury:2022:LAH

[CS22] Prattay Chowdhury and Benjamin Carrion Schafer. Leveraging automatic high-level synthesis resource sharing to maximize dynamical voltage overscaling with error control. ACM Transactions on Design Automation of Electronic Systems, 27(2):14:1–14:18, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3473909.

Cai:2007:WAD

[CSAHR07] Yuan Cai, Marcus T. Schmitz, Bashir M. Al-Hashimi, and Sudhakar M. Reddy. Workloadahead-driven online energy minimization techniques for battery-powered embedded systems with time-constraints. ACM Transactions on Design Automation of Electronic Systems, 12(1):5:1–5:??, January 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2008:TCS

[CSC08] Kuei-Chung Chang, Jih-Sheng Shen, and Tien-Fu Chen. Tailoring circuit-switched networkon-chip to application-specific system-on-chip by two optimization schemes. ACM Transactions on Design Automation of Electronic Systems, 13(1): 12:1–12:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chattopadhyay:2021:CCP

[CSC^{+}21] Saranyu Chattopadhyay, Pranesh Santikellur, Rajat Subhra Chakraborty, Jimson Mathew, and Marco Ottavi. A conditionally chaotic physically unclonable function design framework with high reliability. ACM Transactions on Design Automation of Electronic Systems, 26(6):41:1–41:24, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460004.

Cao:2005:SSL

[CSKR05] Aiqun Cao, Naran Sirisantana, Cheng-Kok Koh, and Kaushik Roy. Synthesis of skewed logic circuits. ACM Transactions on Design Automation of Electronic Systems, 10(2):205–228, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2007:NMA

[CSL^{+}07] Gang Chen, Xiaoyu Song, Feng Liu, Qingping Tan, and Fei He. A note on “A mapping algorithm for computer-assisted exploration in the design of embedded systems”. ACM Transactions on Design Automation of Electronic Systems, 12(4): 52:1–52:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). See [MBB01].

Cicek:2022:EEB

[CSO22] Nihat Mert Cicek, Xipeng Shen, and Ozcan Ozturk. Energy efficient boosting of GEMM accelerators for DNN via reuse. ACM Transactions on Design Automation of Electronic Systems, 27(5): 43:1–43:26, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3503469.

Cong:2005:LSC

[CSX^{+}05] Jason Cong, Joseph R. Shinnerl, Min Xie, Tim Kong, and Xin Yuan. Large-scale circuit placement. ACM Transactions on Design Automation of Electronic Systems, 10(2):389–430, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2013:PDS

[CT13] Yao-Lin Chang and I-Lun Tseng. A parallel dual-scanline algorithm for partitioning parameterized 45-degree polygons. ACM Transactions on Design Automation of Electronic Systems, 18(4):59:1– 59:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Calzada:2024:HIS

[CUA^{+}24] Paul E. Calzada, Md. Sami Ul Islam Sami, Kimia Zamiri Azar, Fahim Rahman, Farimah Farahmandi, and Mark Tehranipoor. Heterogeneous integration supply chain integrity through blockchain and CHSM. ACM Transactions on Design Automation of Electronic Systems, 29(1):9:1–9:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625823.

Chithira:2017:HTS

[CV17] P. R. Chithira and Vinita Vasudevan. A hierarchical technique for statistical path selection and criticality computation. ACM Transactions on Design Automation of Electronic Systems, 23(1): 9:1–9:??, October 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2019:SAT

[CVMP19] Yukai Chen, Sara Vinco, Enrico Macii, and Massimo Poncino. SystemC-AMS thermal modeling for the co-simulation of functional and extra-functional properties. ACM Transactions on Design Automation of Electronic Systems, 24(1): 4:1–4:??, January 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chu:2001:CFS

[CW01] Chris Chu and D. F. Wong. Closed form solutions to simultaneous buffer insertion/sizing and wire sizing. ACM Transactions on Design Automation of Electronic Systems, 6(3): 343–371, July 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cui:2022:ILD

[CWL^{+}22] Lanlan Cui, Fei Wu, Xiaojian Liu, Meng Zhang, Renzhi Xiao, and Changsheng Xie. Improving LDPC decoding performance for 3D TLC NAND flash by LLR optimization scheme for hard and soft decision. ACM Transactions on Design Automation of Electronic Systems, 27(1):5:1–5:20, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3473305.

Chang:1996:USM

[CWW96] Yao-Wen Chang, D. F. Wong, and C. K. Wong. Universal switch modules for FPGA design. ACM Transactions on Design Automation of Electronic Systems, 1(1):80– 101, January 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL http://www.acm.org/

pubs/articles/journals/todaes/
1996-1-1/p80-chang/p80-chang.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1996-1-1/p80-chang/

.

Chen:2016:TTS

[CXH^{+}16] Guoqing Chen, Yi Xu, Xing Hu, Xiangyang Guo, Jun Ma, Yu Hu, and Yuan Xie. TSocket: Thermal sustainable power budgeting. ACM Transactions on Design Automation of Electronic Systems, 21(2): 29:1–29:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2013:NMC

[CXK^{+}13] Xi Chen, Zheng Xu, Hyungjun Kim, Paul Gratz, Jiang Hu, Michael Kishinevsky, and Umit Ogras. In-network monitoring and control policy for DVFS of CMP networks-on-chip and last level caches. ACM Transactions on Design Automation of Electronic Systems, 18(4): 47:1–47:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2022:UTB

[CXLL22] Si Chen, Guoqi Xie, Renfa Li, and Keqin Li. Uncertainty theory based partitioning for cyber-physical systems with uncertain reliability analysis. ACM Transactions on Design Automation of Electronic Systems, 27(3):23:1–23:19, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3490177.

Choudhury:2023:AGC

[CXR^{+}23] Dwaipayan Choudhury, Lizhi Xiang, Aravind Rajam, Anantharaman Kalyanaraman, and Partha Pratim Pande. Accelerating graph computations on 3D NoC-enabled PIM architectures. ACM Transactions on Design Automation of Electronic Systems, 28(3): 30:1–30:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3564290.

Cheng:2023:MLD

[CXS^{+}23] Jiawen Cheng, Yong Xiao, Yun Shao, Guanghai Dong, Songlin Lyu, and Wenjian Yu. Machine-learning-driven architectural selection of adders and multipliers in logic synthesis. ACM Transactions on Design Automation of Electronic Systems, 28(2):20:1–20:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3560712.

Choi:2019:OFT

[CYH19] Junchul Choi, Hoeseok Yang, and Soonhoi Ha. Optimization of fault-tolerant mixedcriticality multi-core systems with enhanced WCRT analysis. ACM Transactions on Design Automation of Electronic Systems, 24(1):6:1–6:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Cui:2024:RAA

[CYLC24] Xiaole Cui, Mingqi Yin, Hanqing Liu, and Xiaoxin Cui. The resistance analysis attack and security enhancement of the IMC LUT based on the complementary resistive switch cells. ACM Transactions on Design Automation of Electronic Systems, 29(1):10:1–10:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616870.

Cheng:2014:ECD

[CYV^{+}14] An-Che Cheng, Chia-Chih (Jack)

Yen, Celina G. Val, Sam Bayless, Alan J. Hu, Iris Hui-Ru Jiang, and Jing-Yang Jou. Efficient coverage-driven stimulus generation using simultaneous SAT solving, with application to SystemVerilog. ACM Transactions on Design Automation of Electronic Systems, 20(1): 7:1–7:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chang:2023:HPM

[CYZL23] Yajing Chang, Yingjian Yan, Chunsheng Zhu, and Yanjiang Liu. A high-performance masking design approach for Saber against high-order sidechannel attack. ACM Transactions on Design Automation of Electronic Systems, 28 (6):91:1–91:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611670.

Chang:2000:TDR

[CZW00] Yao-Wen Chang, Kai Zhu, and D. F. Wong. Timingdriven routing for symmetrical array-based FPGAs. ACM Transactions on Design Automation of Electronic Systems, 5(3):433–450, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p433-chang/p433-
chang.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p433-chang/

.

Chang:2003:AFF

[CZW^{+}03] Yao-Wen Chang, Kai Zhu, Guang-Ming Wu, D. F. Wong, and C. K. Wong. Analysis of FPGA/FPIC switch modules. ACM Transactions on Design Automation of Electronic Systems, 8(1):11–37, January 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Chen:2019:OTL

[CZW19] Yung-Chih Chen, Li-Cheng Zheng, and Fu-Lian Wong. Optimization of threshold logic networks with node merging and wire replacement. ACM Transactions on Design Automation of Electronic Systems, 24(6):67:1–67:18, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3358748.

Chen:2021:RMB

[CZZYW21] Jianli Chen, Ziran Zhu, Wenxing Zhu, and Chang Yao-Wen. A robust modulus-based matrix splitting iteration method for mixed-cell-height circuit legalization. ACM Transactions on Design Automation of Electronic Systems, 26(2):15:1– 15:28, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3423326.

Dasdan:2004:EAF

[Das04] Ali Dasdan. Experimental analysis of the fastest optimum cycle ratio and mean algorithms. ACM Transactions on Design Automation of Electronic Systems, 9(4):385– 418, October 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dasdan:2009:PEA

[Das09] Ali Dasdan. Provably efficient algorithms for resolving temporal and spatial difference constraint violations. ACM Transactions on Design Automation of Electronic Systems, 14(1): 8:1–8:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Darav:2018:ELH

[DBK^{+}18] Nima Karimpour Darav, Ismail S. Bustany, Andrew Kennings, David Westwick, and Laleh Behjat. Eh?Legalizer: a high performance standardcell legalizer observing technology constraints. ACM Transactions on Design Automation of Electronic Systems, 23(4): 43:1–43:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dastidar:2007:VST

[DC07] Tathagato Rai Dastidar and P. P. Chakrabarti. A verification system for transient response of analog circuits. ACM Transactions on Design Automation of Electronic Systems, 12(3):31:1–31:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dong:2023:WCP

[DCC^{+}23] Xiao Dong, Yufei Chen, Jun Chen, Yucheng Wang, Ji Li, Tianming Ni, Zhiguo Shi, Xunzhao Yin, and Cheng Zhuo. Worst-case power integrity prediction using convolutional neural network. ACM Transactions on Design Automation of Electronic Systems, 28(4): 54:1–54:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3564932.

Das:2007:FVT

[DCK07] Dipankar Das, P. P. Chakrabarti, and Rajeev Kumar. Functional verification of task partitioning for multiprocessor embedded systems. ACM Transactions on Design Automation of Electronic Systems, 12(4):44:1– 44:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2009:SBT

[DCK09] Dipankar Das, P. P. Chakrabarti, and Rajeev Kumar. Scenariobased timing verification of multiprocessor embedded applications. ACM Transactions on Design Automation of Electronic Systems, 14(3): 37:1–37:??, May 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2010:TAM

[DCK10] Dipankar Das, P. P. Chakrabarti, and Rajeev Kumar. Thermal analysis of multiprocessor SoC applications by simulation and verification. ACM Transactions on Design Automation of Electronic Systems, 15(2):15:1– 15:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2002:CAI

[DD02] Shantanu Dutt and Wenyong Deng. Cluster-aware iterative improvement techniques for partitioning large VLSI circuits. ACM Transactions on Design Automation of Electronic Systems, 7(1): 91–121, January 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

DaRolt:2013:NDS

[DDFR13] Jean Da Rolt, Giorgio Di Natale, Marie-Lise Flottes, and Bruno Rouzeyre. A novel differential scan attack on advanced DFT structures. ACM Transactions on Design Automation of Electronic Systems, 18(4): 58:1–58:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Doboli:2004:TLL

[DDNAV04] Alex Doboli, Nagu Dhanwada, Adrian Nunez-Aldana, and Ranga Vemuri. A twolayer library-based approach to synthesis of analog systems from VHDL-AMS specifications. ACM Transactions on Design Automation of Electronic Systems, 9(2):238–271, April 2004. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic).

Delshadtehrani:2015:SMR

[DFM15] Leila Delshadtehrani, Hamed Farbeh, and Seyed Ghassem Miremadi. In-scratchpad memory replication: Protecting scratchpad memories in multicore embedded systems against soft errors. ACM Transactions on Design Automation of Electronic Systems, 20(4):61:1– 61:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dasdan:2006:HIT

[DH06] Ali Dasdan and Ivan Hom.

Handling inverted temperature dependence in static timing analysis. ACM Transactions on Design Automation of Electronic Systems, 11(2):306–324, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dubeuf:2016:EPA

[DHB16] Jeremy Dubeuf, David Hely, and Vincent Beroulle. ECDSA passive attacks, leakage sources, and common design mistakes. ACM Transactions on Design Automation of Electronic Systems, 21(2):31:1–31:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Daboul:2023:GIO

[DHNR23] Siad Daboul, Stephan Held, Bento Natura, and Daniel Rotter. Global interconnect optimization. ACM Transactions on Design Automation of Electronic Systems, 28(5): 72:1–72:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3587044.

DosSantos:2000:CMP

[DHV^{+}00] Luiz C. V. Dos Santos, M. J. M. Heijligers, C. A. J. Van Eijk, J. Van Eijndhoven, and J. A. G. Jess. A code-motion pruning technique for global scheduling. ACM Transactions on Design Automation of Electronic Systems, 5(1):1–33, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-1/p1-dos_santos/p1-
dos_santos.pdf;   http://
www.acm.org/pubs/citations/
journals/todaes/2000-5-1/
p1-dos_santos/.

Daboul:2018:AAT

[DHVW18] Siad Daboul, Stephan Held, Jens Vygen, and Sonja Wittke. An approximation algorithm for threshold voltage optimization. ACM Transactions on Design Automation of Electronic Systems, 23(6):68:1– 68:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ding:2023:TMP

[DHW^{+}23] Bo Ding, Jinglei Huang, Junpeng Wang, Qi Xu, Song Chen, and Yi Kang. Task modules partitioning, scheduling and floorplanning for partially dynamically reconfigurable systems with heterogeneous resources. ACM Transactions on Design Automation of Electronic Systems, 28(6):103:1–103:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3625295.

Ding:2023:MAP

[DHX^{+}23] Bo Ding, Jinglei Huang, Qi Xu, Junpeng Wang, Song Chen, and Yi Kang. Memoryaware partitioning, scheduling, and floorplanning for partially dynamically reconfigurable systems. ACM Transactions on Design Automation of Electronic Systems, 28(1):7:1– 7:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3534968.

Duarte:2011:HDP

[DHZ^{+}11] Filipa Duarte, Jos Hulzink, Jun Zhou, Jan Stuijt, Jos Huisken, and Harmke De Groot. A 36μW heartbeat-detection processor for a wireless sensor node. ACM Transactions on Design Automation of Electronic Systems, 16(4):51:1– 51:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Du:2023:TLR

[DHZL23] Yajuan Du, Siyi Huang, Yao Zhou, and Qiao Li. Towards LDPC read performance of 3D flash memories with layerinduced error characteristics. ACM Transactions on Design Automation of Electronic Systems, 28(3):44:1–44:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3585075.

Deb:2021:CRC

[DJP21] Dipika Deb, John Jose, and Maurizio Palesi. COPE: Reducing cache pollution and network contention by intertile coordinated prefetching in NoC-based MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 26(3):17:1–17:31, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3428149.

Das:2008:RSA

[DK08] Sabyasachi Das and Sunil P. Khatri. Resource sharing among mutually exclusive sumof-product blocks for area reduction. ACM Transactions on Design Automation of Electronic Systems, 13(3):51:1– 51:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2016:FBP

[DK16] Shirshendu Das and Hemangee K. Kapoor. A framework for block placement, migration, and fast searching in tiled-DNUCA architecture. ACM Transactions on Design Automation of Electronic Systems, 22(1):4:1–4:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dewan:2022:DAA

[DK22] Monzurul Islam Dewan and Dae Hyun Kim. Design automation algorithms for the NP-separate VLSI design methodology. ACM Transactions on Design Automation of Electronic Systems, 27(5): 53:1–53:20, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3508375.

Darav:2016:EPH

[DKT^{+}16] Nima Karimpour Darav, Andrew Kennings, Aysa Fakheri Tabrizi, David Westwick, and Laleh Behjat. Eh?Placer: a high-performance modern technology-driven placer. ACM Transactions on Design Automation of Electronic Systems, 21(3):37:1–37:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Duan:2015:AAP

[DKZ^{+}15] Qing Duan, Abhishek Koneru, Jun Zeng, Krishnendu Chakrabarty, and Gary Dispoto. Accurate analysis and prediction of enterprise service-level performance. ACM Transactions on Design Automation of Electronic Systems, 20(4):52:1– 52:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dong:2011:PCS

[DL11] Wei Dong and Peng Li. Parallel circuit simulation with adaptively controlled projective integration. ACM Transactions on Design Automation of Electronic Systems, 16(4): 44:1–44:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2017:VBP

[DLC^{+}17] Sourav Das, Dongjin Lee, Wonje Choi, Janardhan Rao Doppa, Partha Pratim Pande, and Krishnendu Chakrabarty. VFI-based power management to enhance the lifetime of highperformance 3D NoCs. ACM Transactions on Design Automation of Electronic Systems, 23(1):7:1–7:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dewan:2024:CAM

[DLK24] Monzurul Islam Dewan, Sheng-En David Lin, and Dae Hyun Kim. Construction of all multilayer monolithic RSMTs and its application to monolithic 3D IC routing. ACM Transactions on Design Automation of Electronic Systems, 29(1):17:1– 17:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3626958.

Dhiman:2010:VSE

[DMR10] Gaurav Dhiman, Giacomo Marchetti, and Tajana Rosing. vGreen: a system for energyefficient management of virtual machines. ACM Transactions on Design Automation of Electronic Systems, 16(1): 6:1–6:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dahiya:2023:MDS

[DMR23] Ayush Dahiya, Poornima Mittal, and Rajesh Rohilla. Modified decoupled sense amplifier with improved sensing speed for low-voltage differential SRAM. ACM Transactions on Design Automation of Electronic Systems, 28(6): 94:1–94:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611672.

Dworak:2012:UIC

[DNA^{+}12] Jennifer Dworak, Kundan Nepal, Nuno Alves, Yiwen Shi, Nicholas Imbriglia, and R. Iris Bahar. Using implications to choose tests through suspect fault identification. ACM Transactions on Design Automation of Electronic Systems, 18(1):14:1–14:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dey:2020:MLA

[DNT20] Sukanta Dey, Sukumar Nandi, and Gaurav Trivedi. Machine learning approach for fast electromigration aware aging prediction in incremental design of large scale on-chip power grid network. ACM Transactions on Design Automation of Electronic Systems, 25(5):42:1– 42:29, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3399677.

deOliveira:2023:MOO

[dONH23] Ricardo Gonzalez de Oliveira, Nicolas Navet, and Achim Henkel. Multi-objective optimization for safety-related available E/E architectures scoping highly automated driving vehicles. ACM Transactions on Design Automation of Electronic Systems, 28(3):

41:1–41:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3582004.

Dandalis:2002:RTP

[DP02] Andreas Dandalis and Viktor K. Prasanna. Run-time performance optimization of an FPGA-based deduction engine for SAT solvers. ACM Transactions on Design Automation of Electronic Systems, 7(4):547– 562, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dandalis:2004:ACE

[DP04] Andreas Dandalis and Viktor K. Prasanna. An adaptive cryptographic engine for Internet protocol security architectures. ACM Transactions on Design Automation of Electronic Systems, 9(3):333–353, July 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dasgupta:2002:MBP

[DPNB02] Parthasarathi Dasgupta, Peichen Pan, Subhas C. Nandy, and Bhargab B. Bhattacharya. Monotone bipartitioning problem in a planar point set with applications to VLSI. ACM Transactions on Design Automation of Electronic Systems, 7(2):231–248, April 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dasdan:1998:TDD

[DRG98] Ali Dasdan, Dinesh Ramanathan, and Rajesh K. Gupta. A timing-driven design and validation methodology for embedded real-time systems. ACM Transactions on Design Automation of Electronic Systems, 3(4):533–553, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p533-dasdan/p533-
dasdan.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-4/p533-dasdan/

.

Davoodi:2005:VSU

[DS05] Azadeh Davoodi and Ankur Srivastava. Voltage scheduling under unpredictabilities: a risk management paradigm. ACM Transactions on Design Automation of Electronic Systems, 10(2):354–368, April 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Davoodi:2006:ETG

[DS06] Azadeh Davoodi and Ankur Srivastava. Effective techniques for the generalized low-power binding problem. ACM Transactions on Design Automation of Electronic Systems, 11(1): 52–69, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Deniz:2012:VCM

[DSH12] Etem Deniz, Alper Sen, and Jim Holt. Verification and coverage of message passing multicore applications. ACM Transactions on Design Automation of Electronic Systems, 17(3): 23:1–23:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2023:CCV

[DSHD23] Sourav Das, Sayandeep Sanyal, Aritra Hazra, and Pallab Dasgupta. CoVerPlan: a comprehensive verification planning framework leveraging PSS specifications. ACM Transactions on Design Automation of Electronic Systems, 28(1):9:1– 9:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3543175.

Dasgupta:2001:SRG

[DSK01] Parthasarathi Dasgupta and Susmita Sur-Kolay. Slicible rectangular graphs and their optimal floorplans. ACM Transactions on Design Automation of Electronic Systems, 6(4):447–470, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Das:2004:MDR

[DSKB04] Sandip Das, Susmita Sur-Kolay, and Bhargab B. Bhattacharya. Manhattan-diagonal routing in channels and switchboxes. ACM Transactions on Design Automation of Electronic Systems, 9(1):75– 104, January 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Darte:2002:CEL

[DSRV02] Alain Darte, Robert Schreiber, B. Ramakrishna Rau, and Frédéric Vivien. Constructing and exploiting linear schedules with prescribed parallelism. ACM Transactions on Design Automation of Electronic Systems, 7(1):159–172, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dontharaju:2009:DAP

[DTC^{+}09] Swapna Dontharaju, Shenchih Tung, James T. Cain, Leonid Mats, Marlin H. Mickle, and Alex K. Jones. A design automation and power estimation flow for RFID systems. ACM Transactions on Design Automation of Electronic Systems, 14(1):7:1–7:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2005:E

[Dut05] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 10(1):1–2, January 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2006:E

[Dut06] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 11(1):1–2, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2007:E

[Dut07] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 12(2):9:1–9:??, April 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2008:Ea

[Dut08a] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 13(1):1:1–1:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2008:Eb

[Dut08b] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 13(2):23:1–23:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2008:E

[Dut08c] Nikil Dutt. Editorial. ACM Transactions on Design Automation of Electronic Systems, 13(3):37:1–37:??, July 2008.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dutt:2002:SBB

[DVA02] Shantanu Dutt, Vinay Verma, and Hasan Arslan. A searchbased bump-and-refit approach to incremental routing for ECO applications in FPGAs. ACM Transactions on Design Automation of Electronic Systems, 7(4):664–693, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

deAbreuMoreira:1997:ADC

[dW97] Dilvan de Abreu Moreira and Les T. Walczowski. AGENTS a distributed client-server system for leaf cell generation. ACM Transactions on Design Automation of Electronic Systems, 2(1):42–61, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-1/p42-moreira/p42-
moreira.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-1/p42-moreira/

.

Dai:2023:SAD

[DY23] Ruochen Dai and Tuba Yavuz. A symbolic approach to detecting hardware Trojans triggered by don’t care transitions. ACM Transactions on Design Automation of Electronic Systems, 28(2):28:1–28:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3558392.

Dong:2018:PAA

[DZ18] Xuan Dong and Lihong Zhang.

PV-aware analog sizing for robust analog layout retargeting with optical proximity correction. ACM Transactions on Design Automation of Electronic Systems, 23(6):71:1– 71:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Dong:2024:DAE

[DZ24] Jinxin Dong and Pingqiang Zhou. Detecting adversarial examples utilizing pixel value diversity. ACM Transactions on Design Automation of Electronic Systems, 29(3): 41:1–41:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3636460.

Duan:2015:DDO

[DZCD15] Qing Duan, Jun Zeng, Krishnendu Chakrabarty, and Gary Dispoto. Data-driven optimization of order admission policies in a digital print factory. ACM Transactions on Design Automation of Electronic Systems, 20(2):21:1–21:??, February 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Deng:2024:ERT

[DZK^{+}24] Libing Deng, Gang Zeng, Ryo Kurachi, Hiroaki Takada, Xiongren Xiao, Renfa Li, and Guoqi Xie. Enhanced realtime scheduling of AVB flows in time-sensitive networking. ACM Transactions on Design Automation of Electronic Systems, 29(2):33:1–33:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3637878.

Delledonne:2018:CDA

[DZS^{+}18] Lorenzo Delledonne, Vittorio Zaccaria, Ruggero Susella, Guido Bertoni, and Filippo Melzani. CASCA: a design automation approach for designing hardware countermeasures against side-channel attacks. ACM Transactions on Design Automation of Electronic Systems, 23(6):69:1–69:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ebrahimi-Azandaryani:2023:ACA

[EAAK^{+}23] Farhad Ebrahimi-Azandaryani, Omid Akbari, Mehdi Kamal, Ali Afzali-Kusha, and Massoud Pedram. Accuracy configurable adders with negligible delay overhead in exact operating mode. ACM Transactions on Design Automation of Electronic Systems, 28(1):13:1– 13:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3549936.

Enrici:2017:MDE

[EAP17] Andrea Enrici, Ludovic Apvrille, and Renaud Pacalet. A modeldriven engineering methodology to design parallel and distributed embedded systems. ACM Transactions on Design Automation of Electronic Systems, 22(2):34:1–34:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Engelke:2009:SSU

[EBR^{+}09] Piet Engelke, Bernd Becker, Michel Renovell, Juergen Schloeffel, Bettina Braitling, and Ilia Polian. SUPERB: Simulator Utilizing Parallel Evaluation of Resistive Bridges. ACM Transactions on Design Automation of Electronic Systems, 14(4):56:1–56:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Edwards:2003:TCC

[Edw03] Stephen A. Edwards. Tutorial: Compiling concurrent languages for sequential processors. ACM Transactions on Design Automation of Electronic Systems, 8(2):141–187, April 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Elangovan:2022:ABA

[EJR22] Reena Elangovan, Shubham Jain, and Anand Raghunathan. Ax-BxP: Approximate blocked computation for precisionreconfigurable deep neural network acceleration. ACM Transactions on Design Automation of Electronic Systems, 27(3): 28:1–28:20, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3492733.

Esbensen:1997:PDI

[EK97] Henrik Esbensen and Ernest S. Kuh. A performance-driven IC/MCM placement algorithm featuring explicit design space exploration. ACM Transactions on Design Automation of Electronic Systems, 2(1):62–80, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-1/p62-esbensen/p62-
esbensen.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-1/p62-esbensen/

.

Ewetz:2016:CRC

[EK16] Rickard Ewetz and Cheng-Kok Koh. Construction of reconfigurable clock trees for MCMM designs using mode separation and scenario compression. ACM Transactions on Design Automation of Electronic Systems, 21(4):57:1– 57:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ebrahimi:2022:PCL

[EKEK22] Zahra Ebrahimi, Dennis Klar, Mohammad Aasim Ekhtiyar, and Akash Kumar. Plasticine: a cross-layer approximation methodology for multi-kernel applications through minimally biased, high-throughput, and energy-efficient SIMD soft multiplier-divider. ACM Transactions on Design Automation of Electronic Systems, 27(2):16:1–16:33, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3486616.

Erb:2014:ELF

[EKS^{+}14] Dominik Erb, Michael A.

Kochte, Matthias Sauer, Stefan Hillebrecht, Tobias Schubert, Hans-Joachim Wunderlich, and Bernd Becker. Exact logic and fault simulation in presence of unknowns. ACM Transactions on Design Automation of Electronic Systems, 19(3): 28:1–28:??, June 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

El-Maleh:2003:TVD

[EMO03] Aiman H. El-Maleh and Yahya E. Osais. Test vector decomposition-based static compaction algorithms for combinational circuits. ACM Transactions on Design Automation of Electronic Systems, 8(4):430–459, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Esmaili:2020:EAS

[ENP20] Amirhossein Esmaili, Mahdi Nazemi, and Massoud Pedram. Energy-aware scheduling of task graphs with imprecise computations and end-to-end deadlines. ACM Transactions on Design Automation of Electronic Systems, 25(1):11:1– 11:21, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3365999.

Erol:2019:KSB

[EO19] Osman Emir Erol and Sule Ozev. Knowledge- and simulation-based synthesis of area-efficient passive loop filter incremental Zoom-ADC for built-in self-test applications. ACM Transactions on Design Automation of Electronic Systems, 24(1):3:1–3:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Emeretlis:2018:SMA

[ETAV18] Andreas Emeretlis, George Theodoridis, Panayiotis Alefragis, and Nikolaos Voros. Static mapping of applications on heterogeneous multicore platforms combining logicbased Benders decomposition with integer linear programming. ACM Transactions on Design Automation of Electronic Systems, 23(2):26:1– 26:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Elshoukry:2007:CPA

[ETR07] Mohammed Elshoukry, Mohammad Tehranipoor, and C. P. Ravikumar. A criticalpath-aware partial gating approach for test power reduction. ACM Transactions on Design Automation of Electronic Systems, 12(2):17:1– 17:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Elmandouh:2018:GFV

[EW18a] Eman M. Elmandouh and Amr G. Wassal. Guiding formal verification orchestration using machine learning methods. ACM Transactions on Design Automation of Electronic Systems, 23(5):62:1–62:??, October 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Eslami:2018:RTC

[EW18b] Fatemeh Eslami and Steven J. E. Wilton. Rapid triggering capability using an adaptive overlay during FPGA debug. ACM Transactions on Design Automation of Electronic Systems, 23(6):72:1–72:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Esper:2023:ASF

[EWT23] Khalil Esper, Stefan Wildermann, and Jürgen Teich. Automatic synthesis of FSMs for enforcing non-functional requirements on MPSoCs using multi-objective evolutionary algorithms. ACM Transactions on Design Automation of Electronic Systems, 28 (6):98:1–98:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3617832.

Elwakil:2012:DRM

[EY12] Mohamed Elwakil and Zijiang Yang. Deterministic replay for message-passing-based concurrent programs. ACM Transactions on Design Automation of Electronic Systems, 17(3): 22:1–22:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ferretti:2023:GNN

[FCZ^{+}23] Lorenzo Ferretti, Andrea Cini, Georgios Zacharopoulos, Cesare Alippi, and Laura Pozzi. Graph neural networks for high-level synthesis design space exploration. ACM Transactions on Design Automation of Electronic Systems, 28 (2):25:1–25:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570925.

Fallahzadeh:2018:TPC

[FG18] Ramin Fallahzadeh and Hassan Ghasemzadeh. Trading off power consumption and prediction performance in wearable motion sensors: an optimal and real-time approach. ACM Transactions on Design Automation of Electronic Systems, 23(5):67:1–67:??, October 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Foroozannejad:2012:PBM

[FHHG12] Mohammad H. Foroozannejad, Trevor Hodges, Matin Hashemi, and Soheil Ghiasi. Postscheduling buffer management trade-offs in streaming software synthesis. ACM Transactions on Design Automation of Electronic Systems, 17(3):27:1–27:??, June 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Feng:2022:TTO

[FHHH22] Lang Feng, Jiayi Huang, Jeff Huang, and Jiang Hu. Toward taming the overhead monster for data-flow integrity. ACM Transactions on Design Automation of Electronic Systems, 27(3):25:1–25:24, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3490176.

Feng:2021:FRT

[FHHR21] Lang Feng, Jeff Huang, Jiang Hu, and Abhijith Reddy. FastCFI: Real-time controlflow integrity using FPGA without code instrumentation. ACM Transactions on Design Automation of Electronic Systems, 26(5):39:1–39:39, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3458471.

Fang:2023:ETC

[FHL^{+}23] Chenlei Fang, Qicheng Huang, Zeye Liu, Ruizhou Ding, and Ronald D. Blanton. Efficient test chip design via smart computation. ACM Transactions on Design Automation of Electronic Systems, 28(2):22:1– 22:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3558393.

Freund:1997:CEA

[FIR^{+}97] L. Freund, M. Israel, F. Rousseau, J. M. Bergé, M. Auguin, C. Belleudy, and G. Gogniat. A codesign experiment in acoustic echo cancellation GMDF. ACM Transactions on Design Automation of Electronic Systems, 2(4):365–383, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-4/p365-freund/p365-
freund.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-4/p365-freund/

.

Feng:2023:GGA

[FLG^{+}23] Lang Feng, Wenjian Liu, Chuliang Guo, Ke Tang, Cheng Zhuo, and Zhongfeng Wang. GANDSE: Generative adversarial network-based design space exploration for neural network accelerator design. ACM Transactions on Design Automation of Electronic Systems, 28(3):35:1–35:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3570926.

Fummi:2009:CMH

[FLPP09] Franco Fummi, Mirko Loghi, Massimo Poncino, and Graziano Pravadelli. A cosimulation methodology for HW/SW validation and performance estimation. ACM Transactions on Design Automation of Electronic Systems, 14(2):23:1– 23:??, March 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fan:2007:ECD

[FLWC07] Hongbing Fan, Jiping Liu, Yu-Liang Wu, and Chak-Chung Cheung. The exact channel density and compound design for generic universal switch blocks. ACM Transactions on Design Automation of Electronic Systems, 12(2): 19:1–19:??, April 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fan:2002:RDG

[FLWW02] Hongbing Fan, Jiping Liu, Yu-Liang Wu, and C. K. Wong. Reduction design for generic universal switch blocks. ACM Transactions on Design Automation of Electronic Systems, 7(4):526–546, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ferres:2023:CFF

[FMR23] Bruno Ferres, Olivier Muller, and Frédéric Rousseau. A chisel framework for flexible design space exploration through a functional approach. ACM Transactions on Design Automation of Electronic Systems, 28(4):63:1–63:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3590769.

Flores:2001:ESM

[FNMS01] Paulo F. Flores, Horácio C. Neto, and João P. Marques-Silva. An exact solution to the minimum size test pattern problem. ACM Transactions on Design Automation of Electronic Systems, 6(4):629– 644, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Faezipour:2009:HPE

[FNP09] Miad Faezipour, Mehrdad Nourani, and Rina Panigrahy. A hardware platform for efficient worm outbreak detection. ACM Transactions on Design Automation of Electronic Systems, 14(4):49:1–49:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fummi:1997:FDT

[FRS97] F. Fummi, U. Rovati, and D. Sciuto. Functional design for testability of controldominated architectures. ACM Transactions on Design Automation of Electronic Systems, 2(2):98–122, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p98-fummi/p98-fummi.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1997-2-2/p98-fummi/

.

Forte:2013:RAA

[FS13] Domenic Forte and Ankur Srivastava. Resource-aware architectures for adaptive particle filter based visual target tracking. ACM Transactions on Design Automation of Electronic Systems, 18(2): 22:1–22:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fujita:2005:ECB

[Fuj05] Masahiro Fujita. Equivalence checking between behavioral and RTL descriptions with virtual controllers and datapaths. ACM Transactions on Design Automation of Electronic Systems, 10(4):610–626, October 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fang:2000:MFP

[FW00] Wen-Jong Fang and Allen C.- H. Wu. Multiway FPGA partitioning by fully exploiting design hierarchy. ACM Transactions on Design Automation of Electronic Systems, 5(1):34–50, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-1/p34-fang/p34-fang.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        todaes/2000-5-1/p34-fang/.

Feng:2005:UDP

[FWCL05] Tao Feng, Li-C Wang, Kwang-Ting (Tim) Cheng, and Chih-Chang (Andy) Lin. Using 2-domain partitioned OBDD data structure in an enhanced symbolic simulator. ACM Transactions on Design Automation of Electronic Systems, 10(4):627–650, October 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Firouzi:2015:AVA

[FYCT15] Farshad Firouzi, Fangming Ye, Krishnendu Chakrabarty, and Mehdi B. Tahoori. Aging- and variation-aware delay monitoring using representative critical path selection. ACM Transactions on Design Automation of Electronic Systems, 20(3): 39:1–39:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fournier:2011:PAC

[FZKS11] Laurent Fournier, Avi Ziv, Ekaterina Kutsy, and Ofer Strichman. A probabilistic analysis of coverage methods. ACM Transactions on Design Automation of Electronic Systems, 16(4):38:1–38:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Fan:2023:PCC

[FZL^{+}23] Shaoze Fan, Shun Zhang, Jianbo Liu, Ningyuan Cao, Xiaoxiao Guo, Jing Li, and Xin Zhang. Power converter circuit design automation using parallel Monte Carlo tree search.

ACM Transactions on Design Automation of Electronic Systems, 28(2):17:1–17:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549538.

Gogniat:2000:CBE

[GABP00] G. Gogniat, M. Auguin, L. Bianco, and A. Pegatoquet. A codesign back-end approach for embedded system design. ACM Transactions on Design Automation of Electronic Systems, 5(3):492–509, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p492-gogniat/p492-
gogniat.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p492-gogniat/

.

Gautier:2022:SMO

[GACK22] Quentin Gautier, Alric Althoff, Christopher L. Crutchfield, and Ryan Kastner. Sherlock: a multi-objective design space exploration framework. ACM Transactions on Design Automation of Electronic Systems, 27(4):33:1–33:20, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3511472.

Gade:2019:EEC

[GADG19] Sri Harsha Gade, M. Meraj Ahmed, Sujay Deb, and Amlan Ganguly. Energy efficient chip-to-chip wireless interconnection for heterogeneous architectures. ACM Transactions on Design Automation of Electronic Systems, 24(5):55:1– 55:27, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3340109.

Goel:2021:MNN

[GAT^{+}21] Abhinav Goel, Sara Aghajanzadeh, Caleb Tung, Shuo-Han Chen, George K. Thiruvathukal, and Yung-Hsiang Lu. Modular neural networks for low-power image classification on embedded devices. ACM Transactions on Design Automation of Electronic Systems, 26(1):1:1–1:35, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408062.

Gorjiara:2007:UFE

[GBC07] Bita Gorjiara, Nader Bagherzadeh, and Pai H. Chou. Ultra-fast and efficient algorithm for energy optimization by gradientbased stochastic voltage and task scheduling. ACM Transactions on Design Automation of Electronic Systems, 12(4):39:1– 39:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gangwar:2007:IIC

[GBK07] Anup Gangwar, M. Balakrishnan, and Anshul Kumar. Impact of intercluster communication mechanisms on ILP in clustered VLIW architectures. ACM Transactions on Design Automation of Electronic Systems, 12(1):1:1–1:??, January 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ghosh:2007:LPT

[GBR07] Swaroop Ghosh, Swarup Bhunia, and Kaushik Roy. Lowpower and testable circuit synthesis using Shannon decomposition. ACM Transactions on Design Automation of Electronic Systems, 12(4):47:1– 47:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ganley:1996:RST

[GC96] Joseph L. Ganley and James P. Cohoon. Rectilinear Steiner trees on a checkerboard. ACM Transactions on Design Automation of Electronic Systems, 1(4):512–522, October 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1996-1-4/p512-ganley/p512-
                                        ganley.pdf;   http://www.

acm.org/pubs/citations/journals/
todaes/1996-1-4/p512-ganley/

.

Gomez:2018:SCP

[GC18] Andres F. Gomez and Victor Champac. Selection of critical paths for reliable frequency scaling under BTI-aging considering workload uncertainty and process variations effects. ACM Transactions on Design Automation of Electronic Systems, 23(3):27:1–27:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gingade:2016:HPM

[GCL^{+}16] Ganesh Gingade, Wenyi Chen, Yung-Hsiang Lu, Jan Allebach, and Hernan Ildefonso Gutierrez-Vazquez. Hybrid power management for office equipment. ACM Transactions on Design Automation of Electronic Systems, 22(1):14:1– 14:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guo:2015:RDS

[GCZ^{+}15] Qi Guo, Tianshi Chen, Zhi-Hua Zhou, Olivier Temam, Ling Li, Depei Qian, and Yunji Chen. Robust design space modeling. ACM Transactions on Design Automation of Electronic Systems, 20(2):18:1–18:??, February 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Goli:2020:PAP

[GD20] Mehran Goli and Rolf Drechsler. PREASC: Automatic portion resilience evaluation for approximating SystemC-based designs using regression analysis techniques. ACM Transactions on Design Automation of Electronic Systems, 25(5):40:1– 40:28, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3388140.

Gade:2022:NHC

[GD22] Sri Harsha Gade and Sujay Deb. A novel hybrid cache coherence with global snooping for many-core architectures. ACM Transactions on Design Automation of Electronic Systems, 27(1):2:1–2:31, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3462775.

Ghosh:2021:PDP

[GDD21] Sumana Ghosh, Soumyajit Dey, and Pallab Dasgupta. Performance-driven post-processing of control loop execution schedules. ACM Transactions on Design Automation of Electronic Systems, 26(2):13:1–13:27, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3421505.

Grosse:2009:MPO

[GDF09] Philippe Grosse, Yves Durand, and Paul Feautrier. Methods for power optimization in SOC-based data flow systems. ACM Transactions on Design Automation of Electronic Systems, 14(3):38:1–38:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guan:2008:SAP

[GDG^{+}08] Nan Guan, Qingxu Deng, Zonghua Gu, Wenyao Xu, and Ge Yu. Schedulability analysis of preemptive and nonpreemptive EDF on partial runtimereconfigurable FPGAs. ACM Transactions on Design Automation of Electronic Systems, 13(4):56:1–56:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Garcia-Dopico:2011:NAV

[GDPRG11] Antonio García-Dopico, Antonio Pérez, Santiago Rodríguez, and María Isabel García. A new algorithm for VHDL parallel simulation. ACM Transactions on Design Automation of Electronic Systems, 16(3): 27:1–27:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Goncalves:2021:SAT

[GdRJM21] Stèphano M. M. Gonçalves, Leomar S. da Rosa Jr, and Felipe S. Marques. SmartDR:

Algorithms and techniques for fast detailed routing with good design rule handling. ACM Transactions on Design Automation of Electronic Systems, 26(2):9:1–9:38, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3417133.

Guo:2017:OBP

[GDTF17] Zimu Guo, Jia Di, Mark M. Tehranipoor, and Domenic Forte. Obfuscation-based protection framework against printed circuit boards unauthorized operation and reverse engineering. ACM Transactions on Design Automation of Electronic Systems, 22(3): 54:1–54:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Galanis:2007:SES

[GDTG07] Michalis D. Galanis, Gregory Dimitroulakos, Spyros Tragoudas, and Costas E. Goutis. Speedups in embedded systems with a highperformance coprocessor datapath. ACM Transactions on Design Automation of Electronic Systems, 12(3):35:1– 35:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Goren:2006:TSG

[GF06] Sezer Gören and F. Joel Ferguson. Test sequence generation for controller verification and test with high coverage. ACM Transactions on Design Automation of Electronic Systems, 11(4):916–938, October 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guan:2010:RFP

[GF10] Xuan Guan and Yunsi Fei.

Register file partitioning and recompilation for register file power reduction. ACM Transactions on Design Automation of Electronic Systems, 15(3): 24:1–24:??, May 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Geelen:2009:SLE

[GFC^{+}09] Bert Geelen, Vissarion Ferentinos, Francky Catthoor, Gauthier Lafruit, Diederik Verkest, Rudy Lauwereins, and Thanos Stouraitis. Spatial locality exploitation for runtime reordering of JPEG2000 wavelet data layouts. ACM Transactions on Design Automation of Electronic Systems, 15(1):8:1– 8:??, December 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ghasemzadeh:2016:HAE

[GFJ16] Hassan Ghasemzadeh, Ramin Fallahzadeh, and Roozbeh Jafari. A hardware-assisted energy-efficient processing model for activity recognition using wearables. ACM Transactions on Design Automation of Electronic Systems, 21(4):58:1– 58:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gasteier:1999:BBC

[GG99] Michael Gasteier and Manfred Glesner. Bus-based communication synthesis on system level. ACM Transactions on Design Automation of Electronic Systems, 4(1):1–11, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-1/p1-gasteier/p1-
gasteier.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1999-4-1/p1-gasteier/

.

Ghosh:2004:COE

[GG04] Arijit Ghosh and Tony Givargis. Cache optimization for embedded processor cores: an analytical approach. ACM Transactions on Design Automation of Electronic Systems, 9(4): 419–440, October 2004. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gong:1997:MRH

[GGB97] Jie Gong, Daniel D. Gajski, and Smita Bakshi. Model refinement for hardware-software codesign. ACM Transactions on Design Automation of Electronic Systems, 2(1):22– 41, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
1997-2-1/p22-gong/p22-gong.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1997-2-1/p22-gong/.

Glebov:2002:FNA

[GGBZ02] A. Glebov, S. Gavrilov, D. Blaauw, and V. Zolotov. False-noise analysis using logic implications. ACM Transactions on Design Automation of Electronic Systems, 7(3):474–498, July 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gupta:2004:CPC

[GGDN04] Sumit Gupta, Rajesh Kumar Gupta, Nikil D. Dutt, and Alexandru Nicolau. Coordinated parallelizing compiler optimizations and high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 9(4):441– 470, October 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gupta:2000:CIP

[GH00] Avaneendra Gupta and John P.

Hayes. CLIP: integer-programming-based optimal layout synthesis of 2D CMOS cells. ACM Transactions on Design Automation of Electronic Systems, 5(3):510–547, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p510-gupta/p510-
gupta.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p510-gupta/

.

Guthaus:2012:HPC

[GHW^{+}12] Matthew R. Guthaus, Xuchu Hu, Gustavo Wilke, Guilherme Flach, and Ricardo Reis. Highperformance clock mesh optimization. ACM Transactions on Design Automation of Electronic Systems, 17(3): 33:1–33:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gupta:2019:DAD

[GHYR19] Ankur Gupta, Juinn-Dar Huang, Shigeru Yamashita, and Sudip Roy. Design automation for dilution of a fluid using programmable microfluidic device-based biochips. ACM Transactions on Design Automation of Electronic Systems, 24(2):21:1–21:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3306492.

Givargis:2006:ZCI

[Giv06] Tony Givargis. Zero cost indexing for improved processor cache performance. ACM Transactions on Design Automation of Electronic Systems, 11(1):3–25, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gopalakrishnan:2007:OPD

[GK07] Sivaram Gopalakrishnan and Priyank Kalla. Optimization of polynomial datapaths using finite ring algebra. ACM Transactions on Design Automation of Electronic Systems, 12(4): 49:1–49:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gopalakrishnan:2009:ATB

[GK09] Sivaram Gopalakrishnan and Priyank Kalla. 2009 ACM TO-DAES best paper award: Optimization of polynomial datapaths using finite ring algebra. ACM Transactions on Design Automation of Electronic Systems, 14(4):47:1–47:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ganeshpure:2014:PDD

[GK14] Kunal Ganeshpure and Sandip Kundu. Performance-driven dynamic thermal management of MPSoC based on task rescheduling. ACM Transactions on Design Automation of Electronic Systems, 19(2): 11:1–11:??, March 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gupta:2005:RAS

[GKM05] Puneet Gupta, Andrew B. Kahng, and Stefanus Mantik. Routing-aware scan chain ordering. ACM Transactions on Design Automation of Electronic Systems, 10(3):546–560, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guo:2022:SSD

[GLD^{+}22] Wenzhong Guo, Sihuang Lian, Chen Dong, Zhenyi Chen, and Xing Huang. A survey on security of digital microfluidic biochips: Technology, attack, and defense. ACM Transactions on Design Automation of Electronic Systems, 27(4): 40:1–40:33, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3494697.

Gong:2012:FNM

[GLY^{+}12] Fang Gong, Xuexin Liu, Hao Yu, Sheldon X. D. Tan, Junyan Ren, and Lei He. A fast non-Monte-Carlo yield analysis and optimization by stochastic orthogonal polynomials. ACM Transactions on Design Automation of Electronic Systems, 17(1):10:1–10:??, January 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Goel:2003:STA

[GM03] Sandeep Kumar Goel and Erik Jan Marinissen. SOC test architecture design for efficient utilization of test bandwidth. ACM Transactions on Design Automation of Electronic Systems, 8(4):399–429, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Garg:2008:SLT

[GM08] Siddharth Garg and Diana Marculescu. System-level throughput analysis for process variation aware multiple voltage-frequency island designs. ACM Transactions on Design Automation of Electronic Systems, 13(4):59:1– 59:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gester:2013:BAD

[GMN^{+}13] Michael Gester, Dirk Müller, Tim Nieberg, Christian Panten, Christian Schulte, and Jens Vygen. BonnRoute: Algorithms and data structures for fast and good VLSI routing. ACM Transactions on Design Automation of Electronic Systems, 18(2):32:1–32:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gopalakrishnan:2023:GMW

[GMS^{+}23] Ramprasath Srinivasa Gopalakrishnan, Meghna Madhusudan, Arvind K. Sharma, Jitesh Poojary, Soner Yaldiz, Ramesh Harjani, Steven M. Burns, and Sachin S. Sapatnekar. A generalized methodology for well island generation and welltap insertion in analog/mixedsignal layouts. ACM Transactions on Design Automation of Electronic Systems, 28(5): 69:1–69:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3580477.

GuerraeSilva:2002:SMA

[GMSSS02] Luís Guerra e Silva, João Marques-Silva, L. Miguel Silveira, and Karem A. Sakallah. Satisfiability models and algorithms for circuit delay computation. ACM Transactions on Design Automation of Electronic Systems, 7(1): 137–158, January 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gnad:2021:VBC

[GNGT21] Dennis R. E. Gnad, Cong Dang Khoa Nguyen, Syed Hashim Gillani, and Mehdi B. Tahoori. Voltage-based covert channels using FPGAs. ACM Transactions on Design Automation of Electronic Systems, 26(6): 43:1–43:25, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460229.

Ge:2022:SBN

[GNQ^{+}22] Mengke Ge, Xiaobing Ni, Xu Qi, Song Chen, Jinglei Huang, Yi Kang, and Feng Wu. Synthesizing brain-network-inspired interconnections for large-scale network-on-chips. ACM Transactions on Design Automation of Electronic Systems, 27(1):9:1–9:30, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3480961.

Goodby:2002:MSP

[GOC02] Laurence Goodby, Alex Orailoğlu, and Paul M. Chau. Microarchitectural synthesis of performance-constrained, lowpower VLSI designs. ACM Transactions on Design Automation of Electronic Systems, 7(1):122–136, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gheorghita:2009:SSB

[GPH^{+}09] Stefan Valentin Gheorghita, Martin Palkovic, Juan Hamers, Arnout Vandecappelle, Stelios Mamagkakis, Twan Basten, Lieven Eeckhout, Henk Corporaal, Francky Catthoor, Frederik Vandeputte, and Koen De Bosschere. System-scenario-based design of dynamic embedded systems. ACM Transactions on Design Automation of Electronic Systems, 14(1): 3:1–3:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gulati:2009:FBH

[GPK^{+}09] Kanupriya Gulati, Suganth Paul, Sunil P. Khatri, Srinivas Patil, and Abhijit Jas. FPGAbased hardware acceleration for Boolean satisfiability. ACM Transactions on Design Automation of Electronic Systems, 14(2):33:1–33:??, March 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guo:2024:YOA

[GPS^{+}24] Nanlin Guo, Fulin Peng, Jiahe Shi, Fan Yang, Jun Tao, and Xuan Zeng. Yield optimization for analog circuits over multiple corners via Bayesian neural networks: Enhancing circuit reliability under environmental variation. ACM Transactions on Design Automation of Electronic Systems, 29(1):12:1–12:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3626321.

Gong:2019:PEH

[GQW19] Yanping Gong, Fengyu Qian, and Lei Wang. Probabilistic evaluation of hardware security vulnerabilities. ACM Transactions on Design Automation of Electronic Systems, 24(2):14:1– 14:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3290405.

Gelosh:2000:MLT

[GS00] Donald S. Gelosh and Dorothy E.

Setliff. Modeling layout tools to derive forward estimates of area and delay at the RTL level. ACM Transactions on Design Automation of Electronic Systems, 5(3):451–491, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p451-gelosh/p451-
gelosh.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p451-gelosh/

.

Gupta:2013:ECR

[GS13] Saket Gupta and Sachin S.

Sapatnekar. Employing circadian rhythms to enhance power and reliability. ACM Transactions on Design Automation of Electronic Systems, 18(3): 38:1–38:??, July 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gao:2018:ECI

[GSD^{+}18] Congming Gao, Liang Shi, Yejia Di, Qiao Li, Chun Jason Xue, Kaijie Wu, and Edwin Sha. Exploiting chip idleness for minimizing garbage collection-induced chip access conflict on SSDs. ACM Transactions on Design Automation of Electronic Systems, 23(2): 15:1–15:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guin:2016:FCS

[GSFT16] Ujjwal Guin, Qihang Shi, Domenic Forte, and Mark M. Tehranipoor. FORTIS: a comprehensive solution for establishing forward trust for protecting IPs and ICs. ACM Transactions on Design Automation of Electronic Systems, 21(4):63:1–63:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gange:2014:SOS

[GSS14] Graeme Gange, Harald Søndergaard, and Peter J. Stuckey. Synthesizing optimal switching lattices. ACM Transactions on Design Automation of Electronic Systems, 20(1):6:1–6:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gebregirogis:2021:ALF

[GT21] Anteneh Gebregirogis and Mehdi Tahoori. Approximate learning and fault-tolerant mapping for energy-efficient neuromorphic systems. ACM Transactions on Design Automation of Electronic Systems, 26(3):21:1–21:23, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436491.

Gupte:2015:FAT

[GVJ15] Adwait Gupte, Sudhanshu Vyas, and Phillip H. Jones. A fault-aware toolchain approach for FPGA fault tolerance. ACM Transactions on Design Automation of Electronic Systems, 20(2):32:1– 32:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Guthaus:2013:RAP

[GWR13] Matthew R. Guthaus, Gustavo Wilke, and Ricardo Reis. Revisiting automated physical synthesis of high-performance clock networks. ACM Transactions on Design Automation of Electronic Systems, 18(2): 31:1–31:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gately:2012:AJO

[GYT12] Matthew B. Gately, Mark B.

Yeary, and Choon Yik Tang. An algorithm for jointly optimizing quantization and multiple constant multiplication. ACM Transactions on Design Automation of Electronic Systems, 17(4):42:1–42:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Gong:2022:AMB

[GYZ^{+}22] Yifan Gong, Geng Yuan, Zheng Zhan, Wei Niu, Zhengang Li, Pu Zhao, Yuxuan Cai, Sijia Liu, Bin Ren, Xue Lin, Xulong Tang, and Yanzhi Wang. Automatic mapping of the best-suited DNN pruning schemes for real-time mobile acceleration. ACM Transactions on Design Automation of Electronic Systems, 27(5): 47:1–47:26, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3495532.

Higgins:2005:SDA

[HA05] Jason T. Higgins and Mark D. Aagaard. Simplifying the design and automating the verification of pipelines with structural hazards. ACM Transactions on Design Automation of Electronic Systems, 10(4): 651–672, October 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hoveida:2017:EMA

[HAB^{+}17] Mohaddeseh Hoveida, Fatemeh Aghaaliakbari, Ramin Bashizade, Mohammad Arjomand, and Hamid Sarbazi-Azad. Efficient mapping of applications for future chipmultiprocessors in dark silicon era. ACM Transactions on Design Automation of Electronic Systems, 22(4):70:1–70:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Heyse:2015:TTM

[HABS15] Karel Heyse, Brahim Al Farisi, Karel Bruneel, and Dirk Stroobandt. TCONMAP: Technology mapping for parameterised FPGA configurations. ACM Transactions on Design Automation of Electronic Systems, 20(4):48:1– 48:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Harris:2005:I

[Har05] Ian G. Harris. Introduction.

ACM Transactions on Design Automation of Electronic Systems, 10(4):587–588, October 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Henkel:2020:ISI

[HAW20] Jörg Henkel, Hussam Amrouch, and Marilyn Wolf. Introduction to the special issue on machine learning for CAD. ACM Transactions on Design Automation of Electronic Systems, 25(5):36:1–36:2, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410864.

Huffmire:2008:DSS

[HBC^{+}08] Ted Huffmire, Brett Brotherton, Nick Callegari, Jonathan Valamehr, Jeff White, Ryan Kastner, and Tim Sherwood. Designing secure systems on reconfigurable hardware. ACM Transactions on Design Automation of Electronic Systems, 13(3):44:1–44:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hatami:2014:MSN

[HBPW14] Nadereh Hatami, Rafal Baranowski, Paolo Prinetto, and Hans-Joachim Wunderlich. Multilevel simulation of nonfunctional properties by piecewise evaluation. ACM Transactions on Design Automation of Electronic Systems, 19(4): 37:1–37:??, August 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hankendi:2017:SCS

[HC17] Can Hankendi and Ayse Kivilcim Coskun. Scale & cap: Scaling-aware resource management for consolidated multithreaded applications. ACM Transactions on Design Automation of Electronic Systems, 22(2):30:1–30:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2018:PSC

[HC18] Sheng-Min Huang and Li-Pin Chang. Providing SLO compliance on NVMe SSDs through parallelism reservation. ACM Transactions on Design Automation of Electronic Systems, 23(3):28:1–28:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2023:RDP

[HC23] Ping-Wei Huang and Yao-Wen Chang. Routabilitydriven power/ground network optimization based on machine learning. ACM Transactions on Design Automation of Electronic Systems, 28(4): 53:1–53:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3587817.

Huang:2001:VSE

[HCC01] Shi-Yu Huang, Kwang-Ting Cheng, and Kuang-Chien Chen. Verifying sequential equivalence using ATPG techniques. ACM Transactions on Design Automation of Electronic Systems, 6(2):244–275, April 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-2/p244-huang/p244-
huang.pdf;   http://www.
acm.org/pubs/citations/journals/

todaes/2001-6-2/p244-huang/

.

Huang:2013:IBM

[HCK13] Po-Chun Huang, Yuan-Hao Chang, and Tei-Wei Kuo. An index-based management scheme with adaptive caching for huge-scale low-cost embedded flash storages. ACM Transactions on Design Automation of Electronic Systems, 18(4): 56:1–56:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2014:GCM

[HCL^{+}14] Po-Chun Huang, Yuan-Hao Chang, Kam-Yiu Lam, Jian-Tao Wang, and Chien-Chin Huang. Garbage collection for multiversion index in flashbased embedded databases. ACM Transactions on Design Automation of Electronic Systems, 19(3):25:1–25:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsiung:1998:IIC

[HCLC98] Pao-Ann Hsiung, Chung-Hwang Chen, Trong-Yen Lee, and Sao-Jie Chen. ICOS: an intelligent concurrent objectoriented synthesis methodology for multiprocessor systems. ACM Transactions on Design Automation of Electronic Systems, 3(2):109–135, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p109-hsiung/p109-
hsiung.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-2/p109-hsiung/

.

Huang:2009:OPR

[HCN09] Shih-Hsu Huang, Chia-Ming Chang, and Yow-Tyng Nieh. Opposite-phase register switching for peak current minimization. ACM Transactions on Design Automation of Electronic Systems, 14(1):14:1– 14:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Han:2011:DIT

[HCRK11] Yiding Han, Koushik Chakraborty, Sanghamitra Roy, and Vilasita Kuntamukkala. Design and implementation of a throughputoptimized GPU floorplanning algorithm. ACM Transactions on Design Automation of Electronic Systems, 16(3): 23:1–23:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hu:2001:ELA

[HCS01] Xiaobo Sharon Hu, Danny Z.

Chen, and Rajeshkumar Sambandam. Efficient listapproximation techniques for floorplan area minimization. ACM Transactions on Design Automation of Electronic Systems, 6(3):372–400, July 2001.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ho:2016:AAD

[HCW^{+}16] Ching-Hsuan Ho, Yung-Chih Chen, Chun-Yao Wang, Ching-Yi Huang, Suman Datta, and Vijaykrishnan Narayanan. Area-aware decomposition for single-electron transistor arrays. ACM Transactions on Design Automation of Electronic Systems, 21(4):70:1– 70:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2016:FTS

[HCZ^{+}16] Jinglei Huang, Song Chen, Wei Zhong, Wenchao Zhang, Shengxi Diao, and Fujiang Lin. Floorplanning and topology synthesis for applicationspecific network-on-chips with RF-interconnect. ACM Transactions on Design Automation of Electronic Systems, 21(3): 40:1–40:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Handique:2022:FLS

[HDB22] Mousum Handique, Jantindra Kumar Deka, and Santosh Biswas. Fault localization scheme for missing gate faults in reversible circuits. ACM Transactions on Design Automation of Electronic Systems, 27(4):39:1–39:29, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3503539.

Holt:2012:FLP

[HDL^{+}12] Jim Holt, Jaideep Dastidar, David Lindberg, John Pape, and Peng Yang. A full lifecycle performance verification methodology for multicore systems-on-chip. ACM Transactions on Design Automation of Electronic Systems, 17(3): 21:1–21:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

He:2020:LHD

[HDZ^{+}20] Xu He, Yu Deng, Shizhe Zhou, Rui Li, Yao Wang, and Yang Guo. Lithography hotspot detection with FFT-based feature extraction and imbalanced learning rate. ACM Transactions on Design Automation of Electronic Systems, 25(2):15:1– 15:21, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3372044.

Huang:2020:TSD

[HFMB20] Qicheng Huang, Chenlei Fang, Soumya Mittal, and R. D. (Shawn) Blanton. Towards smarter diagnosis: a learning-based diagnostic outcome previewer. ACM Transactions on Design Automation of Electronic Systems, 25(5):43:1–43:20, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3398267.

Huang:2007:ESC

[HG07] Po-Kuan Huang and Soheil Ghiasi. Efficient and scalable compiler-directed energy optimization for realtime applications. ACM Transactions on Design Automation of Electronic Systems, 12(3): 27:1–27:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hansson:2009:CTC

[HGBH09] Andreas Hansson, Kees Goossens, Marco Bekooij, and Jos Huisken. CoMPSoC: a template for composable and predictable multi-processor system on chips. ACM Transactions on Design Automation of Electronic Systems, 14(1): 2:1–2:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2016:FOF

[HGLC16] Xing Huang, Wenzhong Guo, Genggeng Liu, and Guolong Chen. FH-OAOS: a fast four-step heuristic for obstacleavoiding octilinear Steiner tree construction. ACM Transactions on Design Automation of Electronic Systems, 21(3):

48:1–48:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hong:2009:RFD

[HH09] Yu-Ru Hong and Juinn-Dar Huang. Reducing fault dictionary size for million-gate large circuits. ACM Transactions on Design Automation of Electronic Systems, 14(2): 27:1–27:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2021:MLE

[HHH^{+}21] Guyue Huang, Jingbo Hu, Yifan He, Jialong Liu, Mingyuan Ma, Zhaoyang Shen, Juejian Wu, Yuanfan Xu, Hengrui Zhang, Kai Zhong, Xuefei Ning, Yuzhe Ma, Haoyu Yang, Bei Yu, Huazhong Yang, and Yu Wang. Machine learning for electronic design automation: a survey. ACM Transactions on Design Automation of Electronic Systems, 26(5):40:1– 40:46, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3451179.

Han:2017:CAB

[HHK^{+}17] Miseon Han, Youngsun Han, Seon Wook Kim, Hokyoon Lee, and Il Park. Content-aware bit shuffling for maximizing PCM endurance. ACM Transactions on Design Automation of Electronic Systems, 22(3): 48:1–48:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ho:2014:USS

[HHL14] Yenpo Ho, Garng M. Huang, and Peng Li. Understanding SRAM stability via bifurcation analysis: Analytical models and scaling trends. ACM Transactions on Design Automation of Electronic Systems, 19(4):41:1–41:??, August 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hong:2023:PMC

[HHX^{+}23] Qinghui Hong, Richeng Huang, Pingdan Xiao, Jun Li, Jingru Sun, and Jiliang Zhang. Programmable in-memory computing circuit of Fast Hartley Transform. ACM Transactions on Design Automation of Electronic Systems, 28(6):100:1–100:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3618112.

Holst:2015:HTL

[HIW15] Stefan Holst, Michael E. Imhof, and Hans-Joachim Wunderlich. High-throughput logic timing simulation on GPGPUs. ACM Transactions on Design Automation of Electronic Systems, 20(3):37:1–37:??, June 2015.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsiao:2008:ISS

[HJ08] Michael S. Hsiao and Robert B.

Jones. Introduction to special section on high-level design, validation, and test. ACM Transactions on Design Automation of Electronic Systems, 13(1):2:1–2:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2023:HAQ

[HJY23] Shanshi Huang, Hongwu Jiang, and Shimeng Yu. Hardwareaware quantization/mapping strategies for compute-in-memory accelerators. ACM Transactions on Design Automation of Electronic Systems, 28(3):34:1–34:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3569940.

Hussain:2018:PPP

[HK18] Siam Umar Hussain and Farinaz Koushanfar. P3: Privacy preserving positioning for smart automotive systems. ACM Transactions on Design Automation of Electronic Systems, 23(6):79:1–79:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsu:2007:ESC

[HKB^{+}07] Chia-Jui Hsu, Ming-Yung Ko, Shuvra S. Bhattacharyya, Suren Ramasubbu, and José Luis Pino. Efficient simulation of critical synchronous dataflow graphs. ACM Transactions on Design Automation of Electronic Systems, 12(3):21:1– 21:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hyun:2023:ROE

[HKJ^{+}23] Daijoon Hyun, Sunwha Koh, Younggwang Jung, Taeyoung Kim, and Youngsoo Shin. Routability optimization of extreme aspect ratio design through non-uniform placement utilization and selective flip-flop stacking. ACM Transactions on Design Automation of Electronic Systems, 28(4): 50:1–50:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3573387.

Ha:2007:PHS

[HKL^{+}07] Soonhoi Ha, Sungchan Kim, Choonseung Lee, Youngmin Yi, Seongnam Kwon, and Young-Pyo Joo. PeaCE: a hardwaresoftware codesign environment for multimedia embedded systems. ACM Transactions on Design Automation of Electronic Systems, 12(3):24:1– 24:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Heo:2015:IAS

[HKL^{+}15] Ingoo Heo, Minsu Kim, Yongje Lee, Changho Choi, Jinyong Lee, Brent Byunghoon Kang, and Yunheung Paek. Implementing an application-specific instruction-set processor for system-level dynamic program analysis engines. ACM Transactions on Design Automation of Electronic Systems, 20(4): 53:1–53:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hu:2007:IHM

[HKV^{+}07] Q. Hu, P. G. Kjeldsberg, A. Vandecappelle, M. Palkovic, and F. Catthoor. Incremental hierarchical memory size estimation for steering of loop transformations. ACM Transactions on Design Automation of Electronic Systems, 12(4): 50:1–50:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsieh:2007:FDC

[HLCH07] Ang-Chih Hsieh, Tzu-Teng Lin, Tsuang-Wei Chang, and Tingting Hwang. A functionalitydirected clustering technique for low-power MTCMOS design— computation of simultaneously discharging current. ACM Transactions on Design Automation of Electronic Systems, 12(3):30:1–30:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2015:OAA

[HLG^{+}15] Xing Huang, Genggeng Liu, Wenzhong Guo, Yuzhen Niu, and Guolong Chen. Obstacleavoiding algorithm in Xarchitecture based on discrete particle swarm optimization for VLSI design. ACM Transactions on Design Automation of Electronic Systems, 20(2):24:1– 24:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hu:2008:PSF

[HLHT08] Yu Hu, Yan Lin, Lei He, and Tim Tuan. Physical synthesis for FPGA interconnect power reduction by dual-vdd budgeting and retiming. ACM Transactions on Design Automation of Electronic Systems, 13(2): 30:1–30:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hosseinabady:2007:LTA

[HLKN07] Mohammad Hosseinabady, Pejman Lotfi-Kamran, and Zainalabedin Navabi. Low test application time resource binding for behavioral synthesis. ACM Transactions on Design Automation of Electronic Systems, 12(2):16:1–16:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2024:ALV

[HLL^{+}24] Bo-Yuan Huang, Steven Lyubomirsky, Yi Li, Mike He, Gus Henry Smith, Thierry Tambe, Akash Gaonkar, Vishal Canumalla, Andrew Cheung, Gu-Yeon Wei, Aarti Gupta, Zachary Tatlock, and Sharad Malik. Application-level validation of accelerator designs using a formal software/hardware interface. ACM Transactions on Design Automation of Electronic Systems, 29(2):35:1– 35:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3639051.

Hung:2023:DDR

[HLW^{+}23] José Romero Hung, Chao Li, Taolei Wang, Jinyang Guo, Pengyu Wang, Chuanming Shao, Jing Wang, Guoyong Shi, Xiangwen Liu, and Hanqing Wu. DRAGON: Dynamic recurrent accelerator for graph online convolution. ACM Transactions on Design Automation of Electronic Systems, 28(1):1:1–1:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3524124.

Huang:2023:SDP

[HLX^{+}23] Kai Huang, Bowen Li, Dongliang Xiong, Haitian Jiang, Xiaowen Jiang, Xiaolang Yan, Luc Claesen, Dehong Liu, Junjian Chen, and Zhili Liu. Structured dynamic precision for deep neural networks quantization. ACM Transactions on Design Automation of Electronic Systems, 28(1):12:1–12:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3549535.

Han:2022:EEF

[HLZ^{+}22] Ding Han, Guohui Li, Quan Zhou, Jianjun Li, Yong Yang, and Xiaofei Hu. An efficient execution framework of twopart execution scenario analysis. ACM Transactions on Design Automation of Electronic Systems, 27(1):3:1–3:24, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3465474.

Hasteer:1998:EEC

[HMB98] G. Hasteer, A. Mathur, and P. Banerjee. Efficient equivalence checking of multi-phase designs using phase abstraction and retiming. ACM Transactions on Design Automation of Electronic Systems, 3(4):600–625, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p600-hasteer/p600-
hasteer.pdf;   http://www.
acm.org/pubs/citations/journals/

todaes/1998-3-4/p600-hasteer/

.

Healy:2011:IMF

[HMLL11] Michael B. Healy, Fayez Mohamood, Hsien-Hsin S. Lee, and Sung Kyu Lim. Integrated microarchitectural floorplanning and run-time controller for inductive noise mitigation. ACM Transactions on Design Automation of Electronic Systems, 16(4):46:1–46:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hu:2020:MLA

[HMMG^{+}20] Yong Hu, Marcel Mettler, Daniel Mueller-Gritschneder, Thomas Wild, Andreas Herkersdorf, and Ulf Schlichtmann. Machine learning approaches for efficient design space exploration of application-specific NoCs. ACM Transactions on Design Automation of Electronic Systems, 25(5):44:1– 44:27, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3403584.

Hu:2014:GLI

[HMO^{+}14] Wei Hu, Dejun Mu, Jason Oberg, Baolei Mao, Mohit Tiwari, Timothy Sherwood, and Ryan Kastner. Gate-level information flow tracking for security lattices. ACM Transactions on Design Automation of Electronic Systems, 20(1):

2:1–2:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2013:SNC

[HMVG13] Chen Huang, Bailey Miller, Frank Vahid, and Tony Givargis. Synthesis of networks of custom processing elements for real-time physical system emulation. ACM Transactions on Design Automation of Electronic Systems, 18(2): 21:1–21:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2007:CSS

[HN07] Shih-Hsu Huang and Yow-Tyng Nieh. Clock skew scheduling with race conditions considered. ACM Transactions on Design Automation of Electronic Systems, 12(4):45:1– 45:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hougardy:2023:FOD

[HNS23] Stefan Hougardy, Meike Neuwohner, and Ulrike Schorr. A fast optimal double-row legalization algorithm. ACM Transactions on Design Automation of Electronic Systems, 28(5): 71:1–71:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3579844.

Hsu:2011:MSS

[HPB11] Chia-Jui Hsu, José Luis Pino, and Shuvra S. Bhattacharyya. Multithreaded simulation for synchronous dataflow graphs. ACM Transactions on Design Automation of Electronic Systems, 16(3):25:1–25:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hong:1999:POU

[HPK99] Inki Hong, Miodrag Potkonjak, and Ramesh Karri. Power optimization using divideand-conquer techniques for minimization of the number of operations. ACM Transactions on Design Automation of Electronic Systems, 4(4):405–429, October 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-4/p405-hong/p405-
hong.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1999-4-4/p405-hong/

.

Hanchate:2006:GTF

[HR06] Narender Hanchate and Nagarajan Ranganathan. A gametheoretic framework for multimetric optimization of interconnect delay, power, and crosstalk noise during wire sizing. ACM Transactions on Design Automation of Electronic Systems, 11(3):711–739, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hassanpourghadi:2021:MLG

[HRC21] Mohsen Hassanpourghadi, Rezwan A. Rasul, and Mike Shuo-Wei Chen. A module-linking graph assisted hybrid optimization framework for custom analog and mixed-signal circuit parameter synthesis. ACM Transactions on Design Automation of Electronic Systems, 26(5):38:1–38:22, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3456722.

Hussain:2018:SSH

[HRK18] Siam Umar Hussain, M. Sadegh Riazi, and Farinaz Koushanfar. SHAIP: Secure Hamming Distance for Authentication of Intrinsic PUFs. ACM Transactions on Design Automation of Electronic Systems, 23(6):75:1– 75:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsiao:2000:DST

[HRP00] Michael S. Hsiao, Elizabeth M. Rudnick, and Janak H. Patel. Dynamic state traversal for sequential circuit test generation. ACM Transactions on Design Automation of Electronic Systems, 5(3):548–565, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p548-hsiao/p548-
hsiao.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p548-hsiao/

.

Han:2018:FCS

[HS18] Inhak Han and Youngsoo Shin.

Folded circuit synthesis: Minarea logic synthesis using dualedge-triggered flip-flops. ACM Transactions on Design Automation of Electronic Systems, 23(5):61:1–61:??, October 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hyun:2019:IAA

[HS19] Daijoon Hyun and Youngsoo Shin. Integrated approach of airgap insertion for circuit timing optimization. ACM Transactions on Design Automation of Electronic Systems, 24(2):24:1–24:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3306494.

Hung:2004:SCR

[HSA^{+}04] William N. N. Hung, Xiaoyu Song, El Mostapha Aboulhamid, Andrew Kennings, and Alan Coppola. Segmented channel routability via satisfiability. ACM Transactions on Design Automation of Electronic Systems, 9(4):517– 528, October 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hsiung:2000:CCM

[Hsi00] Pao-Ann Hsiung. CMAPS: a cosynthesis methodology for application-oriented parallel systems. ACM Transactions on Design Automation of Electronic Systems, 5(1):51– 81, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
2000-5-1/p51-hsiung/p51-
hsiung.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-1/p51-hsiung/

.

Hsiung:2001:PPO

[Hsi01] Pao-Ann Hsiung. POSE: a parallel object-oriented synthesis environment. ACM Transactions on Design Automation of Electronic Systems, 6(1):67–92, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-1/p67-hsiung/p67-
hsiung.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2001-6-1/p67-hsiung/

.

Huang:2022:AHS

[HSP^{+}22] Shanshi Huang, Xiaoyu Sun, Xiaochen Peng, Hongwu Jiang, and Shimeng Yu. Achieving high in situ training accuracy and energy efficiency with analog non-volatile synaptic devices. ACM Transactions on Design Automation of Electronic Systems, 27(4): 37:1–37:19, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3500929.

Hassani:2024:EFA

[HSR^{+}24] Fatemeh Serajeh Hassani, Mohammad Sadrosadati, Nezam Rohbani, Sebastian Pointner, Robert Wille, and Hamid Sarbazi-Azad. An efficient FPGA architecture with turnrestricted switch boxes. ACM Transactions on Design Automation of Electronic Systems, 29(3):42:1–42:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643809.

Huang:2023:STB

[HTC^{+}23] Po-Hsuan Huang, Chia-Heng Tu, Shen-Ming Chung, Pei-Yuan Wu, Tung-Lin Tsai, Yi-An Lin, Chun-Yi Dai, and Tzu-Yi Liao. SecureTVM: a TVM-based compiler framework for selective privacypreserving neural inference. ACM Transactions on Design Automation of Electronic Systems, 28(4):61:1–61:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579049.

Huang:2013:LPA

[HTCP13] Shih-Hsu Huang, Wen-Pin Tu, Chia-Ming Chang, and Song-Bin Pan. Low-power anti-aging zero skew clock gating. ACM Transactions on Design Automation of Electronic Systems, 18(2):27:1–27:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hu:2020:EMN

[Hu20] X. Sharon Hu. Editorial: a message from the new Editorin-Chief. ACM Transactions on Design Automation of Electronic Systems, 25(6):49e:1– 49e:2, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3419376.

Huang:2001:CSP

[Hua01] Ing-Jer Huang. Co-synthesis of pipelined structures and instruction reordering constraints for instruction set processors. ACM Transactions on Design Automation of Electronic Systems, 6(1):93–121, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-1/p93-huang/p93-huang.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/2001-6-1/p93-huang/

.

Huggins:1998:SVP

[HV98] James K. Huggins and David Van Campenhout. Specification and verification of pipelining in the ARM2 RISC microprocessor. ACM Transactions on Design Automation of Electronic Systems, 3(4):563– 580, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
1998-3-4/p563-huggins/p563-
huggins.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-4/p563-huggins/

.

Herrera:2007:FHS

[HV07] Fernando Herrera and Eugenio Villar. A framework for heterogeneous specification and design of electronic embedded systems in SystemC. ACM Transactions on Design Automation of Electronic Systems, 12(3):22:1–22:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hartanto:2001:DSS

[HVF^{+}01] Ismed Hartanto, Srikanth Venkataraman, W. Kent Fuchs, Elizabeth M. Rudnick, Janak H.

Patel, and Sreejit Chakravarty. Diagnostic simulation of stuckat faults in sequential circuits using compact lists. ACM Transactions on Design Automation of Electronic Systems, 6(4):471–489, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hwang:2000:PSS

[HW00] Chi-Hong Hwang and Allen C.-H. Wu. A predictive system shutdown method for energy saving of eventdriven computation. ACM Transactions on Design Automation of Electronic Systems, 5(2):226–241, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-2/p226-hwang/p226-
hwang.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-2/p226-hwang/

.

Hung:2014:AFD

[HW14] Eddie Hung and Steven J. E. Wilton. Accelerating FPGA debug: Increasing visibility using a runtime reconfigurable observation and triggering network. ACM Transactions on Design Automation of Electronic Systems, 19(2):14:1– 14:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2013:OCC

[HWCL13] Chien-Chih Huang, Chin-Long Wey, Jwu-E Chen, and Pei-Wen Luo. Optimal commoncentroid-based unit capacitor placements for yield enhancement of switched-capacitor circuits. ACM Transactions on Design Automation of Electronic Systems, 19(1):7:1–7:??, December 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2015:PDU

[HWCL15] Chien-Chih Huang, Chin-Long Wey, Jwu-E Chen, and Pei-Wen Luo. Performance-driven unit-capacitor placement of successive-approximation-register ADCs. ACM Transactions on Design Automation of Electronic Systems, 21(1):15:1– 15:??, November 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Han:2022:DSL

[HWDQ22] Ming Han, Ye Wang, Jian Dong, and Gang Qu. Doubleshift: a low-power DNN weights storage and access framework based on approximate decomposition and quantization. ACM Transactions on Design Automation of Electronic Systems, 27(2):15:1– 15:16, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3477047.

He:2023:GLP

[HWF^{+}23] Xu He, Yao Wang, Zhiyong Fu, Yipei Wang, and Yang Guo. A general layout pattern clustering using geometric matching-based clip relocation and lower-bound aided optimization. ACM Transactions on Design Automation of Electronic Systems, 28(6): 90:1–90:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3610293.

He:2016:RIM

[HWGY16] Xu He, Yao Wang, Yang Guo, and Evangeline F. Y. Young. Ripple 2.0: Improved movement of cells in routabilitydriven placement. ACM Transactions on Design Automation of Electronic Systems, 22 (1):10:1–10:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

He:2023:SEM

[HWL^{+}23a] Xu He, Yao Wang, Chang Liu, Qiang Wu, Juan Luo, and Yang Guo. A soft-error mitigation approach using pulse quenching enhancement at detailed placement for combinational circuits. ACM Transactions on Design Automation of Electronic Systems, 28(4): 65:1–65:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3595637.

Huang:2023:CFD

[HWL^{+}23b] Yanze Huang, Kui Wen, Limei Lin, Li Xu, and Sun-Yuan Hsieh. Component fault diagnosability of hierarchical cubic networks. ACM Transactions on Design Automation of Electronic Systems, 28(3): 39:1–39:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3577018.

Huang:2014:ICP

[HWX^{+}14] Libo Huang, Zhiying Wang, Nong Xiao, Yongwen Wang, and Qiang Dou. Integrated coherence prediction: Towards efficient cache coherence on NoCbased multicore architectures. ACM Transactions on Design Automation of Electronic Systems, 19(3):24:1–24:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

He:2022:DME

[HXB^{+}22] Jingyu He, Yao Xiao, Corina Bogdan, Shahin Nazarian, and Paul Bogdan. A design methodology for energyaware processing in unmanned aerial vehicles. ACM Transactions on Design Automation of Electronic Systems, 27(1):4:1– 4:20, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3470451.

Huang:2018:DML

[HXC^{+}18] Hantao Huang, Hang Xu, Yuehua Cai, Rai Suleman Khalid, and Hao Yu. Distributed machine learning on smartgateway network toward realtime smart-grid energy management with behavior cognition. ACM Transactions on Design Automation of Electronic Systems, 23(5):56:1– 56:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Huang:2023:RRB

[HXZ^{+}23] Chenglong Huang, Nuo Xu, Junwei Zeng, Wenqing Wang, Yihong Hu, Liang Fang, Desheng Ma, and Yanting Chen. Rescuing ReRAM-based neural computing systems from device variation. ACM Transactions on Design Automation of Electronic Systems, 28(1):6:1– 6:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3533706.

Hoque:2020:HPO

[HYK^{+}20] Tamzidul Hoque, Kai Yang, Robert Karam, Shahin Tajik, Domenic Forte, Mark Tehranipoor, and Swarup Bhunia. Hidden in plaintext: an obfuscation-based countermeasure against FPGA bitstream tampering attacks. ACM Transactions on Design Automation of Electronic Systems, 25(1):4:1–4:32, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3361147.

Huang:2015:CPM

[HYN15] Shih-Hsu Huang, Hua-Hsin Yeh, and Yow-Tyng Nieh. Clock period minimization with minimum leakage power. ACM Transactions on Design Automation of Electronic Systems, 21(1):9:1–9:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Hou:2023:DLF

[HZJC23] Tianshu Hou, Peining Zhen, Zhigang Ji, and Hai-Bao Chen. A deep learning framework for solving stress-based partial differential equations in electromigration analysis. ACM Transactions on Design Automation of Electronic Systems, 28(4): 57:1–57:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567424.

Hong:2022:TNB

[HZL^{+}22] Xin Hong, Xiangzhen Zhou, Sanjiang Li, Yuan Feng, and Mingsheng Ying. A tensor network based decision diagram for representation of quantum circuits. ACM Transactions on Design Automation of Electronic Systems, 27 (6):60:1–60:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3514355.

Huang:2019:ILA

[HZS^{+}19] Bo-Yuan Huang, Hongce Zhang, Pramod Subramanyan, Yakir Vizel, Aarti Gupta, and Sharad Malik. Instruction-level abstraction (ILA): a uniform specification for system-on-chip (SoC) verification. ACM Transactions on Design Automation of Electronic Systems, 24(1):10:1–10:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Inoue:2009:DSD

[IAI^{+}09] Hiroaki Inoue, Tsuyoshi Abe, Kazuhisa Ishizaka, Junji Sakai, and Masato Edahiro. Dynamic security domain scaling on embedded symmetric multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 14(2): 24:1–24:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Issenin:2007:DDR

[IBMD07] Ilya Issenin, Erik Brockmeyer, Miguel Miranda, and Nikil Dutt. DRDU: a data reuse analysis technique for efficient scratch-pad memory management. ACM Transactions on Design Automation of Electronic Systems, 12(2):15:1– 15:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ioannides:2012:CDT

[IE12] Charalambos Ioannides and Kerstin I. Eder. Coveragedirected test generation automated by machine learning — a review. ACM Transactions on Design Automation of Electronic Systems, 17(1): 7:1–7:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ittershagen:2018:IFM

[IGN18] Philipp Ittershagen, Kim Grüttner, and Wolfgang Nebel. An integration flow for mixedcritical embedded systems on a flexible time-triggered platform. ACM Transactions on Design Automation of Electronic Systems, 23(4):51:1– 51:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Indrusiak:2015:FSN

[IHM15] Leandro Soares Indrusiak, James Harbin, and Osmar Marchi Dos Santos. Fast simulation of networks-on-chip with priority-preemptive arbitration. ACM Transactions on Design Automation of Electronic Systems, 20(4):56:1– 56:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ibrahim:2023:OPR

[IIEKS23] Abrar A. Ibrahim, Ahmed M. Y. Ibrahim, Mohamed Watheq El-Kharashi, and Mona Safar. Optimal pattern retargeting in IEEE 1687 networks: a SAT-based upper-bound computation. ACM Transactions on Design Automation of Electronic Systems, 28(4):62:1– 62:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3585074.

Islam:2019:EIT

[IK19] Md Nazmul Islam and Sandip Kundu. Enabling IC traceability via blockchain pegged to embedded PUF. ACM Transactions on Design Automation of Electronic Systems, 24(3): 36:1–36:23, June 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3315669.

Isenberg:2017:PCH

[IPWW17] Tobias Isenberg, Marco Platzner, Heike Wehrheim, and Tobias Wiersema. Proof-carrying hardware via inductive invariants. ACM Transactions on Design Automation of Electronic Systems, 22(4):61:1– 61:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Irwin:2000:E

[Irw00]  Mary Jane Irwin.  Edito-

rial. ACM Transactions on Design Automation of Electronic Systems, 5(3):265–266, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p265-irwin/p265-
irwin.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p265-irwin/

.

Inoue:2008:PVS

[ISE08] Hiroaki Inoue, Junji Sakai, and Masato Edahiro. Processor virtualization for secure mobile terminals. ACM Transactions on Design Automation of Electronic Systems, 13(3): 48:1–48:??, July 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Islam:2021:HLS

[ISK21] Sheikh Ariful Islam, Love Kumar Sah, and Srinivas Katkoori. High-level synthesis of keyobfuscated RTL IP with design lockout and camouflaging. ACM Transactions on Design Automation of Electronic Systems, 26(1):6:1–6:35, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410337.

Ince:2021:FBB

[IYF^{+}21] Mehmet Ince, Ender Yilmaz, Wei Fu, Joonsung Park, Krishnaswamy Nagaraj, Leroy Winemberg, and Sule Ozev. Fault-based built-in self-test and evaluation of phase locked loops. ACM Transactions on Design Automation of Electronic Systems, 26(3):20:1– 20:18, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3427911.

Johnson:1998:MAS

[JB98] E. W. Johnson and J. B. Brockman. Measurement and analysis of sequential design processes. ACM Transactions on Design Automation of Electronic Systems, 3(1):1–20, January 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-1/p1-johnson/p1-johnson.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1998-3-1/p1-johnson/

.

Jamieson:2010:BER

[JBC^{+}10] Peter Jamieson, Tobias Becker, Peter Y. K. Cheung, Wayne Luk, Tero Rissa, and Teemu Pitkänen. Benchmarking and evaluating reconfigurable architectures targeting the mobile domain. ACM Transactions on Design Automation of Electronic Systems, 15(2):14:1– 14:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jagadheesh:2022:NAM

[JBJ22] Samala Jagadheesh, P. Veda Bhanu, and Soumya J. NoC application mapping optimization using reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 27(6):55:1–55:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510381.

Jan:2005:GMR

[JCGP05] Gene Eu Jan, Ki-Yin Chang, Su Gao, and Ian Parberry. A 4-geometry maze router and its application on multiterminal nets. ACM Transactions on Design Automation of Electronic Systems, 10(1): 116–135, January 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jun:2018:RBD

[JCK^{+}18] Jaeyung Jun, Kyu Hyun Choi, Hokwon Kim, Sang Ho Yu, Seon Wook Kim, and Youngsun Han. Recovering from biased distribution of faulty cells in memory by reorganizing replacement regions through universal hashing. ACM Transactions on Design Automation of Electronic Systems, 23(2): 16:1–16:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jiang:2023:ISS

[JCPL23] Iris Hru Jiang, David Chinnery, Gracieli Posser, and Jens Lienig. Introduction to the special section on advances in physical design automation. ACM Transactions on Design Automation of Electronic Systems, 28(5): 68:1–68:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3604593.

Joo:2008:ECP

[JCS^{+}08] Yongsoo Joo, Youngjin Cho, Donghwa Shin, Jaehyun Park, and Naehyuck Chang. An energy characterization platform for memory devices and energy-aware data compression for multilevel-cell flash memory. ACM Transactions on Design Automation of Electronic Systems, 13(3):43:1–43:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jha:2000:HLL

[JD00] Pradip K. Jha and Nikil D. Dutt. High-level library mapping for memories. ACM Transactions on Design Automation of Electronic Systems, 5(3):566–603, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p566-jha/p566-jha.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/2000-5-3/p566-jha/.

Jiao:2018:OER

[JD18] Fengxian Jiao and Sheqin Dong. Ordered escape routing with consideration of differential pair and blockage. ACM Transactions on Design Automation of Electronic Systems, 23(4):46:1–46:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jana:2020:HHC

[JDD20] Rajib Lochan Jana, Soumyajit Dey, and Pallab Dasgupta. A hierarchical HVAC control scheme for energy-aware smart building automation. ACM Transactions on Design Automation of Electronic Systems, 25(4):31:1–31:33, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3393666.

Jiang:2024:RHF

[JDLZ24] Danping Jiang, Zibin Dai, Yanjiang Liu, and Zongren Zhang.

RGMU: a high-flexibility and low-cost reconfigurable Galois field multiplication unit design approach for CGRCA. ACM Transactions on Design Automation of Electronic Systems, 29(2):38:1–38:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639820.

Jones:2008:RFI

[JDT^{+}08] Alex K. Jones, Swapna Dontharaju, Shenchih Tung, Leo Mats, Peter J. Hawrylak, Raymond R. Hoare, James T. Cain, and Marlin H. Mickle. Radio frequency identification prototyping. ACM Transactions on Design Automation of Electronic Systems, 13(2): 29:1–29:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Juan:2014:SPT

[JGM14] Da-Cheng Juan, Siddharth Garg, and Diana Marculescu. Statistical peak temperature prediction and thermal yield improvement for 3D chip multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 19(4): 39:1–39:??, August 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jagannathan:2002:FAC

[JHL02] Ashok Jagannathan, Sung-Woo Hur, and John Lillis. A fast algorithm for contextaware buffer insertion. ACM Transactions on Design Automation of Electronic Systems, 7(1):173–188, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jassi:2018:GGB

[JHMGS18] Munish Jassi, Yong Hu, Daniel Mueller-Gritschneder, and Ulf Schlichtmann. Graphgrammar-based IP-integration (GRIP) — an EDA tool for software-defined SoCs. ACM Transactions on Design Automation of Electronic Systems, 23(3):40:1–40:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jayasinghe:2021:QQB

[JIR^{+}21] Darshana Jayasinghe, Aleksandar Ignjatovic, Roshan Ragel, Jude Angelo Ambrose, and Sri Parameswaran. Quad-Seal: Quadruple balancing to mitigate power analysis attacks with variability effects and electromagnetic fault injection attacks. ACM Transactions on Design Automation of Electronic Systems, 26(5):33:1– 33:36, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3443706.

Jeong:2021:DMB

[JJH21] Eunjin Jeong, Dowhan Jeong, and Soonhoi Ha. Dataflow model-based software synthesis framework for parallel and distributed embedded systems. ACM Transactions on Design Automation of Electronic Systems, 26(5):35:1–35:38, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3447680.

Jayakumar:2010:SIV

[JK10] Nikhil Jayakumar and Sunil P. Khatri. A simultaneous input vector control and circuit modification technique to reduce leakage with zero delay penalty. ACM Transactions on Design Automation of Electronic Systems, 16(1):9:1–9:??, November 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jing:2012:SFE

[JLF^{+}12] Naifeng Jing, Ju-Yueh Lee, Zhe Feng, Weifeng He, Zhigang Mao, and Lei He. SEU fault evaluation and characteristics for SRAM-based FPGA architectures and synthesis algorithms. ACM Transactions on Design Automation of Electronic Systems, 18(1):13:1– 13:??, December 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jafari:2015:LUD

[JLJ15] Fahimeh Jafari, Zhonghai Lu, and Axel Jantsch. Least upper delay bound for VBR flows in networks-on-chip with virtual channels. ACM Transactions on Design Automation of Electronic Systems, 20(3): 35:1–35:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jung:2015:LMS

[JLK15] Dongha Jung, Hokyoon Lee, and Seon Wook Kim. Lowering minimum supply voltage for power-efficient cache design by exploiting data redundancy. ACM Transactions on Design Automation of Electronic Systems, 21(1):11:1–11:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jose:2014:IAH

[JM14] John Jose and Madhu Mutyam.

Implementation and analysis of history-based output channel selection strategies for adaptive routers in mesh NoCs. ACM Transactions on Design Automation of Electronic Systems, 19(4):35:1–35:??, August 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jung:2019:ILP

[JNCS19] Jinwook Jung, Gi-Joon Nam, Woohyun Chung, and Youngsoo Shin. Integrated latch placement and cloning for timing optimization. ACM Transactions on Design Automation of Electronic Systems, 24 (2):22:1–22:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3301613.

Jeong:2017:CSP

[JNS^{+}17] Jae Woong Jeong, Vishwanath Natarajan, Shreyas Sen, Tm Mak, Jennifer Kitchen, and Sule Ozev. A comprehensive BIST solution for polar transceivers using on-chip resources. ACM Transactions on Design Automation of Electronic Systems, 23(1): 2:1–2:??, October 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jung:2017:MSM

[JOH17] Hanwoong Jung, Hyunok Oh, and Soonhoi Ha. Multiprocessor scheduling of a multimode dataflow graph considering mode transition delay. ACM Transactions on Design Automation of Electronic Systems, 22(2):37:1–37:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Johnson:2008:IME

[JP08] F. Ryan Johnson and Joann M. Paul. Interrupt modeling for efficient high-level scheduler design space exploration. ACM Transactions on Design Automation of Electronic Systems, 13(1):10:1–10:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jang:2012:AAA

[JP12] Wooyoung Jang and David Z.

Pan. A3MAP: Architectureaware analytic mapping for networks-on-chip. ACM Transactions on Design Automation of Electronic Systems, 17(3): 26:1–26:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jiang:2006:RCD

[JPCJ06] Iris Hui-Ru Jiang, Song-Ra Pan, Yao-Wen Chang, and Jing-Yang Jou. Reliable crosstalk-driven interconnect optimization. ACM Transactions on Design Automation of Electronic Systems, 11(1): 88–103, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jin:2016:CEE

[JPHL16] Song Jin, Songwei Pei, Yinhe Han, and Huawei Li. A costeffective energy optimization framework of multicore SoCs based on dynamically reconfigurable voltage-frequency islands. ACM Transactions on Design Automation of Electronic Systems, 21(2):27:1– 27:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jun:2019:FTT

[JPM^{+}19] Jaeyung Jun, Yoonah Paik, Gyeong Il Min, Seon Wook Kim, and Youngsun Han. Fault tolerance technique offlining faulty blocks by heap memory management. ACM Transactions on Design Automation of Electronic Systems, 24(4): 47:1–47:25, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3329079.

Johnson:1997:DSM

[JR97] Mark C. Johnson and Kaushik Roy. Datapath scheduling with multiple supply voltages and level converters. ACM Transactions on Design Automation of Electronic Systems, 2(3):227–248, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-3/p227-johnson/p227-
johnson.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-3/p227-johnson/

.

Jeyapaul:2013:EEE

[JS13] Reiley Jeyapaul and Aviral Shrivastava. Enabling energy efficient reliability in embedded systems through smart cache cleaning. ACM Transactions on Design Automation of Electronic Systems, 18(4): 53:1–53:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jalili:2018:ERM

[JSA18] Majid Jalili and Hamid Sarbazi-Azad. Express read in MLC phase change memories. ACM Transactions on Design Automation of Electronic Systems, 23(3):33:1–33:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jin:2009:GND

[JSG09] Zhong-Yi Jin, Curt Schurgers, and Rajesh K. Gupta. A gateway node with duty-cycled radio and processing subsystems for wireless sensor networks. ACM Transactions on Design Automation of Electronic Systems, 14(1):5:1–5:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jiang:2019:EEQ

[JSS^{+}19] Li Jiang, Zhuoran Song, Haiyue Song, Chengwen Xu, Qiang Xu, Naifeng Jing, Weifeng Zhang, and Xiaoyao Liang. Energy-efficient and qualityassured approximate computing framework using a cotraining method. ACM Transactions on Design Automation of Electronic Systems, 24(6): 59:1–59:25, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3342239.

Jone:1998:CAD

[JT98] Wen-Ben Jone and K. S. Tsai. Confidence analysis for defect-level estimation of VLSI random testing. ACM Transactions on Design Automation of Electronic Systems, 3(3):389–407, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1998-3-3/p389-jone/p389-
                                        jone.pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        todaes/1998-3-3/p389-jone/

.

Jones:2008:ISS

[JW08] Alex K. Jones and Robert Walker. Introduction to the special section on demonstrable software systems and hardware platforms II. ACM Transactions on Design Automation of Electronic Systems, 13(3): 38:1–38:??, July 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jone:2003:DTI

[JWL^{+}03] W.-B. Jone, J. S. Wang, Hsueh-I Lu, I. P. Hsu, and J.-Y. Chen. Design theory and implementation for low-power segmented bus systems. ACM Transactions on Design Automation of Electronic Systems, 8(1):38–54, January 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Jiang:2021:PDM

[JYHY21] Chen Jiang, Bo Yuan, Tsung-Yi Ho, and Xin Yao. Placement of digital microfluidic biochips via a new evolutionary algorithm. ACM Transactions on Design Automation of Electronic Systems, 26(6): 42:1–42:22, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460230.

Jiang:2022:ELH

[JYY^{+}22] Yiyang Jiang, Fan Yang, Bei Yu, Dian Zhou, and Xuan Zeng. Efficient layout hotspot detection via neural architecture search. ACM Transactions on Design Automation of Electronic Systems, 27(6): 62:1–62:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3517130.

Jain:2021:TTA

[JZG21] Ayush Jain, Ziqi Zhou, and Ujjwal Guin. TAAL: Tampering attack on any key-based logic locked circuits. ACM Transactions on Design Automation of Electronic Systems, 26(4):28:1–28:22, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3442379.

Jiang:2015:CLF

[JZYZ15] Lei Jiang, Bo Zhao, Jun Yang, and Youtao Zhang. Constructing large and fast on-chip cache for mobile processors with multilevel cell STT–MRAM technology. ACM Transactions on Design Automation of Electronic Systems, 20(4):54:1– 54:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Koblah:2023:SPA

[KAC^{+}23] David Koblah, Rabin Acharya, Daniel Capecci, Olivia Dizon-Paradis, Shahin Tajik, Fatemeh Ganji, Damon Woodard, and Domenic Forte. A survey and perspective on artificial intelligence for security-aware electronic design automation. ACM Transactions on Design Automation of Electronic Systems, 28(2):16:1–16:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3563391.

Kagaris:2005:UMP

[Kag05]  Dimitri Kagaris.  A uni-

fied method for phase shifter computation. ACM Transactions on Design Automation of Electronic Systems, 10(1): 157–167, January 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kamal:2016:YSI

[KAKSP16] Mehdi Kamal, Ali Afzali-Kusha, Saeed Safari, and Massoud Pedram. Yield and speedup improvements in extensible processors by allocating extra cycles to some custom instructions. ACM Transactions on Design Automation of Electronic Systems, 21(2): 28:1–28:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kandemir:2006:REC

[Kan06] Mahmut Taylan Kandemir. Reducing energy consumption of multiprocessor SoC architectures by exploiting memory bank locality. ACM Transactions on Design Automation of Electronic Systems, 11(2): 410–441, April 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Krashinsky:2008:ISV

[KBA08] Ronny Krashinsky, Christopher Batten, and Krste Asanović. Implementing the Scale vectorthread processor. ACM Transactions on Design Automation of Electronic Systems, 13(3): 41:1–41:??, July 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kavousianos:2009:EPS

[KBN09] Xrysovalantis Kavousianos, Dimitris Bakalis, and Dimitris Nikolos. Efficient partial scan cell gating for low-power scanbased testing. ACM Transactions on Design Automation of Electronic Systems, 14(2): 28:1–28:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2015:UIL

[KBV^{+}15] Hyungjun Kim, Siva Bhanu Krishna Boga, Arseniy Vitkovskiy, Stavros Hadjitheophanous, Paul V. Gratz, Vassos Soteriou, and Maria K. Michael. Use it or lose it: Proactive, deterministic longevity in future chip multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 20(4):65:1– 65:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Khordoc:1998:SVA

[KC98] K. Khordoc and E. Cerny. Semantics and verification of action diagrams with linear timing. ACM Transactions on Design Automation of Electronic Systems, 3(1):21–50, January 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/

articles/journals/todaes/
1998-3-1/p21-khordoc/p21-
khordoc.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-1/p21-khordoc/

.

Kim:2010:EEP

[KC10] Jinsik Kim and Pai H. Chou.

Energy-efficient progressive remote update for flash-based firmware of networked embedded systems. ACM Transactions on Design Automation of Electronic Systems, 16(1):7:1– 7:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2013:AMP

[KC13] Sehwan Kim and Pai H. Chou.

Analysis and minimization of power-transmission loss in locally daisy-chained systems by local energy buffering. ACM Transactions on Design Automation of Electronic Systems, 18(3):37:1–37:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kjeldsberg:2004:SRE

[KCA04] P. G. Kjeldsberg, F. Catthoor, and E. J. Aas. Storage requirement estimation for optimized design of data intensive applications. ACM Transactions on Design Automation of Electronic Systems, 9(2):133–158, April 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kritikakou:2013:NOS

[KCKG13] Angeliki Kritikakou, Francky Catthoor, Vasilios Kelefouras, and Costas Goutis. Nearoptimal and scalable intrasignal in-place optimization for non-overlapping and irregular access schemes. ACM Transactions on Design Automation of Electronic Systems, 19(1): 4:1–4:??, December 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kritikakou:2016:ASC

[KCKG16] Angeliki Kritikakou, Francky Catthoor, Vasilios Kelefouras, and Costas Goutis. Array size computation under uniform overlapping and irregular accesses. ACM Transactions on Design Automation of Electronic Systems, 21(2):22:1– 22:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kaur:2024:RRS

[KD24] Jaspinder Kaur and Shirshendu Das. RSPP: Restricted static pseudo-partitioning for mitigation of cross-core covert channel attacks. ACM Transactions on Design Automation of Electronic Systems, 29(2):27:1– 27:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3637222.

Kumar:2008:MSS

[KFH^{+}08] Akash Kumar, Shakith Fernando, Yajun Ha, Bart Mesman, and Henk Corporaal. Multiprocessor systems synthesis for multiple use-cases of multiple applications on FPGA. ACM Transactions on Design Automation of Electronic Systems, 13(3):40:1– 40:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kern:1999:FVH

[KG99] Christoph Kern and Mark R.

Greenstreet. Formal verification in hardware design: a survey. ACM Transactions on Design Automation of Electronic Systems, 4(2):123–193, April 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-2/p123-kern/p123-
kern.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1999-4-2/p123-kern/

.

Kumar:2009:EML

[KG09] Yokesh Kumar and Prosenjit Gupta. External memory layout vs. schematic. ACM Transactions on Design Automation of Electronic Systems, 14(2): 30:1–30:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kamal:2020:ADF

[KGS^{+}20] Nishant Kamal, Ankur Gupta, Ananya Singla, Shubham Tiwari, Parth Kohli, Sudip Roy, and Bhargab B. Bhattacharya. Architectural design of flowbased microfluidic biochips for multi-target dilution of biochemical fluids. ACM Transactions on Design Automation of Electronic Systems, 25(3): 25:1–25:34, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3357604.

Kwon:2010:SPC

[KH10] Seongnam Kwon and Soonhoi Ha. Serialized parallel code generation framework for MPSoC. ACM Transactions on Design Automation of Electronic Systems, 15(2):11:1– 11:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Khatib:2012:MRP

[Kha12] Mohammed G. Khatib. Migrationresistant policies for probewear leveling in MEMS storage devices. ACM Transactions on Design Automation of Electronic Systems, 17(4): 49:1–49:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Khan:2023:HEC

[Kha23] Muhammad Imran Khan. Harmonic estimation and comparative analysis of ultra-high speed flip-flop and latch topologies for low power and high performance future generation micro- /nano electronic systems. ACM Transactions on Design Automation of Electronic Systems, 28(4):64:1–64:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3590770.

Koushanfar:2005:BST

[KHP05] Farinaz Koushanfar, Inki Hong, and Miodrag Potkonjak. Behavioral synthesis techniques for intellectual property protection. ACM Transactions on Design Automation of Electronic Systems, 10(3):523–545, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kuo:2006:DID

[KHW06] Wu-An Kuo, Tingting Hwang, and Allen C.-H. Wu. Decomposition of instruction decoders for low-power designs. ACM Transactions on Design Automation of Electronic Systems, 11(4):880–889, October 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Karri:2001:IRT

[KI01] Ramesh Karri and Balakrishnan Iyer. Introspection: a register transfer level technique for cocurrent error detection and diagnosis in data dominated designs. ACM Transactions on Design Automation of Electronic Systems, 6(4):501– 515, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2003:MDO

[KJKK03] Ki-Wook Kim, Seong-Ook Jung, Taewhan Kim, and Sung-Mo Kang. Minimum delay optimization for domino logic circuits—a coupling-aware approach. ACM Transactions on Design Automation of Electronic Systems, 8(2):203–213, April 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kobayashi:2007:MOS

[KJR^{+}07] Yuki Kobayashi, Murali Jayapala, Praveen Raghavan, Francky Catthoor, and Masaharu Imai. Methodology for operation shuffling and L0 cluster generation for low energy heterogeneous VLIW processors. ACM Transactions on Design Automation of Electronic Systems, 12(4):41:1–41:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Krishna:2004:AHE

[KJT04] C. V. Krishna, Abhijit Jas, and Nur A. Touba. Achieving high encoding efficiency with partial dynamic LFSR reseeding. ACM Transactions on Design Automation of Electronic Systems, 9(4):500–516, October 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2011:CTS

[KK11] Tak-Yung Kim and Taewhan Kim. Clock tree synthesis for TSV-based 3D IC designs. ACM Transactions on Design Automation of Electronic Systems, 16(4):48:1–48:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kang:2014:IRA

[KK14] Minseok Kang and Taewhan Kim. Integrated resource allocation and binding in clock mesh synthesis. ACM Transactions on Design Automation of Electronic Systems, 19(3): 30:1–30:??, June 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2002:LTL

[KKH^{+}02] Ki-Wook Kim, Taewhan Kim, Ting-Ting Hwang, Sung-Mo Kang, and C. L. Liu. Logic transformation for low-power synthesis. ACM Transactions on Design Automation of Electronic Systems, 7(2):265–283, April 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2016:NAP

[KKHK16] Seungwon Kim, Seokhyeong Kang, Ki Jin Han, and Youngmin Kim. Novel adaptive power-gating strategy and tapered TSV structure in multilayer 3D IC. ACM Transactions on Design Automation of Electronic Systems, 21(3): 44:1–44:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kwon:2008:RPP

[KKJ^{+}08] Seongnam Kwon, Yongjoo Kim, Woo-Chul Jeun, Soonhoi Ha, and Yunheung Paek. A retargetable parallel-programming framework for MPSoC. ACM Transactions on Design Automation of Electronic Systems, 13(3):39:1–39:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2012:SAH

[KKK12] Yonghwan Kim, Sanghoon Kwak, and Taewhan Kim. Synthesis of adaptable hybrid adders for area optimization under timing constraint. ACM Transactions on Design Automation of Electronic Systems, 17(4):43:1–43:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2015:MMS

[KKLG15] Myungsun Kim, Jinkyu Koo, Hyojung Lee, and James R.

Geraci. Memory management scheme to improve utilization efficiency and provide fast contiguous allocation without a statically reserved area. ACM Transactions on Design Automation of Electronic Systems, 21(1):4:1–4:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kahng:2015:IMR

[KKLP15] Andrew B. Kahng, Seokhyeong Kang, Jiajia Li, and Jose Pineda De Gyvez. An improved methodology for resilient design implementation. ACM Transactions on Design Automation of Electronic Systems, 20(4):66:1–66:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kastner:2002:IGH

[KKMB02] R. Kastner, A. Kaplan, S. Ogrenci Memik, and E. Bozorgzadeh. Instruction generation for hybrid reconfigurable systems. ACM Transactions on Design Automation of Electronic Systems, 7(4):605–627, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2016:SDM

[KKS16] Sangmin Kim, Seokhyeong Kang, and Youngsoo Shin. Synthesis of dual-mode circuits through library design, gate sizing, and clock-tree optimization. ACM Transactions on Design Automation of Electronic Systems, 21(3): 51:1–51:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kao:2005:EAF

[KL05] Chi-Chou Kao and Yen-Tai Lai. An efficient algorithm for finding the minimal-area FPGA technology mapping. ACM Transactions on Design Automation of Electronic Systems, 10(1):168–186, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Knechtel:2018:MOF

[KLE18] Johann Knechtel, Jens Lienig, and Ibrahim (Abe) M. Elfadel. Multi-objective 3D floorplanning with integrated voltage assignment. ACM Transactions on Design Automation of Electronic Systems, 23(2):22:1– 22:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kuo:2014:RCS

[KLJ14] Hsien-Kai Kuo, Bo-Cheng Charles Lai, and Jing-Yang Jou. Reducing contention in shared last-level cache for throughput processors. ACM Transactions on Design Automation of Electronic Systems, 20(1):12:1– 12:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2017:SBS

[KLK^{+}17] Taehyun Kim, Jongbum Lim, Jinku Kim, Woo-Cheol Cho, Eui-Young Chung, and Hyuk-Jun Lee. Scalable bandwidth shaping scheme via adaptively managed parallel heaps in manycore-based network processors. ACM Transactions on Design Automation of Electronic Systems, 22(4): 59:1–59:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2024:OMP

[KLP^{+}24] Seok Young Kim, Jaewook Lee, Yoonah Paik, Chang Hyun Kim, Won Jun Lee, and Seon Wook Kim. Optimal model partitioning with lowoverhead profiling on the PIMbased platform for deep learning inference. ACM Transactions on Design Automation of Electronic Systems, 29(2):28:1– 28:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3628599.

Kim:2011:MAO

[KLSP11] Yongjoo Kim, Jongeun Lee, Aviral Shrivastava, and Yunheung Paek. Memory access optimization in compilation for coarse-grained reconfigurable architectures. ACM Transactions on Design Automation of Electronic Systems, 16(4):42:1–42:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Keutzer:2009:ATD

[KLSZ09] Kurt Keutzer, Peng Li, Li Shang, and Hai Zhou. ACM Transactions on Design Automation of Electronic Systems (TODAES) special section call for papers: Parallel CAD: Algorithm design and programming. ACM Transactions on Design Automation of Electronic Systems, 15(1):9:1–9:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Keutzer:2011:SSM

[KLSZ11] Kurt Keutzer, Peng Li, Li Shang, and Hai Zhou. A special section on multicore parallel CAD: Algorithm design and programming. ACM Transactions on Design Automation of Electronic Systems, 16(3): 21:1–21:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2015:AIP

[KLV15] Lok-Won Kim, Dong-U Lee, and John Villasenor. Automated iterative pipelining for ASIC design. ACM Transactions on Design Automation of Electronic Systems, 20(2):28:1– 28:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kahng:1997:ARI

[KM97] Andrew B. Kahng and Sudhakar Muddu. Analysis of RC interconnections under ramp input. ACM Transactions on Design Automation of Electronic Systems, 2(2):168–192, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p168-kahng/p168-
kahng.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-2/p168-kahng/

.

Kormicki:1997:PLS

[KMC97] Maciek Kormicki, Ausif Mahmood, and Bradley S. Carlson. Parallel logic simulation on a network of workstations using parallel virtual machine. ACM Transactions on Design Automation of Electronic Systems, 2(2):123–134, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p123-kormicki/p123-
kormicki.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-2/p123-kormicki/

.

Kurimoto:2012:YRI

[KMO^{+}12] Masanori Kurimoto, Jun Matsushima, Shigeki Ohbayashi, Yoshiaki Fukui, Michio Komoda, and Nobuhiro Tsuda. A yield and reliability improvement methodology based on logic redundant repair with a repairable scan flip-flop designed by push rule. ACM Transactions on Design Automation of Electronic Systems, 17(2):17:1–17:??, April 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kritikakou:2018:DDS

[KMR18] Angeliki Kritikakou, Thibaut Marty, and Matthieu Roy. DY-NASCORE: DYNAmic Software COntroller to Increase REsource utilization in mixedcritical systems. ACM Transactions on Design Automation of Electronic Systems, 23(2): 13:1–13:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Karfa:2012:FVC

[KMS12] Chandan Karfa, Chittaranjan Mandal, and Dipankar Sarkar. Formal verification of code motion techniques using data-flow-driven equivalence checking. ACM Transactions on Design Automation of Electronic Systems, 17(3): 30:1–30:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kolson:1996:ORA

[KNDK96] David J. Kolson, Alexandru Nicolau, Nikil Dutt, and Ken Kennedy. Optimal register assignment to loops for embedded code generation. ACM Transactions on Design Automation of Electronic Systems, 1(2):251–279, April 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-2/p251-kolson/p251-
kolson.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-2/p251-kolson/

.

Kulkarni:2006:CTA

[KNRK06] Dhananjay Kulkarni, Walid A.

Najjar, Robert Rinker, and Fadi J. Kurdahi. Compiletime area estimation for LUTbased FPGAs. ACM Transactions on Design Automation of Electronic Systems, 11(1): 104–122, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kashyap:2023:IIM

[KO23] Suhas Krishna Kashyap and Sule Ozev. IMPRoVED: Integrated method to predict PostRouting setup violations in early design stages. ACM Transactions on Design Automation of Electronic Systems, 28(4):49:1–49:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3572546.

Karabacak:2018:RDU

[KOO18] Fatih Karabacak, Umit Ogras, and Sule Ozev. Remote detection of unauthorized activity via spectral analysis. ACM Transactions on Design Automation of Electronic Systems, 23(6):81:1–81:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2009:MLP

[KOS09] Jaehyun Kim, Chungki Oh, and Youngsoo Shin. Minimizing leakage power of sequential circuits through mixed-V_{t} flip-flops and multi-V_{t} combinational gates. ACM Transactions on Design Automation of Electronic Systems, 15(1): 4:1–4:??, December 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kornaros:2013:STC

[KP13] Georgios Kornaros and Dionisios Pnevmatikatos. A survey and taxonomy of on-chip monitoring of multicore systemson-chip. ACM Transactions on Design Automation of Electronic Systems, 18(2):17:1– 17:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kee:2022:LPP

[KP22] Minkwan Kee and Gi-Ho Park.

A low-power programmable machine learning hardware accelerator design for intelligent edge devices. ACM Transactions on Design Automation of Electronic Systems, 27(5): 51:1–51:13, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3531479.

Kukkala:2019:JSF

[KPB19] Vipin Kumar Kukkala, Sudeep Pasricha, and Thomas Bradley. JAMS-SG: a framework for jitter-aware message scheduling for time-triggered automotive networks. ACM Transactions on Design Automation of Electronic Systems, 24(6):63:1– 63:31, November 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3355392.

Kashif:2016:PSR

[KPF16] Hany Kashif, Hiren Patel, and Sebastian Fischmeister. Path selection for real-time communication on priority-aware NoCs. ACM Transactions on Design Automation of Electronic Systems, 21(3):53:1– 53:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kang:2006:STA

[KPR06] Kunhyuk Kang, Bipul C. Paul, and Kaushik Roy. Statistical timing analysis using levelized covariance propagation considering systematic and random variations of process parameters. ACM Transactions on Design Automation of Electronic Systems, 11(4):848–879, October 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kahng:2009:LAA

[KPSW09] Andrew B. Kahng, Chul-Hong Park, Puneet Sharma, and Qinke Wang. Lens aberration aware placement for timing yield. ACM Transactions on Design Automation of Electronic Systems, 14(1):16:1– 16:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kang:2019:TDF

[KQP^{+}19] Ilgweon Kang, Fang Qiao, Dongwon Park, Daniel Kane, Evangeline Fung Yu Young, Chung-Kuan Cheng, and Ronald Graham. Three-dimensional floorplan representations by using corner links and partial order. ACM Transactions on Design Automation of Electronic Systems, 24(1):13:1–13:??, January 2019. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic).

Kundu:2023:MTF

[KR23] Debraj Kundu and Sudip Roy.

Multi-target fluid mixing in MEDA biochips: Theory and an attempt toward waste minimization. ACM Transactions on Design Automation of Electronic Systems, 28(6): 101:1–101:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3622785.

K:2018:AAF

[KRH18] Keerthi K., Chester Rebeiro, and Aritra Hazra. An algorithmic approach to formally verify an ECC library. ACM Transactions on Design Automation of Electronic Systems, 23(5): 63:1–63:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Koch:1998:BBD

[KRK98] Gernot H. Koch, W. Rosenstiel, and U. Kebschull. Breakpoints and breakpoint detection in source-level emulation. ACM Transactions on Design Automation of Electronic Systems, 3(2):209–230, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/

                                        1998-3-2/p209-koch/p209-
                                        koch.pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        todaes/1998-3-2/p209-koch/

.

Kiddie:2015:SEM

[KRL15] Bradley T. Kiddie, William H. Robinson, and Daniel B. Limbrick. Single-event multipletransient characterization and mitigation via alternative standard cell placement methods. ACM Transactions on Design Automation of Electronic Systems, 20(4):60:1– 60:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kandemir:2006:IEB

[KRS06] M. Kandemir, J. Ramanujam, and U. Sezer. Improving the energy behavior of block buffering using compiler optimizations. ACM Transactions on Design Automation of Electronic Systems, 11(1):228–250, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kritikakou:2023:MMS

[KS23] Angeliki Kritikakou and Stefanos Skalistis. Mitigating mode-switch through runtime computation of response time. ACM Transactions on Design Automation of Electronic Systems, 28(5): 81:1–81:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597432.

Kurimoto:2010:PAE

[KSA^{+}10] Masanori Kurimoto, Hiroaki Suzuki, Rei Akiyama, Tadao Yamanaka, Haruyuki Ohkuma, Hidehiro Takata, and Hirofumi Shinohara. Phase-adjustable error detection flip-flops with 2stage hold-driven optimization, slack-based grouping scheme and slack distribution control for dynamic voltage scaling. ACM Transactions on Design Automation of Electronic Systems, 15(2):17:1–17:??, February 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kolhe:2022:BDS

[KSD^{+}22] Gaurav Kolhe, Tyler David Sheaves, Sai Manoj P. D., Hamid Mahmoodi, Setareh Rafatirad, Avesta Sasan, and Houman Homayoun. Breaking the design and security trade-off of look-up-table-based obfuscation. ACM Transactions on Design Automation of Electronic Systems, 27(6): 56:1–56:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510421.

Kadayif:2005:OIT

[KSK^{+}05] I. Kadayif, A. Sivasubramaniam, M. Kandemir, G. Kandiraju, and G. Chen. Optimizing instruction TLB energy using software and hardware techniques. ACM Transactions on Design Automation of Electronic Systems, 10(2):229–257, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Keinert:2009:SAE

[KSS^{+}09] Joachim Keinert, Martin Streubühr, Thomas Schlichter, Joachim Falk, Jens Gladigau, Christian Haubelt, Jürgen Teich, and Michael Meredith. System-CoDesigner — an automatic ESL synthesis approach by design space exploration and behavioral synthesis for streaming applications. ACM Transactions on Design Automation of Electronic Systems, 14(1): 1:1–1:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kagaris:1996:FAM

[KT96] Dimitrios Kagaris and Spyros Tragoudas. A fast algorithm for minimizing FPGA combinational and sequential modules. ACM Transactions on Design Automation of Electronic Systems, 1(3):341–351, July 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-3/p341-kagaris/p341-
kagaris.pdf;   http://www.

acm.org/pubs/citations/journals/
todaes/1996-1-3/p341-kagaris/

.

Kagaris:2001:NHC

[KT01] D. Kagaris and S. Tragoudas.

Von Neumann hybrid cellular automata for generating deterministic test sequences. ACM Transactions on Design Automation of Electronic Systems, 6(3):308–321, July 2001. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kadayif:2013:HSA

[KTKO13] Ismail Kadayif, Mahir Turkcan, Seher Kiziltepe, and Ozcan Ozturk. Hardware/software approaches for reducing the process variation impact on instruction fetches. ACM Transactions on Design Automation of Electronic Systems, 18(4): 54:1–54:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kuchcinski:2003:CDS

[Kuc03] Krzysztof Kuchcinski. Constraintsdriven scheduling and resource assignment. ACM Transactions on Design Automation of Electronic Systems, 8(3):355–383, July 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Krishnaswamy:2008:PTM

[KVMH08] Smita Krishnaswamy, George F.

Viamontes, Igor L. Markov, and John P. Hayes. Probabilistic transfer matrices in symbolic reliability analysis of logic circuits. ACM Transactions on Design Automation of Electronic Systems, 13(1): 8:1–8:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kountouris:2002:ESC

[KW02] Apostolos A. Kountouris and Christophe Wolinski. Efficient scheduling of conditional behaviors for high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 7(3):380–412, July 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Katoen:2016:PMC

[KW16] Joost-Pieter Katoen and Hao Wu. Probabilistic model checking for uncertain scenarioaware data flow. ACM Transactions on Design Automation of Electronic Systems, 22 (1):15:1–15:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kim:2016:IWP

[KYL16] Youngsik Kim, Sungjoo Yoo, and Sunggu Lee. Improving write performance by controlling target resistance distributions in MLC PRAM. ACM Transactions on Design Automation of Electronic Systems, 21(2):23:1–23:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kurimoto:2012:VWR

[KYN^{+}12] Masanori Kurimoto, Takeshi Yamamoto, Satoshi Nakano, Atsuto Hanami, and Hiroyuki Kondo. Verification work reduction methodology in lowpower chip implementation. ACM Transactions on Design Automation of Electronic Systems, 18(1):12:1–12:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Kazerooni-Zand:2023:MBM

[KZKAKP23] Reza Kazerooni-Zand, Mehdi Kamal, Ali Afzali-Kusha, and Massoud Pedram. Memristivebased mixed-signal CGRA for accelerating deep neural network inference. ACM Transactions on Design Automation of Electronic Systems, 28(4): 66:1–66:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3595638.

Liu:2001:ODC

[LAS01] Tai-Hung Liu, Adnan Aziz, and Vigyan Singhal. Optimizing designs containing black boxes. ACM Transactions on Design Automation of Electronic Systems, 6(4):591–601, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2023:EES

[LAYZ23] Chunqiao Li, Chengtao An, Fan Yang, and Xuan Zeng. ESPSim: an efficient scalable power grid simulator based on parallel algebraic multigrid. ACM Transactions on Design Automation of Electronic Systems, 28(1):5:1–5:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529533.

Leupers:2000:GBC

[LB00] Rainer Leupers and Steven Bashford. Graph-based code selection techniques for embedded processors. ACM Transactions on Design Automation of Electronic Systems, 5(4):794–814, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-4/p794-leupers/p794-
leupers.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-4/p794-leupers/

.

Ludwin:2011:EDP

[LB11] Adrian Ludwin and Vaughn Betz. Efficient and deterministic parallel placement for FP-GAs. ACM Transactions on Design Automation of Electronic Systems, 16(3):22:1– 22:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2006:LVA

[LBV^{+}06] Wei Li, Daniel Blakely, Scott Van Sooy, Keven Dunn, David Kidd, Robert Rogenmoser, and Dian Zhou. LVS verification across multiple power domains for a quad-core microprocessor. ACM Transactions on Design Automation of Electronic Systems, 11(2):490–500, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Langevin:1996:RTC

[LC96] M. Langevin and E. Cerny. A recursive technique for computing lower-bound performance of schedules. ACM Transactions on Design Automation of Electronic Systems, 1(4):443–455, October 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-4/p443-langevin/p443-
langevin.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-4/p443-langevin/

.

Liu:2007:IEM

[LC07] Jinfeng Liu and Pai H. Chou.

Idle energy minimization by mode sequence optimization. ACM Transactions on Design Automation of Electronic Systems, 12(4):38:1– 38:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2013:SRB

[LC13] Ren-Jie Lee and Hung-Ming Chen. A study of row-based area-array I/O design planning in concurrent chip-package design flow. ACM Transactions on Design Automation of Electronic Systems, 18(2): 30:1–30:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2014:CPA

[LC14] Seokhyun Lee and Kiyoung Choi. Critical-path-aware highlevel synthesis with distributed controller for fast timing closure. ACM Transactions on Design Automation of Electronic Systems, 19(2):16:1– 16:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2011:GRS

[LCC11] Yih-Lang Li, Yu-Ning Chang, and Wen-Nai Cheng. A gridless routing system with nonslicing floorplanning-based crosstalk reduction on gridless track assignment. ACM Transactions on Design Automation of Electronic Systems, 16(2): 19:1–19:??, March 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2015:EEB

[LCC^{+}15] Jingwei Lu, Pengwen Chen, Chin-Chih Chang, Lu Sha, Dennis Jen-Hsin Huang, Chin-Chi Teng, and Chung-Kuan Cheng. ePlace: Electrostaticsbased placement using Fast Fourier Transform and Nesterov’s method. ACM Transactions on Design Automation of Electronic Systems, 20(2): 17:1–17:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2007:ISS

[LCD07] Jong-Eun Lee, Kiyoung Choi, and Nikil D. Dutt. Instruction set synthesis with efficient instruction encoding for configurable processors. ACM Transactions on Design Automation of Electronic Systems, 12(1): 8:1–8:??, January 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2022:DDN

[LCG^{+}22] Chaojian Li, Wuyang Chen, Yuchen Gu, Tianlong Chen, Yonggan Fu, Zhangyang Wang, and Yingyan Lin. DANCE: DAta-Network Co-optimization for Efficient segmentation model training and inference. ACM Transactions on Design Automation of Electronic Systems, 27(5):

50:1–50:20, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3510835.

Lin:2002:OTB

[LCHT02] Shi-Zheng Eric Lin, Chieh Changfan, Yu-Chin Hsu, and Fur-Shing Tsai. Optimal time borrowing analysis and timing budgeting optimization for latch-based designs. ACM Transactions on Design Automation of Electronic Systems, 7(1):217–230, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2010:ECR

[LCJ^{+}10] Shenghua Liu, Guoqiang Chen, Tom Tong Jing, Lei He, Robi Dutta, and Xian-Long Hong. Effective congestion reduction for IC package substrate routing. ACM Transactions on Design Automation of Electronic Systems, 15(3):27:1–27:??, May 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2022:MEC

[LCJ^{+}22] Sunjung Lee, Jaewan Choi, Wonkyung Jung, Byeongho Kim, Jaehyun Park, Hweesoo Kim, and Jung Ho Ahn. MVP: an efficient CNN accelerator with matrix, vector, and processing-near-memory units. ACM Transactions on Design Automation of Electronic Systems, 27(5): 42:1–42:25, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3497745.

Lee:2009:TSA

[LCK^{+}09] Byunghyun Lee, Ki-Seok Chung, Bontae Koo, Nak-Woong Eum, and Taewhan Kim. Thermal sensor allocation and placement for reconfigurable systems. ACM Transactions on Design Automation of Electronic Systems, 14(4):50:1– 50:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2012:FHA

[LCKT12] Chien-Nan Jimmy Liu, Yen-Lung Chen, Chin-Cheng Kuo, and I-Ching Tsai. A fast heuristic approach for parametric yield enhancement of analog designs. ACM Transactions on Design Automation of Electronic Systems, 17(3): 35:1–35:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2008:EDI

[LCL08] Chao-Hung Lu, Hung-Ming Chen, and Chien-Nan Jimmy Liu. Effective decap insertion in area-array SoC floorplan design. ACM Transactions on Design Automation of Electronic Systems, 13(4):66:1– 66:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2007:CCA

[LCOM07] Hyung Gyu Lee, Naehyuck Chang, Umit Y. Ogras, and Radu Marculescu. On-chip communication architecture exploration: a quantitative evaluation of point-to-point, bus, and network-on-chip approaches. ACM Transactions on Design Automation of Electronic Systems, 12(3):23:1– 23:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2003:TDC

[LCT03] Lei Li, Krishnendu Chakrabarty, and Nur A. Touba. Test data compression using dictionaries with selective entries and fixedlength indices. ACM Transactions on Design Automation of Electronic Systems, 8(4): 470–490, October 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Leung:2012:PVI

[LCY12] Mario K. Y. Leung, Eric K. I. Chio, and Evangeline F. Y. Young. Postplacement voltage island generation. ACM Transactions on Design Automation of Electronic Systems, 17(1): 4:1–4:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2018:RML

[LCYN18] Bo Liu, Gong Chen, Bo Yang, and Shigetoshi Nakatake. Routable and matched layout styles for analog module generation. ACM Transactions on Design Automation of Electronic Systems, 23(4):47:1–47:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2008:FCB

[LCZ^{+}08] Kyungsoo Lee, Naehyuck Chang, Jianli Zhuo, Chaitali Chakrabarti, Sudheendra Kadri, and Sarma Vrudhula. A fuelcell-battery hybrid for portable embedded systems. ACM Transactions on Design Automation of Electronic Systems, 13(1):19:1–19:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liang:2024:DAU

[LCZ^{+}24] Tung-Che Liang, Yi-Chen Chang, Zhanwei Zhong, Yaas Bigdeli, Tsung-Yi Ho, Krishnendu Chakrabarty, and Richard Fair. Dynamic adaptation using deep reinforcement learning for digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 29(2):24:1– 24:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3633458.

Li:2017:ASE

[LD17] Ji Li and Jeffrey Draper. Accelerated soft-error-rate (SER) estimation for combinational and sequential circuits. ACM Transactions on Design Automation of Electronic Systems, 22(3):57:1–57:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2018:PTT

[LDD^{+}18] Dongjin Lee, Sourav Das, Janardhan Rao Doppa, Partha Pratim Pande, and Krishnendu Chakrabarty. Performance and thermal tradeoffs for energy-efficient monolithic 3D network-on-chip. ACM Transactions on Design Automation of Electronic Systems, 23(5): 60:1–60:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2019:IEC

[LDD^{+}19] Dongjin Lee, Sourav Das, Janardhan Rao Doppa, Partha Pratim Pande, and Krishnendu Chakrabarty. Impact of electrostatic coupling on monolithic 3D-enabled network on chip. ACM Transactions on Design Automation of Electronic Systems, 24(6):62:1–62:22, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3357158.

Liao:1999:TCB

[LDK99] Stan Liao, Srinivas Devadas, and Kurt Keutzer. A textcompression-based method for code size minimization in embedded systems. ACM Transactions on Design Automation of Electronic Systems, 4(1):12–38, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-1/p12-liao/p12-liao.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1999-4-1/p12-liao/.

Lezos:2020:LOL

[LDLM20] Christakis Lezos, Grigoris Dimitroulakos, Ioannis Latifis, and Konstantinos Masselos. A locality optimizer for loopdominated applications based on reuse distance analysis. ACM Transactions on Design Automation of Electronic Systems, 25(6):51:1–51:26, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3398189.

Laubeuf:2022:DQR

[LDP^{+}22] Nathan Laubeuf, Jonas Doevenspeck, Ioannis A. Papistas, Michele Caselli, Stefan Cosemans, Peter Vrancx, Debjyoti Bhattacharjee, Arindam Mallik, Peter Debacker, Diederik Verkest, Francky Catthoor, and Rudy Lauwereins. Dynamic quantization range control for analog-in-memory neural networks acceleration. ACM Transactions on Design Automation of Electronic Systems, 27(5):46:1–46:21, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498328.

Luo:2022:FDF

[LDX22] Yukui Luo, Shijin Duan, and Xiaolin Xu. FPGAPRO: a defense framework against crosstalk-induced secret leakage in FPGA. ACM Transactions on Design Automation of Electronic Systems, 27(3): 24:1–24:31, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3491214.

Lin:2012:RSP

[LF12] Hai Lin and Yunsi Fei. Resource sharing of pipelined custom hardware extension for energy-efficient applicationspecific instruction set processor design. ACM Transactions on Design Automation of Electronic Systems, 17(4): 39:1–39:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2009:MAA

[LFG^{+}09] Bo Liu, Francisco V. Fernández, Georges Gielen, R. Castro-López, and E. Roca. A memetic approach to the automatic design of high-performance analog integrated circuits. ACM Transactions on Design Automation of Electronic Systems, 14(3):42:1–42:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Letras:2021:MOO

[LFST21] Martin Letras, Joachim Falk, Tobias Schwarzer, and Jürgen Teich. Multi-objective optimization of mapping dataflow applications to MPSoCs using a hybrid evaluation combining analytic models and measurements. ACM Transactions on Design Automation of Electronic Systems, 26(3):18:1– 18:33, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3431814.

Lee:2012:ECM

[LG12] John Lee and Puneet Gupta.

ECO cost measurement and incremental gate sizing for late process changes. ACM Transactions on Design Automation of Electronic Systems, 18 (1):16:1–16:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2018:LBF

[LG18] Dongwook Lee and Andreas Gerstlauer. Learning-based, fine-grain power modeling of system-level hardware IPs. ACM Transactions on Design Automation of Electronic Systems, 23(3):30:1–30:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lozano:2023:LBP

[LG23] Erika Susana Alcorta Lozano and Andreas Gerstlauer. Learningbased phase-aware multi-core CPU workload forecasting. ACM Transactions on Design Automation of Electronic Systems, 28(2):23:1–23:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3564929.

Livramento:2014:HTD

[LGGJ14] Vinicius S. Livramento, Chrystian Guth, José Luís Güntzel, and Marcelo O. Johann. A hybrid technique for discrete gate sizing based on Lagrangian relaxation. ACM Transactions on Design Automation of Electronic Systems, 19(4): 40:1–40:??, August 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2009:SCD

[LH09] Yen-Chun Lin and Li-Ling Hung. Straightforward construction of depth-size optimal, parallel prefix circuits with fan-out 2. ACM Transactions on Design Automation of Electronic Systems, 14(1): 15:1–15:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2011:GBP

[LH11] Yifang Liu and Jiang Hu.

GPU-based parallelization for fast circuit optimization. ACM Transactions on Design Automation of Electronic Systems, 16(3):24:1–24:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2013:EMA

[LH13] Yu-Min Lee and Pei-Yu Huang.

An efficient method for analyzing on-chip thermal reliability considering process variations. ACM Transactions on Design Automation of Electronic Systems, 18(3):41:1–41:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2014:DCC

[LH14] Chia-Wei Lee and Sun-Yuan Hsieh. Diagnosability of component-composition graphs in the MM* model. ACM Transactions on Design Automation of Electronic Systems, 19(3):27:1–27:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2016:ODM

[LHC16] Zipeng Li, Tsung-Yi Ho, and Krishnendu Chakrabarty. Optimization of 3D digital microfluidic biochips for the multiplexed polymerase chain reaction. ACM Transactions on Design Automation of Electronic Systems, 21(2):25:1– 25:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2024:MFO

[LHC24] Wanqian Li, Yinhe Han, and Xiaoming Chen. Mathematical framework for optimizing crossbar allocation for ReRAMbased CNN accelerators. ACM Transactions on Design Automation of Electronic Systems, 29(1):21:1–21:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631523.

Liu:2005:ETT

[LHCT05] Xiao Liu, Michael S. Hsiao, Sreejit Chakravarty, and Paul J. Thadikaran. Efficient techniques for transition testing. ACM Transactions on Design Automation of Electronic Systems, 10(2):258–278, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2012:HSC

[LHF12] Hai Lin, Tiansi Hu, and Yunsi Fei. A hardware/software cooperative custom register binding approach for register spill elimination in applicationspecific instruction set processors. ACM Transactions on Design Automation of Electronic Systems, 17(4):40:1–40:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2012:RDP

[LHJ12] Jing-Wei Lin, Tsung-Yi Ho, and Iris Hui-Ru Jiang. Reliabilitydriven power/ground routing for analog ICs. ACM Transactions on Design Automation of Electronic Systems, 17(1): 6:1–6:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2015:DES

[LHK^{+}15] Cheng-Yen Lin, Chung-Wen Huang, Chi-Bang Kuan, Shi-Yu Huang, and Jenq-Kuen Lee. The design and experiments of a SID-based power-aware simulator for embedded multicore systems. ACM Transactions on Design Automation of Electronic Systems, 20(2):22:1– 22:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2016:ESM

[LHLP16] Jinyong Lee, Ingoo Heo, Yongje Lee, and Yunheung Paek. Efficient security monitoring with the core debug interface in an embedded processor. ACM Transactions on Design Automation of Electronic Systems, 22(1):8:1–8:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Last:2020:PMC

[LHS20] Felix Last, Max Haeberlein, and Ulf Schlichtmann. Predicting memory compiler performance outputs using feedforward neural networks. ACM Transactions on Design Automation of Electronic Systems, 25(5):39:1–39:19, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3385262.

Li:2021:MNI

[LHS^{+}21] Jun Li, Bowen Huang, Zhibing Sha, Zhigang Cai, Jianwei Liao, Balazs Gerofi, and Yutaka Ishikawa. Mitigating negative impacts of read disturb in SSDs. ACM Transactions on Design Automation of Electronic Systems, 26(1):3:1– 3:24, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3410332.

Lin:1997:STV

[LHW97] Yann-Rue Lin, Cheng-Tsung Hwang, and Allen C.-H. Wu. Scheduling techniques for variable voltage low power designs. ACM Transactions on Design Automation of Electronic Systems, 2(2):81–97, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p81-lin/p81-lin.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1997-2-2/p81-lin/.

Lee:2017:TPT

[LHW^{+}17] Woojoo Lee, Kyuseung Han, Yanzhi Wang, Tiansong Cui, Shahin Nazarian, and Massoud Pedram. TEI-power: Temperature effect inversion-aware dynamic thermal management. ACM Transactions on Design Automation of Electronic Systems, 22(3):51:1–51:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2006:ETO

[LHZ^{+}06] Zuoyuan Li, Xianlong Hong, Qiang Zhou, Jinian Bian, Hannah H. Yang, and Vijay Pitchumani. Efficient thermaloriented 3D floorplanning and thermal via planning for twostacked-die integration. ACM Transactions on Design Automation of Electronic Systems, 11(2):325–345, April 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Long:2000:FFA

[LIA00] David E. Long, Mahesh A. Iyer, and Miron Abramovici. FILL and FUNI: algorithms to identify illegal states and sequentially untestable faults. ACM Transactions on Design Automation of Electronic Systems, 5(3):631–657, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p631-long/p631-
                                        long.pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        todaes/2000-5-3/p631-long/

.

Liu:2022:AAF

[LIK22] Liu Liu, Sibren Isaacman, and Ulrich Kremer. An adaptive application framework with customizable quality metrics. ACM Transactions on Design Automation of Electronic Systems, 27(2):13:1–13:33, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3477428.

Lin:1997:RDH

[Lin97] Youn-Long Lin. Recent developments in high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 2(1):2–21, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-1/p2-lin/p2-lin.pdf;
http://www.acm.org/pubs/
citations/journals/todaes/
1997-2-1/p2-lin/.

Lee:2018:ICA

[LJ18] Pei-Yu Lee and Iris Hui-Ru Jiang. iTimerM: a compact and accurate timing macro model for efficient hierarchical timing analysis. ACM Transactions on Design Automation of Electronic Systems, 23(4): 48:1–48:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2022:NAD

[LJJ^{+}22] Taozhong Li, Naifeng Jing, Jianfei Jiang, Qin Wang, Zhigang Mao, and Yiran Chen. A novel architecture design for output significance aligned flow with adaptive control in ReRAM-based neural network accelerator. ACM Transactions on Design Automation of Electronic Systems, 27(6): 57:1–57:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3510819.

Liu:2023:GFG

[LJL^{+}23] Yiting Liu, Ziyi Ju, Zhengming Li, Mingzhi Dong, Hai Zhou, Jia Wang, Fan Yang, Xuan Zeng, and Li Shang. Graph-Planner: Floorplanning with graph neural network. ACM Transactions on Design Automation of Electronic Systems, 28(2):21:1–21:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3555804.

Lapinskii:2002:CAH

[LJV02] Viktor S. Lapinskii, Margarida F. Jacome, and Gustavo A. De Veciana. Cluster assignment for high-performance embedded VLIW processors. ACM Transactions on Design Automation of Electronic Systems, 7(3):430–454, July 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2024:DRL

[LJZ^{+}24] Chunlin Li, Kun Jiang, Yong Zhang, Lincheng Jiang, Youlong Luo, and Shaohua Wan. Deep reinforcement learningbased mining task offloading scheme for intelligent connected vehicles in UAV-aided MEC. ACM Transactions on Design Automation of Electronic Systems, 29(3):54:1– 54:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3653451.

Lu:2018:FDR

[LKC^{+}18] Guan-Ruei Lu, Chun-Hao Kuo, Kuen-Cheng Chiang, Ansuman Banerjee, Bhargab B. Bhattacharya, Tsung-Yi Ho, and Hung-Ming Chen. Flexible droplet routing in active matrix-based digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 23(3): 37:1–37:??, April 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2019:QEO

[LKH19] Chun-Han Lin, Chih-Kai Kang, and Pi-Cheng Hsiu. Qualityenhanced OLED power savings on mobile devices. ACM Transactions on Design Automation of Electronic Systems, 24(1): 1:1–1:??, January 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2022:DAA

[LKLC22] Jaechul Lee, Cédric Killian, Sebastien Le Beux, and Daniel Chillet. Distance-aware approximate nanophotonic interconnect. ACM Transactions on Design Automation of Electronic Systems, 27(2):17:1– 17:30, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3484309.

Li:2004:PMA

[LKM04] Hao Li, Srinivas Katkoori, and Wai-Kei Mak. Power minimization algorithms for LUT-based FPGA technology mapping. ACM Transactions on Design Automation of Electronic Systems, 9(1):33–51, January 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liao:1998:NVC

[LKTD98] S. Liao, K. Keutzer, S. Tjiang, and S. Devadas. A new viewpoint on code generation for directed acyclic graphs. ACM Transactions on Design Automation of Electronic Systems, 3(1):51–75, January 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-1/p51-liao/p51-liao.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1998-3-1/p51-liao/.

Lee:2015:SLO

[LL15] Jong Chul Lee and Roman Lysecky. System-level observation framework for nonintrusive runtime monitoring of embedded systems. ACM Transactions on Design Automation of Electronic Systems, 20(3):42:1–42:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2019:DDA

[LL19] Sixing Lu and Roman Lysecky.

Data-driven anomaly detection with timing features for embedded systems. ACM Transactions on Design Automation of Electronic Systems, 24(3):33:1– 33:27, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3279949.

Lee:2017:UCP

[LLH^{+}17] Yongje Lee, Jinyong Lee, Ingoo Heo, Dongil Hwang, and Yunheung Paek. Using Core-Sight PTM to integrate CRA monitoring IPs in an ARMbased SoC. ACM Transactions on Design Automation of Electronic Systems, 22(3): 52:1–52:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2003:COV

[LLHT03] Chingren Lee, Jenq Kuen Lee, Tingting Hwang, and Shi-Chun Tsai. Compiler optimization on VLIW instruction scheduling for low power. ACM Transactions on Design Automation of Electronic Systems, 8(2): 252–268, April 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2012:LBC

[LLHT12] Kuan-Yu Lin, Hong-Ting Lin, Tsung-Yi Ho, and Chia-Chun Tsai. Load-balanced clock tree synthesis with adjustable delay buffer insertion for clock skew reduction in multiple dynamic supply voltage designs. ACM Transactions on Design Automation of Electronic Systems, 17(3):34:1–34:??, June 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lim:2014:PMG

[LLK^{+}14] Jieun Lim, Nagesh B. Lakshminarayana, Hyesoon Kim, William Song, Sudhakar Yalamanchili, and Wonyong Sung. Power modeling for GPU architectures using McPAT. ACM Transactions on Design Automation of Electronic Systems, 19(3):26:1–26:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2013:DRN

[LLKC13] Jinho Lee, Dongwoo Lee, Sunwook Kim, and Kiyoung Choi. Deflection routing in 3D network-on-chip with limited vertical bandwidth. ACM Transactions on Design Automation of Electronic Systems, 18(4):50:1–50:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2013:AVC

[LLKY13] Jaekyu Lee, Si Li, Hyesoon Kim, and Sudhakar Yalamanchili. Adaptive virtual channel partitioning for networkon-chip in heterogeneous architectures. ACM Transactions on Design Automation of Electronic Systems, 18(4):48:1– 48:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2018:UHP

[LLL^{+}18] Wuxi Li, Yibo Lin, Meng Li, Shounak Dhar, and David Z. Pan. UTPlaceF 2.0: a high-performance clock-aware FPGA placement engine. ACM Transactions on Design Automation of Electronic Systems, 23(4):42:1–42:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2013:ABF

[LLLC13] Sean Shih-Ying Liu, Wan-Ting Lo, Chieh-Jui Lee, and Hung-Ming Chen. Agglomerativebased flip-flop merging and relocation for signal wirelength and clock tree optimization. ACM Transactions on Design Automation of Electronic Systems, 18(3):40:1–40:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2018:MRB

[LLLL18] Kuen-Wey Lin, Yeh-Sheng Lin, Yih-Lang Li, and Rung-Bin Lin. A maze routing-based methodology with bounded exploration and path-assessed retracing for constrained multilayer obstacle-avoiding rectilinear Steiner tree construction. ACM Transactions on Design Automation of Electronic Systems, 23(4):45:1–45:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liao:2001:CPT

[LLM01] Swanwa Liao, Mario A. Lopez, and Dinesh Mehta. Constrained polygon transformations for incremental floorplanning. ACM Transactions on Design Automation of Electronic Systems, 6(3):322–342, July 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2023:PDW

[LLM^{+}23] Yaguang Li, Yishuang Lin, Meghna Madhusudan, Arvind Sharma, Sachin Sapatnekar, Ramesh Harjani, and Jiang Hu. Performance-driven wire sizing for analog integrated circuits. ACM Transactions on Design Automation of Electronic Systems, 28(2):19:1–19:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3559542.

Lee:2016:DWC

[LLP^{+}16] Sungkwang Lee, Taemin Lee, Hyunsun Park, Junwhan Ahn, Sungjoo Yoo, Youjip Won, and Sunggu Lee. Differential write-conscious software design on phase-change memory: an SQLite case study. ACM Transactions on Design Automation of Electronic Systems, 21(3):47:1–47:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2003:CLF

[LLQ^{+}03] Zhuo Li, Xiang Lu, Wangqi Qiu, Weiping Shi, and D. M. H. Walker. A circuit level fault model for resistive bridges. ACM Transactions on Design Automation of Electronic Systems, 8(4):546–559, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2023:CPN

[LLQD23] Yanjiang Liu, Junwei Li, Tongzhou Qu, and Zibin Dai. CBDC-PUF: a novel physical unclonable function design framework utilizing configurable butterfly delay chain against modeling attack. ACM Transactions on Design Automation of Electronic Systems, 28(5):78:1–78:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3588435.

Li:2010:CPG

[LLYW10] Zhifang Li, Wenjian Luo, Lihua Yue, and Xufa Wang. On the completeness of the polymorphic gate set. ACM Transactions on Design Automation of Electronic Systems, 15(4):32:1– 32:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lopez:1996:EDP

[LM96] Mario A. Lopez and Dinesh P. Mehta. Efficient decomposition of polygons into L-shapes with application to VLSI layouts. ACM Transactions on Design Automation of Electronic Systems, 1(3):371–395, July 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1996-1-3/p371-lopez/p371-
                                        lopez.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1996-1-3/p371-lopez/

.

Lee:2005:PDD

[LM05] Jaehwan John Lee and Vincent John Mooney III. An o(min(m, n)) parallel deadlock detection algorithm. ACM Transactions on Design Automation of Electronic Systems, 10(3):573–586, July 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2019:TMF

[LM19] Xiangwei Li and Douglas L.

Maskell. Time-multiplexed FPGA overlay architectures: a survey. ACM Transactions on Design Automation of Electronic Systems, 24(5):54:1– 54:19, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3339861.

Lyu:2021:MSC

[LM21] Yangdi Lyu and Prabhat Mishra. MaxSense: Sidechannel sensitivity maximization for Trojan detection using statistical test patterns. ACM Transactions on Design Automation of Electronic Systems, 26(3):22:1–22:21, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3436820.

Liu:2016:DAE

[LMA^{+}16] Qixiao Liu, Miquel Moreto, Jaume Abella, Francisco J. Cazorla, and Mateo Valero. DReAM: an approach to estimate per-task DRAM energy in multicore systems. ACM Transactions on Design Automation of Electronic Systems, 22(1):16:1–16:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lari:2012:HPM

[LMB^{+}12] Vahid Lari, Shravan Muddasani, Srinivas Boppu, Frank Hannig, Moritz Schmid, and Jürgen Teich. Hierarchical power management for adaptive tightly-coupled processor arrays. ACM Transactions on Design Automation of Electronic Systems, 18(1):2:1–2:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lukasiewycz:2016:SAO

[LMS16] Martin Lukasiewycz, Philipp Mundhenk, and Sebastian Steinhorst. Security-aware obfuscated priority assignment for automotive CAN platforms. ACM Transactions on Design Automation of Electronic Systems, 21(2):32:1–32:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:1999:PEE

[LMW99] Yau-Tsun Steven Li, Sharad Malik, and Andrew Wolfe. Performance estimation of embedded software with instruction cache modeling. ACM Transactions on Design Automation of Electronic Systems, 4(3):257–279, July 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-3/p257-li/p257-li.
pdf;   http://www.acm.org/

pubs/citations/journals/
todaes/1999-4-3/p257-li/.

Livramento:2016:CTA

[LNG^{+}16] Vinicius Livramento, Renan Netto, Chrystian Guth, José Luís Güntzel, and Luiz C. V. Dos Santos. Clocktree-aware incremental timingdriven placement. ACM Transactions on Design Automation of Electronic Systems, 21(3): 38:1–38:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2023:EGS

[LNPL23] Yi-Chen Lu, Siddhartha Nath, Sai Pentapati, and Sung Kyu Lim. ECO-GNN: Signoff power prediction using graph neural networks with subgraph approximation. ACM Transactions on Design Automation of Electronic Systems, 28(4): 55:1–55:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3569942.

Linehan:2012:MDA

[LOC12] Éamonn Linehan, Eamonn O’Toole, and Siobhán Clarke. Model-driven automation for simulation-based functional verification. ACM Transactions on Design Automation of Electronic Systems, 17(3): 31:1–31:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2008:PVA

[LON08] Fang Liu, Sule Ozev, and Plamen K. Nikolov. Parametric variability analysis for multistage analog circuits using analytical sensitivity modeling. ACM Transactions on Design Automation of Electronic Systems, 13(2):33:1–33:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2003:ACG

[LP03] J.-Y. Lee and I.-C. Park. Address code generation for DSP instruction-set architectures. ACM Transactions on Design Automation of Electronic Systems, 8(3):384–395, July 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lim:2007:ISI

[LP07] Sung Kyu Lim and Massoud Pedram. Introduction to special issue on demonstrable software systems and hardware platforms. ACM Transactions on Design Automation of Electronic Systems, 12(3): 20:1–20:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Latifis:2017:MVC

[LPD^{+}17] Ioannis Latifis, Karthick Parashar, Grigoris Dimitroulakos, Hans Cappelle, Christakis Lezos, Konstantinos Masselos, and Francky Catthoor. A MAT-LAB vectorizing compiler targeting application-specific instruction set processors. ACM Transactions on Design Automation of Electronic Systems, 22(2):32:1–32:28, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2021:RRD

[LPL^{+}21] Anni Lu, Xiaochen Peng, Yandong Luo, Shanshi Huang, and Shimeng Yu. A runtime reconfigurable design of computein-memory-based hardware accelerator for deep learning inference. ACM Transactions on Design Automation of Electronic Systems, 26(6): 45:1–45:18, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460436.

Lee:2022:ION

[LPLK22] Jooyeon Lee, Junsang Park, Seunghyun Lee, and Jaeha Kung. Implication of optimizing NPU dataflows on neural architecture search for mobile devices. ACM Transactions on Design Automation of Electronic Systems, 27(5): 48:1–48:24, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3513085.

Lalgudi:2000:OCE

[LPP00] Kumar N. Lalgudi, Marios C. Papaefthymiou, and Miodrag Potkonjak. Optimizing computations for effective block-processing. ACM Transactions on Design Automation of Electronic Systems, 5(3):604–630, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p604-lalgudi/p604-
                                        lalgudi.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/2000-5-3/p604-lalgudi/

.

Liu:2020:FGA

[LPY^{+}20] Mengyun Liu, Renjian Pan, Fangming Ye, Xin Li, Krishnendu Chakrabarty, and Xinli Gu. Fine-grained adaptive testing based on quality prediction. ACM Transactions on Design Automation of Electronic Systems, 25(5):38:1–38:25, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3385261.

Liu:2022:LOH

[LQD22] Yanjiang Liu, Tongzhou Qu, and Zibin Dai. A lowoverhead and high-security cryptographic circuit design utilizing the TIGFET-based three-phase single-rail pulse register against side-channel attacks. ACM Transactions on Design Automation of Electronic Systems, 27(4):36:1– 36:13, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3498339.

Liu:2024:HPA

[LQL^{+}24] Hongduo Liu, Yijian Qian, Youqiang Liang, Bin Zhang, Zhaohan Liu, Tao He, Wenqian Zhao, Jiangbo Lu, and Bei Yu. A high-performance accelerator for real-time super-resolution on edge FPGAs. ACM Transactions on Design Automation of Electronic Systems, 29(3): 53:1–53:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3652855.

Lu:2024:GPA

[LRHL24] Yi-Chen Lu, Haoxing Ren, Hao-Hsiang Hsiao, and Sung Kyu Lim. GAN-Place: Advancing open source placers to commercial-quality using generative adversarial networks and transfer learning. ACM Transactions on Design Automation of Electronic Systems, 29(2):32:1–32:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3636461.

Liao:2011:AUB

[LS11] Xiongfei Liao and Thambipillai Srikanthan. Accelerating UNISIM-based cycle-level microarchitectural simulations on multicore platforms. ACM Transactions on Design Automation of Electronic Systems, 16(3):26:1–26:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2017:LPC

[LS17] Tiantao Lu and Ankur Srivastava. Low-power clock tree synthesis for 3D-ICs. ACM Transactions on Design Automation of Electronic Systems, 22(3): 50:1–50:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2019:SIP

[LS19] Tengtao Li and Sachin S. Sapatnekar. Stress-induced performance shifts in 3D DRAMs. ACM Transactions on Design Automation of Electronic Systems, 24(5):51:1–51:21, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3331527.

Li:2022:NSI

[LS22] Bo Li and Guoyong Shi. A native SPICE implementation of memristor models for simulation of neuromorphic analog signal processing circuits.

ACM Transactions on Design Automation of Electronic Systems, 27(1):6:1–6:24, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474364.

Last:2023:TPM

[LS23] Felix Last and Ulf Schlichtmann. Training PPA models for embedded memories on a low-data diet. ACM Transactions on Design Automation of Electronic Systems, 28(2):26:1– 26:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3556539.

Liang:2020:SAE

[LSCK20] Tung-Che Liang, Mohammed Shayan, Krishnendu Chakrabarty, and Ramesh Karri. Secure assay execution on MEDA biochips to thwart attacks using real-time sensing. ACM Transactions on Design Automation of Electronic Systems, 25(2):17:1–17:25, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3374213.

Lee:2010:PTP

[LSDV10] Kyoungwoo Lee, Aviral Shrivastava, Nikil Dutt, and Nalini Venkatasubramanian. Partitioning techniques for partially protected caches in resourceconstrained embedded systems. ACM Transactions on Design Automation of Electronic Systems, 15(4):30:1– 30:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2013:LEV

[LSL^{+}13] Jianhua Li, Liang Shi, Qingan Li, Chun Jason Xue, Yiran Chen, Yinlong Xu, and Wei Wang. Low-energy volatile STT–RAM cache design using cache-coherence-enabled adaptive refresh. ACM Transactions on Design Automation of Electronic Systems, 19(1):5:1– 5:??, December 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2014:CRM

[LSPC14] Jongeun Lee, Seongseok Seo, Jongkyung Paek, and Kiyoung Choi. Configurable range memory for effective data reuse on programmable accelerators. ACM Transactions on Design Automation of Electronic Systems, 19(2):13:1–13:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lysecky:2006:WP

[LSV06] Roman Lysecky, Greg Stitt, and Frank Vahid. Warp Processors. ACM Transactions on Design Automation of Electronic Systems, 11(3):659–681, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2021:VAH

[LSZ^{+}21] Xi Li, Soheil Nazar Shahsavani, Xuan Zhou, Massoud Pedram, and Peter A. Beerel. A variation-aware hold time fixing methodology for single flux quantum logic circuits. ACM Transactions on Design Automation of Electronic Systems, 26(6):47:1–47:17, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460289.

Li:2024:PLC

[LSZ^{+}24] Taixin Li, Boran Sun, Hongtao Zhong, Yixin Xu, Vijaykrishnan Narayanan, Liang Shi, Tianyi Wang, Yao Yu, Thomas Kämpfe, Kai Ni, Huazhong Yang, and Xueqing Li. ProtFe: Low-cost secure power sidechannel protection for general and custom FeFET-based memories. ACM Transactions on Design Automation of Electronic Systems, 29(1):3:1– 3:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3604589.

Lu:2011:CBP

[LT11] Jianchao Lu and Baris Taskin.

Clock buffer polarity assignment with skew tuning. ACM Transactions on Design Automation of Electronic Systems, 16(4):49:1–49:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:1999:BBI

[LTH99] Kuen-Jong Lee, Jing-Jou Tang, and Tsung-Chu Huang. BIFEST: a built-in intermediate fault effect sensing and test generation system for CMOS bridging faults. ACM Transactions on Design Automation of Electronic Systems, 4(2):194–218, April 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-2/p194-lee/p194-lee.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1999-4-2/p194-lee/.

Liu:2013:PBA

[LTPR^{+}13] Xue-Xin Liu, Sheldon X.-D. Tan, Adolfo Adair Palma-Rodriguez, Esteban Tlelo-Cuautle, and Guoyong Shi. Performance bound analysis of analog circuits in frequencyand time-domain considering process variations. ACM Transactions on Design Automation of Electronic Systems, 19(1):6:1–6:??, December 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2010:PAL

[LTPT10] Duo Li, Sheldon X.-D. Tan, Eduardo H. Pacheco, and Murli Tirumala. Parameterized architecture-level dynamic thermal models for multicore microprocessors. ACM Transactions on Design Automation of Electronic Systems, 15(2): 16:1–16:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2016:ECM

[LTW^{+}16] Chuangwen Liu, Peishan Tu, Pangbo Wu, Haomo Tang, Yande Jiang, Jian Kuang, and Evangeline F. Y. Young. An effective chemical mechanical polishing fill insertion approach. ACM Transactions on Design Automation of Electronic Systems, 21(3):54:1– 54:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lam:2012:EPL

[LTYW12] Tak-Kei Lam, Wai-Chung Tang, Xiaoqing Yang, and Yu-Liang Wu. ECR: a powerful and low-complexity error cancellation rewiring scheme. ACM Transactions on Design Automation of Electronic Systems, 17(4):50:1–50:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2022:FIA

[LTZ22] Qiang Liu, Honghui Tang, and Peiran Zhang. Fault injection attack emulation framework for early evaluation of IC designs. ACM Transactions on Design Automation of Electronic Systems, 27(1):8:1–8:25, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3480962.

Lysecky:2002:PIB

[LV02] Roman Lysecky and Frank Vahid. Prefetching for improved bus wrapper performance in cores. ACM Transactions on Design Automation of Electronic Systems, 7(1):58– 90, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2014:SIS

[LV14] Lingyi Liu and Shobha Vasudevan. Scaling input stimulus generation through hybrid static and dynamic analysis of RTL. ACM Transactions on Design Automation of Electronic Systems, 20(1):4:1– 4:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lopez-Vallejo:2003:HSP

[LVL03] Marisa López-Vallejo and Juan Carlos López. On the hardware-software partitioning problem: System modeling and partitioning techniques. ACM Transactions on Design Automation of Electronic Systems, 8(3):269–297, July 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2016:TPD

[LVS16] William Lee, Vikas S. Vij, and Kenneth S. Stevens. Timing path-driven cycle cutting for sequential controllers. ACM Transactions on Design Automation of Electronic Systems, 21(4):64:1–64:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2017:DMF

[LW17] Katherine Shu-Min Li and Sying-Jyan Wang. Design methodology of fault-tolerant custom 3D network-on-chip. ACM Transactions on Design Automation of Electronic Systems, 22(4):63:1–63:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2021:DAS

[LW21] Dave Y.-W. Lin and Charles H.-P. Wen. A delay-adjustable, self-testable flip-flop for softerror tolerability and delayfault testability. ACM Transactions on Design Automation of Electronic Systems, 26(6): 50:1–50:12, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3462171.

Li:2007:SBC

[LWC07] Lei Li, Zhanglei Wang, and Krishnendu Chakrabarty. Scan-BIST based on cluster analysis and the encoding of repeating sequences. ACM Transactions on Design Automation of Electronic Systems, 12(1): 4:1–4:??, January 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2018:CMD

[LWC18] Chen-Hsuan Lin, Lu Wan, and Deming Chen. C-Mine: Data mining of logic common cases for improved timing error resilience with energy efficiency. ACM Transactions on Design Automation of Electronic Systems, 23(2):20:1–20:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lin:2023:SRB

[LWG^{+}23] Wenxiong Lin, Haojie Wu, Peng Gao, Wenjun Luo, Shuting Cai, and Xiaoming Xiong. Sequential routing-based timedivision multiplexing optimization for multi-FPGA systems. ACM Transactions on Design Automation of Electronic Systems, 28(6):104:1–104:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3626322.

Liu:2006:CML

[LWH06] Yi-Yu Liu, Kuo-Hua Wang, and Tingting Hwang. Crosstalk minimization in logic synthesis for PLAs. ACM Transactions on Design Automation of Electronic Systems, 11(4): 890–915, October 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2011:SBA

[LWK11] Yu Liu, Kaijie Wu, and Ramesh Karri. Scan-based attacks on linear feedback shift register based stream ciphers. ACM Transactions on Design Automation of Electronic Systems, 16(2):20:1–20:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lee:2023:VEL

[LWX^{+}23] Jayoung Lee, Pengcheng Wang, Ran Xu, Sarthak Jain, Venkat Dasari, Noah Weston, Yin Li, Saurabh Bagchi, and Somali Chaterji. Virtuoso: Energyand latency-aware streamlining of streaming videos on systems-on-chips. ACM Transactions on Design Automation of Electronic Systems, 28(3): 31:1–31:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3564289.

Li:2019:NRM

[LWZ^{+}19] Taozhong Li, Qin Wang, Yongxin Zhu, Jianfei Jiang, Guanghui He, Jing Jin, Zhigang Mao, and Naifeng Jing. A novel resistive memorybased process-in-memory architecture for efficient logic and add operations. ACM Transactions on Design Automation of Electronic Systems, 24(2): 25:1–25:??, March 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lepak:2004:SSI

[LXCH04] Kevin M. Lepak, Min Xu, Jun Chen, and Lei He. Simultaneous shield insertion and net ordering for capacitive and inductive coupling minimization. ACM Transactions on Design Automation of Electronic Systems, 9(3):290–309, July 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Lu:2023:SPI

[LXGM23] Huaixi Lu, Yue Xing, Aarti Gupta, and Sharad Malik. SoC protocol implementation verification using instructionlevel abstraction specifications. ACM Transactions on Design Automation of Electronic Systems, 28(6):89:1–89:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610292.

Liu:2020:AFD

[LXWC20] Mengyun Liu, Lixue Xia, Yu Wang, and Krishnendu Chakrabarty. Algorithmic fault detection for RRAM-based matrix operations. ACM Transactions on Design Automation of Electronic Systems, 25(3): 29:1–29:31, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3386360.

Luo:2024:HTH

[LY24] Yandong Luo and Shimeng Yu.

H3D-transformer: a heterogeneous 3D (H3D) computing platform for transformer model acceleration on edge devices. ACM Transactions on Design Automation of Electronic Systems, 29(3):47:1–47:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3649219.

Lee:2013:RIB

[LYCP13] Jongwon Lee, Jonghee M.

Youn, Doosan Cho, and Yunheung Paek. Reducing instruction bit-width for low-power VLIW architectures. ACM Transactions on Design Automation of Electronic Systems, 18(2): 25:1–25:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2017:ILA

[LYCP17] Derong Liu, Bei Yu, Salim Chowdhury, and David Z. Pan. Incremental layer assignment for timing optimization. ACM Transactions on Design Automation of Electronic Systems, 22(4):75:1–75:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2014:SUM

[LYHL14] Xueliang Li, Guihai Yan, Yinhe Han, and Xiaowei Li. Smart-Cap: Using machine learning for power adaptation of Smartphone’s application processor. ACM Transactions on Design Automation of Electronic Systems, 20(1):8:1–8:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2009:HPO

[LYKW09] Chih-Hung Liu, Shih-Yi Yuan, Sy-Yen Kuo, and Szu-Chi Wang. High-performance obstacle-avoiding rectilinear Steiner tree construction. ACM Transactions on Design Automation of Electronic Systems, 14(3):45:1–45:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2019:SEA

[LYL^{+}19] Jiajun Li, Guihai Yan, Wenyan Lu, Shijun Gong, Shuhao Jiang, Jingya Wu, and Xiaowei Li. SynergyFlow: an elastic accelerator architecture supporting batch processing of largescale deep neural networks. ACM Transactions on Design Automation of Electronic Systems, 24(1):8:1–8:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Li:2023:CAT

[LYL^{+}23] Meng-Jing Li, Yu-Chuan Yen, Yi-Ting Li, Yung-Chih Chen, and Chun-Yao Wang. A constructive approach for threshold function identification. ACM Transactions on Design Automation of Electronic Systems, 28(5): 86:1–86:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3606371.

Lin:2017:HDP

[LYLW17] Ye-Jyun Lin, Chia-Lin Yang, Hsiang-Pang Li, and Cheng-Yuan Michael Wang. A hybrid DRAM/PCM buffer cache architecture for Smartphones with QoS consideration. ACM Transactions on Design Automation of Electronic Systems, 22(2):27:1–27:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liu:2020:APA

[LYM^{+}20] Kang Liu, Haoyu Yang, Yuzhe Ma, Benjamin Tan, Bei Yu, Evangeline F. Y. Young, Ramesh Karri, and Siddharth Garg. Adversarial perturbation attacks on ML-based CAD: a case study on CNN-based lithographic hotspot detection. ACM Transactions on Design Automation of Electronic Systems, 25(5):48:1–48:31, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3408288.

Li:2019:ATR

[LYSO19] Yanjun Li, Ender Yilmaz, Pete Sarson, and Sule Ozev. Adaptive test for RF/analog circuit using higher order correlations among measurements. ACM Transactions on Design Automation of Electronic Systems, 24(4):45:1–45:16, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3308566.

Li:2017:TSL

[LZ17] Xingquan Li and Wenxing Zhu. Two-stage layout decomposition for hybrid e-beam and triple patterning lithography. ACM Transactions on Design Automation of Electronic Systems, 23(1):6:1–6:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Liao:2021:EPA

[LZ21] Tuotian Liao and Lihong Zhang. Efficient parasiticaware g^{m} /I^{D} — based hybrid sizing methodology for analog and RF integrated circuits. ACM Transactions on Design Automation of Electronic Systems, 26(2):10:1–10:31, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3416946.

Luo:2021:TMF

[LZA^{+}21] Yingyi Luo, Joshua C. Zhao, Arnav Aggarwal, Seda Ogrenci-Memik, and Kazutomo Yoshii. Thermal management for FPGA nodes in HPC systems. ACM Transactions on Design Automation of Electronic Systems, 26(2):14:1–14:17, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3423494.

Lin:2023:ISI

[LZR23] Yibo Lin, Avi Ziv, and Haoxing Ren. Introduction to the special issue on machine learning for CAD/EDA. ACM Transactions on Design Automation of Electronic Systems, 28(2):14:1– 14:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3586208.

Li:2023:MLB

[LZY^{+}23] Benzheng Li, Xi Zhang, Hailong You, Zhongdong Qi, and Yuming Zhang. Machine learning based framework for fast resource estimation of RTL designs targeting FPGAs. ACM Transactions on Design Automation of Electronic Systems, 28(2):24:1–24:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3555047.

Le:2023:PBM

[LZZ23] Trung Le, Zhao Zhang, and Zhichun Zhu. Polling-based memory interface. ACM Transactions on Design Automation of Electronic Systems, 28(3): 47:1–47:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3572919.

Lin:2015:SAD

[LZZSV15] Chung-Wei Lin, Bowen Zheng, Qi Zhu, and Alberto Sangiovanni-Vincentelli. Security-aware design methodology and optimization for automotive systems. ACM Transactions on Design Automation of Electronic Systems, 21(1):18:1– 18:??, November 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mehri:2016:GAB

[MA16] Hossein Mehri and Bijan Alizadeh. Genetic-algorithm-based FPGA architectural exploration using analytical models. ACM Transactions on Design Automation of Electronic Systems, 22(1):13:1– 13:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Maric:2014:HCD

[MACV14] Bojan Maric, Jaume Abella, Francisco J. Cazorla, and Mateo Valero. Hybrid cache designs for reliable hybrid high and ultra-low voltage operation. ACM Transactions on Design Automation of Electronic Systems, 20(1):10:1– 10:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Murali:2023:PSR

[MAL23] Gauthaman Murali, Anthony Agnesina, and Sung Kyu Lim. A PPA study of reinforced placement parameter autotuning: Pseudo-3D vs. true-3D placers. ACM Transactions on Design Automation of Electronic Systems, 28(5): 75:1–75:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582007.

Marwedel:2000:GE

[Mar00] Peter Marwedel. Guest Editorial. ACM Transactions on Design Automation of Electronic Systems, 5(4):749–751, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-4/p749-marwedel/p749-
                                        marwedel.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/2000-5-4/p749-marwedel/

.

Mazumdar:2016:CIS

[MAS16] Bodhisatwa Mazumdar, Sk. Subidh Ali, and Ozgur Sinanoglu. A compact implementation of Salsa20 and its power analysis vulnerabilities. ACM Transactions on Design Automation of Electronic Systems, 22(1):11:1– 11:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mahfouzi:2020:SAR

[MAS^{+}20] Rouhollah Mahfouzi, Amir Aminifar, Soheil Samii, Petru Eles, and Zebo Peng. Securityaware routing and scheduling for control applications on Ethernet TSN networks. ACM Transactions on Design Automation of Electronic Systems, 25(1):1:1–1:26, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.

acm.org/doi/abs/10.1145/
3358604.

Murthy:2004:BMP

[MB04] Praveen K. Murthy and Shuvra S. Bhattacharyya. Buffer merging—a powerful technique for reducing memory requirements of synchronous dataflow specifications. ACM Transactions on Design Automation of Electronic Systems, 9(2):212– 237, April 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mariatos:2001:MAC

[MBB01] E. P. Mariatos, A. N. Birbas, and M. K. Birbas. A mapping algorithm for computerassisted exploration in the design of embedded systems. ACM Transactions on Design Automation of Electronic Systems, 6(1):122–147, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-1/p122-mariatos/p122-
mariatos.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2001-6-1/p122-mariatos/

. See note [CSL^{+}07].

Mandal:2020:EAO

[MBD^{+}20] Sumit K. Mandal, Ganapati Bhat, Janardhan Rao Doppa, Partha Pratim Pande, and Umit Y. Ogras. An energyaware online learning framework for resource management in heterogeneous platforms. ACM Transactions on Design Automation of Electronic Systems, 25(3):28:1– 28:26, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3386359.

Mondal:2012:SEP

[MCD12] Arijit Mondal, P. P. Chakrabarti, and Pallab Dasgupta. Symbolicevent-propagation-based minimal test set generation for robust path delay faults. ACM Transactions on Design Automation of Electronic Systems, 17(4):47:1–47:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Muchherla:2008:NEW

[MCMW08] Kishore Kumar Muchherla, Pinhong Chen, Dongsheng Ma, and Janet Meiling Wang. A noniterative equivalent waveform model for timing analysis in presence of crosstalk. ACM Transactions on Design Automation of Electronic Systems, 13(2):25:1–25:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Monjur:2023:HSR

[MCY23] Mohammad Monjur, Joshua Calzadillas, and Qiaoyan Yu. Hardware security risks and threat analyses in advanced manufacturing industry. ACM Transactions on Design Automation of Electronic Systems, 28(5):83:1–83:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603502.

Mao:2016:LBP

[MCZ^{+}16] Fubing Mao, Yi-Chung Chen, Wei Zhang, Hai (Helen) Li, and Bingsheng He. Library-based placement and routing in FP-GAs with support of partial reconfiguration. ACM Transactions on Design Automation of Electronic Systems, 21(4):71:1– 71:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mishra:2008:SDD

[MD08] Prabhat Mishra and Nikil Dutt. Specification-driven directed test generation for validation of pipelined processors. ACM Transactions on Design Automation of Electronic Systems, 13(3):42:1–42:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Marculescu:2013:ESS

[MD13] Diana Marculescu and Chita Das. Editorial to special section on networks on chip: Architecture, tools, and methodologies. ACM Transactions on Design Automation of Electronic Systems, 18(4):45:1–45:??, October 2013. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-

7309 (electronic).

Mathur:1998:RAE

[MDG98] Anmol Mathur, Ali Dasdan, and Rajesh K. Gupta. Rate analysis for embedded systems. ACM Transactions on Design Automation of Electronic Systems, 3(3):408–436, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1998-3-3/p408-mathur/p408-
                                        mathur.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1998-3-3/p408-mathur/

.

Maslov:2007:TSR

[MDM07] D. Maslov, G. W. Dueck, and D. M. Miller. Techniques for the synthesis of reversible Toffoli networks. ACM Transactions on Design Automation of Electronic Systems, 12(4):42:1– 42:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mukherjee:2012:SAA

[MDM^{+}12] Subhankar Mukherjee, Pallab Dasgupta, Siddhartha Mukhopadhyay, Scott Little, John Havlicek, and Srikanth Chandrasekaran. Synchronizing AMS assertions with AMS simulation: From theory to practice. ACM Transactions on Design Automation of Electronic Systems, 17(4): 38:1–38:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mirtar:2015:AAA

[MDR15] Ali Mirtar, Sujit Dey, and Anand Raghunathan. An application adaptation approach to mitigate the impact of dynamic thermal management on video encoding. ACM Transactions on Design Automation of Electronic Systems, 20(4):50:1– 50:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mahmoud:2023:LEP

[MED23] Mervat M. A. Mahmoud, Nahla E. Elashkar, and Heba H. Draz. Low-energy pipelined hardware design for approximate medium filter. ACM Transactions on Design Automation of Electronic Systems, 28(3):42:1–42:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582005.

Mehta:1998:ESR

[Meh98] Dinesh P. Mehta. Estimating the storage requirements of the rectangular and L-shaped corner stitching data structures. ACM Transactions on Design Automation of Electronic Systems, 3(2):272–284, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p272-mehta/p272-
mehta.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-2/p272-mehta/

.

Milder:2012:CGH

[MFHP12] Peter Milder, Franz Franchetti, James C. Hoe, and Markus Püschel. Computer generation of hardware for linear digital signal processing transforms. ACM Transactions on Design Automation of Electronic Systems, 17(2):15:1–15:??, April 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Morgado:2009:GRS

[MFS09] P. Marques Morgado, Paulo F. Flores, and L. Miguel Silveira. Generating realistic stimuli for accurate power grid analysis. ACM Transactions on Design Automation of Electronic Systems, 14(3):40:1–40:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mitra:2015:OWS

[MGR^{+}15] Debasis Mitra, Sarmishtha Ghoshal, Hafizur Rahaman, Krishnendu Chakrabarty, and Bhargab B. Bhattacharya. Offline washing schemes for residue removal in digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 21(1):17:1–17:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Metwalli:2019:SAS

[MHA19] Sara Ayman Metwalli and Yuko Hara-Azumi. SSA-AC: Static significance analysis for approximate computing. ACM Transactions on Design Automation of Electronic Systems, 24(3):34:1–34:17, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3314575.

Ma:2004:SCU

[MHD^{+}04] Yuchun Ma, Xianlong Hong, Sheqin Dong, Yici Cai, Chung-Kuan Cheng, and Jun Gu. Stairway compaction using corner block list and its applications with rectilinear blocks. ACM Transactions on Design Automation of Electronic Systems, 9(2):199–211, April 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Moreno:1996:REU

[MHF96] R. Moreno, R. Hermida, and M. Fernández. Register estimation in unscheduled dataflow graphs. ACM Transactions on Design Automation of Electronic Systems, 1(3):396–403, July 1996. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1996-1-3/p396-moreno/p396-
                                        moreno.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1996-1-3/p396-moreno/

.

Mochocki:2007:TOA

[MHQ07] Bren Mochocki, Xiaobo Sharon Hu, and Gang Quan. Transitionoverhead-aware voltage scheduling for fixed-priority real-time systems. ACM Transactions on Design Automation of Electronic Systems, 12(2):11:1– 11:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Meyer:2014:CEL

[MHT14] Brett H. Meyer, Adam S. Hartman, and Donald E. Thomas. Cost-effective lifetime and yield optimization for NoC-based MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 19(2):12:1– 12:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mittal:2016:STC

[Mit16] Sparsh Mittal. A survey of techniques for cache locking. ACM Transactions on Design Automation of Electronic Systems, 21(3):49:1–49:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Monteiro:2019:OCF

[MJB19] Jucemar Monteiro, Marcelo Johann, and Laleh Behjat. An optimized cost flow algorithm to spread cells in detailed placement. ACM Transactions on Design Automation of Electronic Systems, 24(3):35:1– 35:16, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3317575.

Mittal:2011:TVA

[MJM11] Kartikey Mittal, Arpit Joshi, and Madhu Mutyam. Timing variation-aware scheduling and resource binding in highlevel synthesis. ACM Transactions on Design Automation of Electronic Systems, 16(4): 40:1–40:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Memik:2005:SAO

[MKBS05] Seda Ogrenci Memik, Ryan Kastner, Elaheh Bozorgzadeh, and Majid Sarrafzadeh. A scheduling algorithm for optimization and early planning in high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 10(1):33–57, January 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Majzoobi:2013:LPR

[MKK13] Mehrdad Majzoobi, Joonho Kong, and Farinaz Koushanfar. Low-power resource binding by postsilicon customization. ACM Transactions on Design Automation of Electronic Systems, 18(2):26:1–26:??, March 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Moiseev:2008:TAP

[MKW08] Konstantin Moiseev, Avinoam Kolodny, and Shmuel Wimer. Timing-aware power-optimal ordering of signals. ACM Transactions on Design Automation of Electronic Systems, 13(4):65:1–65:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Moiseev:2009:PDO

[MKW09] Konstantin Moiseev, Avinoam Kolodny, and Shmuel Wimer. Power-delay optimization in VLSI microprocessors by wire spacing. ACM Transactions on Design Automation of Electronic Systems, 14(4):55:1– 55:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mu:2009:AHS

[ML09] Jingqing Mu and Roman Lysecky. Autonomous hardware/software partitioning and voltage/frequency scaling for low-power embedded systems. ACM Transactions on Design Automation of Electronic Systems, 15(1):2:1–2:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Moscola:2008:RCB

[MLC08] James Moscola, John W.

Lockwood, and Young H. Cho. Reconfigurable contentbased router using hardwareaccelerated language parser. ACM Transactions on Design Automation of Electronic Systems, 13(2):28:1–28:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mok:2012:DSL

[MLG12] Santiago Mok, John Lee, and Puneet Gupta. Discrete sizing for leakage power optimization in physical design: a comparative study. ACM Transactions on Design Automation of Electronic Systems, 18(1):15:1– 15:??, December 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Moon:2017:ASP

[MLH^{+}17] Hyungon Moon, Jinyong Lee, Dongil Hwang, Seonhwa Jung, Jiwon Seo, and Yunheung Paek. Architectural supports to protect OS kernels from code-injection attacks and their applications. ACM Transactions on Design Automation of Electronic Systems, 23(1): 10:1–10:??, October 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mukherjee:2008:HLC

[MLMM08] Rajarshi Mukherjee, Song Liu, Seda Ogrenci Memik, and Somsubhra Mondal. A high-level clustering algorithm targeting dual V_{dd} FPGAs. ACM Transactions on Design Automation of Electronic Systems, 13(4): 57:1–57:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mahalat:2022:ICA

[MMM^{+}22] Mahabub Hasan Mahalat, Suraj Mandal, Anindan Mondal, Bibhash Sen, and Rajat Subhra Chakraborty. Implementation, characterization and application of path changing switch based arbiter PUF on FPGA as a lightweight security primitive for IoT. ACM Transactions on Design Automation of Electronic Systems, 27(3):26:1–26:26, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491212.

Marculescu:2000:SSM

[MMP00] Diana Marculescu, Radu Marculescu, and Massoud Pedram. Stochastic sequential machine synthesis with application to constrained sequence generation. ACM Transactions on Design Automation of Electronic Systems, 5(3):658–681, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p658-marculescu/
p658-marculescu.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/2000-5-3/
p658-marculescu/.

Moudallal:2017:GCC

[MN17] Zahi Moudallal and Farid N.

Najm. Generating current constraints to guarantee RLC power grid safety. ACM Transactions on Design Automation of Electronic Systems, 22(4): 66:1–66:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Maleki:2021:EEI

[MNMK^{+}21] Mohammad-Ali Maleki, Alireza Nabipour-Meybodi, Mehdi Kamal, Ali Afzali-Kusha, and Massoud Pedram. An energyefficient inference method in convolutional neural networks based on dynamic adjustment of the pruning level. ACM Transactions on Design Automation of Electronic Systems, 26(6):49:1–49:20, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460972.

Marculescu:2006:CCR

[MOZ06] Radu Marculescu, Umit Y. Ogras, and Nicholas H. Zamora. Computation and communication refinement for multiprocessor SoC design: a systemlevel perspective. ACM Transactions on Design Automation of Electronic Systems, 11(3): 564–592, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mohanty:2007:MBE

[MP07] Sumit Mohanty and Viktor K. Prasanna. A model-based extensible framework for efficient application design using FPGA. ACM Transactions on Design Automation of Electronic Systems, 12(2): 13:1–13:??, April 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mukhopadhyay:2009:IAA

[MPDG09] Rajdeep Mukhopadhyay, S. K. Panda, Pallab Dasgupta, and John Gough. Instrumenting AMS assertion verification on commercial platforms. ACM Transactions on Design Automation of Electronic Systems, 14(2):21:1–21:??, March 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mundhenk:2017:SAN

[MPM^{+}17] Philipp Mundhenk, Andrew Paverd, Artur Mrowca, Sebastian Steinhorst, Martin Lukasiewycz, Suhaib A. Fahmy, and Samarjit Chakraborty. Security in automotive networks: Lightweight authentication and authorization. ACM Transactions on Design Automation of Electronic Systems, 22(2): 25:1–25:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mathaikutty:2007:EMD

[MPSJ07] Deepak Mathaikutty, Hiren Patel, Sandeep Shukla, and Axel Jantsch. EWD: a metamodeling driven customizable multi-MoC system modeling framework. ACM Transactions on Design Automation of Electronic Systems, 12(3): 33:1–33:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Middelhoek:1996:VEF

[MR96] Peter F. A. Middelhoek and Sreeranga P. Rajan. From VHDL to efficient and firsttime-right designs: a formal approach. ACM Transactions on Design Automation of Electronic Systems, 1(2):205–250, April 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-2/p205-middelhoek/
p205-middelhoek.pdf;   http:
//www.acm.org/pubs/citations/

                                        journals/todaes/1996-1-2/
                                        p205-middelhoek/.

Mohanty:2005:EED

[MR05] Saraju P. Mohanty and N. Ranganathan. Energy-efficient datapath scheduling using multiple voltages and dynamic clocking. ACM Transactions on Design Automation of Electronic Systems, 10(2):330–353, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Maestro:2011:MEL

[MRB^{+}11] Juan Antonio Maestro, Pedro Reviriego, Sanghyeon Baeg, Shijie Wen, and Richard Wong. Mitigating the effects of large multiple cell upsets (MCUs) in memories. ACM Transactions on Design Automation of Electronic Systems, 16(4): 45:1–45:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mohanty:2006:IMS

[MRC06] Saraju P. Mohanty, N. Ranganathan, and Sunil K. Chappidi. ILP models for simultaneous energy and transient power minimization during behavioral synthesis. ACM Transactions on Design Automation of Electronic Systems, 11(1):186–212, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Muhammad:2019:RBS

[MRL^{+}19] Shaheer Muhammad, M. Usman Rafique, Shuai Li, Zili Shao, Qixin Wang, and Xue Liu. Reconfigurable battery systems: a survey on hardware architecture and research challenges. ACM Transactions on Design Automation of Electronic Systems, 24(2):19:1– 19:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
ft_gateway.cfm?id=3301301.

Malekpour:2020:HTM

[MRL^{+}20] Amin Malekpour, Roshan Ragel, Tuo Li, Haris Javaid, Aleksandar Ignjatovic, and Sri Parameswaran. Hardware Trojan mitigation in pipelined MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 25(1):6:1– 6:27, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365578.

Moffitt:2008:CDF

[MRMP08] Michael D. Moffitt, Jarrod A.

Roy, Igor L. Markov, and Martha E. Pollack. Constraintdriven floorplan repair. ACM Transactions on Design Automation of Electronic Systems, 13(4):67:1–67:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mehta:2000:UFR

[MS00] Dinesh P. Mehta and Naveed Sherwani. On the use of flexible, rectilinear blocks to obtain minimum-area floorplans in mixed block and cell designs. ACM Transactions on Design Automation of Electronic Systems, 5(1):82–97, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-1/p82-mehta/p82-mehta.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/2000-5-1/p82-mehta/

.

Manolios:2008:AVS

[MS08] Panagiotis Manolios and Sudarshan K. Srinivasan. Automatic verification of safety and liveness for pipelined machines using WEB refinement. ACM Transactions on Design Automation of Electronic Systems, 13(3):45:1–45:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Murugesan:2017:NRM

[MS17] Shanmugakumar Murugesan and Noor Mahammad Sk. A novel range matching architecture for packet classification without rule expansion. ACM Transactions on Design Automation of Electronic Systems, 23(1):8:1–8:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mondal:2021:IFS

[MS21] Ankit Mondal and Ankur Srivastava. Ising-FPGA: a spintronics-based reconfigurable Ising model solver. ACM Transactions on Design Automation of Electronic Systems, 26(1):4:1–4:27, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3411511.

Minakova:2023:MTT

[MS23] Svetlana Minakova and Todor Stefanov. Memory-throughput trade-off for CNN-based applications at the edge. ACM Transactions on Design Automation of Electronic Systems, 28(1):2:1–2:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3527457.

Mehta:2009:ICH

[MSB^{+}09] Gayatri Mehta, Justin Stander, Mustafa Baz, Brady Hunsaker, and Alex K. Jones. Interconnect customization for a hardware fabric. ACM Transactions on Design Automation of Electronic Systems, 14(1):11:1– 11:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mishra:2006:ADL

[MSD06] Prabhat Mishra, Aviral Shrivastava, and Nikil Dutt. Architecture description language (ADL)-driven software toolkit generation for architectural exploration of programmable SOCs. ACM Transactions on Design Automation of Electronic Systems, 11(3):626–658, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Majumder:2007:HPV

[MSKBD07] Subhashis Majumder, Susmita Sur-Kolay, Bhargab B. Bhattacharya, and Swarup Kumar Das. Hierarchical partitioning of VLSI floorplans by staircases. ACM Transactions on Design Automation of Electronic Systems, 12(1): 7:1–7:??, January 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Malik:2009:SCU

[MSR09] Avinash Malik, Zoran Salcic, and Partha S. Roop. SystemJ compilation using the Tandem Virtual Machine approach. ACM Transactions on Design Automation of Electronic Systems, 14(3):34:1–34:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Michael:2002:ATD

[MT02] M. Michael and S. Tragoudas. ATPG tools for delay faults at the functional level. ACM Transactions on Design Automation of Electronic Systems, 7(1):33–57, January 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

More:2015:LAN

[MT15] Ankit More and Baris Taskin.

Locality-aware network utilization balancing in NoCs. ACM Transactions on Design Automation of Electronic Systems, 21(1):6:1–6:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mutyam:2009:SST

[Mut09] Madhu Mutyam. Selective shielding technique to eliminate crosstalk transitions. ACM Transactions on Design Automation of Electronic Systems, 14(3):43:1–43:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Muztoba:2018:IAI

[MVK^{+}18] Md Muztoba, Rohit Voleti, Fatih Karabacak, Jaehyun Park, and Umit Y. Ogras. Instinctive assistive indoor navigation using distributed intelligence. ACM Transactions on Design Automation of Electronic Systems, 23(6):80:1– 80:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Mak:1997:BLM

[MW97] Wai-Kei Mak and D. F. Wong. Board-level multiterminal net routing for FPGAbased logic emulation. ACM Transactions on Design Automation of Electronic Systems, 2(2):151–167, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p151-mak/p151-mak.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1997-2-2/p151-mak/.

Munch:1997:EIB

[MWG97] Michael Münch, Norbert Wehn, and Manfred Glesner. An efficient ILP-based scheduling algorithm for control-dominated VHDL descriptions. ACM Transactions on Design Automation of Electronic Systems, 2(4):344–364, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-4/p344-munch/p344-
munch.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-4/p344-munch/

.

Mohammadzadeh:2021:EOP

[MWK21] Naser Mohammadzadeh, Robert Wille, and Oliver Keszocze. Efficient one-pass synthesis for digital microfluidic biochips.

ACM Transactions on Design Automation of Electronic Systems, 26(4):27:1–27:21, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446880.

Ma:2020:MEF

[MWS^{+}20] Chenlin Ma, Yi Wang, Zhaoyan Shen, Renhai Chen, Zhu Wang, and Zili Shao. MNFTL: an efficient flash translation layer for MLC NAND flash memory. ACM Transactions on Design Automation of Electronic Systems, 25(6):50:1–50:19, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3398037.

Markov:2023:GEI

[MYSZ23] Igor Markov, Fan Yang, Li Shang, and Hai Zhou. Guest Editor’s introduction: Machine learning for VLSI physical design. ACM Transactions on Design Automation of Electronic Systems, 28(4): 48:1–48:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3592606.

Nath:2020:RDB

[NAK20] Arijit Nath, Sukarn Agarwal, and Hemangee K. Kapoor. Reuse distance-based victim cache for effective utilisation of hybrid main memory system. ACM Transactions on Design Automation of Electronic Systems, 25(3):24:1– 24:32, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3380732.

Nourani:2001:ITI

[NCP01] Mehrdad Nourani, Joan Carletta, and Christos Papachristou. Integrated test of interacting controllers and datapaths. ACM Transactions on Design Automation of Electronic Systems, 6(3):401–422, July 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Narang:2023:DPM

[NDA^{+}23] Gaurav Narang, Aryan Deshwal, Raid Ayoub, Michael Kishinevsky, Janardhan Rao Doppa, and Partha Pratim Pande. Dynamic power management in large manycore systems: a learning-to-search framework. ACM Transactions on Design Automation of Electronic Systems, 28(5): 84:1–84:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3603501.

Neuberger:2003:MBU

[NdLCR03] Gustavo Neuberger, Fernanda de Lima, Luigi Carro, and Ricardo Reis. A multiple bit upset tolerant SRAM memory. ACM Transactions on Design Automation of Electronic Systems, 8(4):577–590, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nacul:2006:STC

[NG06] André C. Nácul and Tony Givargis. Synthesis of timeconstrained multitasking embedded software. ACM Transactions on Design Automation of Electronic Systems, 11(4): 822–847, October 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ning:2021:FND

[NGL^{+}21] Xuefei Ning, Guangjun Ge, Wenshuo Li, Zhenhua Zhu, Yin Zheng, Xiaoming Chen, Zhen Gao, Yu Wang, and Huazhong Yang. FTT-NAS: Discovering fault-tolerant convolutional neural architecture. ACM Transactions on Design Automation of Electronic Systems, 26(6):44:1–44:24, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460288.

Naseer:2023:QGA

[NHS23] Mahum Naseer, Osman Hasan, and Muhammad Shafique. QuanDA: GPU accelerated quantitative deep neural network analysis. ACM Transactions on Design Automation of Electronic Systems, 28 (6):95:1–95:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611671.

NS:2024:POA

[NL24] Karthik Somayaji NS and Peng Li. Pareto optimization of analog circuits using reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 29(2):37:1– 37:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3640463.

Nadakuditi:2013:BAS

[NM13] Raj Rao Nadakuditi and Igor L. Markov. On bottleneck analysis in stochastic stream processing. ACM Transactions on Design Automation of Electronic Systems, 18(3): 34:1–34:??, July 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nie:2023:CMD

[NM23] Qi Nie and Sharad Malik. CN-NFlow: Memory-driven data flow optimization for convolutional neural networks. ACM Transactions on Design Automation of Electronic Systems, 28(3):40:1–40:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3577017.

Nahiyan:2020:SCF

[NPH^{+}20] Adib Nahiyan, Jungmin Park, Miao He, Yousef Iskander, Farimah Farahmandi, Domenic Forte, and Mark Tehranipoor. SCRIPT: a CAD framework for power side-channel vulnerability assessment using information flow tracking and pattern generation. ACM Transactions on Design Automation of Electronic Systems, 25(3): 26:1–26:27, May 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3383445.

Narasimhan:2001:FAC

[NR01] M. Narasimhan and J. Ramanujam. A fast approach to computing exact solutions to the resource-constrained scheduling problem. ACM Transactions on Design Automation of Electronic Systems, 6(4):490–500, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Niggemeyer:2003:DAM

[NR03] Dirk Niggemeyer and Elizabeth M. Rudnick. A data acquisition methodology for onchip repair of embedded memories. ACM Transactions on Design Automation of Electronic Systems, 8(4):560–576, October 2003. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic).

Nongpoh:2019:ESE

[NRDB19] Bernard Nongpoh, Rajarshi Ray, Moumita Das, and Ansuman Banerjee. Enhancing speculative execution with selective approximate computing. ACM Transactions on Design Automation of Electronic Systems, 24(2):26:1– 26:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3307651.

Niu:2024:ECS

[NRM^{+}24] Linwei Niu, Danda B. Rawat, Jonathan Musselwhite, Zonghua Gu, and Qingxu Deng. Energyconstrained scheduling for weakly hard real-time systems using standby-sparing. ACM Transactions on Design Automation of Electronic Systems, 29(2):29:1–29:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3631587.

Noltsis:2018:RSC

[NRZ^{+}18] Michail Noltsis, Dimitrios Rodopoulos, Nikolaos Zompakis, Francky Catthoor, and Dimitrios Soudris. Runtime slack creation for processor performance variability using system scenarios. ACM Transactions on Design Automation of Electronic Systems, 23(2): 24:1–24:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nummer:2003:THP

[NS03] Muhammad Nummer and Manoj Sachdev. Testing highperformance pipelined circuits with slow-speed testers. ACM Transactions on Design Automation of Electronic Systems, 8(4):506–521, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nguyen:2017:SAA

[NSCM17] Phuong Ha Nguyen, Durga Prasad Sahoo, Rajat Subhra Chakraborty, and Debdeep Mukhopadhyay. Security analysis of arbiter PUF and its lightweight compositions under predictability test. ACM Transactions on Design Automation of Electronic Systems, 22(2):20:1– 20:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nair:2016:ESP

[NSH^{+}16] Piyoosh Purushothaman Nair, Arnab Sarkar, N. M. Harsha, Megha Gandhi, P. P. Chakrabarti, and Sujoy Ghose. ERfair scheduler with processor suspension for real-time multiprocessor embedded systems. ACM Transactions on Design Automation of Electronic Systems, 22(1):19:1– 19:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nasser:2020:NCM

[NSP^{+}20] Yehya Nasser, Carlo Sau, Jean-Christophe Prévotet, Tiziana Fanni, Francesca Palumbo, Maryline Hélard, and Luigi Raffo. NeuPow: a CAD methodology for high-level power estimation based on machine learning. ACM Transactions on Design Automation of Electronic Systems, 25(5):41:1– 41:29, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3388141.

Narayanaswamy:2016:BRE

[NSS^{+}16] Swaminathan Narayanaswamy, Steffen Schlueter, Sebastian Steinhorst, Martin Lukasiewycz, Samarjit Chakraborty, and Harry Ernst Hoster. On battery recovery effect in wireless sensor nodes. ACM Transactions on Design Automation of Electronic Systems, 21(4):60:1– 60:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Nourani:2005:RHE

[NT05] Mehrdad Nourani and Mohammad H. Tehranipour. RL-Huffman encoding for test compression and power reduction in scan applications. ACM Transactions on Design Automation of Electronic Systems, 10(1): 91–115, January 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Naderan-Tahan:2018:DCE

[NTSA18] Mahmood Naderan-Tahan and Hamid Sarbazi-Azad. Domino Cache: an energy-efficient data cache for modern applications. ACM Transactions on Design Automation of Electronic Systems, 23(3):31:1–31:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ni:2024:ISI

[NWA^{+}24] Tianming Ni, Xiaoqing Wen, Hussam Amrouch, Cheng Zhuo, and Peilin Song. Introduction to the special issue on design for testability and reliability of security-aware hardware. ACM Transactions on Design Automation of Electronic Systems, 29(1):1:1– 1:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631476.

Oh:2019:TAS

[OCK19] Deok Keun Oh, Mu Jun Choi, and Ju Ho Kim. Thermalaware 3D symmetrical buffered clock tree synthesis. ACM Transactions on Design Automation of Electronic Systems, 24(3):28:1–28:22, June 2019.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3313798.

Ostler:2007:IHT

[OCRS07] Chris Ostler, Karam S. Chatha, Vijay Ramamurthi, and Krishnan Srinivasan. ILP and heuristic techniques for system-level design on network processor architectures. ACM Transactions on Design Automation of Electronic Systems, 12(4):48:1– 48:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Osawa:2019:ADR

[OHA19] Hisashi Osawa and Yuko Hara-Azumi. Approximate data reuse-based accelerator design for embedded processor. ACM Transactions on Design Automation of Electronic Systems, 24(5):56:1–56:25, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3342098.

Ozturk:2008:IBE

[OK08] Ozcan Ozturk and Mahmut Kandemir. ILP-based energy minimization techniques for banked memories. ACM Transactions on Design Automation of Electronic Systems, 13(3):50:1–50:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Olney:2020:TFB

[OK20] Brooks Olney and Robert Karam. Tunable FPGA bitstream obfuscation with Boolean satisfiability attack countermeasure. ACM Transactions on Design Automation of Electronic Systems, 25(2): 19:1–19:22, March 2020. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3373638.

Ozturk:2008:APB

[OKC08] Ozcan Ozturk, Mahmut Kandemir, and Guangyu Chen. Access pattern-based code compression for memory-constrained systems. ACM Transactions on Design Automation of Electronic Systems, 13(4):60:1– 60:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Oldja:2022:HSS

[OKJH22] Mari-Liis Oldja, Jangryul Kim, Dowhan Jeong, and Soonhoi Ha. Hierarchical scheduling of an SDF/L graph onto multiple processors. ACM Transactions on Design Automation of Electronic Systems, 27(3): 22:1–22:23, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3489469.

Ogras:2008:AOP

[OM08] Umit Y. Ogras and Radu Marculescu. Analysis and optimization of prediction-based flow control in networks-on-chip. ACM Transactions on Design Automation of Electronic Systems, 13(1):11:1–11:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ochoa-Ruiz:2015:MAR

[ORGD^{+}15] Gilberto Ochoa-Ruiz, Sébastien Guillet, Florent De Lamotte, Eric Rutten, El-Bay Bourennane, Jean-Philippe Diguet, and Guy Gogniat. An MDE approach for rapid prototyping and implementation of dynamic reconfigurable systems. ACM Transactions on Design Automation of Electronic Systems, 21(1):8:1–8:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Obenaus:2003:GFP

[OS03] Stefan Thomas Obenaus and Ted H. Szymanski. Gravity: Fast placement for 3-D VLSI. ACM Transactions on Design Automation of Electronic Systems, 8(3):298–315, July 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Oboril:2015:EIS

[OT15] Fabian Oboril and Mehdi B. Tahoori. Exploiting instruction set encoding for aging-aware microprocessor design. ACM Transactions on Design Automation of Electronic Systems, 21(1):5:1–5:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ozdal:2006:TLB

[OW06] Muhammet Mustafa Ozdal and Martin D. F. Wong. Twolayer bus routing for high-speed printed circuit boards. ACM Transactions on Design Automation of Electronic Systems, 11(1):213–227, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ozdal:2008:ORA

[OWH08] Muhammet Mustafa Ozdal, Martin D. F. Wong, and Philip S. Honsinger. Optimal routing algorithms for rectilinear pin clusters in high-density multichip modules. ACM Transactions on Design Automation of Electronic Systems, 13(4):68:1–68:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2021:LDH

[PA21] Irith Pomeranz and M. Enamul Amyeen. Logic diagnosis with hybrid fail data. ACM Transactions on Design Automation of Electronic Systems, 26(3):19:1–19:13, February 2021. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3433929.

Pomeranz:2017:TMR

[PAV17] Irith Pomeranz, M. Enamul Amyeen, and Srikanth Venkataraman. Test modification for reduced volumes of fail data. ACM Transactions on Design Automation of Electronic Systems, 22(4): 67:1–67:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pan:2012:ERE

[PB12] Zhaoliang Pan and Melvin A. Breuer. Error rate estimation for defective circuits via ones counting. ACM Transactions on Design Automation of Electronic Systems, 17(1): 8:1–8:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Panerati:2014:CEM

[PB14] Jacopo Panerati and Giovanni Beltrame. A comparative evaluation of multi-objective exploration algorithms for highlevel design. ACM Transactions on Design Automation of Electronic Systems, 19(2): 15:1–15:??, March 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Poddar:2022:DDM

[PBF^{+}22] Sudip Poddar, Sukanta Bhattacharjee, Shao-Yun Fang, Tsung-Yi Ho, and B. B. Bhattacharya. Demand-driven multi-target sample preparation on resource-constrained digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 27(1):7:1–7:21, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3474392.

Peng:2024:CTD

[PBH^{+}24] Qingsong Peng, Jingchang Bian, Zhengfeng Huang, Senling Wang, and Aibin Yan. A compact TRNG design for FPGA based on the metastability of RO-driven shift registers. ACM Transactions on Design Automation of Electronic Systems, 29(1):13:1–13:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610295.

Park:2017:HHC

[PBL^{+}17] Jaehyun Park, Seungcheol Baek, Hyung Gyu Lee, Chrysostomos Nicopoulos, Vinson Young, Junghee Lee, and Jongman Kim. HoPE: Hot-cacheline prediction for dynamic early decompression in compressed LLCs. ACM Transactions on Design Automation of Electronic Systems, 22(3):40:1– 40:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pinto:2006:SLD

[PBSV^{+}06] Alessandro Pinto, Alvise Bonivento, Allberto L. Sangiovanni-Vincentelli, Roberto Passerone, and Marco Sgroi. System level design paradigms: Platform-based design and communication synthesis. ACM Transactions on Design Automation of Electronic Systems, 11(3):537–563, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Poddar:2021:RMT

[PBWB21] Sudip Poddar, Tapalina Banerjee, Robert Wille, and Bhargab B. Bhattacharya. Robust multi-target sample preparation on MEDA biochips obviating waste production. ACM Transactions on Design Automation of Electronic Systems, 26(1):7:1–7:29, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3414061.

Pourshirazi:2019:WAL

[PBZM19] Bahareh Pourshirazi, Majed Valad Beigi, Zhichun Zhu, and Gokhan Memik. Writeback-aware LLC management for PCM-Based main memory systems. ACM Transactions on Design Automation of Electronic Systems, 24 (2):18:1–18:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3292009.

Palkovic:2009:TOL

[PCC09] Martin Palkovic, Francky Catthoor, and Henk Corporaal. Trade-offs in loop transformations. ACM Transactions on Design Automation of Electronic Systems, 14(2): 22:1–22:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Panda:2001:DMO

[PCD^{+}01] P. R. Panda, F. Catthoor, N. D. Dutt, K. Danckaert, E. Brockmeyer, C. Kulkarni, A. Vandercappelle, and P. G. Kjeldsberg. Data and memory optimization techniques for embedded systems. ACM Transactions on Design Automation of Electronic Systems, 6(2):149–206, January 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-2/p149-panda/p149-
panda.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2001-6-2/p149-panda/

.

Peng:2017:LSA

[PCT^{+}17] Yin-Chi Peng, Chien-Chih Chen, Hsiang-Jen Tsai, Keng-Hao Yang, Pei-Zhe Huang, Shih-Chieh Chang, Wen-Ben Jone, and Tien-Fu Chen. Leak Stopper: an actively revitalized snoop filter architecture with effective generation control. ACM Transactions on Design Automation of Electronic Systems, 22(3):46:1–46:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Panda:1997:MDO

[PDN97] Preeti Ranjan Panda, Nikil D. Dutt, and Alexandru Nicolau. Memory data organization for improved cache performance in embedded processor applications. ACM Transactions on Design Automation of Electronic Systems, 2(4):384–409, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-4/p384-panda/p384-
panda.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-4/p384-panda/

.

Panda:2000:CVC

[PDN00] Preeti Ranjan Panda, Nikil D. Dutt, and Alexandru Nicolau. On-chip vs. off-chip memory: the data partitioning problem in embedded processor-based systems. ACM Transactions on Design Automation of Electronic Systems, 5(3):682–704, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p682-panda/p682-
panda.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p682-panda/

.

Pasha:2012:SLS

[PDS12] Muhammad Adeel Pasha, Steven Derrien, and Olivier Sentieys. System-level synthesis for wireless sensor node controllers: a complete design flow. ACM Transactions on Design Automation of Electronic Systems, 17(1):2:1–2:??, January 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pedram:1996:PMI

[Ped96]  Massoud Pedram.  Power

minimization in IC design: principles and applications. ACM Transactions on Design Automation of Electronic Systems, 1(1):3–56, January 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-1/p3-pedram/p3-pedram.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1996-1-1/p3-pedram/

.

Pedram:2006:ISI

[Ped06] Massoud Pedram. Introduction to special issue: Novel paradigms in system-level design. ACM Transactions on Design Automation of Electronic Systems, 11(3):535–536, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pedram:2008:E

[Ped08]  Massoud Pedram.  Edito-

rial. ACM Transactions on Design Automation of Electronic Systems, 13(4):55:1– 55:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pedram:2011:CPV

[Ped11] Massoud Pedram. Call for papers: Verification issue and challenges with multicore systems. ACM Transactions on Design Automation of Electronic Systems, 16(2):12:1– 12:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pop:2006:AOD

[PEPP06] Paul Pop, Petru Eles, Zebo Peng, and Traian Pop. Analysis and optimization of distributed real-time embedded systems. ACM Transactions on Design Automation of Electronic Systems, 11(3):593–625, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pilato:2022:ISS

[PFHAH22] Christian Pilato, Zhenman Fang, Yuko Hara-Azumi, and Jim Hwang. Introduction to the special section on highlevel synthesis for FPGA: Nextgeneration technologies and applications. ACM Transactions on Design Automation of Electronic Systems, 27(4): 29:1–29:2, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3519279.

Peter:2015:CBS

[PG15] Steffen Peter and Tony Givargis. Component-based synthesis of embedded systems using satisfiability modulo theories. ACM Transactions on Design Automation of Electronic Systems, 20(4):49:1– 49:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Parulkar:2001:IRC

[PGB01] Ishwar Parulkar, Sandeep K.

Gupta, and Melvin A. Breuer. Introducing redundant computations in RTL data paths for reducing BIST resources. ACM Transactions on Design Automation of Electronic Systems, 6(3):423–445, July 2001. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Poddar:2016:ECS

[PGCB16] Sudip Poddar, Sarmishtha Ghoshal, Krishnendu Chakrabarty, and Bhargab B. Bhattacharya. Error-correcting sample preparation with cyberphysical digital microfluidic lab-on-chip. ACM Transactions on Design Automation of Electronic Systems, 22(1):2:1–2:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pereira:2023:IED

[PGGD23] Danny Pereira, Anirban Ghose, Sumana Ghosh, and Soumyajit Dey. Inferencing on edge devices: a time- and spaceaware co-scheduling approach. ACM Transactions on Design Automation of Electronic Systems, 28(3):38:1–38:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3576197.

Peng:2012:SSE

[PHKW12] Huan-Kai Peng, Hsuan-Ming Huang, Yu-Hsin Kuo, and Charles H.-P. Wen. Statistical soft error rate (SSER) analysis for scaled CMOS designs. ACM Transactions on Design Automation of Electronic Systems, 17(1):9:1–9:??, January 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pees:2000:RCS

[PHM00] Stefan Pees, Andreas Hoffmann, and Heinrich Meyr. Retargetable compiled simulation of embedded processors using a machine description language. ACM Transactions on Design Automation of Electronic Systems, 5(4):815–834, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-4/p815-pees/p815-
pees.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/2000-5-4/p815-pees/

.

Pierre:2016:AVT

[Pie16] Laurence Pierre. Auxiliary variables in temporal specifications: Semantic and practical analysis for system-level requirements. ACM Transactions on Design Automation of Electronic Systems, 21(2):20:1– 20:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pendyala:2020:IAS

[PIK20] Shilpa Pendyala, Sheikh Ariful Islam, and Srinivas Katkoori. Interval arithmetic and selfsimilarity based RTL input vector control for datapath leakage minimization. ACM Transactions on Design Automation of Electronic Systems, 25(6):54:1– 54:26, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3408061.

Pan:2014:SPM

[PJL14] Gung-Yu Pan, Jing-Yang Jou, and Bo-Cheng Lai. Scalable power management using multilevel reinforcement learning for multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 19(4):33:1–33:??, August 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Park:2021:PPD

[PKC^{+}21] Heechun Park, Bon Woong Ku, Kyungwook Chang, Da Eun Shim, and Sung Kyu Lim. Pseudo-3D physical design flow for monolithic 3D ICs: Comparisons and enhancements. ACM Transactions on Design Automation of Electronic Systems, 26(5):37:1–37:25, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3453480.

Paik:2020:GRT

[PKJK20] Yoonah Paik, Seon Wook Kim, Dongha Jung, and Minseong Kim. Generating representative test sequences from real workload for minimizing DRAM verification overhead. ACM Transactions on Design Automation of Electronic Systems, 25(4):

30:1–30:23, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3391891.

Parthasarathy:2003:PTA

[PKP^{+}03] Kumar Parthasarathy, Turker Kuyel, Dana Price, Le Jin, Degang Chen, and Randall Geiger. BIST and production testing of ADCs using imprecise stimulus. ACM Transactions on Design Automation of Electronic Systems, 8(4):522– 545, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pan:1998:OCP

[PL98] Peichen Pan and C. L. Liu. Optimal clock period FPGA technology mapping for sequential circuits. ACM Transactions on Design Automation of Electronic Systems, 3(3):437–462, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-3/p437-pan/p437-pan.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1998-3-3/p437-pan/.

Pinar:2003:CSI

[PL03] Ali Pinar and C. L. Liu.

Compacting sequences with invariant transition frequencies. ACM Transactions on Design Automation of Electronic Systems, 8(2):214–221, April 2003.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pan:2024:RCA

[PLC24] Renjian Pan, Xin Li, and Krishnendu Chakrabarty. Rootcause analysis with semisupervised co-training for integrated systems. ACM Transactions on Design Automation of Electronic Systems, 29(3): 51:1–51:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3649313.

Pan:2024:CEP

[PLH^{+}24] Zijin Pan, Xunyu Li, Weiquan Hao, Runyu Miao, and Albert Wang. On-chip ESD protection design methodologies by CAD simulation. ACM Transactions on Design Automation of Electronic Systems, 29(1):4:1– 4:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3593808.

Prasad:2024:SSE

[PMA24] Govind Prasad, Bipin Mandi, and Maifuz Ali. SEDONUT: a single event double node upset tolerant SRAM for terrestrial applications. ACM Transactions on Design Automation of Electronic Systems, 29(3): 52:1–52:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3651985.

Paul:2010:LOC

[PMB10] Somnath Paul, Hamid Mahmoodi, and Swarup Bhunia. Low-overhead Fmax calibration at multiple operating points using delay-sensitivity-based path selection. ACM Transactions on Design Automation of Electronic Systems, 15(2):19:1–19:??, February 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pagliari:2017:AEE

[PMP17] Daniele Jahier Pagliari, Enrico Macii, and Massimo Poncino. Approximate energyefficient encoding for serial interfaces. ACM Transactions on Design Automation of Electronic Systems, 22(4): 64:1–64:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Park:2015:SGA

[PMS15] Heejong Park, Avinash Malik, and Zoran Salcic. Scheduling globally asynchronous locally synchronous programs for guaranteed response times. ACM Transactions on Design Automation of Electronic Systems, 20(3):40:1–40:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Parane:2020:LDL

[PMT20] Khyamling Parane, Prabhu Prasad B. M., and Basavaraj Talawar. LBNoC: Design of low-latency router architecture with lookahead bypass for network-on-chip using FPGA. ACM Transactions on Design Automation of Electronic Systems, 25(1):9:1– 9:26, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3365994.

Pomeranz:2013:BGM

[Pom13] Irith Pomeranz. Built-in generation of multicycle functional broadside tests with observation points. ACM Transactions on Design Automation of Electronic Systems, 19(1):8:1– 8:??, December 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2014:DTM

[Pom14a]  Irith Pomeranz.  Design-

for-testability for multi-cycle broadside tests by holding of state variables. ACM Transactions on Design Automation of Electronic Systems, 19(2): 19:1–19:??, March 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2014:LPS

[Pom14b] Irith Pomeranz. Low-power skewed-load tests based on functional broadside tests. ACM Transactions on Design Automation of Electronic Systems, 19(2):18:1–18:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2015:ETC

[Pom15a] Irith Pomeranz. Enhanced test compaction for multicycle broadside tests by using state complementation. ACM Transactions on Design Automation of Electronic Systems, 21 (1):13:1–13:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2015:FES

[Pom15b] Irith Pomeranz. FOLD: Extreme static test compaction by folding of functional test sequences. ACM Transactions on Design Automation of Electronic Systems, 20(4):57:1– 57:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2015:GDU

[Pom15c] Irith Pomeranz. A generalized definition of unnecessary test vectors in functional test sequences. ACM Transactions on Design Automation of Electronic Systems, 20(2): 29:1–29:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2016:DTF

[Pom16a] Irith Pomeranz. Design-for-testability for functional broadside tests under primary input constraints. ACM Transactions on Design Automation of Electronic Systems, 21(2): 35:1–35:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2016:DTS

[Pom16b] Irith Pomeranz. N-detection test sets for circuits with multiple independent scan chains. ACM Transactions on Design Automation of Electronic Systems, 21(4):68:1– 68:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2016:PSS

[Pom16c]  Irith Pomeranz.  Periodic

scan-in states to reduce the input test data volume for partially functional broadside tests. ACM Transactions on Design Automation of Electronic Systems, 22(1):7:1–7:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2017:CSL

[Pom17a] Irith Pomeranz. Computation of seeds for LFSR-based ndetection test generation. ACM Transactions on Design Automation of Electronic Systems, 22(2):29:1–29:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2017:GTS

[Pom17b] Irith Pomeranz. Generation of transparent-scan sequences for diagnosis of scan chain faults. ACM Transactions on Design Automation of Electronic Systems, 22(3):43:1–43:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2018:DDP

[Pom18a]  Irith Pomeranz.  Dynami-

cally determined preferred values and a design-for-testability approach for multiplexer select inputs under functional test sequences. ACM Transactions on Design Automation of Electronic Systems, 23(5): 59:1–59:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2018:PIP

[Pom18b] Irith Pomeranz. Partially invariant patterns for LFSRbased generation of closeto-functional broadside tests. ACM Transactions on Design Automation of Electronic Systems, 23(4):53:1–53:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2019:BFB

[Pom19a] Irith Pomeranz. Boundaryfunctional broadside and skewedload tests. ACM Transactions on Design Automation of Electronic Systems, 24(1): 7:1–7:??, January 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2019:ITU

[Pom19b] Irith Pomeranz. Incomplete tests for undetectable faults to improve test set quality. ACM Transactions on Design Automation of Electronic Systems, 24(2):23:1– 23:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3306493.

Pomeranz:2020:TFT

[Pom20] Irith Pomeranz. Target faults for test compaction based on multicycle tests. ACM Transactions on Design Automation of Electronic Systems, 25(2): 18:1–18:14, January 2020. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3375278.

Pomeranz:2021:CTH

[Pom21a]  Irith Pomeranz.  Covering

test holes of functional broadside tests. ACM Transactions on Design Automation of Electronic Systems, 26(3):23:1– 23:15, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3441282.

Pomeranz:2021:EFU

[Pom21b] Irith Pomeranz. Equivalent faults under launch-on-shift (LOS) tests with equal primary input vectors. ACM Transactions on Design Automation of Electronic Systems, 26(4):25:1– 25:15, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3440013.

Pomeranz:2022:IFC

[Pom22] Irith Pomeranz. Increasing the fault coverage of a truncated test set. ACM Transactions on Design Automation of Electronic Systems, 27 (6):54:1–54:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3508459.

Pomeranz:2024:RCS

[Pom24a] Irith Pomeranz. Reduced onchip storage of seeds for builtin test generation. ACM Transactions on Design Automation of Electronic Systems, 29(3): 45:1–45:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3643810.

Pomeranz:2024:TDS

[Pom24b] Irith Pomeranz. Two-dimensional search space for extracting broadside tests from functional test sequences. ACM Transactions on Design Automation of Electronic Systems, 29(3): 48:1–48:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3650207.

Pasricha:2009:SLP

[PPDK09] Sudeep Pasricha, Young-Hwan Park, Nikil Dutt, and Fadi J. Kurdahi. System-level PVT variation-aware power exploration of on-chip communication architectures. ACM Transactions on Design Automation of Electronic Systems, 14(2): 20:1–20:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Papandreou:2015:ERM

[PPP^{+}15] Nikolaos Papandreou, Thomas Parnell, Haralampos Pozidis, Thomas Mittelholzer, Evangelos Eleftheriou, Charles Camp, Thomas Griffin, Gary Tressler, and Andrew Walls. Enhancing the reliability of MLC NAND flash memory systems by read channel optimization. ACM Transactions on Design Automation of Electronic Systems, 20(4):62:1–62:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Prasad:1996:TRP

[PR96] S. C. Prasad and K. Roy.

Transistor reordering for power minimization under delay constraint. ACM Transactions on Design Automation of Electronic Systems, 1(2):280–300, April 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-2/p280-prasad/p280-
prasad.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-2/p280-prasad/

.

Pomeranz:1998:FTG

[PR98] Irith Pomeranz and Sudhakar M. Reddy. Functional test generation for delay faults in combinational circuits. ACM Transactions on Design Automation of Electronic Systems, 3(2):231–248, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-2/p231-pomeranz/p231-
pomeranz.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-2/p231-pomeranz/

.

Pomeranz:2007:FDT

[PR07] Irith Pomeranz and Sudhakar M. Reddy. Forming N-detection test sets without test generation. ACM Transactions on Design Automation of Electronic Systems, 12(2): 18:1–18:??, April 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2009:UST

[PR09] Irith Pomeranz and Sudhakar M. Reddy. Using stuckat tests to form scan-based tests for transition faults in standard-scan circuits. ACM Transactions on Design Automation of Electronic Systems, 15(1):7:1–7:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pomeranz:2011:RSA

[PR11] Irith Pomeranz and Sudhakar M. Reddy. Reducing the switching activity of test sequences under transparentscan. ACM Transactions on Design Automation of Electronic Systems, 16(2):17:1– 17:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Panda:2008:SBV

[PRCK08] S. K. Panda, Arnab Roy, P. P. Chakrabarti, and Rajeev Kumar. Simulation-based verification using Temporally Attributed Boolean Logic. ACM Transactions on Design Automation of Electronic Systems, 13(4):63:1–63:??, September 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Park:2021:HTN

[PRKK21] Naebeom Park, Sungju Ryu, Jaeha Kung, and Jae-Joon Kim. High-throughput nearmemory processing on CNNs with 3D HBM-like memory. ACM Transactions on Design Automation of Electronic Systems, 26(6):48:1–48:20, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460971.

Praveen:2023:DER

[PS23] Pushkar Praveen and R. K.

Singh. Design of enhanced reversible 9T SRAM design for the reduction in sub-threshold leakage current with14nm Fin-FET technology. ACM Transactions on Design Automation of Electronic Systems, 28(6):105:1–105:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616538.

Palchaudhuri:2021:DAT

[PSD21] Ayan Palchaudhuri, Sandeep Sharma, and Anindya Sundar Dhar. Design automation for tree-based nearest neighborhood-aware placement of high-speed cellular automata on FPGA with scan path insertion. ACM Transactions on Design Automation of Electronic Systems, 26(4):31:1– 31:34, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3446206.

Pecenka:2008:ESR

[PSK08] Tomas Pecenka, Lukas Sekanina, and Zdenek Kotasek. Evolution of synthetic RTL benchmark circuits with predefined testability. ACM Transactions on Design Automation of Electronic Systems, 13(3): 54:1–54:??, July 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Passerone:1998:MRS

[PSL^{+}98] C. Passerone, C. Sansoe, L. Lavagno, R. McGeer, J. Martin, R. Passerone, and A. Sangiovanni-Vincentelli. Modeling reactive systems in Java. ACM Transactions on Design Automation of Electronic Systems, 3(4):515–523, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL http://www.acm.org/
                                        pubs/articles/journals/todaes/
                                        1998-3-4/p515-passerone/
                                        p515-passerone.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/todaes/1998-3-4/
                                        p515-passerone/.

Pereira-Santos:2018:RFB

[PSNC18] Leonardo Pereira-Santos, Gabriel Luca Nazar, and Luigi Carro. Repair of FPGA-based real-time systems with variable slacks.

ACM Transactions on Design Automation of Electronic Systems, 23(2):19:1–19:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pandey:2024:NDT

[PSP24] Shailja Pandey, Lokesh Siddhu, and Preeti Ranjan Panda. NeuroCool: Dynamic thermal management of 3D DRAM for deep neural networks through customized prefetching. ACM Transactions on Design Automation of Electronic Systems, 29(1):19:1–19:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3630012.

Padmanaban:2006:IGM

[PT06] Saravanan Padmanaban and Spyros Tragoudas. Implicit grading of multiple path delay faults. ACM Transactions on Design Automation of Electronic Systems, 11(2):346–361, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Paul:2005:HLM

[PTC05] Joann M. Paul, Donald E.

Thomas, and Andrew S. Cassidy. High-level modeling and simulation of single-chip programmable heterogeneous multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 10(3):

431–461, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Potluri:2015:DAT

[PTC^{+}15] Seetal Potluri, A. Satya Trinadh, Sobhan Babu Ch., V. Kamakoti, and Nitin Chandrachoodan. DFT assisted techniques for peak launch-to-capture power reduction during launch-on-shift at-speed testing. ACM Transactions on Design Automation of Electronic Systems, 21(1):14:1– 14:??, November 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Polychronou:2022:CSA

[PTPB22] Nikolaos-Foivos Polychronou, Pierre-Henri Thevenon, Maxime Puys, and Vincent Beroulle. A comprehensive survey of attacks without physical access targeting hardware vulnerabilities in IoT/IIoT devices, and their detection mechanisms. ACM Transactions on Design Automation of Electronic Systems, 27(1):1:1–1:35, January 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3471936.

Pan:2020:ARP

[PTS^{+}20] Renjian Pan, Jun Tao, Yangfeng Su, Dian Zhou, Xuan Zeng, and Xin Li. Analog/RF postsilicon tuning via Bayesian optimization. ACM Transactions on Design Automation of Electronic Systems, 25(1):7:1– 7:17, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365577.

Potkonjak:1999:MAD

[PW99] Miodrag Potkonjak and Wayne Wolf. A methodology and algorithms for the design of hard real-time multitasking ASICs. ACM Transactions on Design Automation of Electronic Systems, 4(4):430–459, October 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-4/p430-potkonjak/
p430-potkonjak.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/1999-4-4/
p430-potkonjak/.

Poon:2005:DPM

[PWY05] Kara K. W. Poon, Steven J. E. Wilton, and Andy Yan. A detailed power model for field-programmable gate arrays. ACM Transactions on Design Automation of Electronic Systems, 10(2):279–302, April 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Pui:2020:LRB

[PY20] Chak-Wa Pui and Evangeline F. Y. Young. Lagrangian relaxation-based time-division multiplexing optimization for multi-FPGA systems. ACM Transactions on Design Automation of Electronic Systems, 25(2):21:1–21:23, March 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3377551.

Qian:2016:PEN

[QBTM16] Zhiliang Qian, Paul Bogdan, Chi-Ying Tsui, and Radu Marculescu. Performance evaluation of NoC-based multicore systems: From traffic analysis to NoC latency modeling. ACM Transactions on Design Automation of Electronic Systems, 21(3):52:1–52:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qin:2012:DTG

[QM12] Xiaoke Qin and Prabhat Mishra. Directed test generation for validation of multicore architectures. ACM Transactions on Design Automation of Electronic Systems, 17(3): 24:1–24:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qiu:2009:CMW

[QS09] Meikang Qiu and Edwin H.-M. Sha. Cost minimization while satisfying hard/soft timing constraints for heterogeneous embedded systems. ACM Transactions on Design Automation of Electronic Systems, 14(2): 25:1–25:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qiu:2011:ATB

[QS11] Meikang Qiu and Edwin H.- M. Sha. 2011 ACM TODAES best paper award. ACM Transactions on Design Automation of Electronic Systems, 16(4): 36:1–36:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qian:2012:FPS

[QSK12] Haifeng Qian, Sachin S. Sapatnekar, and Eren Kursun. Fast Poisson Solvers for thermal analysis. ACM Transactions on Design Automation of Electronic Systems, 17(3): 32:1–32:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qin:2015:CSE

[QSW^{+}15] Ying Qin, Shengyu Shen, Qingbo Wu, Huadong Dai, and Yan Jia. Complementary synthesis for encoder with flow control mechanism. ACM Transactions on Design Automation of Electronic Systems, 21(1):12:1–12:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Qian:2024:ERL

[QZZW24] Yu Qian, Xuegong Zhou, Hao Zhou, and Lingli Wang. An efficient reinforcement learning based framework for exploring logic synthesis. ACM Transactions on Design Automation of Electronic Systems, 29(2):25:1– 25:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3632174.

Rakhmatov:2009:BVM

[Rak09] Daler Rakhmatov. Battery voltage modeling for portable systems. ACM Transactions on Design Automation of Electronic Systems, 14(2):29:1– 29:??, March 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rodrigues:2012:IPP

[RAKK12] Rance Rodrigues, Arunachalam Annamalai, Israel Koren, and Sandip Kundu. Improving performance per watt of asymmetric multi-core processors via online program phase classification and adaptive core morphing. ACM Transactions on Design Automation of Electronic Systems, 18(1):5:1– 5:??, December 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Raval:2019:III

[RB19] Rajkumar K. Raval and Atta Badii. Investigating the impact of image content on the energy efficiency of hardwareaccelerated digital spatial filters. ACM Transactions on Design Automation of Electronic Systems, 24(5):57:1– 57:34, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3341819.

Roy:2021:FVS

[RB21] Pushpita Roy and Ansuman Banerjee. A framework for validation of synthesized MicroElectrode dot array actuations for digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 26(6): 46:1–46:36, November 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3460437.

Ray:2012:ISS

[RBA^{+}12] Sandip Ray, Jayanta Bhadra, Magdy S. Abadir, Li-C. Wang, and Aarti Gupta. Introduction to special section on verification challenges in the concurrent world. ACM Transactions on Design Automation of Electronic Systems, 17(3): 19:1–19:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Roy:2020:HGM

[RBWB20] Pushpita Roy, Ansuman Banerjee, Robert Wille, and Bhargab B. Bhattacharya. Harnessing the granularity of microelectrode-dot-array architectures for optimizing droplet routing in biochips. ACM Transactions on Design Automation of Electronic Systems, 25(1):10:1–10:37, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3365993.

Roy:2007:EPA

[RCD07] Suchismita Roy, P. P. Chakrabarti, and Pallab Dasgupta. Event propagation for accurate circuit delay calculation using SAT. ACM Transactions on Design Automation of Electronic Systems, 12(3):36:1–36:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Resano:2008:ESR

[RCG^{+}08] Javier Resano, Juan Antonio Clemente, Carlos Gonzalez, Daniel Mozos, and Francky Catthoor. Efficiently scheduling runtime reconfigurations. ACM Transactions on Design Automation of Electronic Systems, 13(4):58:1– 58:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Roy:2015:LAM

[RCK^{+}15] Sudip Roy, Partha P. Chakrabarti, Srijan Kumar, Krishnendu Chakrabarty, and Bhargab B. Bhattacharya. Layout-aware mixture preparation of biochemical fluids on applicationspecific digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 20(3):45:1– 45:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ramanathan:2022:CPF

[RCW22] Nadesh Ramanathan, George A.

Constantinides, and John Wickerson. A case for precise, fine-grained pointer synthesis in high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 27(4):30:1–30:26, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491430.

Raval:2010:LPT

[RFB10] R. K. Raval, C. H. Fernandez, and C. J. Bleakley. Lowpower TinyOS tuned processor platform for wireless sensor network motes. ACM Transactions on Design Automation of Electronic Systems, 15(3): 23:1–23:??, May 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Richthammer:2020:SSD

[RFG20] Valentina Richthammer, Fabian Fassnacht, and Michael Glaß.

Search-space decomposition for system-level design space exploration of embedded systems. ACM Transactions on Design Automation of Electronic Systems, 25(2):14:1– 14:32, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3369388.

Rajan:1998:ASD

[RFYL98] S. P. Rajan, M. Fujita, K. Yuan, and M. T-C. Lee. ATM switch design by highlevel modeling, formal verification and high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 3(4):554–562, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p554-rajan/p554-
rajan.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-4/p554-rajan/

.

Rokni:2019:SLF

[RG19] Seyed Ali Rokni and Hassan Ghasemzadeh. Share-n-Learn: a framework for sharing activity recognition models in wearable systems with contextvarying sensors. ACM Transactions on Design Automation of Electronic Systems, 24(4): 39:1–39:27, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3318044.

Ranganathan:2009:VAM

[RGM09] Nagarajan Ranganathan, Upavan Gupta, and Venkataraman Mahalingam. Variation-aware multimetric optimization during gate sizing. ACM Transactions on Design Automation of Electronic Systems, 14(4): 54:1–54:??, August 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Roy:2015:FTE

[RGM15] Debashri Roy, Prasun Ghosal, and Saraju Mohanty. FuzzRoute: a thermally efficient congestion-free global routing method for three-dimensional integrated circuits. ACM Transactions on Design Automation of Electronic Systems, 21(1):1:1–1:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rosales:2014:MHA

[RGT^{+}14] Rafael Rosales, Michael Glass, Jürgen Teich, Bo Wang, Yang Xu, and Ralph Hasholzner. MAESTRO — holistic actororiented modeling of nonfunctional properties and firmware behavior for MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 19(3):23:1–23:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rahman:2024:SES

[RGX^{+}24] Md Moshiur Rahman, Jim Geist, Daniel Xing, Yuntao Liu, Ankur Srivastava, Travis Meade, Yier Jin, and Swarup Bhunia. Security evaluation of state space obfuscation of hardware IP through a red team–blue team practice. ACM Transactions on Design Automation of Electronic Systems, 29(3):50:1–50:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3640461.

Raabe:2008:RDS

[RHA08] Andreas Raabe, Philipp A. Hartmann, and Joachim K. Anlauf. ReChannel: Describing and simulating reconfigurable hardware in systemC. ACM Transactions on Design Automation of Electronic Systems, 13(1):15:1–15:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Raimi:2000:EML

[RHN00] Richard Raimi, Ramin Hojati, and Kedar S. Namjoshi. Environment modeling and language universality. ACM Transactions on Design Automation of Electronic Systems, 5(3):705–725, January 2000. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-3/p705-raimi/p705-
raimi.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2000-5-3/p705-raimi/

.

Ravi:2014:HLT

[RJ14] Srivaths Ravi and Michael Joseph. High-level test synthesis: a survey from synthesis process flow perspective. ACM Transactions on Design Automation of Electronic Systems, 19(4):38:1–38:??, August 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rao:2009:COT

[RJBS09] Rajeev R. Rao, Vivek Joshi, David Blaauw, and Dennis Sylvester. Circuit optimization techniques to mitigate the effects of soft errors in combinational logic. ACM Transactions on Design Automation of Electronic Systems, 15(1): 5:1–5:??, December 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Raghavan:2009:PTG

[RJL^{+}09] Praveen Raghavan, Murali Jayapala, Andy Lambrechts, Javed Absar, and Francky Catthoor. Playing the trade-off game: Architecture exploration using Coffeee. ACM Transactions on Design Automation of Electronic Systems, 14(3): 36:1–36:??, May 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rapp:2024:NAI

[RKKH24] Martin Rapp, Heba Khdr, Nikita Krohmer, and Jörg Henkel. NPU-accelerated imitation learning for thermal optimization of QoS-constrained heterogeneous multi-cores. ACM Transactions on Design Automation of Electronic Systems, 29(1):16:1–16:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3626320.

Ramanujam:2013:DBC

[RL13] Rohit Sunkam Ramanujam and Bill Lin. Destination-based congestion awareness for adaptive routing in 2D mesh networks. ACM Transactions on Design Automation of Electronic Systems, 18(4):60:1– 60:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Reviriego:2009:EED

[RM09] Pedro Reviriego and Juan Antonio Maestro. Efficient error detection codes for multiplebit upset correction in SRAMs with BICS. ACM Transactions on Design Automation of Electronic Systems, 14(1):18:1– 18:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rawat:2023:RSB

[RM23a] Bhawna Rawat and Poornima Mittal. A reconfigurable 7T SRAM bit cell for high speed, power saving and low voltage application. ACM Transactions on Design Automation of Electronic Systems, 28(6):96:1–96:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616872.

Rawat:2023:SNB

[RM23b] Bhawna Rawat and Poornima Mittal. A switching NMOS based single ended sense amplifier for high density SRAM applications. ACM Transactions on Design Automation of Electronic Systems, 28(3): 37:1–37:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3576198.

Reviriego:2010:RAM

[RMB10] Pedro Reviriego, Juan Antonio Maestro, and Chris J. Bleakley. Reliability analysis of memories protected with BICS and a perword parity bit. ACM Transactions on Design Automation of Electronic Systems, 15(2): 18:1–18:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Reddy:2003:TDV

[RMKP03] Sudhakar M. Reddy, Kohei Miyase, Seiji Kajihara, and Irith Pomeranz. On test data volume reduction for multiple scan chain designs. ACM Transactions on Design Automation of Electronic Systems, 8(4):460–469, October 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rahaman:2008:CTB

[RMPJ08] H. Rahaman, J. Mathew, D. K. Pradhan, and A. M. Jabir. Ctestable bit parallel multipliers over GF(2^{m}). ACM Transactions on Design Automation of Electronic Systems, 13(1): 5:1–5:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rokni:2021:TMS

[RNA^{+}21] Seyed Ali Rokni, Marjan Nourollahi, Parastoo Alinia, Iman Mirzadeh, Mahdi Pedram, and Hassan Ghasemzadeh. TransNet: Minimally supervised deep transfer learning for dynamic adaptation of wearable systems. ACM Transactions on Design Automation of Electronic Systems, 26(1):5:1– 5:31, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3414062.

Rahman:2021:SAD

[RNR^{+}21] M. Sazadur Rahman, Adib Nahiyan, Fahim Rahman, Saverio Fazzari, Kenneth Plaks, Farimah Farahmandi, Domenic Forte, and Mark Tehranipoor. Security assessment of dynamically obfuscated scan chain against oracleguided attacks. ACM Transactions on Design Automation of Electronic Systems, 26(4):29:1– 29:27, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3444960.

Roy:2005:FSV

[RPKC05] Arnab Roy, S. K. Panda, Rajeev Kumar, and P. P. Chakrabarti. A framework for systematic validation and debugging of pipeline simulators. ACM Transactions on Design Automation of Electronic Systems, 10(3):462–491, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Roy:2021:MLS

[RPR^{+}21] Urmimala Roy, Tanmoy Pramanik, Subhendu Roy, Avhishek Chatterjee, Leonard F. Register, and Sanjay K. Banerjee. Machine learning for statistical modeling: The case of perpendicular spin-transfer-torque random access memory.

ACM Transactions on Design Automation of Electronic Systems, 26(3):24:1–24:17, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3440014.

Roy:2021:FAA

[RRHB21] Indrani Roy, Chester Rebeiro, Aritra Hazra, and Swarup Bhunia. FaultDroid: an algorithmic approach for fault-induced information leakage analysis. ACM Transactions on Design Automation of Electronic Systems, 26(1):2:1–2:27, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3410336.

Riepe:1998:EBD

[RS98] Michael A. Riepe and Karem A. Sakallah. The edge-based design rule model revisited. ACM Transactions on Design Automation of Electronic Systems, 3(3):463–486, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-3/p463-riepe/p463-
riepe.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1998-3-3/p463-riepe/

.

Riepe:2003:TPN

[RS03] Michael A. Riepe and Karem A.

Sakallah. Transistor placement for noncomplementary digital VLSI cell synthesis. ACM Transactions on Design Automation of Electronic Systems, 8(1):81–107, January 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rosvall:2018:FTA

[RS18] Kathrin Rosvall and Ingo Sander. Flexible and tradeoffaware constraint-based design space exploration for streaming applications on heterogeneous platforms. ACM Transactions on Design Automation of Electronic Systems, 23(2): 21:1–21:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Roop:2001:FST

[RSR01] Partha S. Roop, A. Sowmya, and S. Ramesh. Forced simulation: a technique for automating component reuse in embedded systems. ACM Transactions on Design Automation of Electronic Systems, 6(4):602– 628, October 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Ruan:2005:BEL

[RTNL05] Shanq-Jang Ruan, Kun-Lin Tsai, Edwin Naroska, and Feipei Lai. Bipartitioning and encoding in low-power pipelined circuits. ACM Transactions on Design Automation of Electronic Systems, 10(1): 24–32, January 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Rawat:2003:I

[RW03] Shishpal Rawat and Hans-Joachim Wunderlich. Introduction. ACM Transactions on Design Automation of Electronic Systems, 8(4): 397–398, October 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Saglican:2024:MDV

[SA24] Enes Sağlican and Engin Afacan. MOEA/D vs. NSGA-II: a comprehensive comparison for multi/many objective analog/RF circuit optimization through a generic benchmark. ACM Transactions on Design Automation of Electronic Systems, 29(1):15:1–15:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3626096.

Samavatian:2015:ALL

[SABSA15] Mohammad Hossein Samavatian, Mohammad Arjomand, Ramin Bashizade, and Hamid Sarbazi-Azad. Architecting the last-level cache for GPUs using STT-RAM technology. ACM Transactions on Design Automation of Electronic Systems, 20(4):55:1–55:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Szentimrey:2020:MLC

[SAHF^{+}20] Hannah Szentimrey, Abeer Al-Hyari, Jeremy Foxcroft, Timothy Martin, David Noel, Gary Grewal, and Shawki Areibi. Machine learning for congestion management and routability prediction within FPGA placement. ACM Transactions on Design Automation of Electronic Systems, 25(5):37:1– 37:25, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3373269.

Song:2019:HRB

[SAL19] Yang Song, Olivier Alavoine, and Bill Lin. Harvesting rowbuffer hits via orchestrated last-level cache and DRAM scheduling for heterogeneous multicore systems. ACM Transactions on Design Automation of Electronic Systems, 24(1):5:1–5:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shi:1998:CCT

[SB98] C.-J. Shi and J. A. Brzozowski.

Cluster-cover a theoretical framework for a class of VLSI-CAD optimization problems. ACM Transactions on Design Automation of Electronic Systems, 3(1):76–107, January 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1998-3-1/p76-shi/p76-shi.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        todaes/1998-3-1/p76-shi/.

Shafiee:2018:DFB

[SBB^{+}18] M. Shafiee, N. Beohar, P. Bakliwal, S. Roy, D. Mandal, B. Bakkaloglu, and S. Ozev. A disturbance-free built-in selftest and diagnosis technique for DC–DC converters. ACM Transactions on Design Automation of Electronic Systems, 23(2):25:1–25:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sehgal:2008:PAS

[SBC08] Anuja Sehgal, Sudarshan Bahukudumbi, and Krishnendu Chakrabarty. Power-aware SoC test planning for effective utilization of port-scalable testers. ACM Transactions on Design Automation of Electronic Systems, 13(3):53:1–53:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shojaei:2013:FSM

[SBGD13] Hamid Shojaei, Twan Basten, Marc Geilen, and Azadeh Davoodi. A fast and scalable multidimensional multiple-choice knapsack heuristic. ACM Transactions on Design Automation of Electronic Systems, 18(4):51:1–51:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shrivastava:2006:CFC

[SBH^{+}06] Aviral Shrivastava, Partha Biswas, Ashok Halambi, Nikil Dutt, and Alex Nicolau. Compilation framework for code size reduction using reduced bitwidth ISAs (rISAs). ACM Transactions on Design Automation of Electronic Systems, 11(1):123–146, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Saha:2017:SSS

[SBR^{+}17] Shamik Saha, Prabal Basu, Chidhambaranathan Rajamanikkam, Aatreyi Bal, Koushik Chakraborty, and Sanghamitra Roy. SSAGA: SMs synthesized for asymmetric GPGPU applications. ACM Transactions on Design Automation of Electronic Systems, 22(3):49:1–49:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sha:2020:FPT

[SBY^{+}20] Shi Sha, Ajinkya S. Bankar, Xiaokun Yang, Wujie Wen, and Gang Quan. On fundamental principles for thermalaware design on periodic realtime multi-core systems. ACM Transactions on Design Automation of Electronic Systems, 25(2):23:1–23:23, March 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3378063.

Sapatnekar:2000:PDO

[SC00] Sachin S. Sapatnekar and Weitong Chuang. Power-delay optimizations in gate sizing. ACM Transactions on Design Automation of Electronic Systems, 5(1):98–114, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2000-5-1/p98-sapatnekar/
p98-sapatnekar.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/2000-5-1/
p98-sapatnekar/.

Su:2006:MPF

[SC06] Fei Su and Krishnendu Chakrabarty. Module placement for faulttolerant microfluidics-based biochips. ACM Transactions on Design Automation of Electronic Systems, 11(3):682–710, July 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shenoy:2001:ASL

[SCB01] Nagaraj Shenoy, Alok Choudhary, and Prithviraj Banerjee. An algorithm for synthesis of large time-constrained heterogeneous adaptive systems. ACM Transactions on Design Automation of Electronic Systems, 6(2):207–225, April 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-2/p207-shenoy/p207-
shenoy.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/2001-6-2/p207-shenoy/

.

Su:2008:SNT

[SCCH08] Yu-Shih Su, Po-Hsien Chang, Shih-Chieh Chang, and Tingting Hwang. Synthesis of a novel timing-error detection architecture. ACM Transactions on Design Automation of Electronic Systems, 13(1):14:1– 14:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Schafer:2017:PHL

[Sch17] Benjamin Carrion Schafer. Parallel high-level synthesis design space exploration for behavioral IPs of exact latencies. ACM Transactions on Design Automation of Electronic Systems, 22(4):65:1–65:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Su:2001:IRA

[SCJ01] Chauchin Su, Yue-Tsang Chen, and Shyh-Jye Jou. Intrinsic response for analog module testing using an analog testability bus. ACM Transactions on Design Automation of Electronic Systems, 6(2):226–243, April 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
2001-6-2/p226-su/p226-su.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/2001-6-2/p226-su/.

Sadat:2018:OAL

[SCK18] Sayed Abdullah Sadat, Mustafa Canbolat, and Selçuk Köse. Optimal allocation of LDOs and decoupling capacitors within a distributed on-chip power grid. ACM Transactions on Design Automation of Electronic Systems, 23(4): 49:1–49:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Song:2023:MLA

[SCK^{+}23] Ling-Yen Song, Chih-Yun Chou, Tung-Chieh Kuo, Chien-Nan Liu, and Juinn-Dar Huang. Machine learning assisted circuit sizing approach for low-voltage analog circuits with efficient variation-aware optimization. ACM Transactions on Design Automation of Electronic Systems, 28(2):18:1– 18:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3567422.

Sun:2022:CMO

[SCL^{+}22] Qi Sun, Tinghuan Chen, Siting Liu, Jianli Chen, Hao Yu, and Bei Yu. Correlated multi-objective multi-fidelity optimization for HLS directives design. ACM Transactions on Design Automation of Electronic Systems, 27(4): 31:1–31:27, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3503540.

Schaumont:2006:ICE

[SCV06] Patrick Schaumont, Doris Ching, and Ingrid Verbauwhede. An interactive codesign environment for domain-specific coprocessors. ACM Transactions on Design Automation of Electronic Systems, 11(1): 70–87, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sinha:2009:DIC

[SDP^{+}09] Arnab Sinha, Pallab Dasgupta, Bhaskar Pal, Sayantan Das, Prasenjit Basu, and P. P. Chakrabarti. Design intent coverage revisited. ACM Transactions on Design Automation of Electronic Systems, 14(1): 9:1–9:??, January 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shamshiri:2005:ILT

[SEN05] Saeed Shamshiri, Hadi Esmaeilzadeh, and Zainalabdein Navabi. Instruction-level test methodology for CPU core self-testing. ACM Transactions on Design Automation of Electronic Systems, 10(4): 673–689, October 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sen:2011:COV

[Sen11]  Alper Sen.  Concurrency-

oriented verification and coverage of system-level designs. ACM Transactions on Design Automation of Electronic Systems, 16(4):37:1–37:??, October 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Srivastav:2015:DUL

[SESN15] Meeta Srivastav, Mohammed Ehteshamuddin, Kyle Stegner, and Leyla Nazhandali. Design of ultra-low power scalablethroughput many-core DSP applications. ACM Transactions on Design Automation of Electronic Systems, 20(3): 34:1–34:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Schwarzer:2019:CDA

[SFM^{+}19] Tobias Schwarzer, Joachim Falk, Simone Müller, Martin Letras, Christian Heidorn, Stefan Wildermann, and Jürgen Teich. Compilation of dataflow applications for multi-cores using adaptive multi-objective optimization. ACM Transactions on Design Automation of Electronic Systems, 24(3):29:1– 29:23, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3310249.

Schneider:2014:QNE

[SGC^{+}14] Reinhard Schneider, Dip Goswami, Samarjit Chakraborty, Unmesh Bordoloi, Petru Eles, and Zebo Peng. Quantifying notions of extensibility in FlexRay schedule synthesis. ACM Transactions on Design Automation of Electronic Systems, 19(4): 32:1–32:??, August 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Schirner:2010:FAP

[SGD10] Gunar Schirner, Andreas Gerstlauer, and Rainer Dömer. Fast and accurate processor models for efficient MPSoC design. ACM Transactions on Design Automation of Electronic Systems, 15(2):10:1– 10:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sinha:2014:FAI

[SGGR14] Roopak Sinha, Alain Girault, Gregor Goessler, and Partha S. Roop. A formal approach to incremental converter synthesis for system-on-chip design. ACM Transactions on Design Automation of Electronic Systems, 20(1):13:1–13:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sosic:1996:UAF

[SGJ96] Rok Sosic̄, Jun Gu, and Robert R. Johnson. The Unison algorithm: fast evaluation of Boolean expressions. ACM Transactions on Design Automation of Electronic Systems, 1(4):456–477, October 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-4/p456-sosic/p456-
sosic.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-4/p456-sosic/

.

Sankar:2024:TTA

[SGJN24] Syam Sankar, Ruchika Gupta, John Jose, and Sukumar Nandi. TROP: TRust-aware OPportunistic routing in NoC with hardware Trojans. ACM Transactions on Design Automation of Electronic Systems, 29(2):34:1–34:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639821.

Saluja:2008:SBA

[SGK08] Nikhil Saluja, Kanupriya Gulati, and Sunil P. Khatri. SATbased ATPG using multilevel compatible don’t-cares. ACM Transactions on Design Automation of Electronic Systems, 13(2):24:1–24:??, April 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sinha:2021:DSO

[SHBD21] Mitali Sinha, Gade Sri Harsha, Pramit Bhattacharyya, and Sujay Deb. Design space optimization of shared memory architecture in acceleratorrich systems. ACM Transactions on Design Automation of Electronic Systems, 26(4):30:1– 30:31, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3446001.

Shi:2017:TAA

[SHD17] Guoyong Shi, Hanbin Hu, and Shuwen Deng. Topological approach to automatic symbolic Macromodel generation for analog integrated circuits. ACM Transactions on Design Automation of Electronic Systems, 22(3):47:1–47:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shi:2020:ASF

[Shi20] Guoyong Shi. Automatic stageform circuit reduction for multistage opamp design equation generation. ACM Transactions on Design Automation of Electronic Systems, 25(1):2:1– 2:26, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3363499.

Song:2019:COR

[SHL^{+}19] Youngsoo Song, Daijoon Hyun, Jingon Lee, Jinwook Jung, and Youngsoo Shin. Cut optimization for redundant via insertion in self-aligned double patterning. ACM Transactions on Design Automation of Electronic Systems, 24(6):61:1– 61:21, November 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3355391.

Su:1998:EFL

[SHLL98] Alan Su, Yu-Chin Hsu, Ta-Yung Liu, and Mike Tien-Chien Lee. Eliminating false loops caused by sharing in control path. ACM Transactions on Design Automation of Electronic Systems, 3(3):487–495, July 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-3/p487-su/p487-su.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1998-3-3/p487-su/.

Srivastav:2012:DEE

[SHN12] Meeta Srivastav, M. B. Henry, and Leyla Nazhandali. Design of energy-efficient, adaptable throughput systems at near/ sub-threshold voltage. ACM Transactions on Design Automation of Electronic Systems, 18(1):3:1–3:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sarrafzadeh:2002:GE

[SJ02] Majid Sarrafzadeh and Rajeev Jayaraman. Guest editorial. ACM Transactions on Design Automation of Electronic Systems, 7(4):499–500, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sivakumar:2023:SAL

[SJ23] S. Sivakumar and John Jose.

Self adaptive logical split cache techniques for delayed aging of NVM LLC. ACM Transactions on Design Automation of Electronic Systems, 28 (6):97:1–97:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616871.

Song:2023:VEE

[SJL23] Zhuoran Song, Naifeng Jing, and Xiaoyao Liang. E^{2}-VOR: an end-to-end en/decoder architecture for efficient video object recognition. ACM Transactions on Design Automation of Electronic Systems, 28(1):10:1–10:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3543852.

Sivakumar:2024:ELP

[SJN24] S. Sivakumar, John Jose, and Vijaykrishnan Narayanan. Enhancing lifetime and performance of MLC NVM caches using embedded trace buffers. ACM Transactions on Design Automation of Electronic Systems, 29(3):58:1–58:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3659102.

Su:2006:AMS

[SKCM06] Qing Su, Jamil Kawa, Charles Chiang, and Yehia Massoud. Accurate modeling of substrate resistive coupling for floating substrates. ACM Transactions on Design Automation of Electronic Systems, 11(1): 44–51, January 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Steinhorst:2016:CPC

[SKM^{+}16] Sebastian Steinhorst, Matthias Kauer, Arne Meeuw, Swaminathan Narayanaswamy, Martin Lukasiewycz, and Samarjit Chakraborty. Cyber-physical co-simulation framework for smart cells in scalable battery packs. ACM Transactions on Design Automation of Electronic Systems, 21(4):62:1– 62:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Siddhu:2021:LAD

[SKP21] Lokesh Siddhu, Rajesh Kedia, and Preeti Ranjan Panda. Leakage-aware dynamic thermal management of 3D memories. ACM Transactions on Design Automation of Electronic Systems, 26(2):12:1– 12:31, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3419468.

Shiri:2022:EEE

[SKR^{+}22] Aidin Shiri, Uttej Kallakuri, Hasib-Al Rashid, Bharat Prakash, Nicholas R. Waytowich, Tim Oates, and Tinoosh Mohsenin. E2HRL: an energyefficient hardware accelerator for hierarchical deep reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 27(5): 45:1–45:19, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3498327.

Singh:2012:ATA

[SKS12] Amit Kumar Singh, Akash Kumar, and Thambipillai Srikanthan. Accelerating throughputaware runtime mapping for heterogeneous MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 18(1):9:1–9:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shalu:2018:DDS

[SKS^{+}18] Shalu, Srijan Kumar, Ananya Singla, Sudip Roy, Krishnendu Chakrabarty, Partha P. Chakrabarti, and Bhargab B. Bhattacharya. Demand-driven single- and multitarget mixture preparation using digital microfluidic biochips. ACM Transactions on Design Automation of Electronic Systems, 23(4):55:1–55:??, July 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Seo:2018:NIS

[SL18] Minjun Seo and Roman Lysecky. Non-intrusive in-situ requirements monitoring of embedded system. ACM Transactions on Design Automation of Electronic Systems, 23(5): 58:1–58:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sha:2022:DMB

[SLC^{+}22] Zhibing Sha, Jun Li, Zhigang Cai, Min Huang, Jianwei Liao, and Francois Trahay. Degraded mode-benefited I/O scheduling to ensure I/O responsiveness in RAID-enabled SSDs. ACM Transactions on Design Automation of Electronic Systems, 27(6):64:1–64:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3522755.

Shamsi:2019:IPS

[SLP^{+}19] Kaveh Shamsi, Meng Li, Kenneth Plaks, Saverio Fazzari, David Z. Pan, and Yier Jin. IP protection and supply chain security through logic obfuscation: a systematic overview. ACM Transactions on Design Automation of Electronic Systems, 24(6):65:1–65:36, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3342099.

Sjovall:2022:HLS

[SLV^{+}22] Panu Sjövall, Ari Lemmetti, Jarno Vanne, Sakari Lahti, and Timo D. Hämäläinen. Highlevel synthesis implementation of an embedded real-time HEVC intra encoder on FPGA for media applications. ACM Transactions on Design Automation of Electronic Systems, 27(4):35:1–35:34, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3491215.

Shi:2012:HND

[SLXZ12] Liang Shi, Jianhua Li, Chun Jason Xue, and Xuehai Zhou. Hybrid nonvolatile disk cache for energy-efficient and highperformance systems. ACM Transactions on Design Automation of Electronic Systems, 18(1):8:1–8:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sudarsanam:2000:SRA

[SM00] Ashok Sudarsanam and Sharad Malik. Simultaneous reference allocation in code generation for dual data memory bank ASIPs. ACM Transactions on Design Automation of Electronic Systems, 5(2):242–264, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-2/p242-sudarsanam/
                                        p242-sudarsanam.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/todaes/2000-5-2/
                                        p242-sudarsanam/.

Sayed:2019:CAP

[SMBT19] Nour Sayed, Longfei Mao, Rajendra Bishnoi, and Mehdi B. Tahoori. Compiler-assisted and profiling-based analysis for fast and efficient STT-MRAM onchip cache design. ACM Transactions on Design Automation of Electronic Systems, 24(4): 41:1–41:25, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3321693.

S:2022:EEE

[SMS22] Skandha Deepsita S., Dhayala Kumar M., and Noor Mahammad SK. Energy efficient error resilient multiplier using low-power compressors. ACM Transactions on Design Automation of Electronic Systems, 27(3):21:1–21:26, May 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3488837.

Suhaib:2005:XIM

[SMSB05] Syed M. Suhaib, Deepak A.

Mathaikutty, Sandeep K. Shukla, and David Berner. XFM: an incremental methodology for developing formal models. ACM Transactions on Design Automation of Electronic Systems, 10(4):589– 609, October 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shi:2007:CSO

[SMYH07] Yiyu Shi, Paul Mesa, Hao Yu, and Lei He. Circuit-simulated obstacle-aware Steiner routing. ACM Transactions on Design Automation of Electronic Systems, 12(3):28:1–28:??, August 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singh:2010:AJE

[SN10] Montek Singh and Steven M. Nowick. ACM Journal on Emerging Technologies in Computing Systems. ACM Transactions on Design Automation of Electronic Systems, 16(1):11:1–11:??, November 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Saxena:2002:ESL

[SNH02] Vikram Saxena, Farid N. Najm, and Ibrahim N. Hajj. Estimation of state line statistics in sequential circuits. ACM Transactions on Design Automation of Electronic Systems, 7(3):455–473, July 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singh:2012:TRT

[SNL12] Padmaraj Singh, Vijaykrishnan Narayanan, and David L. Landis. Targeted random test generation for power-aware multicore designs. ACM Transactions on Design Automation of Electronic Systems, 17(3): 25:1–25:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Su:2006:CTD

[SOC06] Fei Su, Sule Ozev, and Krishnendu Chakrabarty. Concurrent testing of digital microfluidics-based biochips.

ACM Transactions on Design Automation of Electronic Systems, 11(2):442–464, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Suresh:2015:AGU

[SOS15] Chandra K. H. Suresh, Sule Ozev, and Ozgur Sinanoglu. Adaptive generation of unique IDs for digital chips through analog excitation. ACM Transactions on Design Automation of Electronic Systems, 20(3): 46:1–46:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singhal:2003:SOA

[SPA^{+}03] Vigyan Singhal, Carl Pixley, Adnan Aziz, Shaz Qadeer, and Robert Brayton. Sequential optimization in the absence of global reset. ACM Transactions on Design Automation of Electronic Systems, 8(2):222– 251, April 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sharma:2015:AIE

[SPC^{+}15] Namita Sharma, Preeti Ranjan Panda, Francky Catthoor, Praveen Raghavan, and Tom Vander Aa. Array interleaving — an energy-efficient data layout transformation. ACM Transactions on Design Automation of Electronic Systems, 20(3):44:1–44:??, June 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sanz:2008:CSS

[SPG^{+}08] Concepción Sanz, Manuel Prieto, José Ignacio Gómez, Antonis Papanikolaou, Miguel Miranda, and Francky Catthoor. Combining system scenarios and configurable memories to tolerate unpredictability. ACM Transactions on Design Automation of Electronic Systems, 13(3):49:1–49:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singh:2002:ECC

[SPMS02] Amit Singh, Ganapathy Parthasarathy, and Małgorzata Marek-Sadowska. Efficient circuit clustering for area and power reduction in FPGAs. ACM Transactions on Design Automation of Electronic Systems, 7(4): 643–663, October 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Salcic:2017:NHH

[SPT^{+}17] Zoran Salcic, Heejong Park, Jürgen Teich, Avinash Malik, and Muhammad Nadeem. Noc-HMP: a heterogeneous multicore processor for embedded systems designed in SystemJ. ACM Transactions on Design Automation of Electronic Systems, 22(4):73:1–73:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sun:2024:LLD

[SQL^{+}24] Rihui Sun, Pengfei Qiu, Yongqiang Lyu, Jian Dong, Haixia Wang, Dongsheng Wang, and Gang Qu. Lightning: Leveraging DVFSinduced transient fault injection to attack deep learning accelerator of GPUs. ACM Transactions on Design Automation of Electronic Systems, 29(1):14:1–14:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3617893.

Salamy:2012:ISA

[SR12] Hassan Salamy and J. Ramanujam. An ILP solution to address code generation for embedded applications on digital signal processors. ACM Transactions on Design Automation of Electronic Systems, 17(3): 28:1–28:??, June 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Subramaniam:2015:FPM

[SRC15] Anupama R. Subramaniam, Janet Roveda, and Yu Cao. A finite-point method for efficient gate characterization under multiple input switching. ACM Transactions on Design Automation of Electronic Systems, 21(1):10:1–10:??, November 2015. CODEN ATASFO.

ISSN 1084-4309 (print), 1557-7309 (electronic).

Senapati:2023:TTA

[SRKS23] Debabrata Senapati, Kousik Rajesh, Chandan Karfa, and Arnab Sarkar. TMDS: Temperature-aware makespan minimizing DAG scheduler for heterogeneous distributed systems. ACM Transactions on Design Automation of Electronic Systems, 28(6): 99:1–99:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3616869.

Smirnov:2019:AOV

[SRTG19] Fedor Smirnov, Felix Reimann, Jürgen Teich, and Michael Glaß. Automatic optimization of the VLAN partitioning in automotive communication networks. ACM Transactions on Design Automation of Electronic Systems, 24(1): 9:1–9:??, January 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sun:2011:GDD

[SS11] Wei-Tsun Sun and Zoran Salcic. GALS-Designer: a design framework for GALS software systems. ACM Transactions on Design Automation of Electronic Systems, 16(4): 39:1–39:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sinha:2014:DGP

[SS14] Sharad Sinha and Thambipillai Srikanthan. Dataflow graph partitioning for area-efficient high-level synthesis with systems perspective. ACM Transactions on Design Automation of Electronic Systems, 20(1): 5:1–5:??, November 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

S:2024:SAS

[SS24] Deepanjali S. and Noor Mahammad SK. Scalable and accelerated self-healing control circuit using evolvable hardware. ACM Transactions on Design Automation of Electronic Systems, 29(2):31:1– 31:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3634682.

Saha:2017:STS

[SSC17] Sangeet Saha, Arnab Sarkar, and Amlan Chakrabarti. Spatiotemporal scheduling of preemptive real-time tasks on partially reconfigurable systems. ACM Transactions on Design Automation of Electronic Systems, 22(4):71:1–71:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shin:2010:PGC

[SSCS10] Youngsoo Shin, Jun Seomun, Kyu-Myung Choi, and Takayasu Sakurai. Power gating: Circuits, design methodologies, and best practice for standard-cell VLSI designs. ACM Transactions on Design Automation of Electronic Systems, 15(4):28:1– 28:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Saladi:2012:CAC

[SSG12] Kalyan Saladi, Harikumar Somakumar, and Mahadevan Ganapathi. Concurrencyaware compiler optimizations for hardware description languages. ACM Transactions on Design Automation of Electronic Systems, 18(1):10:1– 10:??, December 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singh:2003:MST

[SSGS03] Adit D. Singh, Markus Seuring, Michael Gössel, and Egor S. Sogomonyan. Multimode scan: Test per clock BIST for IP cores. ACM Transactions on Design Automation of Electronic Systems, 8(4): 491–505, October 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sanchez:2023:CSE

[SSK^{+}23] Daniela Sánchez, Lorenzo Servadei, Gamze Naz Kiprit, Robert Wille, and Wolfgang Ecker. A comprehensive survey on electronic design automation and graph neural networks: Theory and applications. ACM Transactions on Design Automation of Electronic Systems, 28(2):15:1– 15:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3543853.

Song:2017:STV

[SSL17] Yang Song, Kambiz Samadi, and Bill Lin. A single-tier virtual queuing memory controller architecture for heterogeneous MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 22(3): 56:1–56:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Singh:2022:PFE

[SSN22] Somesh Singh, Tejas Shah, and Rupesh Nasre. ParTBC: Faster estimation of top-k betweenness centrality vertices on GPU. ACM Transactions on Design Automation of Electronic Systems, 27(2):12:1– 12:25, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3486613.

Suresh:2016:AVD

[SSO16] Chandra K. H. Suresh, Ozgur Sinanoglu, and Sule Ozev. Adapting to varying distribution of unknown response bits. ACM Transactions on Design Automation of Electronic Systems, 21(2):33:1–33:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sundararajan:2004:NAI

[SSP04] Vijay Sundararajan, Sachin S. Sapatnekar, and Keshab K. Parhi. A new approach for integration of min-area retiming and min-delay padding for simultaneously addressing shortpath and long-path constraints. ACM Transactions on Design Automation of Electronic Systems, 9(3):273–289, July 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sahoo:2019:FMV

[SSS^{+}19] Debiprasanna Sahoo, Swaraj Sha, Manoranjan Satpathy, Madhu Mutyam, S. Ramesh, and Partha Roop. Formal modeling and verification of a victim DRAM cache. ACM Transactions on Design Automation of Electronic Systems, 24(2):20:1–20:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3306491.

Shi:1999:SSL

[ST99] C.-J. Richard Shi and Michael W.

Tian. Simulation and sensitivity of linear analog circuits under parameter variations by Robust interval analysis. ACM Transactions on Design Automation of Electronic Systems, 4(3):280–312, July 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-3/p280-shi/p280-shi.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1999-4-3/p280-shi/.

Sun:2015:NUB

[STGR15] Jin Sun, Claudio Talarico, Priyank Gupta, and Janet Roveda. A new uncertainty budgeting-based method for robust analog/mixed-signal design. ACM Transactions on Design Automation of Electronic Systems, 21(1):16:1– 16:??, November 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Somashekar:2016:NEG

[STJG16] Ahish Mysore Somashekar, Spyros Tragoudas, Rathish Jayabharathi, and Sreenivas Gangadhar. Non-enumerative generation of path delay distributions and its application to critical path selection. ACM Transactions on Design Automation of Electronic Systems, 22(1):17:1–17:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shen:2013:AAP

[STL^{+}13] Hao Shen, Ying Tan, Jun Lu, Qing Wu, and Qinru Qiu. Achieving autonomous power management using reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 18(2): 24:1–24:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shen:2012:FSF

[STWX12] Ruijing Shen, Sheldon X.-D. Tan, Hai Wang, and Jinjun Xiong. Fast statistical full-chip leakage analysis for nanometer VLSI systems. ACM Transactions on Design Automation of Electronic Systems, 17(4): 51:1–51:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shiue:2001:DMD

[SUC01] Wen-Tsong Shiue, Sathishkumar Udayanarayanan, and Chaitali Chakrabarti. Data memory design and exploration for low-power embedded systems. ACM Transactions on Design Automation of Electronic Systems, 6(4): 553–568, October 2001. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Stitt:2007:BS

[SV07] Greg Stitt and Frank Vahid.

Binary synthesis. ACM Transactions on Design Automation of Electronic Systems, 12(3): 34:1–34:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Stitt:2011:TWD

[SV11] Greg Stitt and Frank Vahid.

Thread warping: Dynamic and transparent synthesis of thread accelerators. ACM Transactions on Design Automation of Electronic Systems, 16(3): 32:1–32:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

S:2016:EAD

[SV16] Ramprasath S. and Vinita Vasudevan. Efficient algorithms for discrete gate sizing and threshold voltage assignment based on an accurate analytical statistical yield gradient. ACM Transactions on Design Automation of Electronic Systems, 21(4):66:1–66:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Santos:2017:SMH

[SVK17] Rui Santos, Shyamsundar Venkataraman, and Akash Kumar. Scrubbing mechanism for heterogeneous applications in reconfigurable devices. ACM Transactions on Design Automation of Electronic Systems, 22(2):33:1–33:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Song:1999:CDP

[SW99] Xiaoyu Song and Yuke Wang. On the crossing distribution problem. ACM Transactions on Design Automation of Electronic Systems, 4(1):39– 51, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
1999-4-1/p39-song/p39-song.
pdf;   http://www.acm.org/
pubs/citations/journals/
todaes/1999-4-1/p39-song/.

Sabade:2004:BTM

[SW04] Sagar S. Sabade and Duncan M. Walker. IDDX-based test methods: a survey. ACM Transactions on Design Automation of Electronic Systems, 9(2):159–198, April 2004. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Schafer:2012:DCH

[SW12] Benjamin Carrion Schafer and Kazutoshi Wakabayashi. Divide and conquer high-level synthesis design space exploration. ACM Transactions on Design Automation of Electronic Systems, 17(3):29:1– 29:??, June 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Spieck:2023:LBM

[SWT23] Jan Spieck, Stefan Wildermann, and Jürgen Teich. A learning-based methodology for scenario-aware mapping of soft real-time applications onto heterogeneous MPSoCs. ACM Transactions on Design Automation of Electronic Systems, 28(1):4:1–4:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3529230.

Shao:2006:LST

[SXX^{+}06] Zili Shao, Bin Xiao, Chun Xue, Qingfeng Zhuge, and Edwin H.- M. Sha. Loop scheduling with timing and switching-activity minimization for VLIW DSP. ACM Transactions on Design Automation of Electronic Systems, 11(1):165–185, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Shi:2013:OSC

[SXZV13] Yiyu Shi, Jinjun Xiong, Vladimir Zolotov, and Chandu Visweswariah. Order statistics for correlated random variables and its application to atspeed testing. ACM Transactions on Design Automation of Electronic Systems, 18(3):

42:1–42:??, July 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sham:2007:ARD

[SY07] Chiu-Wing Sham and Evangeline F. Y. Young. Area reduction by deadspace utilization on interconnect optimized floorplan. ACM Transactions on Design Automation of Electronic Systems, 12(1): 3:1–3:??, January 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sohrabizadeh:2022:AES

[SYGC22] Atefeh Sohrabizadeh, Cody Hao Yu, Min Gao, and Jason Cong. AutoDSE: Enabling software programmers to design efficient FPGA accelerators. ACM Transactions on Design Automation of Electronic Systems, 27(4):32:1–32:27, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3494534.

Shi:2022:CHD

[SYH^{+}22] Xiao Shi, Hao Yan, Qiancun Huang, Chengzhen Xuan, Lei He, and Longxing Shi. A compact high-dimensional yield analysis method using low-rank tensor approximation. ACM Transactions on Design Automation of Electronic Systems, 27(2):19:1–19:23, March 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3483941.

Shih:2014:COR

[SYHL14] Wen-Li Shih, Yi-Ping You, Chung-Wen Huang, and Jenq Kuen Lee. Compiler optimization for reducing leakage power in multithread BSP programs. ACM Transactions on Design Automation of Electronic Systems, 20(1):9:1–9:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sham:2009:CPE

[SYL09] Chiu-Wing Sham, Evangeline F. Y. Young, and Jingwei Lu. Congestion prediction in early stages of physical design. ACM Transactions on Design Automation of Electronic Systems, 14(1):12:1–12:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sun:2012:PTA

[SYX12] Guangyu Sun, Huazhong Yang, and Yuan Xie. Performance/ thermal-aware design of 3Dstacked L2 caches for CMPs. ACM Transactions on Design Automation of Electronic Systems, 17(2):13:1–13:??, April 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sham:2008:OWR

[SYZ08] Chiu-Wing Sham, Evangeline F. Y. Young, and Hai Zhou. Optimizing wirelength and routability by searching alternative packings in floorplanning. ACM Transactions on Design Automation of Electronic Systems, 13(1):21:1– 21:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Su:2017:EMC

[SZB17] Hang Su, Dakai Zhu, and Scott Brandt. An elastic mixed-criticality task model and early-release EDF scheduling algorithms. ACM Transactions on Design Automation of Electronic Systems, 22(2): 28:1–28:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Sun:2012:STD

[SZV^{+}12] Jin Sun, Rui Zheng, Jyothi Velamala, Yu Cao, Roman Lysecky, Karthik Shankar, and Janet Roveda. A self-tuning design methodology for powerefficient multi-core systems. ACM Transactions on Design Automation of Electronic Systems, 18(1):4:1–4:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Thakur:2024:VLL

[TAP^{+}24] Shailja Thakur, Baleegh Ahmad, Hammond Pearce, Benjamin Tan, Brendan Dolan-Gavitt, Ramesh Karri, and Siddharth Garg. VeriGen: a large language model for Verilog code generation. ACM Transactions on Design Automation of Electronic Systems, 29(3): 46:1–46:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3643681.

Tadros:2020:TFT

[TB20] Ramy N. Tadros and Peter A.

Beerel. A theoretical foundation for timing synchronous systems using asynchronous structures. ACM Transactions on Design Automation of Electronic Systems, 25(2):16:1– 16:28, March 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3373355.

Tang:2017:PPE

[TBCH17] Li Tang, Richard F. Barrett, Jeanine Cook, and X. Sharon Hu. PeaPaw: Performance and energy-aware partitioning of workload on heterogeneous platforms. ACM Transactions on Design Automation of Electronic Systems, 22(3): 41:1–41:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tong:2013:TCT

[TBZ13] Jason G. Tong, Marc Boulé, and Zeljko Zilic. Test compaction techniques for assertion-based test generation. ACM Transactions on Design Automation of Electronic Systems, 19(1):9:1–9:??, December 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tiruvuri:1998:ELB

[TC98] Giri Tiruvuri and Moon Chung. Estimation of lower bounds in scheduling algorithms for high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 3(2):162–180, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1998-3-2/p162-tiruvuri/p162-
                                        tiruvuri.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1998-3-2/p162-tiruvuri/

.

Tsai:2014:PAE

[TCL14] Kun-Lin Tsai, Hao-Tse Chen, and Yo-An Lin. Power and area efficiency NoC router design for application-specific SoC by using buffer merging and resource sharing. ACM Transactions on Design Automation of Electronic Systems, 19(4): 36:1–36:??, August 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Thadikaran:1997:ACB

[TCP97] Paul Thadikaran, Sreejit Chakravarty, and Janak Patel. Algorithms to compute bridging fault coverage of IDDQ test sets. ACM Transactions on Design Automation of Electronic Systems, 2(3):281–305, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-3/p281-thadikaran/
p281-thadikaran.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/1997-2-3/
p281-thadikaran/.

Tseng:2020:MAU

[TCW20] Tien-Hung Tseng, Chung-Han Chou, and Kai-Chiang Wu. Making aging useful by recycling aging-induced clock skew. ACM Transactions on Design Automation of Electronic Systems, 25(2):13:1– 13:24, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3363186.

Tragoudas:2003:PDF

[TD03] S. Tragoudas and N. Denny.

Path delay fault testing using test points. ACM Transactions on Design Automation of Electronic Systems, 8(1):1–10, January 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Taktak:2008:TAD

[TDE08] Sami Taktak, Jean-Lou Desbarbieux, and Emmanuelle Encrenaz. A tool for automatic detection of deadlock in wormhole networks on chip. ACM Transactions on Design Automation of Electronic Systems, 13(1):6:1–6:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Taskin:2009:CTR

[TDF^{+}09] Baris Taskin, Joseph Demaio, Owen Farell, Michael Hazeltine, and Ryan Ketner. Custom topology rotary clock router with tree subnetworks. ACM Transactions on Design Automation of Electronic Systems, 14(3):44:1–44:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Truong:2018:LSE

[TEK18] Anh Truong, S. Rasoul Etesami, and Negar Kiyavash. Learning from sleeping experts: Rewarding informative, available, and accurate experts. ACM Transactions on Design Automation of Electronic Systems, 23(6):77:1–77:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tessier:2002:FPA

[Tes02] Russell Tessier. Fast placement approaches for FPGAs. ACM Transactions on Design Automation of Electronic Systems, 7(2):284–305, April 2002.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tang:2024:MIP

[TFW24] Ke Tang, Lang Feng, and Zhongfeng Wang. Mixed integer programming based placement refinement by RSMT model with movable pins. ACM Transactions on Design Automation of Electronic Systems, 29(2):36:1–36:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3639365.

Tu:2014:PPP

[THC^{+}14] Chia-Heng Tu, Hui-Hsin Hsu, Jen-Hao Chen, Chun-Han Chen, and Shih-Hao Hung. Performance and power profiling for emulated Android systems. ACM Transactions on Design Automation of Electronic Systems, 19(2):10:1– 10:??, March 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tsai:2013:ROC

[THL^{+}13] Mei-Hsiang Tsai, Po-Yang Hsu, Hung-Yi Li, Yi-Huang Hung, and Yi-Yu Liu. Routability optimization for crossbarswitch structured ASIC design. ACM Transactions on Design Automation of Electronic Systems, 18(3):39:1–39:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Taouil:2015:YIW

[THM15] Mottaqiallah Taouil, Said Hamdioui, and Erik Jan Marinissen. Yield improvement for 3D wafer-to-wafer stacked ICs using wafer matching. ACM Transactions on Design Automation of Electronic Systems, 20(2):19:1–19:??, February 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tu:2012:MFS

[THT12] Chia-Heng Tu, Shih-Hao Hung, and Tung-Chieh Tsai. MCEmu: a framework for software development and performance analysis of multicore systems. ACM Transactions on Design Automation of Electronic Systems, 17(4):36:1–36:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tseng:1999:TLL

[TJ99] Jyh-Mou Tseng and Jing-Yang Jou. Two-level logic minimization for low power. ACM Transactions on Design Automation of Electronic Systems, 4(1):52–69, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-1/p52-tseng/p52-tseng.
pdf;   http://www.acm.org/

pubs/citations/journals/
todaes/1999-4-1/p52-tseng/

.

Topaloglu:2018:ETS

[TK18] Rasit O. Topaloglu and Farinaz Koushanfar. Editorial for TODAES special issue on Internet of Things system performance, reliability, and security. ACM Transactions on Design Automation of Electronic Systems, 23(6):74:1–74:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tang:2007:PDF

[TKVN07] Weiyu Tang, Arun Kejariwal, Alexander V. Veidenbaum, and Alexandru Nicolau. A predictive decode filter cache for reducing power consumption in embedded processors. ACM Transactions on Design Automation of Electronic Systems, 12(2):14:1–14:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tu:2019:AOS

[TL19] Chia-Heng Tu and Te-Sheng Lin. Augmenting operating systems with OpenCL accelerators. ACM Transactions on Design Automation of Electronic Systems, 24(3):30:1– 30:29, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3315569.

Tan:2016:ESE

[TLCF16] Jingweijia Tan, Zhi Li, Mingsong Chen, and Xin Fu. Exploring soft-error robust and energy-efficient register file in GPGPUs using resistive memory. ACM Transactions on Design Automation of Electronic Systems, 21(2):34:1–34:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Thorolfsson:2010:LPH

[TMDF10] Thorlindur Thorolfsson, Samson Melamed, W. Rhett Davis, and Paul D. Franzon. Lowpower hypercube divided memory FFT engine using 3D integration. ACM Transactions on Design Automation of Electronic Systems, 16(1):5:1– 5:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Thornton:1999:BSC

[TN99] M. A. Thornton and V. S. S. Nair. Behavioral synthesis of combinational logic using spectral-based heuristics. ACM Transactions on Design Automation of Electronic Systems, 4(2):219–230, April 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-2/p219-thornton/p219-
thornton.pdf;   http://www.
acm.org/pubs/citations/journals/

todaes/1999-4-2/p219-thornton/

.

Tseng:2008:PPD

[TP08] I-Lun Tseng and Adam Postula. Partitioning parameterized 45-degree polygons with constraint programming. ACM Transactions on Design Automation of Electronic Systems, 13(3):52:1–52:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Trinadh:2017:ODC

[TPC^{+}17] A. Satya Trinadh, Seetal Potluri, Sobhan Babu Ch., V. Kamakoti, and Shiv Govind Singh. Optimal don’t care filling for minimizing peak toggles during at-speed stuck-at testing. ACM Transactions on Design Automation of Electronic Systems, 23(1):5:1–5:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Teman:2016:PAP

[TRM^{+}16] Adam Teman, Davide Rossi, Pascal Meinerzhagen, Luca Benini, and Andreas Burg. Power, area, and performance optimization of standard cell memory arrays through controlled placement. ACM Transactions on Design Automation of Electronic Systems, 21(4): 59:1–59:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Thanvantri:1996:OFS

[TS96] Venkat Thanvantri and Sartaj Sahni. Optimal folding of standard and custom cells. ACM Transactions on Design Automation of Electronic Systems, 1(1):123–143, January 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-1/p123-thanvantri/
p123-thanvantri.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/1996-1-1/
p123-thanvantri/.

Thakur:1996:SPF

[TW96] Shashidhar Thakur and D. F. Wong. Series-parallel functions and FPGA logic module design. ACM Transactions on Design Automation of Electronic Systems, 1(1):102–122, January 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1996-1-1/p102-thakur/p102-
thakur.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1996-1-1/p102-thakur/

.

Tannir:2016:AMN

[TWL16] Dani Tannir, Ya Wang, and Peng Li. Accurate modeling of nonideal low-power PWM DC–DC converters operating in CCM and DCM using enhanced circuit-averaging techniques. ACM Transactions on Design Automation of Electronic Systems, 21(4):61:1– 61:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tan:2023:IPC

[TWM^{+}23] Jingweijia Tan, Weiren Wang, Maodi Ma, Xiaohui Wei, and Kaige Yan. Improving the performance of CNN accelerator architecture under the impact of process variations. ACM Transactions on Design Automation of Electronic Systems, 28(5):85:1–85:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3604236.

Tomiyama:1997:CPT

[TY97] Hiroyuki Tomiyama and Hiroto Yasuura. Code placement techniques for cache miss rate reduction. ACM Transactions on Design Automation of Electronic Systems, 2(4):410–429, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-4/p410-tomiyama/p410-
tomiyama.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-4/p410-tomiyama/

.

Tan:2019:EMI

[TY19] Jingweijia Tan and Kaige Yan. Efficiently managing the impact of hardware variability on GPUs’ streaming processors. ACM Transactions on Design Automation of Electronic Systems, 24(1):12:1–12:??, January 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tzeng:2008:VPS

[TYH08] Chao-Wen Tzeng, Jheng-Syun Yang, and Shi-Yu Huang. A versatile paradigm for scan chain diagnosis of complex faults using signal processing techniques. ACM Transactions on Design Automation of Electronic Systems, 13(1): 9:1–9:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Tan:2020:EEG

[TYSF20] Jingweijia Tan, Kaige Yan, Shuaiwen Leon Song, and Xin Fu. Energy-efficient GPU L2 cache design using instructionlevel data locality similarity. ACM Transactions on Design Automation of Electronic Systems, 25(6):52:1–52:18, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3408060.

Torabi:2017:FHA

[TZ17] Mohammad Torabi and Lihong Zhang. A fast hierarchical adaptive analog routing algorithm based on integer linear programming. ACM Transactions on Design Automation of Electronic Systems, 22(3): 55:1–55:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Torabi:2020:LAA

[TZ20] Mohammad Torabi and Lihong Zhang. LDE-aware analog layout migration with OPCinclusive routing. ACM Transactions on Design Automation of Electronic Systems, 25 (6):49:1–49:22, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3398190.

Tang:2022:ETE

[TZZH22] Yue Tang, Xinyi Zhang, Peipei Zhou, and Jingtong Hu. EF-Train: Enable efficient ondevice CNN training on FPGA through data reshaping for online adaptation or personalization. ACM Transactions on Design Automation of Electronic Systems, 27(5): 49:1–49:36, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3505633.

Uysal:2022:SCN

[UE22] Necati Uysal and Rickard Ewetz. Synthesis of clock networks with a mode-reconfigurable topology. ACM Transactions on Design Automation of Electronic Systems, 27(4):38:1– 38:22, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3503538.

Utyamishev:2023:MPP

[UPV23] Dmitry Utyamishev and Inna Partin-Vaisband. Multiterminal pathfinding in practical VLSI systems with deep neural networks. ACM Transactions on Design Automation of Electronic Systems, 28(4): 51:1–51:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3564930.

Vatanparvar:2017:ASR

[VA17a] Korosh Vatanparvar and Mohammad Abdullah Al Faruque. Application-specific residential microgrid design methodology. ACM Transactions on Design Automation of Electronic Systems, 22(3):44:1–44:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vatanparvar:2017:EVO

[VA17b] Korosh Vatanparvar and Mohammad Abdullah Al Faruque. Electric vehicle optimized charge and drive management. ACM Transactions on Design Automation of Electronic Systems, 23(1):3:1–3:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

VanCampenhout:1998:HLD

[VAAH^{+}98] D. Van Campenhout, H. Al-Asaad, J. P. Hayes, T. Mudge, and R. B. Brown. High-level design verification of microprocessors via error modeling. ACM Transactions on Design Automation of Electronic Systems, 3(4):581–599, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p581-campenhout/
p581-campenhout.pdf;   http:
//www.acm.org/pubs/citations/
journals/todaes/1998-3-4/
p581-campenhout/.

Vahid:1999:PCT

[Vah99]  Frank Vahid.  Procedure

cloning: a transformation for improved system-level functional partitioning. ACM Transactions on Design Automation of Electronic Systems, 4(1):70–96, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-1/p70-vahid/p70-vahid.
pdf;   http://www.acm.org/

                                        pubs/citations/journals/
                                        todaes/1999-4-1/p70-vahid/

.

Vahid:2002:PSP

[Vah02] Frank Vahid. Partitioning sequential programs for CAD using a three-step approach. ACM Transactions on Design Automation of Electronic Systems, 7(3):413–429, July 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vinco:2019:CLV

[VBP^{+}19] Sara Vinco, Nicola Bombieri, Daniele Jahier Pagliari, Franco Fummi, Enrico Macii, and Massimo Poncino. A crosslevel verification methodology for digital IPs augmented with embedded timing monitors. ACM Transactions on Design Automation of Electronic Systems, 24(3):27:1– 27:23, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3308565.

VanAchteren:2003:SSD

[VCLD03] Tanja Van Achteren, Francky Catthoor, Rudy Lauwereins, and Geert Deconinck. Search space definition and exploration for nonuniform data reuse opportunities in datadominant applications. ACM Transactions on Design Automation of Electronic Systems, 8(1):125–139, January 2003.

CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Venkatasubramanian:2016:PID

[VEO16] Ramachandran Venkatasubramanian, Robert Elio, and Sule Ozev. Process independent design methodology for the active RC and single-inverter-based rail clamp. ACM Transactions on Design Automation of Electronic Systems, 21(3): 50:1–50:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Venieris:2023:MMW

[VFML23] Stylianos I. Venieris, Javier Fernandez-Marques, and Nicholas D. Lane. Mitigating memory wall effects in CNN engines with on-the-fly weights generation. ACM Transactions on Design Automation of Electronic Systems, 28(6):92:1–92:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3611673.

Valencia:2019:CPA

[VGG19] Juan Valencia, Dip Goswami, and Kees Goossens. Comparing platform-aware control design flows for composable and predictable TDM-based execution platforms. ACM Transactions on Design Automation of Electronic Systems, 24(3):32:1– 32:26, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3315572.

Vanna-Iampikul:2023:GBM

[VILSL23] Pruek Vanna-Iampikul, Yi-Chen Lu, Da Eun Shim, and Sung Kyu Lim. GNNbased multi-bit flip-flop clustering and post-clustering design optimization for energyefficient 3D ICs. ACM Transactions on Design Automation of Electronic Systems, 28(5): 76:1–76:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3588570.

Vanbroekhoven:2007:PDS

[VJBC07] Peter Vanbroekhoven, Gerda Janssens, Maurice Bruynooghe, and Francky Catthoor. A practical dynamic single assignment transformation. ACM Transactions on Design Automation of Electronic Systems, 12(4):40:1– 40:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vemuri:2002:ERO

[VKKR02] Ranga Vemuri, Srinivas Katkoori, Meenakshi Kaul, and Jay Roy. An efficient register optimization algorithm for high-level synthesis from hierarchical behavioral specifications. ACM Transactions on Design Automation of Electronic Systems, 7(1):189–216, January 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vemuri:2002:BBL

[VKT02] Navin Vemuri, Priyank Kalla, and Russell Tessier. BDDbased logic synthesis for LUTbased FPGAs. ACM Transactions on Design Automation of Electronic Systems, 7(4):501– 525, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

VanPraet:2001:PMC

[VLGG01] J. Van Praet, D. Lanneer, W. Geurts, and G. Goossens. Processor modeling and code selection for retargetable compilation. ACM Transactions on Design Automation of Electronic Systems, 6(3):277–307, July 2001. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vahid:1998:FPI

[VLH98] Frank Vahid, Thuy Dm Le, and Yu-Chin Hsu. Functional partitioning improvements over structural partitioning for packaging constraints and synthesis: tool performance. ACM Transactions on Design Automation of Electronic Systems, 3(2):181–208, April 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/

                                        1998-3-2/p181-vahid/p181-
                                        vahid.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/1998-3-2/p181-vahid/

.

Vicente:2004:APT

[VLH04] Juan D. Vicente, Juan Lanchares, and Román Hermida. Annealing placement by thermodynamic combinatorial optimization. ACM Transactions on Design Automation of Electronic Systems, 9(3):310–332, July 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

VanEijk:2000:CAC

[VMP^{+}00] Koen Van Eijk, Bart Mesman, Carlos A. Alba Pinto, Qin Zhao, Marco Bekooij, Jef Van Meerbergen, and Jochen Jess. Constraint analysis for code generation: basic techniques and applications in FACTS. ACM Transactions on Design Automation of Electronic Systems, 5(4):774–793, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-4/p774-van_eijk/p774-
                                        van_eijk.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        todaes/2000-5-4/p774-van_
                                        eijk/.

Vegesna:2019:NRM

[VNS19] S. M. Srinivasavarma Vegesna, Ashok Chakravarthy Nara, and Noor Mahammad Sk. A novel rule mapping on TCAM for power efficient packet classification. ACM Transactions on Design Automation of Electronic Systems, 24(5):48:1– 48:23, October 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3328103.

Voeten:2001:FLT

[Voe01] Jeroen Voeten. On the fundamental limitations of transformational design. ACM Transactions on Design Automation of Electronic Systems, 6(4): 533–552, October 2001. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Verbeek:2012:EFS

[VS12a] Freek Verbeek and Julien Schmaltz. Easy formal specification and validation of unbounded Networks-on-Chips architectures. ACM Transactions on Design Automation of Electronic Systems, 17(1): 1:1–1:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Verbeek:2012:TFV

[VS12b] Freek Verbeek and Julien Schmaltz. Towards the formal verification of cache coherency at the architectural level. ACM Transactions on Design Automation of Electronic Systems, 17(3):20:1–20:??, June 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Vijayan:2020:RIH

[VTC20] Arunkumar Vijayan, Mehdi B. Tahoori, and Krishnendu Chakrabarty. Runtime identification of hardware Trojans by feature analysis on gate-level unstructured data and anomaly detection. ACM Transactions on Design Automation of Electronic Systems, 25(4):33:1–33:23, September 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3391890.

Tsao:2002:UDC

[wATkK02] Chung wen Albert Tsao and Cheng kok Koh. UST/DME: a clock tree router for general skew constraints. ACM Transactions on Design Automation of Electronic Systems, 7(3):359–379, July 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:1998:MEV

[WAZ98] Li-C. Wang, Magdy S. Abadir, and Jing Zeng. On measuring the effectiveness of various design validation approaches for PowerPC microprocessor embedded arrays. ACM Transactions on Design Automation of Electronic Systems, 3(4):524–532, October 1998. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1998-3-4/p524-wang/p524-
wang.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1998-3-4/p524-wang/

.

Wang:2016:ERL

[WB16] Hongfei Wang and R. D. (Shawn)

Blanton. Ensemble reduction via logic minimization. ACM Transactions on Design Automation of Electronic Systems, 21(4):67:1–67:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2024:FDC

[WB^{+}24] Dan Wu, Peng Chen *, Thilini Kaushalya Bandara, Zhaoying Li, and Tulika Mitra. Flip: Data-centric edge CGRA accelerator. ACM Transactions on Design Automation of Electronic Systems, 29(1):22:1– 22:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3631118.

Wu:2006:MWR

[WC06] Zhong-Zhen Wu and Shih-Chieh Chang. Multiple wire reconnections based on implication flow graph. ACM Transactions on Design Automation of Electronic Systems, 11(4):

939–952, October 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2010:SCR

[WC10] Yu-Ze Wu and Mango C.- T. Chao. Scan-cell reordering for minimizing scan-shift power based on nonspecified test cubes. ACM Transactions on Design Automation of Electronic Systems, 16(1):10:1– 10:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2015:BST

[WCB15] Ran Wang, Krishnendu Chakrabarty, and Sudipta Bhawmik. Built-in self-test and test scheduling for interposer-based 2.5D IC. ACM Transactions on Design Automation of Electronic Systems, 20(4):58:1–58:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2003:RBP

[WCC03] Guang-Ming Wu, Yun-Chih Chang, and Yao-Wen Chang. Rectilinear block placement using B*-trees. ACM Transactions on Design Automation of Electronic Systems, 8(2):188– 202, April 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2014:CTS

[WCCC14] Chun-Kai Wang, Yeh-Chi Chang, Hung-Ming Chen, and Ching-Yu Chin. Clock tree synthesis considering slew effect on supply voltage variation. ACM Transactions on Design Automation of Electronic Systems, 20(1):3:1–3:??, November 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2024:SAM

[WCX^{+}24] Ying Wu, Chuangtao Chen, Weihua Xiao, Xuan Wang, Chenyi Wen, Jie Han, Xunzhao Yin, Weikang Qian, and Cheng Zhuo. A survey on approximate multiplier designs for energy efficiency: From algorithms to circuits. ACM Transactions on Design Automation of Electronic Systems, 29(1):23:1–23:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3610291.

Wang:2024:MLC

[WCZ^{+}24] Jianfeng Wang, Zhonghao Chen, Jiahao Zhang, Yixin Xu, Tongguang Yu, Ziheng Zheng, Enze Ye, Sumitha George, Huazhong Yang, Yongpan Liu, Kai Ni, Vijaykrishnan Narayanan, and Xueqing Li. A module-level configuration methodology for programmable camouflaged logic. ACM Transactions on Design Automation of Electronic Systems, 29(2):39:1–39:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3640462.

Wen:2022:MCT

[WDC^{+}22] Chenyi Wen, Xiao Dong, Baixin Chen, Umamaheswara Rao Tida, Yiyu Shi, and Cheng Zhuo. Magnetic core TSVinductor design and optimization for on-chip DC-DC converter. ACM Transactions on Design Automation of Electronic Systems, 27(5): 52:1–52:23, September 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3507700.

Wang:2023:DDD

[WDD^{+}23] Junpeng Wang, Haitao Du, Bo Ding, Qi Xu, Song Chen, and Yi Kang. DDAM: Data distribution-aware mapping of CNNs on processing-in-memory systems. ACM Transactions on Design Automation of Electronic Systems, 28(3): 36:1–36:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3576196.

Wang:2017:WPL

[WDLD17] Shuai Wang, Guangshan Duan, Yupeng Li, and Qianhao Dong. Word- and partition-level write variation reduction for improving non-volatile cache lifetime. ACM Transactions on Design Automation of Electronic Systems, 23(1):4:1–4:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2021:CPO

[WDLX21] Nan Wu, Lei Deng, Guoqi Li, and Yuan Xie. Core placement optimization for multichip many-core neural network systems with reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 26(2):11:1– 11:27, February 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3418498.

Wang:2016:MSM

[WDZG16] Chao Wang, Chuansheng Dong, Haibo Zeng, and Zonghua Gu. Minimizing stack memory for hard real-time applications on multicore platforms with partitioned fixedpriority or EDF scheduling. ACM Transactions on Design Automation of Electronic Systems, 21(3):46:1–46:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2020:MFS

[WFSS20] Yi Wang, Paul D. Franzon, David Smart, and Brian Swahn. Multi-fidelity surrogatebased optimization for electromagnetic simulation acceleration. ACM Transactions on Design Automation of Electronic Systems, 25(5):45:1– 45:21, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3398268.

Wang:2019:CPR

[WFT^{+}19] Chengning Wang, Dan Feng, Wei Tong, Jingning Liu, Zheng Li, Jiayi Chang, Yang Zhang, Bing Wu, Jie Xu, Wei Zhao, Yilin Li, and Ruoxi Ren. Cross-point resistive memory: Nonideal properties and solutions. ACM Transactions on Design Automation of Electronic Systems, 24(4):46:1– 46:37, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3325067.

Wang:2011:RCM

[WG11] Xiaofang Wang and Pallav Gupta. Resource-constrained multiprocessor synthesis for floating-point applications on FPGAs. ACM Transactions on Design Automation of Electronic Systems, 16(4):41:1– 41:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2007:ETR

[WGDK07] Gang Wang, Wenrui Gong, Brian Derenzi, and Ryan Kastner. Exploring time/resource trade-offs by solving dual scheduling problems with the ant colony optimization. ACM Transactions on Design Automation of Electronic Systems, 12(4):46:1–46:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Won:2016:RSC

[WGSH16] Jae-Yeon Won, Paul V. Gratz, Srinivas Shakkottai, and Jiang Hu. Resource sharing centric dynamic voltage and frequency scaling for CMP cores, uncore, and memory. ACM Transactions on Design Automation of Electronic Systems, 21(4):69:1– 69:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2017:MSS

[WGT^{+}17] Ya Wang, Di Gao, Dani Tannir, Ning Dong, G. Peter Fang, Wei Dong, and Peng Li. Multiharmonic small-signal modeling of low-power PWM DC-DC converters. ACM Transactions on Design Automation of Electronic Systems, 22(4): 68:1–68:??, July 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2005:EAV

[WH05] Zhong Wang and Xiaobo Sharon Hu. Energy-aware variable partitioning and instruction scheduling for multibank memory architectures. ACM Transactions on Design Automation of Electronic Systems, 10(2): 369–388, April 2005. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2019:ITD

[WH19] Hongfei Wang and Kun He. Improving test and diagnosis efficiency through ensemble reduction and learning. ACM Transactions on Design Automation of Electronic Systems, 24(5): 49:1–49:26, October 2019. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3328754.

Wang:2020:IBT

[WH20] Chih-Hao Wang and Tong-Yu Hsieh. An implication-based test scheme for both diagnosis and concurrent error detection applications. ACM Transactions on Design Automation of Electronic Systems, 25(1):3:1– 3:27, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3364681.

Wang:2012:BOD

[WHRC12] An-Ping Wang, Jiwon Hahn, Mahshid Roumi, and Pai H. Chou. Buffer optimization and dispatching scheme for embedded systems with behavioral transparency. ACM Transactions on Design Automation of Electronic Systems, 17(4):41:1– 41:??, October 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2013:EWD

[WHXZ13] Guanying Wu, Xubin He, Ningde Xie, and Tong Zhang. Exploiting workload dynamics to improve SSD read latency via differentiated error correction codes. ACM Transactions on Design Automation of Electronic Systems, 18(4): 55:1–55:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2019:RTS

[WJG^{+}19] Kankan Wang, Xu Jiang, Nan Guan, Di Liu, Weichen Liu, and Qingxu Deng. Real-time scheduling of DAG tasks with arbitrary deadlines. ACM Transactions on Design Automation of Electronic Systems, 24(6):66:1–66:22, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3358603.

Witharana:2024:ICT

[WJM24] Hasini Witharana, Aruna Jayasena, and Prabhat Mishra. Incremental concolic testing of register-transfer level designs. ACM Transactions on Design Automation of Electronic Systems, 29(3):55:1–55:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3655621.

Wu:2007:EPM

[WJY^{+}07] Wei Wu, Lingling Jin, Jun Yang, Pu Liu, and Sheldon X.-D. Tan. Efficient power modeling and software thermal sensing for runtime temperature monitoring. ACM Transactions on Design Automation of Electronic Systems, 12(3): 26:1–26:??, August 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2011:ALR

[WJYZ11] Shaoxi Wang, Xinzhang Jia, Arthur B. Yeh, and Lihong Zhang. Analog layout retargeting using geometric programming. ACM Transactions on Design Automation of Electronic Systems, 16(4):50:1– 50:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Weng:2012:TOS

[WKC12] Shih-Hung Weng, Yu-Min Kuo, and Shih-Chieh Chang. Timing optimization in sequential circuit by exploiting clockgating logic. ACM Transactions on Design Automation of Electronic Systems, 17(2):

16:1–16:??, April 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2018:ACS

[WKL^{+}18] Shao-Chung Wang, Li-Chen Kan, Chao-Lin Lee, Yuan-Shin Hwang, and Jenq-Kuen Lee. Architecture and compiler support for GPUs using energy-efficient affine register files. ACM Transactions on Design Automation of Electronic Systems, 23(2):18:1–18:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wolinski:2009:ADA

[WKR09] Christophe Wolinski, Krzysztof Kuchcinski, and Erwan Raffin. Automatic design of application-specific reconfigurable processor extensions with UPaK synthesis kernel. ACM Transactions on Design Automation of Electronic Systems, 15(1):1:1–1:??, December 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2012:TAS

[WL12] Chin-Hsien Wu and Hsin-Hung Lin. Timing analysis of system initialization and crash recovery for a segment-based flash translation layer. ACM Transactions on Design Automation of Electronic Systems, 17(2): 14:1–14:??, April 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2002:PDP

[WLC02] Guang-Ming Wu, Jai-Ming Lin, and Yao-Wen Chang. Performance-driven placement for dynamically reconfigurable FPGAs. ACM Transactions on Design Automation of Electronic Systems, 7(4): 628–642, October 2002. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2009:PCV

[WLCJ09] Meng-Chen Wu, Ming-Ching Lu, Hung-Ming Chen, and Jing-Yang Jou. Performanceconstrained voltage assignment in multiple supply voltage SoC floorplanning. ACM Transactions on Design Automation of Electronic Systems, 15(1):3:1– 3:??, December 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2020:HER

[WLH20] Hongfei Wang, Jianwen Li, and Kun He. Hierarchical ensemble reduction and learning for resource-constrained computing. ACM Transactions on Design Automation of Electronic Systems, 25(1):12:1– 12:21, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3365224.

Wang:2011:OAE

[WLL^{+}11] Yi Wang, Hui Liu, Duo Liu, Zhiwei Qin, Zili Shao, and Edwin H.-M. Sha. Overheadaware energy optimization for real-time streaming applications on multiprocessor System-on-Chip. ACM Transactions on Design Automation of Electronic Systems, 16(2): 14:1–14:??, March 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2016:PAC

[WLLH16] Po-Hsun Wu, Mark Po-Hung Lin, Xin Li, and Tsung-Yi Ho. Parasitic-aware commoncentroid FinFET placement and routing for current-ratio matching. ACM Transactions on Design Automation of Electronic Systems, 21(3): 39:1–39:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Witharana:2021:DTG

[WLM21] Hasini Witharana, Yangdi Lyu, and Prabhat Mishra. Directed test generation for activation of security assertions in RTL models. ACM Transactions on Design Automation of Electronic Systems, 26(4):26:1– 26:28, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3441297.

Wu:2008:CPR

[WLT08] Meng-Chiou Wu, Rung-Bin Lin, and Shih-Cheng Tsai. Chip placement in a reticle for multiple-project wafer fabrication. ACM Transactions on Design Automation of Electronic Systems, 13(1): 22:1–22:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wan:2019:DRP

[WLZ^{+}19] Bo Wan, Xi Li, Bo Zhang, Caixu Zhao, Xianglan Chen, Chao Wang, and Xuehai Zhou. DCW: a reactive and predictable programming framework for LET-based distributed real-time systems. ACM Transactions on Design Automation of Electronic Systems, 24(3):37:1–37:35, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3317574.

Weerasena:2024:SEO

[WM24] Hansika Weerasena and Prabhat Mishra. Security of electrical, optical, and wireless onchip interconnects: a survey. ACM Transactions on Design Automation of Electronic Systems, 29(2):40:1–40:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3631117.

Wang:2024:SSG

[WML^{+}24] Bo Wang, Sheng Ma, Shengbai Luo, Lizhou Wu, Jianmin Zhang, Chunyuan Zhang, and Tiejun Li. SparGD: a sparse GEMM accelerator with dynamic dataflow. ACM Transactions on Design Automation of Electronic Systems, 29(2):26:1– 26:??, March 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3634703.

Wang:2016:HDT

[WMT^{+}16] Hai Wang, Jian Ma, Sheldon X.-D. Tan, Chi Zhang, He Tang, Keheng Huang, and Zhenghong Zhang. Hierarchical dynamic thermal management method for highperformance many-core microprocessors. ACM Transactions on Design Automation of Electronic Systems, 22(1): 1:1–1:??, December 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wolf:1996:OOC

[Wol96] Wayne Wolf. Object-oriented cosynthesis of distributed embedded systems. ACM Transactions on Design Automation of Electronic Systems, 1(3): 301–314, July 1996. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL http://www.acm.org/
pubs/articles/journals/todaes/
1996-1-3/p301-wolf/p301-

                                        wolf.pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        todaes/1996-1-3/p301-wolf/

.

Wang:2008:LAS

[WPHL08] Sying-Jyan Wang, Kuo-Lin Peng, Kuang-Cyun Hsiao, and Katherine Shu-Min Li. Layoutaware scan chain reorder for launch-off-shift transition test coverage. ACM Transactions on Design Automation of Electronic Systems, 13(4):64:1– 64:??, September 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2023:IBS

[WPL23] Jun-Sheng Wu, Chi-An Pan, and Yi-Yu Liu. ILP-based substrate routing with mismatched via dimension consideration for wire-bonding FBGA package design. ACM Transactions on Design Automation of Electronic Systems, 28(5): 77:1–77:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3579843.

Wang:2019:LRA

[WPR^{+}19] Naixing Wang, Irith Pomeranz, Sudhakar M. Reddy, Arani Sinha, and Srikanth Venkataraman. Layout resynthesis by applying design-for-manufacturability guidelines to avoid low-coverage areas of a cell-based design. ACM Transactions on Design Automation of Electronic Systems, 24(4): 42:1–42:19, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3325066.

Wang:2016:ADB

[WQC^{+}16] Yi Wang, Zhiwei Qin, Renhai Chen, Zili Shao, and Laurence T. Yang. An adaptive demand-based caching mechanism for NAND flash memory storage systems. ACM Transactions on Design Automation of Electronic Systems, 22 (1):18:1–18:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2022:LPE

[WS22] Zi Wang and Benjamin Carrion Schafer. Learning from the past: Efficient high-level synthesis design space exploration for FPGAs. ACM Transactions on Design Automation of Electronic Systems, 27(4): 34:1–34:23, July 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3495531.

Wurth:1999:FMO

[WSEA99] Bernd Wurth, Ulf Schlichtmann, Klaus Eckl, and Kurt J. Antreich. Functional multipleoutput decomposition with application to technology mapping for lookup table-based FP-GAs. ACM Transactions on Design Automation of Electronic Systems, 4(3):313–350, July 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1999-4-3/p313-wurth/p313-
wurth.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1999-4-3/p313-wurth/

.

Wang:2018:VAG

[WSH^{+}18] Chen Wang, Yanan Sun, Shiyan Hu, Li Jiang, and Weikang Qian. Variation-aware global placement for improving timing-yield of carbonnanotube field effect transistor circuit. ACM Transactions on Design Automation of Electronic Systems, 23(4): 44:1–44:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2016:OAW

[WSRH16] Yu-Wei Wu, Yiyu Shi, Sudip Roy, and Tsung-Yi Ho. Obstacleavoiding wind turbine placement for power loss and wake effect optimization. ACM Transactions on Design Automation of Electronic Systems, 22(1):5:1–5:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Winograd:2018:PGU

[WSS^{+}18] Ted Winograd, Gaurav Shenoy, Hassan Salmani, Hamid Mahmoodi, Setareh Rafatirad, and Houman Homayoun. Programmable gates using hybrid CMOS–STT design to prevent IC reverse engineering. ACM Transactions on Design Automation of Electronic Systems, 23(6):76:1–76:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wei:2014:TSE

[WSV^{+}14] Yaoguang Wei, Cliff Sze, Natarajan Viswanathan, Zhuo Li, Charles J. Alpert, Lakshmi Reddy, Andrew D. Huber, Gustavo E. Tellez, Douglas Keller, and Sachin S. Sapatnekar. Techniques for scalable and effective routability evaluation. ACM Transactions on Design Automation of Electronic Systems, 19(2): 17:1–17:??, March 2014. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2023:BIH

[WSY23] Zilu Wang, Xinming Shi, and Xin Yao. A brain-inspired hardware architecture for evolutionary algorithms based on memristive arrays. ACM Transactions on Design Automation of Electronic Systems, 28(5):82:1–82:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3598421.

Wang:2013:CTM

[WTL^{+}13] Hai Wang, Sheldon X.-D. Tan, Duo Li, Ashish Gupta, and Yuan Yuan. Composable thermal modeling and simulation for architecture-level thermal designs of multicore microprocessors. ACM Transactions on Design Automation of Electronic Systems, 18(2): 28:1–28:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2012:CMI

[WTR12] Hai Wang, Sheldon X.-D. Tan, and Ryan Rakib. Compact modeling of interconnect circuits over wide frequency band by adaptive complex-valued sampling method. ACM Transactions on Design Automation of Electronic Systems, 17(1): 5:1–5:??, January 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wei:2023:APP

[WTW^{+}23] Min Wei, Xingyu Tong, Yuan Wen, Jianli Chen, Jun Yu, Wenxing Zhu, and Yao-Wen Chang. Analytical placement with 3D Poisson’s equation and ADMM-based optimization for large-scale 2.5D heterogeneous FPGAs. ACM Transactions on Design Automation of Electronic Systems, 28(5): 70:1–70:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3582554.

Wu:2009:EER

[Wu09] Chin-Hsien Wu. An energyefficient I/O request mechanism for multi-bank flashmemory storage systems. ACM Transactions on Design Automation of Electronic Systems, 14(1):6:1–6:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2002:BSF

[WV02] Haibo Wang and Sarma B. K.

Vrudhula. Behavioral synthesis of field programmable analog array circuits. ACM Transactions on Design Automation of Electronic Systems, 7(4):563– 604, October 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:1999:PRP

[WVYG99] Qi Wang, Sarma B. K. Vrudhula, Gary Yeap, and Shantanu Ganguly. Power reduction and power-delay trade-offs using logic transformations. ACM Transactions on Design Automation of Electronic Systems, 4(1):97–121, January 1999. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1999-4-1/p97-wang/p97-wang.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        todaes/1999-4-1/p97-wang/.

Wu:2004:BBA

[WWC04] Lieh-Ming Wu, Kuochen Wang, and Chuang-Yi Chiu. A BNFbased automatic test program generator for compatible microprocessor verification. ACM Transactions on Design Automation of Electronic Systems, 9(1):105–132, January 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2018:MTI

[WWCT18] Shengcheng Wang, Ran Wang, Krishnendu Chakrabarty, and Mehdi B. Tahoori. Multicast testing of interposer-based 2.5D ICs: Test-architecture design and test scheduling. ACM Transactions on Design Automation of Electronic Systems, 23(3):35:1–35:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2012:ESF

[WWFT12] Meng-Huan Wu, Peng-Chih Wang, Cheng-Yang Fu, and Ren-Song Tsay. An extended SystemC framework for efficient HW/SW co-simulation. ACM Transactions on Design Automation of Electronic Systems, 17(2):11:1–11:??, April 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2008:PVA

[WWG08] Huaizhi Wu, Martin D. F.

Wong, and Wilsin Gosti. Postplacement voltage assignment under performance constraints. ACM Transactions on Design Automation of Electronic Systems, 13(3):46:1–46:??, July 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wu:2012:LST

[WWW^{+}12] Shianling Wu, Laung-Terng Wang, Xiaoqing Wen, Wen-Ben Jone, Michael S. Hsiao, Fangfang Li, James Chien-Mo Li, and Jiun-Lang Huang. Launch-on-shift test generation for testing scan designs containing synchronous and asynchronous clock domains. ACM Transactions on Design Automation of Electronic Systems, 17(4):48:1–48:??, October 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2019:RSE

[WXH^{+}19] Hai Wang, Tao Xiao, Darong Huang, Lang Zhang, Chi Zhang, He Tang, and Yuan Yuan. Runtime stress estimation for three-dimensional IC reliability management using artificial neural network. ACM Transactions on Design Automation of Electronic Systems, 24(6):69:1–69:29, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3363185.

Wang:2006:PDT

[WY06] Chi-Shong Wang and Chingwei Yeh. Performance-driven technology mapping with MSG partition and selective gate duplication. ACM Transactions on Design Automation of Electronic Systems, 11(4): 953–973, October 2006. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2010:CDF

[WYC10] Renshen Wang, Evangeline Young, and Chung-Kuan Cheng. Complexity of 3-D floorplans by analysis of graph cuboidal dual hardness. ACM Transactions on Design Automation of Electronic Systems, 15(4):33:1– 33:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2007:DIC

[WYIG07] Chao Wang, Zijiang Yang, Franjo Ivančić, and Aarti Gupta. Disjunctive image computation for software verification. ACM Transactions on Design Automation of Electronic Systems, 12(2):10:1– 10:??, April 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Wang:2023:TPI

[WZH^{+}23] Senling Wang, Xihong Zhou, Yoshinobu Higami, Hiroshi Takahashi, Hiroyuki Iwata, Yoichi Maeda, and Jun Matsushima. Test point insertion for multi-cycle power-on selftest. ACM Transactions on Design Automation of Electronic Systems, 28(3):46:1–46:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3563552.

Wu:2021:DHC

[WZL^{+}21] Chin-Hsien Wu, Hao-Wei Zhang, Chia-Wei Liu, Ta-Ching Yu, and Chi-Yen Yang. A dynamic Huffman coding method for reliable TLC NAND flash memory. ACM Transactions on Design Automation of Electronic Systems, 26(5):34:1–34:25, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3446771.

Xama:2020:MLB

[XAG^{+}20] Nektar Xama, Martin Andraud, Jhon Gomez, Baris Esen, Wim Dobbelaere, Ronny Vanhooren, Anthony Coyette, and Georges Gielen. Machine learning-based defect coverage boosting of analog circuits under measurement variations.

ACM Transactions on Design Automation of Electronic Systems, 25(5):47:1–47:27, October 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3408063.

Xiang:2018:FTU

[XCF18] Dong Xiang, Krishnendu Chakrabarty, and Hideo Fujiwara. Faulttolerant unicast-based multicast for reliable network-on-chip testing. ACM Transactions on Design Automation of Electronic Systems, 23(6):73:1– 73:??, December 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xiang:2012:SFF

[XCW12] Dong Xiang, Zhen Chen, and Laung-Terng Wang. Scan flipflop grouping to compress test data and compact test responses for launch-on-capture delay testing. ACM Transactions on Design Automation of Electronic Systems, 17(2): 18:1–18:??, April 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xiao:2016:HTL

[XFJ^{+}16] K. Xiao, D. Forte, Y. Jin, R. Karri, S. Bhunia, and M. Tehranipoor. Hardware Trojans: Lessons learned after one decade of research. ACM Transactions on Design Automation of Electronic Systems, 22(1):6:1–6:??, December 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2020:MCM

[XGC^{+}20] Qi Xu, Hao Geng, Song Chen, Bei Yu, and Feng Wu. Memristive crossbar mapping for neuromorphic computing systems on 3D IC. ACM Transactions on Design Automation of Electronic Systems, 25(1):8:1– 8:19, January 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3365576.

Xiao:2024:IRI

[XGWL24] Jie Xiao, Yingying Ge, Ru Wang, and Jungang Lou. ICP-RL: Identifying critical paths for fault diagnosis using reinforcement learning. ACM Transactions on Design Automation of Electronic Systems, 29(1):11:1– 11:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3610294.

Xiang:2024:TCL

[Xia24] Dong Xiang. Test compression for launch-on-capture transition fault testing. ACM Transactions on Design Automation of Electronic Systems, 29(1):7:1–7:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3597433.

Xing:2023:BPB

[XJF^{+}23] Wei W. Xing, Xiang Jin, Tian Feng, Dan Niu, Weisheng Zhao, and Zhou Jin. BoA-PTA: a Bayesian Optimization Accelerated PTA solver for SPICE simulation. ACM Transactions on Design Automation of Electronic Systems, 28(2):27:1– 27:??, March 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3555805.

Xu:1997:LDR

[XK97] Min Xu and Fadi J. Kurdahi. Layout-driven RTL binding techniques for high-level synthesis using accurate estimators. ACM Transactions on Design Automation of Electronic Systems, 2(4):312–343, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        1997-2-4/p312-xu/p312-xu.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        todaes/1997-2-4/p312-xu/.

Xiang:2013:TCS

[XLCL13] Dong Xiang, Jianbo Li, Krishnendu Chakrabarty, and Xijiang Lin. Test compaction for small-delay defects using an effective path selection scheme. ACM Transactions on Design Automation of Electronic Systems, 18(3):44:1–44:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2016:ACS

[XLL^{+}16] Chang Xu, Guojie Luo, Peixin Li, Yiyu Shi, and Iris Hui-Ru Jiang. Analytical clustering score with application to postplacement register clustering. ACM Transactions on Design Automation of Electronic Systems, 21(3):41:1–41:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xue:2017:FCT

[XLNB17] Yuankun Xue, Ji Li, Shahin Nazarian, and Paul Bogdan. Fundamental challenges toward making the IoT a reachable reality: a model-centric investigation. ACM Transactions on Design Automation of Electronic Systems, 22(3): 53:1–53:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2015:DCD

[XLS15] Tong Xu, Peng Li, and Savithri Sundareswaran. Decoupling capacitance design strategies for power delivery networks with power gating. ACM Transactions on Design Automation of Electronic Systems, 20(3): 38:1–38:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xie:2018:TER

[XLY^{+}18] Guoqi Xie, Zhetao Li, Na Yuan, Renfa Li, and Keqin Li. Toward effective reliability requirement assurance for automotive functional safety. ACM Transactions on Design Automation of Electronic Systems, 23(5): 65:1–65:??, October 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2015:ICF

[XNZ^{+}15] Cong Xu, Dimin Niu, Yang Zheng, Shimeng Yu, and Yuan Xie. Impact of cell failure on reliable cross-point resistive memory design. ACM Transactions on Design Automation of Electronic Systems, 20(4):63:1– 63:??, September 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xydis:2012:CLE

[XPSE12] Sotirios Xydis, Kiamal Pekmestzi, Dimitrios Soudris, and George Economakos. Compiler-in-the-loop exploration during datapath synthesis for higher quality delay-area trade-offs. ACM Transactions on Design Automation of Electronic Systems, 18(1):11:1–11:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xie:2021:DFM

[XPX^{+}21] Guoqi Xie, Hao Peng, Xiongren Xiao, Yao Liu, and Renfa Li. Design flow and methodology for dynamic and static energyconstrained scheduling framework in heterogeneous multicore embedded devices. ACM Transactions on Design Automation of Electronic Systems, 26(5):36:1–36:18, June 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3450448.

Xie:2018:ADI

[XPZ^{+}18] Mimi Xie, Chen Pan, Mengying Zhao, Yongpan Liu, Chun Jason Xue, and Jingtong Hu. Avoiding data inconsistency in energy harvesting powered embedded systems. ACM Transactions on Design Automation of Electronic Systems, 23(3): 38:1–38:??, April 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2019:ESC

[XRS^{+}19] Xiaolin Xu, Fahim Rahman, Bicky Shakya, Apostol Vassilev, Domenic Forte, and Mark Tehranipoor. Electronics supply chain integrity enabled by blockchain. ACM Transactions on Design Automation of Electronic Systems, 24(3):31:1– 31:25, June 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3315571.

Xiang:2016:NUB

[XS16] Dong Xiang and Kele Shen. A new unicast-based multicast scheme for network-on-chip router and interconnect testing. ACM Transactions on Design Automation of Electronic Systems, 21(2):24:1–24:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2016:HSL

[XT16] Yang Xu and Jürgen Teich. Hierarchical statistical leakage analysis and its application. ACM Transactions on Design Automation of Electronic Systems, 21(4):65:1– 65:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xiang:2005:AIP

[XTW05] Hua Xiang, Xiaoping Tang, and Martin D. F. Wong. An algorithm for integrated pin assignment and buffer planning. ACM Transactions on Design Automation of Electronic Systems, 10(3):561–572, July 2005. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xian:2024:WJP

[XXC^{+}24] Juming Xian, Yan Xing, Shuting Cai, Weijun Li, Xiaoming Xiong, and Zhengfa Hu.

WCPNet: Jointly predicting wirelength, congestion and power for FPGA using multitask learning. ACM Transactions on Design Automation of Electronic Systems, 29(3): 57:1–57:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3656170.

Xu:2016:PPA

[XYG^{+}16] Xiaoqing Xu, Bei Yu, Jhih-Rong Gao, Che-Lun Hsu, and David Z. Pan. PARR: Pinaccess planning and regular routing for self-aligned double patterning. ACM Transactions on Design Automation of Electronic Systems, 21(3): 42:1–42:??, July 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Xu:2009:STA

[XZC09] Qiang Xu, Yubin Zhang, and Krishnendu Chakrabarty. SOC test-architecture optimization for the testing of embedded cores and signal-integrity faults on core-external interconnects. ACM Transactions on Design Automation of Electronic Systems, 14(1):4:1–4:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2000:TLB

[Yan00]  Jin-Tai Yan.  Three-layer

bubble-sorting-based nonManhattan channel routing. ACM Transactions on Design Automation of Electronic Systems, 5(3):726–734, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p726-yan/p726-yan.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        todaes/2000-5-3/p726-yan/.

Yan:2008:TDO

[Yan08] Jin-Tai Yan. Timing-driven octilinear Steiner tree construction based on Steiner-point reassignment and path reconstruction. ACM Transactions on Design Automation of Electronic Systems, 13(2): 26:1–26:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2011:ICA

[Yan11] Jin-Tai Yan. IO connection assignment and RDL routing for flip-chip designs. ACM Transactions on Design Automation of Electronic Systems, 16(4): 47:1–47:??, October 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2016:PDA

[Yan16] Jin-Tai Yan. Performancedriven assignment of buffered I/O signals in area-I/O flipchip designs. ACM Transactions on Design Automation of Electronic Systems, 21(2): 21:1–21:??, January 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2017:LAE

[Yan17] Jin-Tai Yan. Layer assignment of escape buses with consecutive constraints in PCB designs. ACM Transactions on Design Automation of Electronic Systems, 22(3):45:1– 45:??, May 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2018:DCR

[Yan18]  Jin-Tai Yan.  Direction-

constrained rectangle escape routing. ACM Transactions on Design Automation of Electronic Systems, 23(3):34:1– 34:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2019:TSN

[Yan19] Jin-Tai Yan. Two-sided net untangling with internal detours for single-layer bus routing. ACM Transactions on Design Automation of Electronic Systems, 24(6):68:1–68:23, November 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3363184.

Yan:2020:SLO

[Yan20]  Jin-Tai Yan.  Single-layer

obstacle-aware substrate routing via iterative pin reassignment and wire assignment. ACM Transactions on Design Automation of Electronic Systems, 25(2):22:1–22:21, March 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3378162.

Yamin:2023:UAE

[YB23] Nuzhat Yamin and Ganapati Bhat. Uncertaintyaware energy harvest prediction and management for IoT devices. ACM Transactions on Design Automation of Electronic Systems, 28(5): 87:1–87:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3606372.

Yellu:2021:STA

[YBM^{+}21] Pruthvy Yellu, Landon Buell, Miguel Mark, Michel A. Kinsy, Dongpeng Xu, and Qiaoyan Yu. Security threat analyses and attack models for approximate computing systems: From hardware and micro-architecture perspectives. ACM Transactions on Design Automation of Electronic Systems, 26(4):32:1– 32:31, April 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3442380.

Yang:2018:UUE

[YBS^{+}18] Kun Yang, Ulbert Botero, Haoting Shen, Damon L. Woodard, Domenic Forte, and Mark M. Tehranipoor. UCR: an unclonable environmentally sensitive chipless RFID tag for protecting supply chain. ACM Transactions on Design Automation of Electronic Systems, 23(6):74:1–74:??, December 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yao:2003:FRC

[YCCG03] Bo Yao, Hongyu Chen, Chung-Kuan Cheng, and Ronald Graham. Floorplan representations: Complexity and connections. ACM Transactions on Design Automation of Electronic Systems, 8(1):55– 80, January 2003. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2000:ERC

[YCHT00] Cheng-Hsing Yang, Sao-Jie Chen, Jan-Ming Ho, and Chia-Chun Tsai. Efficient routability check algorithms for segmented channel routing. ACM Transactions on Design Automation of Electronic Systems, 5(3):735–747, January 2000. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/todaes/
                                        2000-5-3/p735-yang/p735-
                                        yang.pdf;   http://www.acm.
                                        org/pubs/citations/journals/
                                        todaes/2000-5-3/p735-yang/

.

Yang:2020:HSS

[YCL^{+}20] Yajun Yang, Zhang Chen, Yuan Liu, Tsung-Yi Ho, Yier Jin, and Pingqiang Zhou. How secure is split manufacturing in preventing hardware Trojan? ACM Transactions on Design Automation of Electronic Systems, 25(2):20:1– 20:23, March 2020. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/abs/10.1145/3378163.

Yang:2023:ATF

[YCL^{+}23] Chun-Chieh Yang, Yi-Ru Chen, Hui-Hsin Liao, Yuan-Ming Chang, and Jenq-Kuen Lee. Auto-tuning fixed-point precision with TVM on RISC-V packed SIMD extension. ACM Transactions on Design Automation of Electronic Systems, 28(3):33:1–33:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3569939.

Young:2016:PSS

[YD16] Evangeline Young and Azadeh Davoodi. Preface to special section on new physical design techniques for the next generation of integration technology. ACM Transactions on Design Automation of Electronic Systems, 21(3):36:1–36:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2017:CCS

[YFT17] Kun Yang, Domenic Forte, and Mark M. Tehranipoor. CDTA: a comprehensive solution for counterfeit detection, traceability, and authentication in the IoT supply chain. ACM Transactions on Design Automation of Electronic Systems, 22(3): 42:1–42:??, May 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2018:RRE

[YFT18] Kun Yang, Domenic Forte, and Mark Tehranipoor. ReSC: an RFID-Enabled solution for defending IoT supply chain. ACM Transactions on Design Automation of Electronic Systems, 23(3):29:1–29:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yuan:2010:HSP

[YGH^{+}10] Mingxuan Yuan, Zonghua Gu, Xiuqiang He, Xue Liu, and Lei Jiang. Hardware/software partitioning and pipelined scheduling on runtime reconfigurable FPGAs. ACM Transactions on Design Automation of Electronic Systems, 15(2):13:1– 13:??, February 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2004:FVE

[YGZ04] Jun Yang, Rajiv Gupta, and Chuanjun Zhang. Frequent value encoding for low power data buses. ACM Transactions on Design Automation of Electronic Systems, 9(3):354–384, July 2004. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yalcin:1997:EPC

[YH97] Hakan Yalcin and John P. Hayes. Event propagation conditions in circuit delay computation. ACM Transactions on Design Automation of Electronic Systems, 2(3):249–280, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-3/p249-yalcin/p249-
yalcin.pdf;   http://www.
acm.org/pubs/citations/journals/
todaes/1997-2-3/p249-yalcin/

.

Yu:2009:APG

[YHH09] Hao Yu, Joanna Ho, and Lei He. Allocating power ground vias in 3D ICs for simultaneous power and thermal integrity. ACM Transactions on Design Automation of Electronic Systems, 14(3):41:1–41:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

You:2007:CCP

[YHL07] Yi-Ping You, Chung-Wen Huang, and Jenq Kuen Lee. Compilation for compact powergating controls. ACM Transactions on Design Automation of Electronic Systems, 12(4):51:1– 51:??, September 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2011:MUT

[YHL^{+}11] Guihai Yan, Yinhe Han, Hui Liu, Xiaoyao Liang, and Xiaowei Li. MicroFix: Using timing interpolation and delay sensors for power reduction. ACM Transactions on Design Automation of Electronic Systems, 16(2):16:1–16:??, March 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2014:WLL

[YKCG14] Yoon Seok Yang, Reeshav Kumar, Gwan Choi, and Paul V. Gratz. WaveSync: Lowlatency source-synchronous bypass network-on-chip architecture. ACM Transactions on Design Automation of Electronic Systems, 19(4):34:1– 34:??, August 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

You:2006:CLP

[YLL06] Yi-Ping You, Chingren Lee, and Jenq Kuen Lee. Compilers for leakage power reduction. ACM Transactions on Design Automation of Electronic Systems, 11(1):147–164, January 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yoon:2013:ACC

[YLP^{+}13] Jonghee W. Yoon, Jongeun Lee, Sanghyun Park, Yongjoo Kim, Jinyong Lee, Yunheung Paek, and Doosan Cho. Architecture customization of onchip reconfigurable accelerators. ACM Transactions on Design Automation of Electronic Systems, 18(4):52:1–52:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2023:CCE

[YLY^{+}23] Dingcheng Yang, Haoyuan Li, Wenjian Yu, Yuanbo Guo, and Wenjie Liang. CNN-Cap: Effective convolutional neural network-based capacitance models for interconnect capacitance extraction. ACM Transactions on Design Automation of Electronic Systems, 28(4): 56:1–56:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3564931.

Yang:2017:ELD

[YLZ^{+}17] Yunfeng Yang, Wai-Shing Luk, Hai Zhou, David Z. Pan, Dian Zhou, Changhao Yan, and Xuan Zeng. An effective layout decomposition method for DSA with multiple patterning in contact-hole generation. ACM Transactions on Design Automation of Electronic Systems, 23(1):11:1–11:??, October 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yonga:2015:ABE

[YMB15] Franck Yonga, Michael Mefenza, and Christophe Bobda. ASPbased encoding model of architecture synthesis for smart cameras in distributed networks. ACM Transactions on Design Automation of Electronic Systems, 20(2):27:1– 27:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2013:RAG

[YMC^{+}13] Tan Yan, Qiang Ma, Scott Chilstedt, Martin D. F. Wong, and Deming Chen. A routing algorithm for graphene nanoribbon circuit. ACM Transactions on Design Automation of Electronic Systems, 18(4):61:1–61:??, October 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yu:2010:EPE

[YP10] Chenjie Yu and Peter Petrov. Energy- and performanceefficient communication framework for embedded MPSoCs through application-driven release consistency. ACM Transactions on Design Automation of Electronic Systems, 16(1): 8:1–8:??, November 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2017:EEE

[YPCF17] Kaige Yan, Lu Peng, Mingsong Chen, and Xin Fu. Exploring energy-efficient cache design in emerging mobile platforms. ACM Transactions on Design Automation of Electronic Systems, 22(4):58:1–58:??, July 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yu:2011:MQS

[YRH11] Yue Yu, Shangping Ren, and Xiaobo Sharon Hu. A metric for quantifying similarity between timing constraint sets in real-time systems. ACM Transactions on Design Automation of Electronic Systems, 16(3): 34:1–34:??, June 2011. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2018:HEP

[YSF^{+}18] Kun Yang, Haoting Shen, Domenic Forte, Swarup Bhunia, and Mark Tehranipoor.

Hardware-enabled pharmaceutical supply chain security. ACM Transactions on Design Automation of Electronic Systems, 23(2):23:1–23:??, January 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:1997:HFM

[YTHC97] Cheng-Hsing Yang, Chia-Chun Tsai, Jan-Ming Ho, and Sao-Jie Chen. Hmap: a fast mapper for EPGAs using extended GBDD hash tables. ACM Transactions on Design Automation of Electronic Systems, 2(2):135–150, January 1997. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL

http://www.acm.org/pubs/
articles/journals/todaes/
1997-2-2/p135-yang/p135-
yang.pdf;   http://www.acm.
org/pubs/citations/journals/
todaes/1997-2-2/p135-yang/

.

Yan:2014:EFG

[YVC14] Jackey Z. Yan, Natarajan Viswanathan, and Chris Chu. An effective floorplan-guided placement algorithm for largescale mixed-size designs. ACM Transactions on Design Automation of Electronic Systems, 19(3):29:1–29:??, June 2014. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yan:2009:TAS

[YW09] Tan Yan and Martin D. F.

Wong. Theories and algorithms on single-detour routing for untangling twisted bus. ACM Transactions on Design Automation of Electronic Systems, 14(3):46:1–46:??, May 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2009:MCS

[YWGI09] Zijiang Yang, Chao Wang, Aarti Gupta, and Franjo Ivanvčić. Model checking sequential software programs via mixed symbolic analysis. ACM Transactions on Design Automation of Electronic Systems, 14(1):10:1–10:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2003:CRD

[YWK^{+}03] X. Yang, M. Wang, R. Kastner, S. Ghiasi, and M. Sarrafzadeh. Congestion reduction during placement with provably good approximation bound. ACM Transactions on Design Automation of Electronic Systems, 8(3):316–333, July 2003. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yu:2010:PSA

[YWW10] Cheng-Juei Yu, Yi-Hsin Wu, and Sheng-De Wang. An inplace search algorithm for the resource constrained scheduling problem during high-level synthesis. ACM Transactions on Design Automation of Electronic Systems, 15(4):29:1– 29:??, September 2010. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yang:2024:FEA

[YXG^{+}24] Bo Yang, Qi Xu, Hao Geng, Song Chen, Bei Yu, and Yi Kang. Floorplanning with edge-aware graph attention network and hindsight experience replay. ACM Transactions on Design Automation of Electronic Systems, 29(3): 56:1–56:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3653453.

Yuh:2007:TFU

[YYC07] Ping-Hung Yuh, Chia-Lin Yang, and Yao-Wen Chang. Temporal floorplanning using the three-dimensional transitive closure subGraph. ACM Transactions on Design Automation of Electronic Systems, 12(4):37:1–37:??, September 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yuh:2009:TTB

[YYC09] Ping-Hung Yuh, Chia-Lin Yang, and Yao-Wen Chang. T-trees: a tree-based representation for temporal and three-dimensional floorplanning. ACM Transactions on Design Automation of Electronic Systems, 14(4):51:1– 51:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yu:2016:EOA

[YYG^{+}16] Bei Yu, Kun Yuan, Jhih-Rong Gao, Shiyan Hu, and David Z. Pan. EBL overlapping aware stencil planning for MCC system. ACM Transactions on Design Automation of Electronic Systems, 21(3):43:1–43:??, July 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yi:2015:ESF

[YYL^{+}15] Qiuping Yi, Zijiang Yang, Jian Liu, Chen Zhao, and Chao Wang. Explaining software failures by cascade fault localization. ACM Transactions on Design Automation of Electronic Systems, 20(3): 41:1–41:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Yuh:2009:LAT

[YYLL09] Ping-Hung Yuh, Chia-Lin Yang, Chi-Feng Li, and Chung-Hsiang Lin. Leakage-aware task scheduling for partially dynamically reconfigurable FPGAs. ACM Transactions on Design Automation of Electronic Systems, 14(4):52:1–52:??, August 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2017:OIM

[ZABGZ17] Qingling Zhao, Zaid Al-Bayati, Zonghua Gu, and Haibo Zeng. Optimized implementation of multirate mixedcriticality synchronous reactive models. ACM Transactions on Design Automation of Electronic Systems, 22(2): 23:1–23:??, March 2017. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zanini:2012:OTC

[ZAJ^{+}12] Francesco Zanini, David Atienza, Colin N. Jones, Luca Benini, and Giovanni De Micheli. Online thermal control methods for multiprocessor systems. ACM Transactions on Design Automation of Electronic Systems, 18(1):6:1–6:??, December 2012. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2013:SRE

[ZAZ13] Baoxian Zhao, Hakan Aydin, and Dakai Zhu. Shared recovery for energy efficiency and reliability enhancements in realtime applications with precedence constraints. ACM Transactions on Design Automation of Electronic Systems, 18(2): 23:1–23:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2023:SLR

[ZBG^{+}23] Shanglin Zhou, Mikhail A. Bragin, Deniz Gurevin, Lynn Pepin, Fei Miao, and Caiwen Ding. Surrogate Lagrangian relaxation: a path to retrainfree deep neural network pruning. ACM Transactions on Design Automation of Electronic Systems, 28(6):102:1–102:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3624476.

Zoni:2018:CSC

[ZBPF18] Davide Zoni, Alessandro Barenghi, Gerardo Pelosi, and William Fornaciari. A comprehensive side-channel information leakage analysis of an in-order RISC CPU microarchitecture. ACM Transactions on Design Automation of Electronic Systems, 23(5):57:1–57:??, October 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhu:2006:CZD

[ZCG06] Haikun Zhu, Chung-Kuan Cheng, and Ronald Graham. On the construction of zerodeficiency parallel prefix circuits with minimum depth. ACM Transactions on Design Automation of Electronic Systems, 11(2):387–409, April 2006. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2023:MSF

[ZF23] Zhiqiang Zhao and Zhuo Feng.

A multilevel spectral framework for scalable vectorless power/thermal integrity verification. ACM Transactions on Design Automation of Electronic Systems, 28(1):11:1– 11:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3529534.

Zhou:2022:QCT

[ZFL22] Xiangzhen Zhou, Yuan Feng, and Sanjiang Li. Quantum circuit transformation: a Monte Carlo tree search framework. ACM Transactions on Design Automation of Electronic Systems, 27(6):59:1–59:??, November 2022. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3514239.

Zeng:2011:LDP

[ZFLS11] Zhiyu Zeng, Zhuo Feng, Peng Li, and Vivek Sarin. Localitydriven parallel static analysis for power delivery networks. ACM Transactions on Design Automation of Electronic Systems, 16(3):28:1–28:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zheng:2023:BVD

[ZGB^{+}23] Su Zheng, Hao Geng, Chen Bai, Bei Yu, and Martin D. F.

Wong. Boosting VLSI design flow parameter tuning with random embedding and multiobjective trust-region Bayesian optimization. ACM Transactions on Design Automation of Electronic Systems, 28(5): 74:1–74:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3597931.

Zhao:2024:DPD

[ZGB^{+}24] Aidong Zhao, Tianchen Gu, Zhaori Bi, Fan Yang, Changhao Yan, Xuan Zeng, Zixiao Lin, Wenchuang Hu, and Dian Zhou. D^{3}PBO: Dynamic domain decomposition-based parallel Bayesian optimization for large-scale analog circuit sizing. ACM Transactions on Design Automation of Electronic Systems, 29(3):44:1–44:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3643811.

Zhao:2018:TSB

[ZHC^{+}18] Hengyang Zhao, Qi Hua, Hai-Bao Chen, Yaoyao Ye, Hai Wang, Sheldon X.-D. Tan, and Esteban Tlelo-Cuautle. Thermal-sensor-based occupancy detection for smart buildings using machinelearning methods. ACM Transactions on Design Automation of Electronic Systems, 23(4): 54:1–54:??, July 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2021:DPR

[ZHC^{+}21] Ying Zhang, Xinpeng Hong, Zhongsheng Chen, Zebo Peng, and Jianhui Jiang. A deterministic-path routing algorithm for tolerating many faults on very-large-scale networkon-chip. ACM Transactions on Design Automation of Electronic Systems, 26(1):8:1– 8:26, January 2021. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3414060.

Zhang:2023:SKR

[ZHC^{+}23] Ruisi Zhang, Shehzeen Hussain, Huili Chen, Mojan Javaheripi, and Farinaz Koushanfar. Systemization of knowledge: Robust deep learning using hardware–software codesign in centralized and federated settings. ACM Transactions on Design Automation of Electronic Systems, 28 (6):88:1–88:??, November 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3616868.

Zhou:2023:FAO

[ZHJ^{+}23] Yuhao Zhou, Zhenxue He, Jianhui Jiang, Jia Liu, Juncai He, Tao Wang, Limin Xiao, and Xiang Wang. Fast area optimization approach for XNOR/ OR-based fixed polarity Reed– Muller logic circuits based on multi-strategy wolf pack algorithm. ACM Transactions on Design Automation of Electronic Systems, 28(3):45:1– 45:??, May 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        doi/10.1145/3587818.

Zhang:2023:CCM

[ZHL^{+}23] Qing Zhang, Huajie Huang, Jizuo Li, Yuhang Zhang, and Yongfu Li. CmpCNN: CMP modeling with transfer learning CNN architecture. ACM Transactions on Design Automation of Electronic Systems, 28(4):58:1–58:??, July 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3569941.

Zamora:2007:SLP

[ZHM07] Nicholas H. Zamora, Xiaoping Hu, and Radu Marculescu. System-level performance/power analysis for platform-based design of multimedia applications. ACM Transactions on Design Automation of Electronic Systems, 12(1):2:1–2:??, January 2007. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2008:NER

[Zho08] Hai Zhou. A new efficient retiming algorithm derived by formal manipulation. ACM Transactions on Design Automation of Electronic Systems, 13(1):7:1–7:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zamora:2008:EMU

[ZHOM08] Nicholas H. Zamora, Xiaoping Hu, Umit Y. Ogras, and Radu Marculescu. Enabling multimedia using resource-constrained video processing techniques: a node-centric perspective. ACM Transactions on Design Automation of Electronic Systems, 13(1):18:1–18:??, January 2008. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhu:2009:ESA

[ZHTC09] Yi Zhu, Yuanfang Hu, Michael B.

Taylor, and Chung-Kuan Cheng. Energy and switch area optimizations for FPGA global routing architectures. ACM Transactions on Design Automation of Electronic Systems, 14(1):13:1–13:??, January 2009. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2015:RLP

[ZK15] Renyuan Zhang and Mineo Kaneko. Robust and low-power digitally programmable delay element designs employing neuron-MOS mechanism. ACM Transactions on Design Automation of Electronic Systems, 20(4):64:1–64:??, September 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2016:SRE

[ZKS^{+}16] Yue Zhao, Taeyoung Kim, Hosoon Shin, Sheldon X.-D. Tan, Xin Li, Haibao Chen, and Hai Wang. Statistical rareevent analysis and parameter guidance by elite learning sample selection. ACM Transactions on Design Automation of Electronic Systems, 21(4):56:1– 56:??, September 2016. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zheng:2019:HEB

[ZLG^{+}19] Jianwei Zheng, Chao Lu, Jiefeng Guo, Deming Chen, and Donghui Guo. A hardwareefficient block matching algorithm and its hardware design for variable block size motion estimation in ultrahigh-definition video encoding. ACM Transactions on Design Automation of Electronic Systems, 24(2):15:1– 15:??, March 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

                                        URL https://dl.acm.org/
                                        ft_gateway.cfm?id=3290408.

Zeng:2013:IPD

[ZLL13] Zhiyu Zeng, Suming Lai, and Peng Li. IC power delivery: Voltage regulation and conversion, system-level cooptimization and technology implications. ACM Transactions on Design Automation of Electronic Systems, 18(2): 29:1–29:??, March 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2016:CFS

[ZLL^{+}16] Daming Zhang, Shuangchen Li, Yongpan Liu, Xiaobo Sharon Hu, Xinyu He, Yining Zhang, Pei Zhang, and Huazhong Yang. A C2RTL framework supporting partition, parallelization, and FIFO sizing for streaming applications. ACM Transactions on Design Automation of Electronic Systems, 21(2):19:1–19:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2015:RBA

[ZLQ15] Jiliang Zhang, Yaping Lin, and Gang Qu. Reconfigurable binding against FPGA replay attacks. ACM Transactions on Design Automation of Electronic Systems, 20(2): 33:1–33:??, February 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2015:LRR

[ZLW^{+}15] Qi Zhang, Xuandong Li, Linzhang Wang, Tian Zhang, Yi Wang, and Zili Shao. Lazy-RTGC: a real-time lazy garbage collection mechanism with jointly optimizing average and worst performance for NAND flash memory storage systems. ACM Transactions on Design Automation of Electronic Systems, 20(3): 43:1–43:??, June 2015. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2015:LDP

[ZLY^{+}15] Ye Zhang, Wai-Shing Luk, Yunfeng Yang, Hai Zhou, Changhao Yan, David Z. Pan, and Xuan Zeng. Layout decomposition with pairwise coloring and adaptive multi-start for triple patterning lithography. ACM Transactions on Design Automation of Electronic Systems, 21(1):2:1–2:??, November 2015. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhu:2007:HMF

[ZM07] Xinping Zhu and Sharad Malik. A hierarchical modeling framework for on-chip communication architectures of multiprocessing SoCs. ACM Transactions on Design Automation of Electronic Systems, 12(1): 6:1–6:??, January 2007. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2024:EEE

[ZMLH24] Yunping Zhao, Sheng Ma, Hengzhu Liu, and Libo Huang. EPHA: an energy-efficient parallel hybrid architecture for ANNs and SNNs. ACM Transactions on Design Automation of Electronic Systems, 29(3): 43:1–43:??, May 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3643134.

Zuluaga:2016:SSN

[ZMP16] Marcela Zuluaga, Peter Milder, and Markus Püschel. Streaming sorting networks. ACM Transactions on Design Automation of Electronic Systems, 21(4):55:1–55:??, September 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zimmermann:2019:ADL

[ZMS^{+}19] Thomas Zimmermann, Mathias Mora, Sebastian Steinhorst, Daniel Mueller-Gritschneder, and Andreas Jossen. Analysis of dissipative losses in modular reconfigurable energy storage systems using SystemC TLM and SystemC-AMS. ACM Transactions on Design Automation of Electronic Systems, 24(4):40:1– 40:33, July 2019. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/abs/10.1145/3321387.

Zhao:2013:PSA

[ZMTC13] Wei Zhao, Junxia Ma, Mohammad Tehranipoor, and Sreejit Chakravarty. Power-safe application of tdf patterns to flip-chip designs during wafer test. ACM Transactions on Design Automation of Electronic Systems, 18(3):43:1–43:??, July 2013. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2008:HTC

[ZP08] Xiangrong Zhou and Peter Petrov. Heterogeneously tagged caches for low-power embedded systems with virtual memory support. ACM Transactions on Design Automation of Electronic Systems, 13(2): 32:1–32:??, April 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2023:MMR

[ZPLI23] Zhonghua Zhou, Yuxuan Pan, Guy G. F. Lemieux, and André Ivanov. MEDUSA: a multi-resolution machine learning congestion estimation method for 2D and 3D global routing. ACM Transactions on Design Automation of Electronic Systems, 28(5): 73:1–73:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3590768.

Zhao:2002:TMA

[ZS02] Min Zhao and Sachin S. Sapatnekar. Technology mapping algorithms for domino logic.

ACM Transactions on Design Automation of Electronic Systems, 7(2):306–335, April 2002. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2016:PPG

[ZS16] Le Zhang and Vivek Sarin.

Parallel power grid analysis based on enlarged partitions. ACM Transactions on Design Automation of Electronic Systems, 21(2):26:1–26:??, January 2016. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2018:RRD

[ZSY18] Hongxia Zhou, Chiu-Wing Sham, and Hailong Yao. Revisiting routability-driven placement for analog and mixedsignal circuits. ACM Transactions on Design Automation of Electronic Systems, 23(2): 17:1–17:??, January 2018. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2010:CSD

[ZSZ10] Yufu Zhang, Ankur Srivastava, and Mohamed Zahran. On-chip sensor-driven efficient thermal profile estimation algorithms. ACM Transactions on Design Automation of Electronic Systems, 15(3):25:1–25:??, May 2010. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:1998:ORR

[ZW98] Hai Zhou and D. F. Wong. Optimal river routing with crosstalk constraints. ACM Transactions on Design Automation of Electronic Systems, 3(3):496–514, July 1998. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (elec-

tronic). URL http://www.
acm.org/pubs/articles/journals/
todaes/1998-3-3/p496-zhou/
p496-zhou.pdf;   http://
www.acm.org/pubs/citations/
journals/todaes/1998-3-3/
p496-zhou/.

Zhu:2011:MPL

[ZWD11] Yuhao Zhu, Bo Wang, and Yangdong Deng. Massively parallel logic simulation with GPUs. ACM Transactions on Design Automation of Electronic Systems, 16(3):29:1– 29:??, June 2011. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zeng:2023:AMM

[ZXC^{+}23] Junwei Zeng, Nuo Xu, Yabo Chen, Chenglong Huang, Zhiwei Li, and Liang Fang. AIMCU-MESO: an in-memory computing unit constructed by MESO device. ACM Transactions on Design Automation of Electronic Systems, 28(1):8:1– 8:??, January 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

URL https://dl.acm.org/
doi/10.1145/3539575.

Zhou:2008:AAS

[ZYDP08] Xiangrong Zhou, Chenjie Yu, Alokika Dash, and Peter Petrov. Application-aware snoop filtering for low-power cache coherence in embedded multiprocessors. ACM Transactions on Design Automation of Electronic Systems, 13(1): 16:1–16:??, January 2008. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2009:TAR

[ZYP09] Xiangrong Zhou, Chenjie Yu, and Peter Petrov. Temperatureaware register reallocation for register file power-density minimization. ACM Transactions on Design Automation of Electronic Systems, 14(2): 26:1–26:??, March 2009. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhu:2017:CCA

[ZYPC17] Di Zhu, Siyu Yue, Massoud Pedram, and Lizhong Chen. CALM: Contentionaware latency-minimal application mapping for flattened butterfly on-chip networks. ACM Transactions on Design Automation of Electronic Systems, 22(2):21:1–21:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhou:2012:ONC

[ZYS12] Pingqiang Zhou, Ping-Hung Yuh, and Sachin S. Sapatnekar. Optimized 3D networkon-chip design using simulated allocation. ACM Transactions on Design Automation of Electronic Systems, 17(2): 12:1–12:??, April 2012. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhai:2018:ENG

[ZYW^{+}18] Jinyuan Zhai, Changhao Yan, Sheng-Guo Wang, Dian Zhou, Hai Zhou, and Xuan Zeng. An efficient non-Gaussian sampling method for high sigma SRAM yield analysis. ACM Transactions on Design Automation of Electronic Systems, 23(3):36:1–36:??, April 2018. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhao:2013:CSL

[ZYZ^{+}13] Bo Zhao, Jun Yang, Youtao Zhang, Yiran Chen, and Hai Li. Common-source-line array: an area efficient memory architecture for bipolar nonvolatile devices. ACM Transactions on Design Automation of Electronic Systems, 18(4): 57:1–57:??, October 2013. CO-DEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhang:2024:HTS

[ZZ24] Yuan Zhang and Jiliang Zhang.

A high throughput STR-based TRNG by jitter precise quantization superposing. ACM Transactions on Design Automation of Electronic Systems, 29(1):6:1–6:??, January 2024. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3606373.

Zhang:2017:RTV

[ZZCY17] Xianwei Zhang, Youtao Zhang, Bruce R. Childers, and Jun Yang. On the restore time variations of future DRAM memory. ACM Transactions on Design Automation of Electronic Systems, 22(2):26:1– 26:??, March 2017. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic).

Zhu:2023:DSE

[ZZL^{+}23] Binwu Zhu, Xinyun Zhang, Yibo Lin, Bei Yu, and Martin Wong. DRC-SG 2.0: Efficient design rule checking script generation via key information extraction. ACM Transactions on Design Automation of Electronic Systems, 28(5):80:1–80:??, September 2023. CODEN ATASFO. ISSN 1084-4309 (print), 1557-7309 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3594666.